A Bibliography of Publications of Alan Mathison Turing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

06 February 2025 Version 1.273

Abstract This bibliography records publications of Alan Mathison Turing (1912–

1954).

Title word cross-reference

$1 [Fis15, CAC14b]. 1 [PSS11, WWG12]. $139.99 [Ano20]. $16.95 [Sal12]. $16.96 [Kru05]. $17.95 [Hai16]. $19.99 [Jon17]. 2 [Fai10b]. $21.95 [Sal12]. $22.50 [LH83]. $24.00/$34 [Kru05]. $24.95 [Sal12, Ano04, Kru05]. $25.95 [KP02]. $26.95 [Kru05]. $29.95 [Ano20, CK12b]. 3 [Ano11c]. $54.00 [Kru05]. $69.95 [Kru05]. $75.00 [Jon17, Kru05]. $9.95 [CK02]. \frac{1}{2} [Sha14]. H [Wri16]. λ [Tur37a]. λ − K [Tur37c]. M [Wri16]. O(z) [Fef95]. p [Tur37c]. × [Jon17].

-computably [Fai10b]. -conversion [Tur37c]. -D [WWG12]. -definability [Tur37a]. -function [Tur37c].

. [Nic17]. Życie [Hod02b].

0-19-825079-7 [Hod06a]. 0-19-825080-0 [Hod06a]. 0-19-853741-7 [Rus89].

1 [Ano12g]. 1-84046-250-7 [CK02]. 100 [Ano20, FB17, Gin19]. 10011-4211 [Kru05]. 10th [Ano51]. 11th [Ano51]. 12th [Ano51]. 1942 [Tur42b]. 1945 [TDCKW84]. 1947 [CV13b, Tur47, Tur95a]. 1949 [Ano49, TNPY49]. 1950s [Ell19]. 1951 [Ano51]. 1988 [Man90]. 1995 [Fef99, Sha00].

2 [DH10]. 2.0 [Wat12o]. 20 [CV13b]. 2001 [Don01a]. 2002 [Wel02]. 2003 [Kov03]. 2004 [Pip04]. 2005 [Bro05]. 2006 [Mai06, Mai07]. 2008 [Wil10]. 2011 [Str11]. 2012 [Gol12, Smi14]. 20th [Kru05]. 25th [TDCKW84]. 26 [DB04]. 27 [TNPY49]. 2952222 [CBB12].

3 [Ano20, Mar11c, Mar11d]. 320pp [Sal12]. 32nd [WTP^{+}06]. 38th [BFG^{+}12].

4 [Chr22b, Fai12, Mar11a]. 423pp [CK12b]. 432pp [Sal12].

5 [Cra10b, Man90]. 50-pound [Ano19d]. 505 [Boo52]. 50th [Fis17, Set17]. 53 [AH85]. 53/7/77 [AH85]. 55.00 [Rus89]. 5th [DIMV11].

60 [FOO71]. 632pp [Sal12]. 66 [Goo92]. 6th [DMV12].

7 [Sal12]. 77 [AH85].

8 [Dal12b, Gee12a]. 8th [CDL12].

9/11 [Kru05]. 93 [Ano88]. 978 [Ano20, Chr22b, Dal12b, Fai12, Gee12a, Sal12]. 978-0-19-871918-2 [Hai16]. 978-0-19-874782-6 [Jon17]. 978-0-19-874783-3 [Ano20, Jon17]. 978-0-19-960915-4 [Fai12]. 978-0-19-963979-3 [Sal12]. 978-0-262-03454-8 [Nic17]. 978-0-691-15564-7 [Sal12]. 978-0-691-15574-6 [Shi14, vL13]. 978-0-713-99750-7 [Sal12]. 978-1-3988-1244-4 [Chr22b]. 978-1-4000-7599-7 [Sal12]. 978-1-906124-90-8 [Dal12b, Gee12a]. 978-3-319-53278-3 [Ano20]. 9th [Ano51, ACL12].

A-5 [Chr22b]. A. [Bod49, Bri90, CD86, Fie15, Goo79b, Har47, Kid96, TDCKW84, Tur72, TWCD86, Tur01a, TB12]. A.C.E. [Ano46]. A.L.I.C.E. [Wal95, Wal09]. Abstract [DL06]. Abuse [Kru05]. Abyss [Ken17]. accelerated [PR10]. Accelerating [CS11b, Kur04]. access [KvLP88, Mai06]. accidentally [McG12]. account [DT12, Pap12]. ACE [AWL^{+}88, Tur45, Ano11a, Ano12d, CK17, CD86, Cop12a, Dow12b, Har47, TWCD86, Wil80, Ano13]. achievement [Jam06]. achievements [Hae12]. Acid [LE91]. ACM [Ano99, Ash87, Fis15, Owe12, Set17, CAC14b]. across [BSK^{+}15]. Active [BB16]. activity [Dav13, Ell13]. Actor [Hew13]. actress [Kah84]. Actually [Hai14]. Ad [Cha94]. Ada [Swa13b]. adaptivity [Sie13]. Add [Fra06]. Adding [Ano09a, Mai06]. additional [AH85]. Addressing [Day21]. adventure [Lom05]. African [CFK^{+}91]. After [Daw16, Hod04b, Mur12, Coo12b, CP00, Dav13, Gal06, Par14]. Again [Cas01, Res17]. Against [LA12, DB04]. Agar [CK02]. Age [Hal13, Kov03, MBC06, Pri21, Cop12b, Got96, Hal14, SG17, Bol84, Hod06a, Sal12, Bea84, Cer85, Hai16, Joh15, Sut85]. Aged [Sha14]. Agencies [Kru05]. Agent [Cas01]. Agnes [Bur11]. AI [SCT^{+}17, Cop23, Cop09, Cro94, Lev17, Yap12]. aid [PA11b]. al [CFK^{+}91, Cop23]. al-Khwarizmi [CFK^{+}91]. Alan [Ano99, Ano20, CK84, Chr16, Chr22a, Chr24, Coo06a, Cop25, Dys12a, Gin19, GKO95, Ham16, Hod12c, Hof85, Kru05, Lie11, Lip11, May61, MMB13, Swa13a, TDCKW84, AB00, AW77, AH85, Ano96, Ano00a, Ano00b, Ano09b, Ano12d, Ano12b, Ano12a, Ano12c, Ano12h, Ano13, Ano15a, Ano19a, Ano19b, Ano19c, Ano19d, Ano20, Ano21, App12, Asp80, AB12, AB14, Bar98, Bau12, Ben12, Blu14, Bre12a, Bre12c, Bro09, CK12a, Cap05, Cas01, Cas13, Che93, Chr10, Chr13, CM96, CS12, CBB12, Coo12b, Coo12c, Coo12d, CV13a, CvL13, CP96, CP99, Cop05a, CP12b, CGLWVR12, Cop12a, CL17b, Cop18, CP23, Cor07, Cow19, Dav13, Daw16, DC12, DC13, Don14, Dow13, Dys12a, Ell13, Ell19, FH15, FB17, Fre86, Fri05, GMC12, Gam13]. Alan [Gee12b, Ghe11, Gla01, Gla03, Gla04, GR12, Gla12, Gol12, Gon23b, GKO95, Got96, Gou99, GC12b, GC12a, GC12c, GC12d, GG13, Hae12, Har12a, Hel17, Hen11, Hid12, Hil93, Hil91, Hoc87, HG89, Hod83a, Hod83b, Hod85, Hod88, Hod89a, Hod89b, Hod92, Hod94a, Hod94b, Hod95a, Hod95b, Hod97a, Hod97b, HP00, Hod00, Hod01, Hod02a, Hod02b, Hod03a, Hod03b, Hod04a, Hod04b, Hod08a, Hod08b, Hod09, Hod12d, Hod12b, Hod12c, Hod14, Hou12, Hym12, Irv04, IM13, Jac12, Kah12, Kie12, LCKBJ12, Lea05, Lea07, Lea12, Lei01, Lem04, Lem12, Lie11, Liv02, Lol13, Lov04, Mac12a, Mac12b, Mar13b, MD11, Mei12a, Mic08, MC96, MJ84, Müh09, Nan03, Nau09, New55, New12, New03, Num05, OF03, O’N23, O’R12, Odi12, Pap12]. Alan [Par14, Pat04, Pat07, Pea19, Pet08, Pic03a, Pit14, Ran72a, Ran72b, Rob97, Sal04, Sal12, Sau93, Sev12, Sha14, Sie12, Sol87, Sor05, Str15, Swa13b, Swi19, Tau56, Ter11, Teu04a, Teu12, The87, THWV88, Tsa19, Tur42b, Tur59, Tur00, TP06, Tur12, Tur15b, Tur15a, Tur21c, Und13, Unk84, Vin13, Vos13, Web12, Wel12, Whi87, Whi91, Yan12, Zab95, Zab12, Zab17, Zas18, de 12, vL13, And08, Ano14, Asp84, Avi14, Chr15, Dal12b, Ers84, Fai12, Gee12a, Hof83, Kid06, Lav12, Lea19, LH83, Lov04, Rid84, Shi14, Shu87, Smi14]. Alana [Hod02b]. AlanTuring.net [CP01]. Algebraic [Cha95]. Algebras [HTG12]. ALGOL [FOO71, FOO71]. Algorithm [Cai12, BFP07, CWS^{+}24]. Algorithmic [DH10, Dow14a]. Algorithms [Gur95, Par17, SGV94]. Alignment [Don14]. alikes [BA05]. Alisa [Ano20, Gin19]. All-Against-All [LA12]. alle [Dys12a]. allegations [Irv04]. Allen [GC12e, Sal12, SCT^{+}17]. aller [GKO95]. Allgemeine [Tur60a]. Allied [Kah84]. Allies [AWL^{+}88]. Almost [Tau61b, Tur35a]. Always [OSZ03]. Am [Hod94c]. America [Kru05, DB04]. Americas [Kru05]. amplitude [Dut10]. analog [Cor17]. Analyses [WS00]. Analysis [Cuc12, KW12, Kle95, Tau63b, AB12, AB14, Blo98, CP10, DDL01, Ghe11, Mad12, Sie14]. Analyst [Wil71]. Anatomy [Wal95, Wal09]. ancestry [GC12e]. Andrew [Asp84, CK84, Hof83, LH83, Rid84, Sal12, Shi14, Shu87, vL13]. Anecdotes [SHH81, THWV88]. Anerkennung [Hod12c]. angle [Pro17c]. Anhang [Tur60a]. Animal [Mur12, Poo92]. Annals [Boo52]. Anniversary [CFK^{+}91, May01, TDCKW84, Fis17]. annotated [Lip11, Pet08, Wil10]. Annual [ACL12]. anticipation [CP96, CP23, Dow14a, Goo00]. Ants [HL02]. apology [Ano09b, Nau09]. appalling [Bro09]. Appel [Shi14, vL13]. appendix [Tur60a]. Apple [Pat04, Lem04, Lem12, Vin13, Cas16]. Application [Chu13, EH91, Tur36, Tur21a, The87, Tur37b]. Applications [ACL12, BAC14, Kle95, Kru05, Tur41a, Zab12, Zab17, DIMV11, DMV12, Soa16, Gas16]. Applied [BBST53, GGZ06]. approach [GAM11]. Approaches [DP02, BBLT06]. approximations [Tur38b]. Arbeiten [Hod12c, ST12]. Archimedes [Bra13]. architect [Got96, GW14]. Architecture [Mak95]. Archive [Ano10a, CP01]. Arcturus [Chr22b]. Arid [KW12]. Arise [FGG^{+}24]. Art [Gol12, GF91]. Article [Goo92, The87]. articles [FF63]. Artificial [Bry22, CP04, Cop04, Cop05b, Edm03, Fur12, Rou18, Wie12, Yan12, AB00, Bod17, Moo03b, Web12, FRT14]. artikkel [The87]. Artilect [DH09]. Artin [Boo06a]. Artistic [Mas12]. arvoitus [HP00]. Asimov [CFK^{+}91]. Aspects [FGG^{+}24, The87]. aspekter [The87]. Asperger [Jam06, OF03]. Aspray [CFK^{+}91]. assessment [de 12]. Association [Sof83]. astronomy [FF91]. Astrophysics [Tau63c]. Asymptotically [OSZ03]. Atanasoff [Ano96, Smi10]. attribute [EH91]. Auction [Ano15a]. Aufholjagd [Hod94d]. Australian [CFK^{+}91]. Author [Pit14]. authorship [Fie15]. Automata [Dow12a, IT12, Mar13a, Tau63b, Tur60a, DIMV11, DMV12, Sha09a, CFK^{+}91]. Automaten [Tur60a]. Automatic [And08, Ano49, AWL^{+}88, Jon04, Kid06, New49, Cop05a, Tur45]. Automatism [Eri03]. Automaton [MC12b, DDL01]. Autonomy [Cas01]. aux [Bia79]. av [The87]. Ave [Kru05]. Avenue [Kru05]. avtomatov [Tur60a]. Award [Ano99, Ash87, Fis17, Mic15, Var17, Ano14, Fie15, Fis15, Lip12, CAC14b]. awarded [Gee11]. Axes [Whi12].

B [And08, EW17, Fai12, Joh15]. B. [Hod06a, Sal12, TW12]. Babbage [OS65, SHH81, Das14, Jon16, Swa13b, Swa17, THWV88]. Baby [Cop11b, Cop17b]. Back [Res17, Coa13, Coo12d, Moo15]. Bacteria [Mar13a]. Bad [Pip04, Pip05, Hai17]. Ballesteros [Hid12, Hid12]. Balliett [Kru05]. Bamford [Kru05]. Banburisms [Sim17b]. Banburismus [Sim17a].

Bank [Pea19, Tsa19]. Banknote [Ano19a, Hum14]. Barrier [NA06, CS19]. barriers [BBLT06]. Barry [Chr15, Löw16]. Based [Cai12, Mar11a, EH91, Tur38c, Tur39, Tur65]. Basic [Kru05, Dav65]. Basis [Dys12a, Fre86, Nan03, Tur90, Tur52]. Basque [JTS97]. Baton [Chr24]. battle [SM07]. Bayes [EW17, Goo00, McG11, Sim17a]. Bayesian [Fie06, Fie06]. Bayley [TB12]. BBC [Jon04, Syk92]. BCS [Don01a]. Be [Cha16, Tsa19]. beaten [Hej07]. Beautiful [Gon24b, Vos13]. Beaver [Bra95]. Became [CL17b]. become [Fie06]. becomes [Ano19d]. Before [CFK^{+}91, RA04, RA03]. began [Das14]. Begegnung [GKO95]. Behavior [Dre10, Shi04]. Behaviorist [Wha09]. Behind [RA04, Rue07, Hod12c, RA03]. Beijing [ACL12]. Being [Pel09, Dav13]. belated [Ano09b]. Belief [PR17]. Berechenbarkeit [ST12]. Berners [Jor07]. Berners-Lee [Jor07]. Bernhardt [Nic17, Tho18]. berühmt [Hod12c]. Berührungspunkte [GR12]. Better [BBF03, Ack14, Wel02]. Between [Gla04, Dys12a, Emm13, GKO95, Kah84, LL12, Sen21]. Beyond [Has95, Hod12e, Kan12, MC12a, Pau91, Roc12, Sav24, Sie95, Bra13, Fre12c, SCT^{+}17, GKO95]. Bicentenary [CFK^{+}91]. Bifurcation [RMP11, Dil05]. Big [Wat12b, Coo12d, Str99]. biggest [Bie12, Cop17g]. Bilayer [DKM^{+}24]. bilinear [TT56]. Bill [Hou12]. binario [Hid12]. biochemical [GAM11]. Biografie [Ano12b]. biographer [McG12]. Biographies [Chr13, Wei88]. Biography [Hod04a, CFK^{+}91, Hod12b, Smi10, Tur15a, Ano12b]. bioinformatics [GMC12]. Biological [DP02, Mit12, Mei12b, SNUM03]. biologischen [Mei12b]. Biology [Mur93, Sau93, GMC12, HL02, Man90, Mis09, Zas18]. biomathematics [GMC12]. Biomedical [Mur12]. Biometrika [Goo92]. Birth [Ber16, Hod06a, Nic17, Tho18]. bis [Hod12c]. Bit [Cas06a, Hej07, Hid12]. Bletchley [Chr21, Sev12, Ano10a, Ano11b, Cop06a, Cop17a, Gee11, Goo79a, GMT^{+}12, GW14, HS93, Pri21, Sal04, Smi15b, Smi15a, Tur20]. blossoms [Han12]. Blue [Kru05]. Blueprint [Cas06a]. boat [DB04]. Body [Cla72, Hof83]. Bokulich [Ano20, Gin19]. Bold [Pic11]. Bolter [Bea84, Cer85, Sut85]. Bombe [CFK^{+}91, Chr22b, DK90, Gla03, Tur03, Tur21b, Dea98, Bur11, Car10, Tur04]. Bombes [CVC17]. Book [Ano04, Asp84, Bea84, Bla14, CK02, Cha94, Chr15, Coo06a, CV13c, Dal12b, Ers84, Fai12, Fef99, Gee12a, Gin19, GC12e, Hai16, Hay17, Hod06a, Hod06b, Hof83, Jon17, KP02, Lav12, Lea19, Lov04, Nic17, Nof17, OS65, Pet18, Rid84, Rus89, Shi14, Smi02, Tho18, Tur40, Wil10, vL13, Avi14, Hod14, KH19, Lip11, Par14, Smi14]. Books [CK02, CK12b, Día12, LH83, Por19, Smi14]. Boosts [Fis15]. Bot [Bie12]. boundaries [Fie15]. Bounded [Gur95]. Bowen [Ano20, Jon17, Pet18]. Box [EG12]. Boy [Sha14]. boyish [Bie12]. brackets [Tur42a]. Brain [Cop12a, Dal12a, Mur12, Rue07, Ano12j, JTS97, PA13, PA13, Fai12, Smi14]. Brains [IM13, RA03, RA04, Sie12]. Brake [Hic08]. breachable [CS19]. Break [NA06]. Breaker [Ano15a, Cow19, Hen11]. Breaking [Bat17, GMT^{+}12, Hil00b, LJWH97, RA03, RA04, Whi87, WH87a, WH87b, WH88a, WH88b, Rob17, Rob97]. breakthroughs [Haw05]. Brett [Kru05]. Bridge [BCT10]. bridging [Dys12a]. brief [CFK^{+}91]. Bright [Ken17]. Bringing [Cop09, Ste12b]. Bristow [Pit14]. Britain [Cha16, Cop23, Sum14]. British [Gee12a, Ano15a, Lav80, WCK89]. Broadcasts [Jon04]. Broadway [Kru05]. broke [Hea15, Lem04]. Broken [AWL^{+}88, Tur18, Mac12b]. Brooker [Met19]. Brown [Nau09]. Brückenschlag [Dys12a]. Buifendam [AWL^{+}88]. Build [And08, Cop12a, Kid06, Cop05a]. Building [Lav12, Len95, Len09, LCKBJ12, DB04]. Built [Lav80]. Bursts [ÇG12]. Burton [Gee12a]. Business [CFK^{+}91, Wat12b]. Busy [Bra95]. byte [Cle17, CFK^{+}91].

C [AWL^{+}88, PC88]. CA [CS11a, USE83]. Cabinet [Coa13]. cafe [Ive15]. calculability [Sie14]. Calculable [Kar95]. Calculating [Ano88, AWL^{+}88, CFK^{+}91, Jon04, Jon16]. calculation [Tur43]. calculations [Tur53]. Calculator [Tur46, Tur05b, AWL^{+}88]. Caldwell [Kru05]. Cambridge [Ano04, CFK^{+}91, CDL12, Kru05, Nic17, Flo17, HM02]. Campbell [Gee12a, CFK^{+}91]. Campbell-Kelly [Gee12a, CFK^{+}91]. Can [Coo06c, Den04, May52, NA06, Tur91, Wat95, Wat09, Wie12, Tur60a, TvN99]. Canada [Sof83, Kru05]. cancellation [Boo52, Tur50c]. canciones [Hid12]. Cantor [Lom05]. Capsule [TDCKW84]. Career [Pri24]. Carpenter [AWL^{+}88]. Cartesian [Eri03]. Case [TDCKW84, Zde03, Che93]. Cash [Tur01c]. castration [Dav13]. catalogue [AH85]. Catalytic [CP12a]. Catching [Hod94d]. Cathedral [Bla14, CK12b, Día12, Dys12c, GC12e, Sal12]. Cayley [dC11a]. Ce [Mar13b]. celebrate [Fis17]. Celebration [Owe12, Set17, Dic13, Hae12]. Cellular [Dow12a, LGB11, Mar13a, DDL01]. Centenary [CDL12, Ano12e, Ano12j, Bre12b, Coo12e, Dys12b, GMC12, Hae12, Rei12, Sal12, Wel12, BAC14, Owe12]. Central [Zab95]. centuries [McG11]. Century [Rus89, B^{+}11, GGZ06, Her88, Her95, Tay98, Wol17, MHR80]. certainty [GKO95]. Ceruzzi [TDCKW84]. chain [Tia11]. Challenge [SCT^{+}17]. Cham [Ano20]. changé [VB15]. changed [Haw05, VB15]. changing [Coo08]. Channel [CEL10, Whi12]. characters [KAB99]. Charles [OS65, Wil10, SHH81, Swa13b]. Chasing [Kru05]. Checking [Tur49]. Chemical [Fre86, HSD09, Nan03, Tur90, WS00, Dav13, McG12, Poo92, Tur52]. Chemotaxis [FGG^{+}24]. chess [CP17a, Lev88]. Child [Pro17a, Ste12b]. Child-Machine [Ste12b]. China [ACL12, BAC14, Set17]. Chlorite [LE91]. Chris [Gee12a, Nic17, Tho18]. Christof [Kru05, Lov04]. Christos [Ano04]. Church [AD12, BA05, CS19, Cot03, Dav06a, Deu85, Dow12a, Gal06, NT42, Par17, Pic11, Sie14, Tay98, Tim04, Tur42a, Yao03, Zie09]. CiE [BBLT06, CLS07, CDL12]. Cipher [AWL^{+}88]. ciphers [GMT^{+}12]. Circular [LA12]. circulation [Ano21]. claims [McG12, Par14]. Clarendon [Hod06a]. clash [WB12]. Classic [Lew21]. Classical [Kru05, Yao03]. Classics [Man90]. classification [Mon19]. Claude [SG17]. clock [Sut12]. Closed [LKE93]. Closing [Den12a]. cloth [KP02, Nic17]. Co [Fie15, Gla04]. Co-authorship [Fie15]. Co-operation [Gla04]. Co. [CFK^{+}91]. Coalgebraic [Jac11]. Coat [Mur12, BSK^{+}15]. Code [Ano15a, Cow19, GC12a, RA03, RA04, Whi87, WH87b, Bre12a, Bre12b, Ell19, Hen11, LJWH97, Mac12b, McG11, Moo14, Rob17, Rob97, SM07, WH87a, WH88a, WH88b]. Code-Making [GC12a]. Codebreaker [And08, Chr24, Dav13, Hil00a, Kid06, McG12, Bro13, Cop05a, Swa13a]. Codebreakers [Chr21, HS93, Tur20]. Codebreaking [GC12a, Cop06a]. codes [DB04, Hea15, Hil00b, WB12]. Coding [Joy00, OG12, Whi12]. Cognition [Har12a]. Cognitive [AWL^{+}88, Wel04, Wie12, Wel06b]. cold [Sen21]. Collaboration [Bro05, MJ09]. Collected [AWL^{+}88, Kid96, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Tur01a]. Collection [Ano19b, Ano19c, MHR80, FF63]. collections [Ano20]. colorful [KAB99]. Colors [BT12]. Colossal [Hai17]. COLOSSUS [Ran76, Cop06a, Pri21, Shi12]. Comes [MBC06]. Coming [Wat12l]. Commemorative [Gee12b]. Commentaries [AWL^{+}88]. Commentary [Luc95, Luc09, Zab12]. Comments [Tro93, Tro95, Wil71]. Common [Flo17, Lev17, FRT14]. common-sense [FRT14]. Communication [Che93]. communications [Kah84]. Commuting [TT56]. Companion [Chr16]. Company [KP02]. Comparison [LL12, WS00, Lie11, PC88]. compendium [Lev88]. Compiled [TB12]. Complete [CP12a, Pit23]. Complexity [Axe12, Ben95, MC12b, MD11, Mar11a, NW12, HS14, Ste90, Zie09]. compliance [Sum14]. composer [Ano12e]. Computability [AB12, AB14, BBLT06, Coo06b, CLS07, CDL12, Dow14a, Gas16, Kle95, Soa07, Tur37a, dLMSS56, Che93, CP10, Lip11, Pet08, Por19, ST12, SS91, Soa14, Soa16, SS15, Löw16, Nof17]. Computable [Chu13, Fai10a, FHM14, OG12, Tur36, Tur21a, Dav65, Ghe11, The87, Tur37b, Zen13, Coo08]. computably [Fai10b]. computadora [Lea12]. Computation [ACL12, Aho12, Ano49, AWL^{+}88, Bac12, Baj12, BAC14, Bee95, Buz12, Con12, Coo12a, Dah95, Den12a, DW12, Den12b, Den12c, DC11b, Dre10, DL06, EGW04, Fra12, Fre12a, Gel12, GC12b, Hew13, Jac11, Min67, Min72, Mit12, QSW11, Ros12, Sie95, Weg12, Ano04, Blu14, Mar11b, Pap03, Pos23a, Pos23b, Wel06b, Zen13, CLS07]. Computational [Aho12, CM10, DC12, Mar11a, Müh09, MJ09, Tra12, Wha09, Wie12, BBLT06, Coo08, DC13, HS14, The87, Zie09]. Computationalism [Sch02]. Computations [Fen95]. Compute [Coo06c, CS11b]. Computer [Ano51, Ano12e, Ano16, Bea84, Ber16, Bia79, BFG^{+}12, Bri90, CK02, Cer85, CP99, CP04, CP12b, Cop11a, Cop11b, Cop12a, CP17a, CL17c, CL17a, CH83a, CH83b, Cow19, Cuf24, Dav95a, Dav18, Deu85, Eps95, EBR09, Eps09, Eva81, Fly02, Gee12a, Har12b, Hod06a, KP02, Ken89, Kil14b, Lap96, LPAA22, Lev88, Lew21, Met19, Mic80, Nic17, Spr12, Sut85, Tho18, TDCKW84, Tur72, Wat12a, WTP^{+}06, WCK89, dSAL^{+}13, Aga01, Ano96, Ano13, Asp80, BB12a, BB12b, Bre12c, Bro97, BDD15, CK12b, Cop05a, CSS17, Cor17, Das14, Dav00, Dav12, Dew89, Dew93, DT12, Dys12a, Fie15, Goo84, Got96, Hai14, HP20, HH84, Hol90, HH90, Ire17, JTS97, Kil14a, Lea05, Lea07, Lea12, Lie11, dBPZM10, Shi12, Smi10, Smi05, Str99, Sum14, Tur50b, Tur51b, Bol84, BTHS12]. Computer [Dys12a, Spr12, Smi02, And08, Coo06a, Kid06]. computer-science [Bre12c]. Computerizing [Bee95]. Computers [BBST53, Bia79, Dav95b, Day13, DB05, Dys12a, FF63, Gon24a, Goo79a, IM13, Lav80, Lie11, NA06, Tau63b, Tim04, Wat12b, Wat12c, Cop06a, Cor17, Jac12, LCKBJ12, Ran72a, Ran72b, Ran17a, Sch04a, CFK^{+}91, Lav12]. Computes [CDL12]. Computing [And08, Blu04, Bow53a, Bra13, Bul15, CFK^{+}91, CH16, Cop04, Cop05a, Cop06b, Fef99, Hin17, Kid06, Kov03, MHR80, New49, O’N23, Par12, Ros12, Sha00, Swa13b, Ted15, Tur45, TNPY49, Tur50a, Tur95b, T^{+}06, Tur09, Wat12d, Wil97, Ano19d, CS11a, CP17b, Cop17h, Dys12b, Hen11, HP15, Jac12, LTM^{+}51, Mei12a, Mis09, Wat12m, Yan12, Zie09, CFK^{+}91, CP01, Cas13, Luc95, Luc09]. Concept [Ano19a, TDCKW84, Kan12, Pro17d]. conceptions [Coo08]. Concepts [CM96, Mak95, PR10]. concerning [Irv04]. Concise [Mar11b]. Concurrent [HP88b, HP88a]. Condemned [Cow19]. conditional [FRT14]. Conference [ACL12, Ano49, Ano51, BBLT06, BFG^{+}12, CS11a, CLS07, CDL12, DMV12, MBS11, Sof83, USE83, WTP^{+}06, DIMV11, Set17]. Conferences [WCK89]. conferenza [Odi12]. Configurations [Jea12]. Confirmed [Ano15b]. confirms [Irv04]. Conflicting [Zde03]. Confluence [Gan95]. Conjecture [Boo06a, Mon19]. Connecting [AS08b]. Connection [Bea89, Goo00]. Connectionism [CM96, CP17b, Teu04b, CP96, CP23]. Connections [AS08a]. Consciousness [Wat12f, Wat12m]. Consensus [Dav06a]. Consequence [Szu12]. Conservative [DL06]. Consideration [Fre12a]. Considered [Fre12a, HF95]. Constructibility [Edm03]. Construction [Wil80, Gon23a]. constructivity [Asp80]. contact [GR12]. Contemporaries [Lav12, LCKBJ12, Dal12b, Gee12a]. Contemporary [Kru05]. Contest [Loe95, Loe09, Bro97]. Context [Hod12f, Sch12a]. Continuous [Tra12, Tau62]. Contribution [EW17]. Contributions [Gla04]. Controlled [AWL^{+}88]. Controversia [And84]. controversy [McG11, And84]. conventions [Tur51a]. Convergence [RV12]. Conversation [Hut95, Hut09, LC01, TP06]. conversations [WS16]. conversion [Tur37c]. convicted [Dav13]. conviction [Ell13]. Cooper [Chr15, Löw16]. Copeland [Fai12, Hai16, Hod06a, Jon17, Pet18, Sal12, And08, Ano20, Joh15, TW12]. corneae [BSK^{+}15]. Corporation [Tur01c]. Corps [Hod94e]. Corrected [Cuf24]. Correction [CP23, Pos23a, Tro95, Tur37b]. Correspondence [AMKM66, BSPI65, HLOS65, AH85]. Coruña [DMV12]. Costs [ZM08]. could [BB12a, BB12b, Jac12]. Counterpoint [Shu87]. Counting [Kör96]. counts [Sch04a]. Coupling [FGG^{+}24]. course [Sha09a, Sor05]. Cover [Mau09]. Coverand [Cas06a]. CPU [Reg23]. cracked [Ell19, McG11, Moo14]. cracker [McG12]. Cream [Sch04b]. Created [Isa14, Lew21, Haw05, Tia11]. Creativity [BBF03]. Cribs [Gla03, Tur03]. Crime [Cha16, Cop17c]. Critique [Gla03, Tur03]. croqué [Lem04]. croquée [Lem12]. Cross [DKM^{+}24, Tia11]. Cross-Diffusion [DKM^{+}24, Tia11]. Crossley [CFK^{+}91]. crunched [Lem12]. cryptanalysis [Goo00]. Cryptography [Kru05, Tur41a, Zab12, Zab17, Blö12, Joy00]. Cryptologia [DKK^{+}98]. Cryptologic [Kru05]. Cryptology [Bau00, Chr10, Kah84, Kru05]. CSAC [AH85]. Cult [ZM08]. Culture [Bea84, Cer85, Sut85, Bol84, Emm13, Ive15, LC01]. Current [BFG^{+}12, WTP^{+}06]. cyberculture [TJC03]. Cybernetics [ER68]. Cybersecurity [Hel17]. cybertragicommedia [Pat07]. Cyberworld [Pip04, Pip05]. Cylons [Mar13c]. Czech [BFG^{+}12, WTP^{+}06].

D [TB12, Ano11c, WWG12]. Daisies [Swi04, GMC12]. damit [Dys12a]. data [DB05]. datafaglige [The87]. dauerte [Hod12c]. David [Bea84, Cer85, Fef99, Sut85]. Davidson [Sha12]. Davis [KP02, Kil14a, Kil14b]. Dawn [Day12a, Pri21, Wat12d, Dys12b, Wat12m]. Day [DW16]. Days [Bul15, Hus91]. Dayton [Gla01, Tur01c]. DC [Tur42b]. Deadheads [Wat12e]. Deal [Par12]. Death [CFK^{+}91, Moo14, Ter11, Hod02b, McG12, Syk92]. Deavours [CFK^{+}91]. Debate [DH09, Pro17b, Tro93, Tro95]. Debs [Smi15b]. Decidability [Whi17]. Deciphering [Bau00]. decoded [Tur15a, Chr16, Ham16]. Decomposability [Fai10b]. Defeat [Tur20]. Defeated [Chr22b, Tur21b]. defending [Sha12]. Defense [Bac90]. Defiance [Sum14]. Definability [CL02, Tur37a]. Defined [Edm03]. definition [HMRC88]. degrees [Fai10b, Mon19]. Delay [Hod94l]. Delays [LGB11]. DELILAH [TB12, Cop17d, Cop25]. delirio [Paz03]. delirium [Paz03, PC06]. dell’incomputabile [Cap05]. demasiado [Lea12]. demise [Ste90]. denies [Irv04]. denken [Tur60a]. dens [The87]. Dense [Fu12]. Dependent [RV12, AKS11]. Depth [Ben95]. depths [Sim17a]. Dermot [Chr16, Chr21, Chr22b, Ham16, Chr22a]. Deromanticizing [Day12b]. Descartes [Abr11]. Description [TB12, Nau93]. Deserts [She12, Smi14]. Design [Bro05, CG87, Hol86, HSD09, Liv02, HP20, HMRC88, LTS^{+}21, Tau63b]. desktop [GC12e]. deterministic [Wel06a]. DEUCE [Wil80]. Deutschland [GR12]. developed [Kah84]. Development [AWL^{+}88, Tur45, Tur72, Daw16, DT12, HS14, Mad12, Poo92, TDCKW84]. Developments [Ano88, AWL^{+}88, Dow14c, Dow14b]. Deviant [CP10]. Devil [Par12]. Dewdney [Bri90, Ken89]. Dial [Kru05]. Dialogue [GF91, GG12]. Did [CP00, Hod08b, OF03, Poo91, Poo92, BDD15, Fie06, Hai14, McG12]. didn’t [HP20, Par14]. Diego [USE83]. Dies [Met19]. difference [Sen21]. differential [Dow13]. différentielles [Dow13]. Diffusion [AWL^{+}88, BVE11, CEL10, DKM^{+}24, FGG^{+}24, KW12, Kon12, AKS11, Ano15b, CWS^{+}24, Dut10, LGS22, Tia11]. Digital [AWL^{+}88, BBST53, Bla14, Bow53a, Bul15, CK12b, Día12, Dys12c, Gon24a, Isa14, Mai07, Pri21, Swa13b, TDCKW84, Wat12f, Wat12g, GC12e, Ran72a, Ran72b, Ran17a, Sal12, Smi10, Wat12m, Cas16]. dígito [Hid12]. Dijkstra [Day12a]. Dimensional [Ano89, BVE11, IST^{+}10, UST^{+}10]. directions [Sch02]. disabled [Mai06]. disciplinary [Fie15]. Discipline [Ted15]. Discourse [Zde03]. Discover [Poo92, Poo91]. Discovery [dSAL^{+}13, AS08a, Mei12b, Rob12, Soa14]. Discrete [RMP11, LTS^{+}21]. Discussion [TNPY49, The87, CBB12]. Displacement [LP11]. Disputes [Gon24a]. dissertation [Tur38c]. distribution [Leh70]. Diverse [BSK^{+}15]. Division [Tur45]. DNA [CS11a, CS11a, QSW11]. Do [vEB12, CS11b]. doch [Hod12c]. Does [Fra06, Hut84]. Doing [Har12a, Las09, Las95]. Domains [LGB11]. Donald [CFK^{+}91]. Done [Hod04b]. Doran [AWL^{+}88]. Dotcom [Wat12h]. dots [Tur42a]. Doubleday [Kru05]. down [McG11]. Dr. [Gan54, Syk92]. Dream [Hin17]. dreamers [VB15]. dreams [Lev06]. drift [Mai06]. Driscoll [Bur11]. drøfting [The87]. DSLTrans [BLA^{+}11]. dubious [Hid12]. Dudley [THWV88]. dudoso [Hid12]. Dummies [vEB12]. During [RA04, Kah84]. Dustin [Kru05]. Dusting [Fre12b]. Dutch [Day13]. Duxford [CK02]. Dynamical [Del06]. Dynamics [LGB11, LGS22]. Dyson [CK12b, Día12, GC12e, Sal12].

E. [TDCKW84]. Early [Bul15, Cop23, Goo79a, Hus91, Lav80, MJ84, Par12, Pri24, WCK89, Ano46, Web12]. Easy [Har12a]. eboluzioaz [JTS97]. Eckert [Ano96]. eclectic [Odi12]. eclettico [Odi12]. Ecological [Wel04]. Economy [Don01a]. Ed [Kru05, Shi14, AWL^{+}88, Hod06a, Rus89, vL13]. edge [Hol18]. Edited [Ano04, And08, Chr15, Dal12b, Lov04, TW12]. edition [Sal12]. Editor [MMB13, EH91, CAC14a, Str65, Var14]. eds [Ano20, AWL^{+}88, Gin19, Nof17]. Education [Kru05]. effective [Sie14]. Efficient [AG11, QSW11]. Eight [Mah10, Cha94]. Eindhoven [MBS11]. einem [Tur60a]. einfache [FOO71]. Einführung [ST12]. Einstein [Hol18, Sen21, Sha14]. Elastic [Liv02]. Electric [New49]. Electrode [LOM^{+}01]. Electronic [Cop12a, Fai12, Tur46, Tur72, Tur05b, Smi14, Tur50b, Tur51b]. elimination [Grc11]. Elsevier [Avi14]. elusive [Moo03b]. Embedding [Edm95, Edm09]. Embeddings [OG12]. emerged [McG11]. Emergence [Coo06b, MJ09]. Emotions [Sha14]. emphasis [GMT^{+}12]. empirical [Goo00]. Emulated [Reg23]. encodings [CP10]. Encounter [Liv02]. Encounters [Bul21, Cra10a]. encrypting [Cop17d]. Encyclopedia [CFK^{+}91, CF98]. end [Ive15]. Enduring [For12]. Engine [And08, Kid06, Löw16, Cop05a, Tur45]. Engineering [Day12a, MBS11, Smi05]. Engineers [Kah84, Ano96]. England [Pea19, Tsa19]. enhancement [Mei12b]. ENIAC [TDCKW84]. Enigma [AWL^{+}88, Bro13, CK84, Chr24, Ell19, Hof85, Ran12, Shu87, Cap05, Hod83a, Hod83b, Hod85, Hod88, Hod89a, Hod89b, Hod92, HP00, Hod00, Hod01, Hod03b, Hod12d, Hod14, Sal12, Bur11, Cap05, Cas06b, CV13c, Cop04, Dav13, DB04, Goo00, Gre17, Hod94a, Hod94b, Hod02b, Hod03b, Joy00, Mah10, McG12, McG11, Par14, RA03, RA04, SM07, TDCKW84, Tur40, Tur99, Tur18, Tur21b, Unk84, Asp84, Chr22b, Hof83, LH83, Rid84, Ers84]. énigme [Hod88, Hod01]. Enjoying [Sch04b]. Enlightenment [Gör95a]. Enough [CFK^{+}91, DK90, KGJ^{+}24, Len95, Len09, RS03, Dea98]. Ensure [KGJ^{+}24]. Entdeckung [Mei12b]. Entertaining [Cra10a]. Entscheidungsproblem [The87, Tur36, Tur37b, Tur21a, Chu13, Whi17]. Entzifferte [Bau00]. enumerable [Fai10b]. Environment [EH91, KW12, PA11a, CG87]. epic [Rob12]. Epilogue [Hod94g]. Epistemology [Bea89]. Epstein [WWG12]. Equation [Cha95, Dut10]. equations [Dow13, Dow13]. Equivalence [Tur35a]. Erfinder [Lie11]. Ergodic [Tau61b]. Errata [Ano88]. Error [Tur35b, ZM08]. Errors [Bod49, Tur48c]. Ershov [CL02, Fai10a, Fai10b, Fai11]. Erzählung [Hoc87]. Esprit [Hod94e]. Essays [MHR80, Ano20]. Essential [Cop04, Hod06a, Hod06b]. Essentials [Rue07]. Estimation [OSZ03]. eta [JTS97]. Europe [BBLT06, CLS07, CDL12, Löw16]. every [Pro17c]. Everyday [Cra10a]. Evidence [RAM95]. Evolution [Weg12, JTS97]. evolutionary [Lei01, Yan12]. EVZI [Ste12a]. Exact [PSS11]. Examination [Cop00a]. Exclusion [Mai07]. Excursions [Bri90, Ken89, Dew89, Dew93, Hol18]. Exercise [BT12]. exhibition [Mac12b]. Exhibits [Ano02]. Expanded [Cuf24, Blu14]. Experiment [Gon22, Gon24b, WS00, Gon23a, Vos13]. Experimental [HSD09, RAM95, Dav13]. experiments [SNUM03]. Explained [Ano10b, Sha14]. explains [Sen21]. explanations [Lei01]. Explorations [Ano20, FB17, Gin19]. explores [Mac12b]. Exploring [GC12e, Zen13]. extensions [Tur38a]. Externalist [Sch12c].

F.R.S. [Gan54]. Face [Tsa19, Ano19d, Pea19]. Facing [Ran00]. Factor [EW17, Wri16]. Facts [Ano19d]. failure [Coa13]. Fallibility [Kah12]. Families [OG12, Fai10a]. famous [Cop17e, Hod12c]. fanden [Hod12c]. Farben [BT12]. Fast [KvLP88, SGV94]. Faster [Bow53a, Wel02]. Fastest [Cop12a]. Father [Ano19d, CFK^{+}91, CP12b, O’N23, Dys12a]. fatherhood [Cor17]. Faults [ÇG12, Ran00]. Fauvel [CFK^{+}91]. favour [Moo15]. feature [Ell19]. Features [Sch04b]. Featuring [Bul21]. February [CV13b, Tur47, Tur95a]. Feeling [Har12a]. Feigenbaum [TDCKW84]. Ferranti [Tur00]. Fibonacci [Swi04]. Fifteenth [CFK^{+}91]. Fifth [Kru05, TDCKW84]. Fiftieth [May01]. Fifty [Ano21]. Figuring [Cra10a]. Files [Ano02]. film [Hod14]. filter [Zas18]. Filtering [Cai12]. Filtering-Based [Cai12]. fingerprints [GAM11]. Finite [Fin95, Min67, Min72, Tur38b, Whi12]. Finnish [HP00]. First [Ano16, Chr10, Chr24, Day13, Lav12, Ash87, CP17a, Ire17, LCKBJ12, PA13, Tay98, Zas18]. fish [Cop17g, HM96, KA96]. Five [Jon04]. Flat [KW12]. flower [Han12]. Flowers [Hai17]. Floyd [Ano20, Gin19]. fly [Gub86]. focus [Por19]. Folk [CM96]. Following [Pro04]. fondamentale [Bia79]. food [Tia11]. Fools [DV09]. footsteps [Yan12]. foresaw [Sie12]. forest [Rob97]. Foreword [Teu12]. Forgotten [CP99]. Formal [Dow12a, NT42, Sha09a, WB12]. Formation [BB16, DKM^{+}24, FGG^{+}24, KW12, Kon12, KGJ^{+}24, RMP11, WS00, BGL24, Daw16, DDL01, GAM11, LGS22, Mei12b, Rei12, SNUM03]. Forming [Coo12a]. forms [Tur48b]. formula [GC17a]. Foundations [Kru05, Sch12c, Ghe11, Zie09]. four [IST^{+}10, Smi14, UST^{+}10, VB15, Kru05]. four-dimensional [IST^{+}10, UST^{+}10]. Fourier [DDL01]. fractal [dBPZM10]. fragility [Lei01]. frame [Cro94]. Free [LKE93, Pro17b, Llo12]. freedom [Lom05]. French [Ano96, Bia79, Bre12c, Dow13, Gou99, Hod88, Hod01, Lea07, Lem04, Lem12, Lom05, Mar13b, VB15]. Frenkel [CFK^{+}91]. fridge [Sen21]. Friendly [PA11a]. Frontiers [SG18]. FRS [AW77, AH85]. Fuller [CFK^{+}91]. Function [Tur35b, Leh70, Leh56, Tur37c, Tur43, Tur53]. Functions [Tau61b, Zde03, Dav65]. Fundamental [Bac12, Bia79]. funds [Gee11]. Fusion [MJ09]. Future [CH16, Lew21, Moo03a, Web12].

Galilean [Gon23a]. Game [Bra95, Cop05b, Reg23, Pro15, WS16, Cho09, Hod14, Las09, Las95, Lon09, Pic03b, Cho95]. Games [BBST53, ŁW11, Tau63c]. Gandy [Dah95]. gap [Dys12a]. Gardner [AWL^{+}88]. garuna [JTS97]. garunaren [JTS97]. Gaussian [Grc11, Tur35b]. Gave [Hod06a, Jac12]. Gay [Cha16, Mar13c, Ell13, Mac12a]. Gaylons [Mar13c]. Geeks [Isa14]. Geheimnisse [Bau00]. Geheimschreiber [Joy00]. Geist [Mei12a, Hod94f]. geistige [Dys12a]. gelang [Dys12a]. gene [Blo98]. genealogia [Cap05]. genealogy [Cap05]. General [CH83a, CH83b, NA06, Szu12, GMT^{+}12, LTM^{+}51, Tur60a]. generalizations [Nor14]. Generation [TDCKW84]. genesis [Das14]. génies [VB15, Ano12b]. genio [Rig91]. Genius [Ano12b, OS65, Phi65, Hai17, Hen11, Hil93, Hil17, Mac12b, Mic84, Par14, Rig91]. Geniuses [Isa14, Pri21, VB15]. Genome [Kov03]. Geometrical [DL06]. Geometry [DKM^{+}24, Lon09, Tau62]. George [CK12b, Día12, GC12e, Sal12]. Gerdes [CFK^{+}91]. German [AWL^{+}88, Ano12b, Bau00, BT12, Blö12, Bre12a, Bul21, Dys12a, FOO71, Fur12, GR12, Gla12, GKO95, Hil00b, Hoc87, Hod94d, Hod94e, Hod94f, Hod94g, Hod94h, Hod94i, Hod94j, Hod94k, Hod94c, Hod94l, Hod94m, Hod12c, Lie11, Mei12a, Mei12b, OW12, Pil12, ST12, Spr12, Tur60a, Tur87]. Germany [GR12]. Gets [Wat12a, Ano09b]. Gettier [PR17]. Gewissheit [GKO95]. Ghost [Cha94, Mei12a]. Giovanni [Nof17]. given [Dav13]. Giza [JTS97]. gjennomgang [The87]. Glue [PSS11]. glycolysis [LGS22]. glycolytic [Dut10]. Gnirut [Pla09]. Go [Wat12c]. Goal [Hau03]. goals [Hol86]. God [Haw05]. Gödel [Bre13, Hol18, Jor07, Ken17, Lom05]. Goes [Ano15a, Lew78]. Going [Mau09]. Goldreich [Kru05]. Good [OSZ03, Pip04, Pip05, Ano12j]. Google [Fis15]. Gordon [Nau09, GW14]. Got [Poo92, Poo91]. Gower [Tia11]. Grace [Wat12j]. Gradient [LKE93]. Gradient-Free [LKE93]. Grammar [Mar13c, EH91]. grands [Ano96]. Graphs [dC11a]. Great [ALdlP20, Kru05, Lap96, Lov04, RM00b, RM01, Rue07, Cop17h, RM00a, Rob97, Teu04a, Ano96]. Group [Isa14, Sof83, Tau61b, Tur38a]. groups [Boo52, Tur38b, Tur50c]. Grow [Swi04, GMC12]. Growing [LGB11]. guarantee [Kah84]. Guard [THWV88]. guide [CBSW17, Mar11b, Ano20, Jon17, Pet18, Hay17, KH19]. guided [Lip11, Pet08].

H [Ano04]. Hackers [Isa14]. Hacking [Hea15, Pat04, Pat07]. Hairs [Ano06b]. Haldane [EW17]. Half [Rus89, Her88, Her95]. Half-Century [Rus89, Her88, Her95]. Hall [Kru05, Don14]. hallmark [Shi04]. Halting [Cha95, Fra06, Hut84, Wel06a]. Hamkins [Wel06a]. Hamkins-Kidder [Wel06a]. handbook [Tur50b, Tur51b]. Hard [Har12a]. Hardback [Jon17, Hod06a, vL13]. Hardcover [Ano20, CK12b]. Hark [Kru05]. Harmful [Fre12a, HF95]. Hartree [Ano88, AWL^{+}88, Gon24a]. Harvard [AWL^{+}88]. hbk [Shi14]. Heads [Wat12e]. Hedy [Kah84]. held [Ano51, Man90]. Help [Coo06c, DW16, ST12]. Helped [Tur20]. Helquist [Kru05]. Henry [CFK^{+}91]. heretical [Tur96]. Herken [Rus89]. Hero [O’N23, BDD15]. heute [Hod12c]. heutigen [Dys12a]. hexagonal [OS91]. Hidden [SW10, Smi15a]. Hierarchy [CL02, Fai10a, Fai10b, Fai11]. High [Ano49, GvN51, vNG47, Bro97, Ive15, Jam06]. high-end [Ive15]. Higher [Nor14]. Hilbert [B^{+}11, Cop17e]. Hillston [BTHS12]. him [HP20]. himself [McG12, Par14, Pro15, McG12]. hinter [Hod12c]. histoire [VB15]. historic [Gee11, Lip11, Pet08]. Historical [Hai17, HP20]. History [AWL^{+}88, CFK^{+}91, CP01, Cop06b, Cop11a, Cop11b, CL17a, Cuf24, DKK^{+}98, Eva81, Fef99, MHR80, Sha00, Wil97, Ano46, Day12b, Goo79b, Haw05, HHW08, Mah10, Smi15a, TJC03, VB15, CFK^{+}91, TDCKW84]. Hitler [Cop17g, Hea15, Moo14, Rob17]. HL [Hou12]. Hodges [LH83, Sal12, Shu87, TDCKW84, Asp84, CK84, Hof83, Rid84]. Hold [Loe95, Loe09]. hole [BGL24]. Holling [Tia11]. Homage [Cas01]. hombre [Lea12]. Home [Hod97b, THWV88, Hod97a]. homme [Lea07, Lem04, Lem12]. homosexual [Dav13]. homosexuality [Ell19]. honor [Hym12]. honors [Sor05]. honours [Ano12e]. Hopf [Dil05, RMP11]. Horizon [Syk92]. hot [Sen21]. House [Kru05]. HP [Ano10a]. Hub [Gar95, Gar09]. Human [Bry22, Cop05b, Hic08, Mau09, Pel09, Ste17, JTS97, KH19]. Human-Like [Bry22]. hunted [McG11]. Husson [TDCKW84]. Hut [Mah10]. Hybrid [SCPC23]. Hydrodynamics [Tau63c]. Hyperbolic [Mar13a]. Hypercomputation [Cot03, Dav04]. Hypercomputational [Sta04]. Hypotheses [Zie09]. Hypothesis [Boo06a, EW17, GAM11, Boo06b, Tim04].

Ian [Kru05]. IBM [TDCKW84]. ICL [CFK^{+}91]. Icon [CK02, O’N23]. Idea [Wol16, Gou99, Str99]. Ideas [CP99, Dow14c, EGW04, Gan95, Hod06a, Lew21, Nof17, Coo12d, Daw16, Dow14b, Rob97, SS15, Wol16]. idée [Gou99]. Identification [Tra03, CWS^{+}24]. Idiotic [Sch04b]. IEE [Don01a]. If [Bri95, Bri09]. II [Bre12a, Goo92, Kru05, Bre12a, GvN51, Goo79b, Kah84, Sal04, Tau61b, Tia11, Tur51b]. III [Tau63a]. Illogical [Hel17]. Illustrated [LH83]. Illustrator [Kru05]. im [Lie11]. Image [CFK^{+}91]. Images [Mak95]. Imagine [Emm13]. Imitate [Pel09]. Imitation [Cop05b, Hod14, Hai17, Pro15, WS16, Cho09, Las09, Las95, Lon09, Pic03b, Cho95]. Immortal [Jea12]. Impact [Ano88, AWL^{+}88, CvL13, Hop12, Nof17, CBB12, Pap12, SS15, Chr15, Ano14, Avi14]. Implementation [RTM04, SGV94]. Implementing [dBPZM10]. implications [Lei01]. Impossible [ApS65, BT12, Lon09, SHH81, AMKM66, BSPI65, HLOS65, Str65]. imprint [DT12]. Improvements [Tru11]. Inaugural [Ano51]. Including [Cop00a]. Inclusion [Mai07]. Incomplete [BLA^{+}11, Dal12a]. Incompleteness [Fra06, Sut13]. Incomputability [Coo12b, Soa07]. Incomputable [Coo12c, Cap05, Coo12d, Coo12e]. independent [CWS^{+}24, Ive15]. Indianapolis [Lip11]. Indistinguishable [Bre13]. Individuals [HTG12]. Inevitable [Kah12]. inference [Fie06]. Inferentialism [CM10]. Infinite [CEL10, Min67, Min72, Whi12]. Infinitum [Cha94]. Infinity [Cha94]. influence [Abr11, Day12c, Gla12]. Influences [Dav95a]. Influential [Rog10]. Info [DC12, DC13]. Info-Computational [DC12, DC13]. Inform [San05]. Information [Baj12, Hai16, Joh15, Roc12, SG18, Cop12b, Sal12, SG17, Cer04]. Informationierung [Pil12]. Informatique [Bia79, Bre12c, Lea07]. informatization [Pil12]. Informing [Coo12a]. ing [Vos13]. ingénieurs [Ano96]. inhibition [Mei12b, Mei12b]. inhumane [Nau09]. Inmos [AWL^{+}88]. Innovation [AWL^{+}88, Don01a, Jon16]. Innovators [Isa14]. input [IST^{+}10]. insect [BSK^{+}15]. insights [Poo92, Var17]. Inspired [Har12b, Hod14, Zas18]. Instabilities [KGJ^{+}24, Dil05, RR12]. instability [AKS11]. Instrument [CL17b]. Instruments [Ano88, AWL^{+}88]. insubstantiality [Lei01]. integers [Haw05]. Integration [Ste17]. intellectual [TJC03]. Intelligence [AS08b, Bry22, Chu95, Chu09, CP04, Cop04, Cop05b, Edm95, Edm03, Edm09, EG12, Fur12, GC12c, Kru05, Müh09, MJ09, RS03, Rou18, Sav24, Ste03, Ste17, Tur50a, Tur87, Tur95b, Tur09, Tur20, AB00, Ano12j, Cro94, Gon23a, GW14, Hod85, Hod88, Hod01, MM69a, MM72, MM69b, Moo03b, Num05, Pro17d, Shi04, Tur92c, Yan12, FRT14, Tur87, Shu87, Cas13, Luc95, Luc09]. Intelligence* [Ste00]. Intelligent [Cop17f, Gon23b, Tur48a, Tur69, Tur96]. Intelligenz [Fur12]. Intelligenza [Tur94, Num05]. Interaction [KW12]. Interactive [Gol95]. Intercepts [Don14]. Interfaces [Gar95, Gar09]. Interlocutor [SW10]. Internal [Sha54]. International [Ano20, CS11a, DMV12, MBS11, DIMV11]. Internet [Ben97]. Interpretation [Pro06, Pos23a, Pos23b]. interpretive [CG87]. Interrogative [ŁW11]. Interrupted [GC12d]. Introducing [Sim17b, WW17, Bro97]. Introduction [DW12, HH84, Hol90, HH90, MC12a, Wat12i, MMB13, ST12, TW12]. Introductory [Goo92, Hod04a, Hod12b]. invención [Lea12]. invent [Hai14]. inventa [Lea07]. inventé [Ano96]. Invented [CP00, Ano96, Lea07, Smi10, SG17]. Invention [Coo06a, Pri10, Cop17h, Lea05, Lea12]. inventor [Lie11]. Inverting [GvN51, vNG47]. Invited [BTHS12]. Iodide [LE91]. Iraq [Kru05].

Irascible [Phi65, OS65]. Irony [Gon23b]. Irruption [Coc12]. ISBN [Ano20, CK02, Chr22b, Dal12b, Fai12, Gee12a, Hai16, Hod06a, Jon17, Nic17, Rus89, Shi14, vL13]. ISBN-13 [Dal12b]. Isolated [ÇG12]. Issue [Cop00a, MMB13, TDCKW84, GMC12, Teu12]. Issues [EBR09, Nau09]. Italian [Cap05, Hod03b, Num05, Odi12, Rig91]. Italy [CLS07]. IV [Hod12c, CK02, Hod12c, Tau62].

J [Bea84, Cer85, Kru05, Tur60a]. J. [EW17, Sut85, Tur60a]. Jack [And08, Ano20, Fai12, Hai16, Hod06a, Joh15, Jon17, Pet18, Sal12, TW12]. James [Kru05]. Jan [Chr15]. Jane [BTHS12]. January [BFG^{+}12, Jon17, WTP^{+}06]. Japanese [Don14]. Jenseits [GKO95]. job [Dav13]. Jobs [ZM08]. John [Lie11, Ano96, Asp80, Fie15, IM13, Lie11, Mad12, Müh09, Pri24, Smi10, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Johnson [Gee12a]. Jon [CK02]. Jonathan [Ano20, Jon17, Pet18]. Journal [BTHS12, TDCKW84]. journey [HP15]. Judge [Bri95, Bri09]. Juliet [Ano20, Gin19]. July [Ano51, BBLT06, Man90, Sof83]. Jumps [CZ12, Fai10a, Fai10b, Fai11]. June [Ano49, BBLT06, CLS07, CDL12, Swa13a]. Jungle [Roc12]. Juni [Str11]. just [Mac12a, Smi14]. Justice [Las09, Las95, ZM08]. Justified [PR17].

Kahan [Ano99]. Kann [Tur60a]. Kasparov [Ano12f]. Keep [Sch12b]. Keira [Bro13]. Kelly [Gee12a, CFK^{+}91]. key [ER68]. Keyboards [CFK^{+}91]. Keynote [Lis12]. Khwarizmi [CFK^{+}91]. Kidder [Wel06a]. kill [McG12, Par14]. Knew [Coo06a, Lea12, Lea05]. Knightley [Bro13]. Knowledge [Gol95, GF91]. Knuth [CFK^{+}91]. Kode [Bre12a]. Konrad [Lie11]. Kozaczuk [AWL^{+}88]. Kruh [CFK^{+}91]. Kryptografie [Blö12]. Kryptologie [Bau00]. Künstliche [Fur12].

Laboratory [AWL^{+}88, Fef99, Sha00, Wil80]. Lady [SHH81, Gon24a]. laid [Dys12a]. laissé [Mar13b]. Lamarr [Kah84]. Land [Fef95]. Lane [Sal12, GC12e]. Lane/Pantheon [GC12e]. lang [Mei12b]. lang-reichweitiger [Mei12b]. lange [Hod12c]. Language [BLA^{+}11, CH83a, CH83b, Fen95, HC88, MBS11, EH91, HH84, Hol86, HC87, HP88a, HP88b, HMRC88, Hol90, HH90, DIMV11, DMV12]. language-based [EH91]. Languages [Dow12a, Sha09a]. Laplace [Lon09]. Large [AWL^{+}88, Tur49]. Large-Scale [AWL^{+}88]. larger [Lei01]. Last [Cle17, Moo15]. LATA [DIMV11, DMV12]. Later [SCA00, SCA03, Sea95, Sea09, Daw16, Mei12b]. laureates [Fis17]. L’aventure [Lom05]. Lavington [Dal12b, Gee12a]. Law [Kur04, THWV88]. Lazy [Cha94]. leaders [Kah84]. Leading [HSD09]. Learner [PA11a]. Learner- [PA11a]. Learning [EG12, dBPZM10]. least [CWS^{+}24]. leaves [Und13]. Lebens [Bre12a]. Lecture [Bro05, CV13b, Don01a, Kov03, Mai06, Mai07, Pip04, Rob97, Tur47, Tur95a, Wel02, TW05, TW12, BTHS12, Don01b, Hel17, Pip05, Ran00]. lectures [Har47, Odi12, Ash87]. Ledger [TDCKW84]. Lee [Jor07]. Leeuwen [Chr15]. left [Mar13b, Tur35a]. legacies [FRT14]. Legacy [Ano20, Cho12, CS12, Cuc12, DC12, DC13, Dow14c, Fef99, FB17, For12, Gin19, Kar95, Kru05, Lov04, Müh09, Sha00, Swa13a, Ano11a, AB12, AB14, Ben97, BC17, CM96, Dow14b, HfB98, MC96, Por19, Teu04a, Und13]. legte [Dys12a]. Leibniz [KP02, Kil14a, Kil14b, Dav00, Dav12, Dav18, MC12a]. Leistungen [Hod12c]. Lengyel [WWG12]. leopard [Poo91, Poo92]. Lesehilfe [ST12]. Leslie [Tia11]. let [Hum14]. letra [Hid12]. Letter [Str65, Var14]. Letters [CAC14a]. li [Tur60a, Tur60b, TvN99]. li.vi.5 [CFK^{+}91]. Liberate [Ano10a]. liberté [Lom05]. Library [CFK^{+}91]. Lie [Tur38b]. Lieutenant [TB12]. Life [Bra95, Bre12a, Bre12b, Cop04, CB17, Cop09, FH15, GC12d, Hod03b, Kru05, Lov04, Swa13a, Ter11, Bod17, GC12c, Hod02b, Mac12b, Pap12, Syk92, Teu04a]. Lifestyle [Mai06]. light [New03]. Like [Bry22, DDL01, Spr12]. Limit [Sie95, Zab95, CS19]. Limitations [Mar11d]. Limited [Chr22b]. Limits [Kar95]. lineages [BSK^{+}15]. Lines [Hod12a]. Literature [AWL^{+}88]. Little [Ive15]. Lives [ZM08, B^{+}11, Wol16]. LLMs [Sav24]. Local [Tur51a, Mei12b]. location [Ive15]. Loebner [Zde03]. Logic [Bea89, Bee95, Bul21, CLS07, Cop04, Dav95a, Dav95b, Dow14c, Hel17, Pri10, vL13, App12, Asp80, Dow14b, Dys12a, GGZ06, Shi14, Tur38c, Tur39, Tur65, Tur01a, Tur01b, Tau61a]. Logical [Ben95, Hod08a, Tur60a, CS19, BBLT06]. logiceskaja [Tur60a]. Logician [GG13]. logics [GGZ06]. Logik [Dys12a]. logische [Tur60a]. lokaler [Mei12b]. London [Chr22b, Shu87, CV13b, Swa13a, Tur47, Tur95a]. lone [Zas18]. loner [Mic84]. Long [RMP11, Mei12b]. long-range [Mei12b]. Long-Term [RMP11]. Look [Hau03, War12, BA05]. look-alikes [BA05]. Lost [Cop18, Cop25, Dav13]. Lovelace [BBF03, Gon24a, SHH81, Swa13b, Swa17]. lover [Bro13]. Loving [Wat12j]. low [Fai10a, Fai10b]. Ludwig [GKO95, GKO95]. lunaires [VB15]. lunar [VB15]. lyrics [Hid12].

M [Ano99, AWL^{+}88, Bod49, CD86, Chu13, Fie15, Gan54, Gla03, Gla04, Goo79b, Goo92, Har47, Kid96, Lol13, May61, Sor05, The87, Tur59, Tur72, TWCD86, Tur01a, Tur12, TB12, Tur15b, Zab12]. M. [Boo52, TDCKW84]. MA [Ano04, Nic17]. macchina [Cap05]. macchine [Num05]. Machine [AWL^{+}88, AG11, BCT10, Bow53b, CK02, ÇG12, Chr22b, Cop00a, CL17b, Cur65, EG12, For12, Gon23b, Har12b, Hod08b, Hof83, Len95, Len09, Mar11c, Mar11d, MM69a, MM72, MM69b, PA11a, PA11b, Rus89, Ste12b, Ste17, THWV88, TNPY49, T^{+}06, Tur21b, Aga01, Ano12j, Bia79, Blo98, Bro97, CK12a, Cap05, CD77, Cas06b, Coo12f, Cop17h, FOO71, Gon23a, Gou99, Gre17, HP20, Hal14, Hod12c, LTM^{+}51, Lis12, Mal87, Mei12a, Nau93, Num05, Par14, ST23, SGV94, Sie12, Smi02, Tur60a, Tur91, TG95, TvN99, Wat12m, Wel14, Bia79, Blo98, Cas06a, Coo12a, CP00, Cop00b, Hal13, Has95, Her88, Her95, Her98, Hod95a, Jac12, dBPZM10, Pet08, RTM04, Sha54, Ste90, Wel04, Cha94, Con22, Lip11]. Machine-Inspired [Har12b]. Machinery [Cas13, Luc95, Luc09, Tur48a, Tur50a, Tur95b, Tur09, Cop17f, Dys12a, Tur69, Tur96, AWL^{+}88]. Machines [Ano88, Ano89, AWL^{+}88, Axe12, BLvT11, vEB12, Bow53a, Bul21, CFK^{+}91, Cla72, CM10, CS11b, Den04, Har03, Hin17, Hod12e, Hop84, IT12, Jea12, Jon04, Jon16, Mar11a, May52, Mic08, MC96, Min67, Min72, NW12, New49, Pri10, RS03, Rou18, Wat12j, Wat12k, Wat95, Wat09, dC11a, And64, Bat17, BB94, Gam13, IST^{+}10, KvLP88, Lev06, III14, PR10, Pro17a, Sha09b, Sze94, TPD85, UST^{+}10, Web12, dLMSS56, And84, Arb95, AWL^{+}88, Hop12, Mar13a, Wel06a]. madman [Lev06]. Maestro [CFK^{+}91]. Major [Kle95]. Make [Ano89, BDD15]. Makers [AG02, Wol16]. Making [CK02, Eva81, GC12a, Sch04b, Tra03, Aga01, Smi02]. Malonic [LE91]. Man [Bea84, Cer85, CFK^{+}91, Coo06a, Jef49, Lea12, Pau91, Sut85, Bol84, Cap05, GG12, Hod12c, Lea05, Lea07, Lem04, Lem12, Moo14, Smi10, Tur17, Reg23]. Manchester [Ano51, Ano12g, Cop11a, Cop11b, LTM^{+}51, O’N23, Swi19, Tur50b, Tur51b, Lea19]. Manipulation [Con12]. Mann [Hod12c]. Männer [Hod94h]. Manual [AWL^{+}88, Tur00]. Manuscript [Bau12]. Many [CFK^{+}91, Hod12c]. máquina [Tur74]. maquinas [TPD85, And84]. March [DMV12]. Marian [Chr24]. Mark [Ano12g, Ano20, Jon17, Pet18, Tur51b]. Marling [CFK^{+}91]. Marriages [Mur12]. Mars [Smi05]. Marshall [Don14]. Martin [Gee12a, Kil14a, Kil14b, KP02, Mon19]. martyr [Mac12b]. Marvelous [Wat12k]. Masani [AWL^{+}88]. Maschine [Hod12c, Tur60a]. Maschinen [Dys12a]. mashina [Tur60b, TvN99]. masina [Tur60a]. Master [And08, Kid06, Cop05a]. Matching [LA12]. matematiche [B^{+}11]. matematico [Odi12]. material [AH85, Man90]. Materialism [Coc12]. Math [Cra10a, Emm13, Sie12]. Mathematical [Ano88, AWL^{+}88, Cas06b, CV13b, Dav95a, Dav95b, GGZ06, LGS22, Lom05, Mur93, Mur12, Pic03a, Ste94, Tur47, Tur95a, Tur01b, Asp80, Haw05, Hil93, RR12, B^{+}11, Tur01a]. Mathematician [Ano12h, Dav13, Ell19, MS17, Rid84, Rue07, AW77, McG12, Odi12, Zas18]. Mathematicians [Grc11]. Mathematics [Bee95, Bee04, BH03, Hal13, Hal14, Rue07, Boo52, Daw16, Emm13, FF91, KAB99, Tur92b, WW17, Tur45]. mathématique [Lom05]. Mathison [AW77, AH85, Ano00a, Che93, CV13a, New55, Tau56, Ter11]. Matrices [GvN51, vNG47, TT56]. Matrix [Bod49, Tur48c]. Matter [vEB12, Hut84, Jon16]. Mauchly [Ano96]. Maurice [Ben12]. Maverick [AWL^{+}88]. Max [CV13a, GG13, GG17]. Maxima [Bau00]. Maximen [Bau00]. May [ACL12, DIMV11, McG12, Poo92]. McBain. [Kru05]. McCorduck [TDCKW84]. McLean [AWL^{+}88]. McNeil [CFK^{+}91]. Mean [Hod08b]. Meaning [Bra95, Gon24a]. means [Mic15]. meccanica [Tur94]. Mechanical [Jef49, Tur92c, HHW08]. Mechanics [DP02, Tau61a]. Mechanism [BB16, Cop00a, LL12]. Mechanisms [Dah95, LGS22]. Mechanization [Bee04]. Medal [Fie15]. Media [HS82]. Meeting [Hil17, GKO95, Man90, PA13]. Meets [BLvT12, SG18, Blu04, Mis09]. mela [Pat07]. Membrane [GS12, Zas18]. membranes [TCP^{+}18]. Memoir [Hus91]. Memorial [Don01b, Ran00]. Memories [Cuf24]. Memory [Cop11a, Dre10, Tur05a]. Men [Cha16, Hod94h]. Mental [Mak95]. Mentes [TPD85, And84]. Mentor [GG13, GG17]. Merin [WTP^{+}06]. metaheuristics [Yan12]. Metamathematics [Bea89]. Meteorology [Tau63c]. Method [Boo06a, HSD09, Tru11, Tur43]. Methoden [Bau00]. Methodological [EBR09]. Methods [Bau00, GMT^{+}12, Tur51a]. Meyer [Bur11]. MI5 [Irv04]. Micro [Eva81]. Microelectronics [AWL^{+}88]. Microemulsions [BVE11]. Might [Lip12, Kah12]. Military [Hod03a]. million [Fis15, CAC14b]. Milner [BLvT12]. mimic [Sie12]. Mimulus [SCPC23]. Mind [AWL^{+}88, Cla72, CM10, Cop00a, Cop05b, Hof83, Jef49, Mck95, Mck09, Mic08, TNPY49, T^{+}06, Cho12, HHW08, Sie12, SG17, Spr17, Wel06b]. Mind-Body [Cla72]. Minds [And64, Har03, Rue07, HM02, And84, TPD85]. Minimum [Liv02]. Minimum-Weight [Liv02]. Minister [Nau09]. ministers [Coa13]. Miracles [Ter11]. misfit [Mic84]. Misidentification [SW10]. mistaken [Cro94]. Mistakes [Sch04b]. Mk [Tur00]. Mlýn [BFG^{+}12]. mm [Jon17]. Model [Ano10b, Ano15b, Dah95, DC11b, Hew13, KW12, MBC06, SCPC23, Tra12, AKS11, Ano12j, Dal12a, Dut10, FHM14, Jac12, LGS22, RR12, Spr17, Tia11, Nor14]. Modeling [LE91]. Modelle [Mei12b]. Modelling [LP11, Mur12]. Models [ACL12, BAC14, DC11b, EGW04, Sta04, Wie12, DDL01, GS12, SNUM03, Wel14, Mei12b]. Modern [And08, Bia79, CK02, Cop06b, CP12b, Dav95b, DW16, Kid06, LPAA22, O’N23, dSAL^{+}13, Aga01, Ano19d, Cop05a, Cor17, Mad12, Smi02]. Modern-Day [DW16]. modernes [Bia79]. Modest [Pic11]. Molecular [MBC06, CS11a]. moments [CP17a]. monument [GC17b]. Morphogen [LGB11]. Morphogen-Regulated [LGB11]. Morphogenesis [Coo12a, Fre86, Ric06, SNUM03, Tur90, Tur92a, WBM17, Kid96, Tur52, Nan03]. Most [Rog10, Rob17]. Mother [Sha14]. Motivating [Tay98]. Mouse [Ano06b]. Moving [Fre12c, Hau03, SCT^{+}17]. Mozet [Tur60a]. Mozhet [Tur60b, TvN99]. MR [CBB12]. MR2919681 [Smi14]. MR2963548 [Smi14]. MR3185259 [Smi14]. MR3618873 [KH19]. Ms [CFK^{+}91]. Much [Coo06a, Lea12, Lea05]. Muddled [TDCKW84]. Multigenerational [SCPC23]. Multitape [IT12, SGV94]. Münster [CBB12, Gla12]. Murdered [Pit14, Par14]. Murphy [THWV88]. Museum [Swa13a]. Music [Ano16, CL17a, CL17c, Hid12]. música [Hid12]. Musical [CL17b]. Musings [Ner14]. Musterbildung [Mei12b]. My [Hum95, Hum09, May01]. myslit [Tur60a, Tur60b, TvN99]. Myth [Dav04].

Nachwort [Hod94g]. Named [Ste94]. nanopatterns [BSK^{+}15]. nanoscale [TCP^{+}18]. nanotechnology [Wel02]. Narration [Hoc87]. Narrow [Cop00a]. National [Fef99, Sha00, Tur01c, Wil80]. Natural [DC11b, Gel12, Whi12, Hod97c]. naturalized [Sch88]. naturally [Mon19]. Nature [Chu95, Chu09, Coo06c, DC12, DC13, Zen13]. naval [Goo00, Don14, Mah10]. Navy [Gla03, Tur03]. Nazis [Hea15, Tur20]. near [Dil05]. neat. [HP20]. Negative [PSS11]. neither [Irv04]. Netherlands [MBS11]. Nets [CP12a, SS91]. networks [RR12, Web12]. Neue [Hod94h]. Neumann [CFK^{+}91, Fie15, Tur60a, Ano96, Asp80, CK12b, HP20, IM13, Lie11, Müh09, Sch88, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Tur60a]. neural [SS91, Web12]. neurons [CP17b]. Newman [CV13a, GG13, GG17]. News [Ack14, Fis15, Fis17, McG12, Sav24, CAC14b]. Newton [Blu04]. NJ [Kru05]. no [Sha12]. no-priority [Sha12]. Non [Wel06a]. Non-deterministic [Wel06a]. Nonexponentially [Fu12]. Nonlinear [KW12]. nor [Irv04]. Norbert [AWL^{+}88]. Normal [Bec12, BFP07, Turxx]. Norman [Ano12h]. Norris [AWL^{+}88]. Norton [KP02]. Norwegian [The87]. notable [Wol16]. Note [CZ12, New49, Sch04b, Tsa19, Ano19d, Ano21, Ell19, Pea19, Turxx]. Notebook [Ano15a, Cop18]. Notes [Tur05a, Hut84]. nous [Mar13b]. Novel [FGG^{+}24, Ano04, HM92, Pap03]. November [Tur42b]. NP [Fu12]. NPL [Ano11a]. Nuclear [Hel17]. numberings [Fai10a]. Numbers [Bec12, Chu13, Pau91, Tur36, Tur21a, BFP07, Bra13, The87, Tur37b, Turxx]. Numeracy [Pau91]. Numerical [Cuc12, GvN51, HAC^{+}85, Tau63b, Wil71, vNG47, Dow13, Mad12]. numérique [Dow13]. NUPT [Ste12a]. NY [Kru05].

O.B.E [Gan54]. OBE [AW77]. Objection [Gon24a, Pic03a]. Obscaja [Tur60a]. Observatory [THWV88]. Observers [Sut13]. occurring [Mon19]. Ockham [Res17]. October [MBS11, Swa13a, TNPY49]. Oded [Kru05]. ODFL [Ste12a]. Off [Fre12b, Tur48c, Hej07, Bod49]. offer [Poo92]. Oh [Wei88]. Ohio [Tur01c]. Old [Pro17b]. Omnibus [Bri90, Ken89, Dew89, Dew93]. Once [Cha16, CH16]. one [LC01]. Online [CF98]. Only [KH19]. ont [VB15]. Ontario [Sof83]. OO [BB94]. Opening [Den12b]. operate [Jac12]. operating [HP88a]. Operation [AWL^{+}88, Gla04]. Operations [Pri10]. Operators [Tau63a, Ire17, Tau61b]. Opposition [Dav06a]. Optimal [OSZ03]. option [HM92]. Oracles [BCT10]. ordenadorea [JTS97]. Order [GvN51, vNG47]. Orders [DJ12]. ordinals [Tur38c, Tur39, Tur65]. ordinateur [Ano96]. ordinateurs [Bia79]. organisational [Smi15a]. Organizational [AWL^{+}88]. Origin [Ano46, Dav95b]. Original [Tur72, Kan12]. Origins [Bla14, CK12b, CFK^{+}91, Día12, Dys12c, MD11, Mic80, Swa13b, Asp80, GC12e, Kah84, Ran72a, Ran72b, Ran17a, Sal12]. Other [AWL^{+}88, CD86, Sch04b, Tau62, Blu14, CK12a, CD77, Smi15b, TWCD86]. Out-of-the-Box [EG12]. Outlaw [Hod94m]. Output [PR10]. Overcoming [THWV88]. Oxford [Ano20, Fai12, Hai16, Hod06a, Jon17, Rus89, Sal12, Smi14, vL13, Man90].

P [Ano96, TDCKW84, Rou18]. P-Type [Rou18]. P. [TDCKW84]. Pac [Reg23]. Pac-Man [Reg23]. Page [Hod97b, Hod97a]. Pages [Fai12, Ano04, Ano20, Chr22b, Gee12a, KP02, LH83]. Pantheon [CK12b, Día12, GC12e]. Papadimitriou [Ano04]. Paper [Day13, Lip11, Pet08, Zas18]. Paperback [Ano20, Jon17, Sal12, CK02, Chr22b, Hai16, Hod06a]. Paperbacks [Shu87]. Papers [CD86, Lap96, Lew21, AW77, AH85, Ano11b, Dav65, ER68, TWCD86, AWL^{+}88, MBS11]. para [Hid12]. parade [Ano12k]. paradigms [Coo08, GS12]. paradox [Gör91, Gör95b, HS82]. Parallel [Dah95, IST^{+}10, UST^{+}10]. parameter [CWS^{+}24]. paranormal [Lea17]. Pardon [Dav13, Hou12, Coa13, Ell13]. Pardoned [Cha16]. Park [Ano11b, Cop06a, Cop17a, Gee11, GMT^{+}12, GW14, HS93, Pri21, Sal04, Sev12, Smi15b, Smi15a, Tur20, Chr21]. Parsing [ERB08, EBR09]. Part [Cop11a, Cop11b, Cra10b, DH10, Mar11a, Mar11c, Mar11d]. Pasadena [CS11a]. Pascal [Jon16, PC88]. Pass [EG12, Len95, Len09, Rap03, Wie12, BB12a, BB12b]. Passed [Chr24, Hum95, Hum09, Var14]. passes [Zas18]. Passing [Mau09, Zde03]. Passive [BB16]. Paternity [LPAA22]. path [Hej07]. Pattern [Ano09a, BB16, Cai12, DKM^{+}24, FGG^{+}24, KW12, Kon12, KGJ^{+}24, LA12, Rei12, RMP11, WS00, BGL24, Daw16, DDL01, Dut10, LGS22, Mei12b, SNUM03]. Patterning [Ano06b, LTS^{+}21]. Patterns [Ano01, Ano11c, BVE11, CEL10, HSD09, LGB11, LOM^{+}01, Mur12, She12, AS08a, CWS^{+}24, Dil05, GAM11, HM96, KA96, OS91, Poo92, Tia11]. PC [GC12e]. Pearcey [CFK^{+}91]. Pearson [Kru05]. Peirce [Pos23a, Pos23b]. pencil [Bat17]. Pennings [AWL^{+}88]. pensar [Tur74]. pensé [Gou99]. pensée [Bre12c]. People [DKK^{+}98, Lav80, Wat95, Wat09, Jam06, Wol16]. Perceptions [Pra95]. performance [GC17a]. Peril [Bry22]. Periodic [Tau61b, LGS22]. periodicity [Tur35a]. persecuted [Ell19]. Person [BB12a, BB12b]. Personal [Rue07, Wat12a, DT12, Nau93, Pap12, Wol16, de 12]. personality [Bie12]. Perspective [Bre13, Wel04, Mis09, Sch12b]. perspectives [Wol16]. Petal [SCPC23]. Petre [CFK^{+}91]. Petri [CP12a]. Petzold [Wil10]. Phenotypes [SCPC23]. philosopher [Hod97c]. Philosophers [RM00b, RM01, RM00a]. Philosophical [EBR09, FB17, Ano20, Gin19]. Philosophy [Cop04, CF98, DC12, DC13, Gör95a, Rob97]. Photography [Con22]. Phyllotaxis [Swi04]. Physical [AD12, BCT10, Ben95, Cot03, Dow12a, Fef99, Hod08a, Pic11, Sha00, Szu12, Wil80, Zie09]. Physically [Zie09, ST23]. Physically-relativized [Zie09]. Physics [Ano88, AWL^{+}88, Bac12, Fin95, FF91, Yao03, Zie09]. Pilot [Ano13, Ano11a, Wil80]. Pioneer [Bod17, Hai16, Joh15, Mac12b, Met19, Sal12, Cop12b, Smi10]. Pioneers [Wei88]. Plane [Ano89]. Plant [KW12]. Plato [PR17]. Platonists [CM10]. play [SG17, WH88b]. played [Kah84]. Playing [Cha94]. Pleasures [Kör96]. plus [Cop04, HP88a, HP88b, PC88]. Point [Gon23b]. points [GR12]. poisoned [McG12, Vin13]. Polish [Hod02b, RA04]. Polyamide [TCP^{+}18]. Polymers [QSW11]. pomme [Lem04, Lem12]. Popperian [Bea89]. Populations [HTG12]. Portrait [AWL^{+}88, Ano21]. positive [Mai06]. possibilities [Web12]. Posthumous [Ell13]. Posts [Hau03]. Postscript [Hod94i]. Postskriptum [Hod94i]. postwar [Sum14]. Potential [Ano01, Sie12]. pound [Ano19d, Ano21]. Powerful [LP11]. Pp [CK02, Hod06a, Nic17, Rus89, Shi14, vL13, Boo52, Hai16, Jon17, Kru05]. Practical [Gör95a, SW10, Tur48b, Gou99]. Practice [BFG^{+}12, WTP^{+}06]. pratique [Gou99]. pre [Cor17]. pre-war [Cor17]. Predator [RMP11, AKS11]. predicted [Zas18]. Preface [GMC12]. Prefiguring [TJC03]. Prehistory [TDCKW84]. Prentice [Kru05]. presentation [Lis12]. presented [Man90]. Press [Ano04, Ano20, Fai12, Hai16, Hod06a, Jon17, Kru05, Nic17, Rus89, Sal12, Shi14, Smi14, vL13]. Press/Random [Kru05]. Prestigious [Ano14]. Pretext [Kru05]. Prey [RMP11, AKS11]. Price [Con95]. Prime [Nau09]. Primes [Bul15]. Princeton [Sal12, Shi14, vL13, App12]. Principle [Szu12, Wie12, Deu85, Tim04]. Principles [AD12, Dah95, Dow12a, LTS^{+}21]. priority [Sha12]. prize [Fis15, CAC14b, Ano90]. probabilistic [dLMSS56]. Probability [Cha95, OSZ03, Tur41a, Zab12, Zab17, Goo79b]. Problem [Cla72, Fra06, Har12a, Hut84, Boo52, Cop17e, Cro94, Tur50c]. Problems [Tra12, Dav65, GGZ06, III14, Tur54]. Proceedings [Sof83, USE83, PA13, ACL12, AWL^{+}88, BFG^{+}12, CLS07, CDL12, DIMV11, DMV12, WTP^{+}06, BBLT06, CS11a]. Process [Fra12, HTG12]. Processes [Tur48c, Bod49]. processing [DB05]. Prodigy [CFK^{+}91]. produces [Poo92]. Prof [CV13c, Tur15a, Tur40, Chr16, Ham16]. Prof. [BTHS12]. Program [ApS65, Day21, Hum95, Hum09, MJ84, SHH81, TDCKW84, AMKM66, BSPI65, CPR11, HLOS65, Nau93, Str65]. Programmability [Con95]. Programme [Bea89]. Programmer [Tur50b, Tur51b]. Programming [Bul21, EH91, HC88, Lis12, Met19, PA11a, Pri10, CS11a, Day12c, HH84, Hol86, HC87, HP88a, HP88b, HMRC88, Hol90, HH90, dBPZM10, Tur51a]. Progressions [DJ12]. Project [Cop25, Ano46, Sch88]. Promise [Bry22]. Proof [MJ84]. Proofs [Gol95]. Propeller [Wat12e]. properties [UST^{+}10]. Proposal [Tur45, Tur72]. Proposed [Tur46, Tur05b]. propositions [Dav65]. PROSE [Ano14]. Prospective [Ano88, AWL^{+}88]. protagonists [B^{+}11]. Proving [CPR11]. proximity [Fie15]. Psychology [CM96]. psychotic [Lom05]. psychotiques [Lom05]. Publications [May61]. Publisher [Wil10]. Publishing [Ano20, Chr22b]. Puede [Tur74]. punishment [Cop17c]. Pure [Tur92b]. purification [TCP^{+}18]. Purpose [CH83a, CH83b]. Pushdown [IT12]. Pushing [Coo12d]. put [Hum14]. puzzle [Lei01].

qu’Alan [Mar13b]. Quantum [AD12, Bre13, Deu85, Dow12a, Her98, Jac11, Tau61a, Tim04, Joy00]. Quantumland [Buh14]. quatre [VB15]. que [Lea12]. queer [Vos13]. Queries [Tro93, Tro95]. Quest [Eps95, EBR09, Eps09, Lev17]. Question [Day21, Fre12a]. Qui [Ano96, Lea07, Lem04, Lem12, VB15]. quirky [Odi12]. rêveurs [VB15]. races [Hod94j]. Radiolaria [Ric17]. raised [CAC14b]. Ramsey [CZ12]. Random [Kru05, KvLP88, Mai06]. Randomness [Bec12, DH10, Gol95, Dow14a, Dow17]. range [Mei12b]. ratio [AKS11]. ratio-dependent [AKS11]. Rational [DL06]. Re [Cop00a]. Re-Examination [Cop00a]. Reaction [Ano09a, Ano15b, BVE11, CEL10, DKM^{+}24, FGG^{+}24, LE91, CWS^{+}24, Dut10, LGS22]. Reaction-Diffusion [BVE11, CEL10, Ano15b, Dut10]. Reactive [BLvT11]. Read [AWL^{+}88]. reading [ST12]. Readings [Bul21]. Real [CLS07, Tur18, Bra13, Lev17]. Reality [Pra95, Coo12e]. reals [Blu04]. reasoning [FRT14]. Reasons [O’N23]. receive [Mic15]. Receives [Dav13]. Rechenmaschine [Mei12a]. recital [Hid12]. Reckoners [TDCKW84]. Reckoning [Jon16]. Recognition [Cai12]. recognized [Hod12c]. Reconciliation [Hod94k]. Reconciling [BB94]. Recorded [Ano16]. Recreations [Ste94]. recursive [Sie14, Wri16]. Reducibility [Dav06b, DH10]. Reducible [Fu12]. Reduction [Axe12]. Reflections [Den12c, Hil00a, Tur21c, Hai17, HP20, Wel12, Chr22a]. Register [Tur01c]. Regulated [LGB11]. reichweitiger [Mei12b]. Rejewski [Chr24]. Relais [Hod94j]. Relation [Bea89]. Relative [Chr22a, Tur21c]. Relativistic [NA06]. Relativity [Sha14]. relativized [Zie09]. Relay [Hod94j]. Relays [TDCKW84]. remarkable [Jam06, Web12]. Remarks [Goo92]. REME [TB12]. remembered [New12]. Reminiscences [Hil00a]. removal [BGL24]. Rennen [Hod94j]. Repetitions [Tur41b]. replicating [RTM04]. Reply [May01]. Report [AW77, AWL^{+}88, CD86, TB12, GMT^{+}12, Tur42b, TWCD86, Ano49]. Representations [vEB12]. Reproduces [SCPC23]. Reproduction [Arb95]. Republic [BFG^{+}12, WTP^{+}06]. Repurposing [Hic08]. Research [TDCKW84]. Resisting [ÇG12]. résolution [Dow13]. Resonances [Flo17, Gon23a]. Resources [Gur95]. Response [BTHS12, Hod12f, LGB11]. restoration [Gee11]. Restoring [Ano16]. Results [Har12b]. Rethinking [Wel06b]. Returns [Kur04]. revealed [Pea19]. Reveals [BVE11]. revelations [Ran17b]. reverse [PA11b]. Reversible [AG11, Axe12, DL06]. Review [Ano04, Ano20, Asp84, Bea84, Bla14, Bod49, CK84, CK02, CK12b, Cha94, Chr13, Chr15, Chr21, Chr22b, Chr22a, Chr24, Chu13, Coo06a, Dal12b, Día12, Ers84, Fai12, Fef99, Gas16, Gee12a, Gin19, GC12e, Ham16, Hay17, Hod06b, Hof83, Hof85, Jon17, KP02, Ken89, Kil14a, Kil14b, Lea19, Mac12b, Nic17, Nof17, OS65, Pet18, Rid84, Rus89, Shi14, Smi02, Sut85, Tho18, Wil10, Avi14, Cer85, Chr16, Con22, Joh15, Kid06, KH19, Lip11, Sha00, Smi14, The87]. Reviews [Ano06a, AWL^{+}88, Bri90, CFK^{+}91, Hai16, Hod06a, Kru05, Lov04, TDCKW84, vL13, Sal12, Smi14, Ano88]. Revised [Cop11a, Cop11b, Cuf24, MBS11]. Revision [Chr10]. revisited [Cor17, Shi12, Sim17a]. Revolution [Isa14, Nof17, HP15, SS15, AWL^{+}88]. Revolutions [CK02]. Rich [CS12, Und13]. Richard [Kru05]. Richards [Ric06]. riddle [McG12]. Riemann [Boo06a, Boo06b, Leh70, Leh56, Tur53]. Right [Tra03, Tra12, Tur35a]. rigor [Lom05]. rigueur [Lom05]. Rings [Tau63a]. Risk [Buz12]. River [Kru05]. Road [Dav18, KP02, Kil14b, Dav00, Dav12, Kil14a]. Robert [Gas16]. Robin [Ano20, Jon17, Pet18]. Robots [CFK^{+}91, Pro04]. Robust [Cai12, BGL24]. Roger [Gee12a, Pit14]. role [Day12b, Gon23a, Kah84, Mei12b]. Rolf [Rus89]. Rolle [Mei12b]. roots [Leh56]. Rounding [Tur48c, Bod49]. Rounding-Off [Tur48c, Bod49]. Routine [Tur49]. Routledge [Ano12h, CF98]. Rowland [AWL^{+}88]. Royal [Dav13]. Rule [Kru05, Pro04, McG11]. Rule-Following [Pro04]. Rules [Pic03b, Jac12]. Ruminations [Pau91]. Running [Gla03, Tur03]. Runs [Reg23]. Russian [McG11, TvN99].

S [Chr15, EW17, Gla03, TDCKW84, Tur03]. S. [TDCKW84]. sabía [Lea12]. Saddle [Kru05]. Saga [AWL^{+}88]. Said [Den12a, Pro15]. San [USE83]. Sara [May61]. saved [Ano11b]. Saving [Mur12]. say [Kah12]. Says [Tsa19, Pit14]. Scale [AWL^{+}88]. scandal [Rig91]. scandalo [Rig91]. Scheutz [THWV88]. Scholarship [Day21]. Scholastic [Kru05]. school [Bro97]. Schriften [Tur87]. Schuster [Kru05, LH83]. Sci [Avi14]. Science [AG02, AWL^{+}88, Ber16, Bia79, BFG^{+}12, CK02, CP99, Dav95a, Har12b, Ken89, Lap96, Lew21, Pri10, SCT^{+}17, Swa13a, SG18, Ted15, Wel04, WTP^{+}06, Asp80, Bre12c, Bro97, BDD15, Das14, Dew89, Dew93, DT12, Fie15, HH84, Hol90, HH90, LC01, Lea07, Smi05, Wel06b, Nic17, Bri90, Tho18]. Sciences [Mur12, Fly02]. Scientific [Hin17, Pri24, Tra12, HM02]. Scientists [Rog10, Rob12]. scrambler [Kah84]. script [Bre12a, Bre12b]. Second [BBLT06, Wat12l, Sha09a, RA03, RA04]. Secrecy [TB12]. Secret [Cop25, Lew78, Tur20, DB04, Hea15, Rob17]. Secrets [Bau00, Cop04, Cop06a]. secure [Kah84]. Security [Hel17, Pip04, Pip05]. Segarra [Hid12, Hid12]. seine [Hod12c]. Selbstverstärkung [Mei12b]. Selected [MBS11]. Selections [DKK^{+}98]. Self [Arb95, RTM04, Mei12b]. self-enhancement [Mei12b]. Self-replicating [RTM04]. Self-Reproduction [Arb95]. semi [Boo52, Tur50c]. semi-groups [Boo52, Tur50c]. Seminal [Cop04]. semiotic [Pos23a, Pos23b]. Sense [Flo17, FRT14, Lev17]. Sensible [Hut95, Hut09]. September [CS11a]. Sequence [AWL^{+}88]. ser [Boo52]. Serendipitous [dSAL^{+}13]. series [TW05, TW12]. Service [Tur87]. services [Hod12c]. set [BSK^{+}15, Jac12]. Sets [Fu12, OG12, Tau61a, Fai10a]. Sex [Cha16]. shadow [New03]. Shannon [Smi05, SG17, SG18]. Shapes [PSS11]. Shaping [Ted15]. Shells [DP02]. shops [Ive15]. Shore [Hod94c]. Short [Gla03, Hut84, Tur03]. Shroud [Con22]. Siena [CLS07]. signaling [RR12]. Significance [DC11b, ZM08, CAC14a]. Simon [Dal12b, Gee12a, Kru05, LH83]. Simple [AG11, FOO71, HP20, Poo92]. Simplicity [Hin17, MC12b]. Simulating [LP11]. Simulation [Hut95, Hut09, FOO71, FRT14, GC17a, KvLP88, FOO71]. Simulator [Cur65, Her98]. Single [PSS11]. Sinkov [Chr10]. Skill [Gör95a]. skin [HM96, KA96]. SLAIS [MMB13]. Slave [CFK^{+}91]. SLE [MBS11]. Small [NW12]. Smaller [Wel02]. Smart [Len95, Len09]. śmierć [Hod02b]. Soare [Gas16]. sobre [And84]. Social [Edm95, Edm09, EG12, Mur12, Par17, Smi15a]. Socialization [RS03]. Society [Gee12a, Tur47, Tur95a, Mac12b, Cer04, CV13b]. Socrates [RM00b, RM01]. SOFSEM [BFG^{+}12, WTP^{+}06]. Software [Day12a, Kah12, Sof83, MBS11]. Solution [RV12, Dow13, Wri16]. solutions [LGS22]. Solvable [Tur54]. Some [Rue07, Tur53, UST^{+}10, Wel12, Wil71, Jam06, Lei01, Wol16]. Sommaruga [Nof17]. Son [CFK^{+}91]. songs [Hid12]. Sorry [San05]. Space [Gol12, Mck95, Mck09, RMP11, Sze94]. Space- [RMP11]. Spain [DIMV11, DMV12]. Spake [CFK^{+}91]. Spanish [And84, Hid12, Lea12, Paz03, TPD85]. spätere [Mei12b]. Special [Teu12, GMC12, Por19, MMB13]. species [Tia11]. Spectral [Hod12a]. spectrum [Kah84]. Speech [TB12, Cop17d]. Speed [Ano49]. sphere [BGL24]. Spider [Tur04]. Spillman [Kru05]. Spindleruv [BFG^{+}12]. spirit [Hod94e, Hod94f]. spiritual [Dys12a]. Spot [SCPC23]. Spots [Poo92, Poo91]. spread [Kah84]. Sprevak [Ano20, Jon17, Pet18]. Springer [Ano20, Kru05]. Springer-Verlag [Kru05]. squares [CWS^{+}24]. Stability [KW12, WWG12]. stage [Sol87]. Stamp [Gee12b]. Standard [Gar95, Gar09, ZM08, Moo03b]. Starch [LE91]. Starts [Ano12c]. State [HS82, Pri24, OS91]. Statement [Den12a, Den12b]. States [Bre13, Sha54, Tro93, Tro95]. Station [Tur20]. Stationary [RV12]. Statistical [Goo92, ZM08, Goo79b, GMT^{+}12]. Statistics [Tur41b, Goo79b]. Status [Moo03a]. Statutory [Hou12]. Step [Sie14]. still [Dal12a, Gal06]. Stochastic [HTG12]. Stored [TDCKW84]. Storia [Hod03b]. stories [Smi15b]. Story [Chr24, Cop25, Hod03b, Lav80, Lew78, Tur18, Whi87, DB04, Hea15, HS93, Tur21c, Chr22a]. Strahm [Nof17]. Strand [LP11]. strange [Syk92]. stravagante [Odi12]. Strawberries [Sch04b]. Street [Kru05]. striped [OS91]. stronger [Gam13]. Structural [DP02]. Structure [Kop95, Zas18]. Structures [DP02, LE91, LKE93, RV12, RAM95, TCP^{+}18, WWG12]. Struggle [And08, Cop12a, Kid06, Cop05a]. students [Bro97]. Studies [Goo79b]. study [GC17a, New03]. style [Nau93]. sublogarithmic [Sze94]. submarines [McG11]. Subway [Ste94]. succeeded [Dys12a]. Summer [Sof83]. Summit [CFK^{+}91]. Sunflowers [Ano12l]. super [GS12, Sie13]. super-Turing [GS12, Sie13]. Supplementary [AH85]. Surface [KW12]. Surfaces [LOM^{+}01]. Surprises [Hut95, Hut09]. Survey [Gol95, NW12, Rus89, Her88, Her95]. Swansea [BBLT06]. Switzerland [Ano20]. Symbol [Con12]. Symmetry [MC12b]. Symposium [AWL^{+}88, Bow53a, Den12c]. Synchronizability [IT12]. Synchronization [IT12]. syndrome [Jam06, OF03]. Synopsis [Lav12]. System [KW12, LE91, LKE93, RMP11, TB12, vL13, Dut10, GS12, Poo92, Shi14, Tur42a, WWG12]. Systems [CEL10, Del06, DKM^{+}24, HS82, LP11, Tur38c, Tur39, Tur65, Wie12, App12, CWS^{+}24, HP88a, LTS^{+}21, SNUM03].

Take [Pro17b]. Takes [Wha09]. Talking [RS03]. TAMC [ACL12]. Tape [Axe12, EH91]. tapes [IST^{+}10]. Tarragona [DIMV11]. Taster [Yap12]. taxi [Cle17]. Teacher [PA11a]. Teacher-Friendly [PA11a]. Team [Hod94e]. Technical [CFK^{+}91, Mis09, TB12]. Technology [AWL^{+}88, DKK^{+}98, Don01a, GF91, Gör95a, Wil97, Kah84, Mai06, CFK^{+}91]. Telecollaboration [Bro05]. teleprinter [GMT^{+}12]. tell [HP20]. Temperature [PSS11]. ten [Coo12d]. teorija [Tur60a]. Term [RMP11]. termination [CPR11]. terrible [Tur17]. Test [Ano90, Bac90, BBF03, Cra10b, Dew92, EG12, Eri03, EW17, Fre12b, Gon24b, Ham90, Hod09, Mas12, Sch12c, VFR^{+}12, War12, Wie12, Ack14, Bie12, Cro94, Fre12c, HF95, Hod12a, Lev17, Llo12, Nau86, Pat07, Pro17c, PA11b, Shi04, CAC14a, Zas18, BB12a, BB12b, BBF03, CP95, Cop03, CP04, CP09, Edm03, ERB08, EBR09, Gar95, Gar09, Gon22, Hod95b, Hor95, Hor09, Hum95, Hum09, KK09, Len95, Len09, Loe95, Loe09, Moo03a, Moo03b, Pat04, Pla09, Pro06, Rap03, SCA00, SCA03, SCT^{+}17, Sea95, Sea09, Sha12, Tra03, Var14, Wha09, Zde03]. Testery [Rob17]. Testing [PA11a, Sav24]. Tests [Pav17, SW10, Ste00, Ste03]. Teuscher [Kru05, Lov04]. Text [Day21, CFK^{+}91]. Textbook [Chr10]. Their [Ano88, AWL^{+}88, DJ12, IM13]. Them [Lav80, Rue07]. Theorem [Fra06, NT42, Zab95]. Theorems [CZ12]. Theoretical [HL02, Man90]. Theorie [Mei12b, Tur60a]. Theories [Roc12]. Theory [ACL12, AD12, BAC14, BFG^{+}12, CFK^{+}91, CM10, Deu85, Dow12a, Gas16, Sha14, Tau61a, Tau61b, Tau63b, Tur60a, WTP^{+}06, WS00, WBM17, Blu14, Cas06b, DIMV11, DMV12, Dow14a, FHM14, Han12, Joy00, Mar11b, McG11, Mei12b, Moo15, NT42, PA13, Ric17, Sha09a, Soa16, Ste90, Tur48b, Tur96, Zie09, Tau63c, PA13]. There [Par17]. these [Gal06]. Thesis [AD12, Cot03, Dav06a, Dow12a, Fef06, Ner14, Pic11, Szu12, App12, BA05, CS19, Gal06, Sha12, Tay98, Par17, Yao03, vL13, Shi14]. Things [Kru05]. Think [Den04, May52, Wat95, Wat09, Tur60a, Tur91, TvN99]. Thinker [Kru05, Teu04a, Lov04]. Thinking [Aho12, Eps95, EBR09, Eps09, Jon16, Jor07, Nau86]. Thinning [Cai12]. Third [CLS07, MBS11]. Thomas [CFK^{+}91, Nof17]. Thomason [Kru05]. Thought [Bow53a, Gon22, Gon24b, MC96, Bre12c, Coc12, FF63, Gou99, HP20, Hol18]. Thousands [Cha16]. Three [BVE11, Bul21, Por19, Sal12, Tia11]. Three-Dimensional [BVE11]. Thus [CFK^{+}91]. Time [Axe12, RV12, RMP11, Rog10, Whi12, CWS^{+}24, Hod12a]. Time-Dependent [RV12]. Time-Discrete [RMP11]. Times [Bau12, LH83, Wel06a]. Titanic [Coo12f]. todas [Hid12]. Today [Ano12c, Dys12a, Hod12c]. Todd [Mad12]. Toister [Con22]. Tomography [BVE11]. Tony [Met19]. Too [Coo06a, Lea12, Lea05]. Topics [Tau62, LTM^{+}51]. Toronto [Sof83]. Total [Sch12c]. Tour [Ano06a, Rue07, Lip11, Pet08]. toxic [McG12]. Tracks [Ano89]. Trade [Chr22b]. trail [HL02]. Transfinite [Wel14]. Transformation [BLA^{+}11]. Transformations [Bul15, TT56]. Transgressive [SCPC23]. Transient [LKE93]. Transients [RMP11].

Transition [OS91]. Translation [CFK^{+}91]. Transport [BB16]. Trap [Bry22]. treasury [FF91]. Treatise [CV13c, Tur40, Tur99]. Treatment [Bro09, Nau09]. Trends [BFG^{+}12, WTP^{+}06]. Trieste [PA13]. triumphant [McG11]. trousers [Tur17]. True [PR17]. Truly [Sch12c]. Truth [Hod94f]. Tumble [Pit23]. Tumours [Mur12]. Tunny [Cop17g, GMT^{+}12]. Turin [Con22]. Turing [AW77, AH85, And08, Ano99, Ano12d, Ano12b, Ano12c, Ano13, Ano14, Ano20, Arb95, Ash87, Asp84, AWL^{+}88, Avi14, Blo98, Bre12a, Bri90, CK84, CK02, CFK^{+}91, Cha94, Chr15, Chr16, Chr22b, Chr22a, Con22, Coo06a, CDL12, CP00, Cop25, Dal12b, Deu85, Don01a, Dys12a, EH91, Ers84, Fai12, Fie15, Gan54, Gee12a, Gin19, GR12, Gol12, Goo92, GKO95, Ham16, Hay17, Her88, Her95, Hid12, Hoc87, Hod94a, Hod94b, Hod06a, Hod12c, Hof83, Hof85, Hou12, Jon17, Ken89, Kid06, KH19, Kru05, Lav12, Lea19, LH83, Lie11, Lip11, Lov04, Mei12a, MMB13, Owe12, Par17, Pet18, RTM04, Rid84, Rus89, Sal12, Set17, Sev12, Shi14, Shu87, Smi14, CAC14b, Sut85, Swa13a, Swi19, Tim04, THWV88, TDCKW84, Und13, Wel06a, Yao03, Zie09]. Turing [vL13, Abr11, Ack14, Aga01, AB00, AKS11, ALdlP20, Ano89, Ano96, Ano00a, Ano00b, Ano01, Ano02, Ano06a, Ano06b, Ano09a, Ano10b, Ano11c, Ano11b, Ano12b, Ano12a, Ano12e, Ano12f, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano15a, Ano15b, Ano19a, Ano19b, Ano19c, Ano19d, Ano20, Ano21, App12, AD12, Asp80, AB12, AB14, AG11, Axe12, Bac90, BLvT11, BLvT12, BB12a, WBM17, BVE11, BAC14, Bar98, BLA^{+}11, Bau12, Bea89, BFP07, Bec12, BCT10, BA05, Ben97, Ben12, Ber16, BB94, Bia79, Bie12, BSK^{+}15, Blö12, Blu04, Blu14, vEB12, Bod49, Bol84, Boo06a, Boo06b, Boo52, BGL24, BB12b, BC17, Bra13, Bre12b, Bre12c, Bre13, BBF03, Bro97, Bro05, Bro13, BB16, Bro09, Bry22, Buh14, BDD15, Bul21, CK12a, Cap05, ÇG12]. Turing [CZ12, CD77, CD86, CD17, Car10, Cas06a, Cas01, Cas13, Cer04, CWS^{+}24, CEL10, Che93, Cho95, Cho09, Cho12, Chr10, Chr13, Chr16, Chr21, Chu13, CP12a, CM96, CS12, Cla72, CBB12, Cle17, Coa13, Coc12, CM10, CL02, Coo12b, Coo12c, Coo12d, Coo12e, Coo12a, Coo12f, CV13a, CvL13, CV13b, CV13c, CH16, CP95, CP96, CP99, Cop00a, CP00, Cop00b, CP01, Cop03, CP04, Cop04, Cop05a, CP09, CP10, CP12b, CS11b, CGLWVR12, Cop12a, Cop12b, CL17b, CBSW17, Cop17h, CL17a, Cop18, CS19, CP23, Cop23, CH83a, CH83b, CG87, Cor07, Cor17, Cot03, Cow19, Cra10b, Cro94, Cuc12, Cur65, Dal12a, Dav13, Dav00, Dav06a, Dav06b, Dav12, Dav18, Daw16, Day12a, Day12b, Day12c, Day13, Day21, DW16, DK90, Del06, Dew89]. Turing [Dew92, Dew93, DT12, Dic13, DKM^{+}24, Dil05, DC11b, DC12, DC13, Don01b, Don14, DDL01, Dow12a, Dow13, DH10, Dow14a, Dow14b, Dow14c, Dow17, Dow12b, Dre10, DJ12, DL06, Dut10, Dys12a, Dys12b, Dys12c, EGW04, Edm03, EG12, Ell13, Ell19, EH91, ERB08, EBR09, FH15, Fai10a, Fai10b, Fai11, Fef95, Fef06, FOO71, Fis15, Fis17, FB17, Flo17, For12, FGG^{+}24, Fre86, FRT14, Fre12b, Fre12c, Fri05, Fu12, Fur12, Gal06, GMC12, Gam13, Gar95, Gar09, GAM11, Gee12b, GS12, Ghe11, Gla01, Gla03, Gla04, GR12, Gla12, Gon22, Gon23a, Gon23b, Gon24a, Gon24b, Goo79b, Goo84, Goo00, Gör91, GKO95, Gör95b, Got96, Gou99, GC17a, GC12b, GC12a, GC12c, GC12d, GG12, GG13, GG17, GC17b, Gub86, Hae12]. Turing [Hai14, Hai17, HP20, Hal13, Hal14, Ham16, Ham90, HL02, Han12, Har03, Har12a, HM92, Har12b, Har47, Has95, HfB98, HF95, Hej07, Hel17, Hen11, Her98, Hew13, Hic08, Hid12, Hil93, Hil91, Hin17, Hoc87, HG89, Hod83a, Hod83b, Hod85, Hod88, Hod89a, Hod89b, Hod92, Hod95a, Hod95b, Hod97a, Hod97b, Hod97c, Hod99, HP00, Hod00, Hod01, Hod02a, Hod02b, Hod03a, Hod03b, Hod04a, Hod04b, Hod08a, Hod08b, Hod09, Hod12d, Hod12b, Hod12e, Hod12c, Hod12f, Hod12a, Hod14, HM96, HH84, Hol86, HC87, HP88a, HP88b, HC88, HMRC88, Hol90, HS14, Hop84, Hop12, Hor95, Hor09, HSD09, HAC^{+}85, HH90, Hum14, Hum95, Hum09, Hym12, IT12, Irv04, IM13, IST^{+}10, Jac12, Jac11, Jea12, Jor07, Kah12, KP02, Kan12, Kar95, KvLP88, KW12, Ken17, Kid96, Kie12, Kle95]. Turing [KA96, Kon12, Kov03, KGJ^{+}24, Dea98, KK09, LP11, Las98, LL12, LCKBJ12, Lea05, Lea07, Lea12, Lea17, LGB11, Lei01, Lem04, Lem12, Len95, Len09, LE91, LKE93, Lev17, Lev06, LTS^{+}21, LOM^{+}01, Lie11, Lip12, Lis12, LGS22, Liv02, Llo12, Loe95, Loe09, Lol13, Lom05, Lon09, Lov04, Luc95, Luc09, ŁW11, Mac12a, Mac12b, MBC06, MC12a, Mai06, Mai07, Mal87, Mar13a, Mar13b, MD11, Mar11a, Mar11c, Mar11d, Mas12, May61, May01, Mei12a, Mei12b, Mic15, Mic80, Mic08, III14, MC96, Mon19, Moo15, Moo03a, Moo03b, MJ84, Müh09, MS17, Mur12, Nan03, Nau09, Nau86, Nau93, NW12, NA06, Ner14, New55, New12, New03, Nic17, Nof17, Nor14, Num05, OF03, O’N23, O’R12, OG12, Odi12, OW12, OSZ03, OS91, Pap03, Pap12, PR17, Par14]. Turing [Pat04, Pat07, PSS11, Pav17, Paz03, PC06, Pea19, dBPZM10, PC88, Pet08, Pic03a, Pic03b, Pic11, Pip04, Pip05, Pit14, Pit23, Poo91, Poo92, Por19, Pos23a, Pos23b, PR10, Pra01, PA11a, Pra95, Pro06, Pro15, Pro17d, Pro17b, Pro17c, PA11b, QSW11, RV12, Ran72a, Ran72b, Ran00, Ran12, Ran17a, Rap03, RM00b, RM01, RR12, Reg23, Rei12, Res17, Ric06, Ric17, Rig91, Rob97, RMP11, RAM95, Sal04, Sal12, ST23, Sau93, Sav24, SCA00, SCA03, Sch04a, Sch12a, Sch12b, Sch88, SCT^{+}17, SGV94, ST12, Sch12c, Sea95, Sea09, SW10, Sha14, Sha12, Sha09b, Sha54, She12, Shi04, Shi12, SS91, Sie95, Sie13, Sie12, SCPC23, Smi02, Smi05, Soa14, Soa16, Sol87, SS15, Sor05, Spr17, CAC14a, Ste00, Ste03, Ste12b, Ste17]. Turing [Ste90, Ste94, Str99, Str15, Sut13, Swa13b, Swa17, Swi04, Syk92, Sze94, SG18, Szu12, TCP^{+}18, Tau56, Tay98, Ter11, Teu04a, Teu04b, Teu12, The87, Tho18, Tia11, Tim04, Tra03, Tro93, Tro95, Tru11, Tsa19, Tur40, Tur42b, Tur59, Tur72, TWCD86, TG95, Tur99, Tur00, Tur01a, TW05, Tur12, TB12, TW12, Tur15b, Tur15a, Tur21c, UST^{+}10, Unk84, VFR^{+}12, Var14, Var17, Vin13, Vos13, War12, WS16, Web12, WWG12, Wel12, Wel02, Wel04, Wel06b, Wha09, Whi87, Whi91, WW17, Wie12, Wil80, WB12, Wol17, WS00, Wri16, Yan12, Yap12, Zab95, Zab12, Zab17, Zas18, Zde03, dC11a, de 12, dSAL^{+}13, Bow53b, Ano90, HS82, Ano09b, Chr22a, Kil14a, Kil14b, May61, Wil10, Ano04, Bea84, Bla14, CK12b, Cer85, Chr24, Día12, Fef99, Gas16, GC12e]. Turing [Hod06b, Joh15, Sha00, Hai16]. Turing-Complete [Pit23]. Turing-like [DDL01]. Turing-Powerful [LP11]. Turing-Type [LOM^{+}01].

Turing-Universal [DL06, QSW11]. Turinga [Hod02b]. Turingmaschine [FOO71]. Turings [Gla12, Mei12b, ST12]. Turmites [Ano89]. Tutte [Hai17]. twentieth [B^{+}11, MHR80]. twenty [Ash87, Tay98]. twenty-first [Tay98]. Two [Ano89, Bau12, HS82, Pra95, Sha54, Ste00, Ste03, Ano20, McG11, AWL^{+}88]. Two-Dimensional [Ano89]. Type [LOM^{+}01, Rou18, Tia11, Tur48b]. types [NT42]. tyranny [Sut12].

U [Gla03, Tur03, DB04]. U-boat [DB04]. Überleitung [Hod94k]. Ufer [Hod94c]. Ugly [Pip04, Pip05]. UK [BBLT06, CDL12, Jon17, Ano19d, Gla04, Man90]. ultra [DB04, GW14, Lew78, Ran17b]. unbreakable [Kah84]. Uncertainty [Buz12]. Uncomputable [CS11b]. undecidable [Dav65]. Understanding [Nau93, Cro94, Zen13, Poo92]. Underworld [Wat12g]. unfinished [Sch88]. uniform [OS91]. unique [Ive15]. United [Tro93, Tro95]. unity [Lei01]. Universal [AG11, CK02, CL17b, Dav18, Deu85, DL06, KP02, Kil14b, NW12, QSW11, Rus89, Sha54, Aga01, CK12a, Cho12, Cop17h, Dav00, Dav12, FOO71, HP20, Kil14a, Mei12a, Nau93, ST23, Smi02, Wat12m, Arb95, Blo98, CP00, Her88, Her95, RTM04]. Universality [Del06, Mar11d, PSS11, Sut13]. Universe [Dys12c, MC12b, CSS17, HP15, Sen21, Zen13, Sal12, CK12b, Día12, GC12e, Bla14]. universelle [FOO71]. universellen [Mei12a]. University [Ano51, Ano20, CFK^{+}91, Fai12, Hai16, Jon17, Kru05, Rus89, Sal12, Shi14, Smi14, vL13]. UNIX [CH83a, CH83b]. Unknown [ALdlP20, WS16]. Unmögliche [BT12]. Unorganised [Rou18]. unorganized [Web12]. unpublished [BFP07]. Unsolvability [Fra06]. unsolvable [Dav65, Tur54]. until [Hod12c]. Untold [Chr24, DB04]. Unwin [Shu87]. uomo [Cap05]. Upper [Kru05]. USA [CS11a, Kru05]. use [Hod03a, Tur42a]. USENIX [Sof83]. Using [PA11a, Rou18, GAM11, HH84, HP88a, HP88b, Hol90, HH90]. Utopia [Gon23b].

V [Ano99, Lie11, Tau63b]. valid [Gal06]. validating [Ric17]. Vardi [Var17]. Vater [Dys12a]. Vegetative [KW12]. Venice [Ive15]. Venus [Smi05]. verbal [Shi04]. Vergleich [Lie11]. Verification [OW12]. Verifikation [OW12]. Verifying [LP11]. Verlag [Kru05]. Vermeer [Kru05]. Versus [Tim04, Ano12f, Cop00a, Hew13]. very [Jam06]. Verzögerung [Hod94l]. VI [Tau63c]. via [Day21, FRT14]. victim [Mac12a]. video [Ano13]. viele [Hod12c]. Viewpoint [BDD15, Mic15]. Views [Cop00a]. viii [Hod06a]. Vintage [Lav80, Sal12]. Vision [Ber16, Nic17, Tho18, PA11b]. Visionary [Cow19]. Visions [Sum14]. Visit [Gla01, Tro93, Tro95, Tur01c]. Vita [Hod03b]. Vite [B^{+}11]. Vogelfrei [Hod94m]. vol [Boo52]. Volume [Kru05, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Voynich [Bau12]. vs [Bur11].

W [KP02, Shi14, vL13]. Wager [KK09]. Wahrheit [Hod94f]. Wahrnehmung [BT12]. walk [Mai06]. walked [Hol18]. Walks [Jac11]. Wanted [DW16]. War [AWL^{+}88, Goo79b, Goo92, Kah84, Kru05, Lew78, O’N23, Pri21, Sal04, Wat12c, BH03, Coa13, Cor17, DB04, Sch04a, RA03, RA04]. wartime [Hen11]. Was [AWL^{+}88, CFK^{+}91, DK90, Par14, Pit14, Tur18, Bro09, Dav13, Ell19, HP20, Hod12c, Dea98, Mac12a]. Washington [Tur42b]. Watching [Swi04, GMC12]. Water [KW12, Zas18, TCP^{+}18]. waterside [Ive15]. Watson [CFK^{+}91]. Wavefronts [CEL10]. way [Poo92]. Weak [IT12]. Weaving [Wat12n]. Web [Jor07, Wat12o, Wat12n]. Weight [Liv02]. Welchman [GW14]. Well [DJ12]. Well-Orders [DJ12]. Welt [Pil12]. Were [Bri95, Bri09, Ire17]. West [Kru05]. Western [Bea84, Cer85, Sut85, Bol84]. Wheatstone [BCT10]. Where [Mis09, Blu04]. Who [Chr24, Coo06a, DV09, Hau03, Lav80, Lea12, Sha14, Ell19, Hea15, Lea05, Lea07, Lem04, Lem12, Moo14, Smi10, VB15, Ano96]. whole [CSS17]. Whom [DV09]. Wide [Cop00a]. wie [Spr12]. Wiener [AWL^{+}88]. Wiles [B^{+}11]. Wiley [Lip11, Wil10]. Wilkes [Ben12]. Wilkinson [TW05, TW12]. Will [EG12, Pro17b, Tsa19, Llo12]. William [Ano99]. Wilson [Ano20, Jon17, Pet18]. Wine [Yap12]. winners [Fie15]. Wins [Ano14, Bie12]. Winter [USE83]. wir [Spr12]. Wirkung [Gla12]. wisdom [Rob97]. without [Kon12]. Wittgenstein [GKO95, Pra95]. women [Hea15]. Womersley [Pri24]. Won [Lip12, Var17]. word [Boo52, III14, Tur50c]. words [Sal12]. Work [Ano14, Chr15, CvL13, Goo79a, Goo92, Wil80, Avi14, CB17, Goo79b, Hod12c, Lei01, Mac12b, Sch04a, Sch12b, ST12, Vos13, Web12]. Working [Hil91, Pri24]. Works [AWL^{+}88, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Kid96, Tur01a, de 12]. World [Ano16, AWL^{+}88, CLS07, CDL12, CH16, Cop12a, Goo92, Gur95, Lav12, Mai07, RA03, RA04, FF91, Ire17, Jac12, LCKBJ12, Pil12, Goo79b, Kah84, Sal04]. Worth [LL12]. Worthy [AWL^{+}88]. Would [Hod04b, Var14, Var17, McG11]. Writing [LL12, Whi91]. Writings [Cop04, Tur87]. WW2 [Don14].

X [Tur18]. xii [KP02]. xiv [Rus89]. xv [Jon17, vL13]. XXIst [GGZ06]. XXXVII [Goo79b].

Yanai [Con22]. Yates [Fef99]. Year [Ano12a, Gol12, Hae12, Ano12c, Und13]. Years [Bau12, SCA00, SCA03, Sea95, Sea09, Ash87, Gal06, MMB13]. yes [Pit14]. York [KP02, Kru05].

Z [Tur18]. Zeitgeist [CD17]. Zero [Cha94, Gol95]. Zero-Knowledge [Gol95]. zeros [Leh70]. zeta [Leh70, Leh56, Tur43, Tur53]. zeta-function [Leh70, Leh56, Tur43, Tur53]. zur [Mei12b]. Zuse [Lie11, Lie11, MC12a]. zwischen [Dys12a, GKO95].

References

Akman:2000:ATA

[AB00] Varol Akman and Patrick Blackburn, editors. Alan Turing and artificial intelligence. Kluwer Academic Publishers, Dordrecht, The Netherlands, 2000. CODEN JLLIEN. ISSN 0925-8531 (print), 1572-9583 (electronic). ii + 391–509 pp. J. Logic Lang. Inform. 9 (2000), no. 4.

Avigad:2012:CAL

[AB12] Jeremy Avigad and Vasco Brattka. Computability and analysis: the legacy of Alan Turing. arxiv.org, June 2012. URL http://

adsabs.harvard.edu/abs/2012arXiv1206.3431A.

Avigad:2014:CAL

[AB14] Jeremy Avigad and Vasco Brattka. Computability and analysis: the legacy of Alan Turing. In Downey [Dow14c], chapter 2, pages 1–47. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Abramson:2011:DIT

[Abr11] Darren Abramson. Descartes’ influence on Turing. Studies in History and Philosophy of Science Part A, 42(4):544–551, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000689.

Ackerman:2014:BTT

[Ack14] Evan Ackerman. A better test than Turing [news]. IEEE Spectrum, 51(10):20–21, October 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Agrawal:2012:TAM

[ACL12] Manindra Agrawal, S. Barry Cooper, and Angsheng Li, editors. Theory and Applications of Models of Computation: 9th Annual Conference, TAMC 2012, Beijing, China, May 16– 21, 2012. Proceedings, volume 7287 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-29951-2. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/
content/2738n2330255x474/;   http://www.springerlink.
com/content/978-3-642-29951-3/.

Arrighi:2012:PCT

[AD12] Pablo Arrighi and Gilles Dowek. The physical Church–Turing thesis and the principles of quantum theory. International Journal of Foundations of Computer Science (IJFCS), 23(5):1131–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002. URL http:

//www.gbv.de/dms/goettingen/360423353.pdf;   http://
www.loc.gov/catdir/enhancements/fy0611/2001048396-d.
html.

Axelsen:2011:SEU

[AG11] Holger Bock Axelsen and Robert Glück. A simple and efficient universal reversible Turing machine. In Dediu et al. [DIMV11], pages 117–128. ISBN 3-642-21253-0 (softcover).

LCCN ???? URL http://www.springerlink.com/content/
j2778305m6846x87/.

Agar:2001:TUM

[Aga01] Jon Agar. Turing and the universal machine: the making of the modern computer. Revolutions in science. Icon, Cambridge, UK, 2001. ISBN 1-84046-250-7. iv + 153 pp. LCCN QA76.2.T87 A43 2001.

Alton:1985:SCP

[AH85] Jeannine Alton and Peter Harper. Supplementary catalogue of papers and correspondence of Alan Mathison Turing, FRS (1912–1954) material additional to CSAC 53/7/77. Technical Report CSAC 104/1/85, Contemporary Scientific Archives Centre, London, UK, 1985. 16 pp.

Aho:2012:CCT

[Aho12] Alfred V. Aho. Computation and computational thinking. The Computer Journal, 55(7):832–835, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/832.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Aly:2011:TIR

[AKS11] Shaban Aly, Imbunm Kim, and Dongwoo Sheen. Turing instability for a ratio-dependent predator–prey model with diffusion. Applied Mathematics and Computation, 217(17):7265–7281, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300311001949.

Anguera:2020:TGU

[ALdlP20] Aurea Anguera, Juan A. Lara, and F. David de la Peña. Turing: The great unknown. Foundations of Science, 25(4):1203– 1225, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-019-09596-6.

ApSimon:1966:CIP

[AMKM66] H. G. ApSimon, W. D. Maurer, P. J. H. King, and C. H. R. Morris. Correspondence: An impossible program. The Computer Journal, 8(4):329–330, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic).  URL http://comjnl.oxfordjournals.org/
content/8/4/329.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/329.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/330.tif. See [Str65, ApS65, HLOS65,

BSPI65].

Anderson:1964:MM

[And64] Alan Ross Anderson, editor. Minds and machines. Contemporary prospects in philosophy series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1964. viii + 114 pp. LCCN Q335.5 .A5.

Anderson:1984:CSM

[And84] Alan Ross Anderson, editor. Controversia sobre mentes y máquinas. (Spanish) [Controversy over Minds and Machines]. Tusquets Editores, Barcelona, Spain, 1984. ISBN 84-7223-624-2. 171 pp. LCCN ???? Translated by Francisco Martín.

Anderson:2008:ATA

[And08] David Anderson. Alan Turing’s Automatic Computing Engine: The Master Codebreaker’s Struggle to Build the Modern Computer, edited by B. Jack Copeland. History and Philosophy of Logic, 29(4):389–396, November 2008. ISSN 0144-5340 (print), 1464-5149 (electronic).

Anonymous:1946:CPO

[Ano46] Anonymous. A.C.E. project – origin and early history. Technical Report DSIR 10/385, National Physical Laboratory,

Teddington, UK, 1946. URL http://www.AlanTuring.net/
aceearlyhistory.

Anonymous:1949:RCH

[Ano49] Anonymous, editor. Report on a Conference on High Speed Automatic Computation, June 1949. University Mathematical Laboratory, Cambridge University, Cambridge, UK, 1949. LCCN ???? Inaugural conference of the EDSAC computer at the Mathematical Laboratory, Cambridge, UK.

Anonymous:1951:MUC

[Ano51] Anonymous, editor. Manchester University Computer: Inaugural Conference held at the University on the 9th, 10th, 11th and 12th July, 1951. Tillotsons, Bolton, UK, 1951. LCCN ????

Anonymous:1988:ERH

[Ano88] Anonymous. Errata: Reviews: Hartree: Calculating Machines: Recent and Prospective Developments and Their Impact on Mathematical Physics and Calculating Instruments and Machines, 10(1) 93. Annals of the History of Computing, 10(3):234, July/September 1988. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a3234.pdf;   http://www.computer.
org/annals/an1988/a3234abs.htm. See [AWL^{+}88].

Anonymous:1989:TDT

[Ano89] Anonymous. Two-dimensional Turing machines and turmites make tracks on a plane. Scientific American, 261(3):124–??, September 1989. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1990:TTP

[Ano90] Anonymous. ‘Turing Test’ Prize. Science, 248(4963):1610, June 29, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/248/
4963/1610.3.full.pdf.

Anonymous:1996:QIO

[Ano96] Anonymous. Qui a inventé l’ordinateur?: grands ingénieurs: Alan Turing, John Mauchly, John P. Eckert, John Atanasoff, John von Neumann. (French) [Who invented the computer? Great engineers: Alan Turing, John Mauchly, John P. Eckert, John Atanasoff, John von Neumann], volume 36 of Les Cahiers de Science et vie (Paris). Excelsior, Paris, France, 1996. ISBN ???? ISSN 1157-4887. 96 pp. LCCN ????

Anonymous:1999:AAM

[Ano99] Anonymous. ACM Alan M. Turing Award: William V. Kahan. World Wide Web document, February 8, 1999. URL http://

www.acm.org/awards/turing_citations/kahan.html.

Anonymous:2000:AMT

[Ano00a] Anonymous. Alan Mathison Turing. World-Wide Web

site., 2000. URL http://www-groups.dcs.st-and.ac.uk/~
history/Mathematicians/Turing.html.

Anonymous:2000:AT

[Ano00b] Anonymous. Alan Turing (1912–1954). Minds and Machines, 10(4):461–??, November 2000. CODEN MMACEO. ISSN 0924-6495 (print), 1572-8641 (electronic). URL http://link.

springer.com/article/10.1023/A:1017354226375.

Anonymous:2001:PTP

[Ano01] Anonymous. Potential for Turing patterns. Science, 291(5512): 2271, March 23, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/291/5512/2271.4.full.

Anonymous:2002:ETF

[Ano02] Anonymous. Exhibits: The Turing files. Science, 297(5578):19, July 5, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
297/5578/19.3.full.pdf.

Anonymous:2004:BRT

[Ano04] Anonymous. Book review: Turing (a novel about computation): Edited by Christos H. Papadimitriou. The MIT Press. Cambridge, MA. (2003) 284 pages. $24.95. Computers and Mathematics with Applications, 47(8–9):1486, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122104901514.

Anonymous:2006:RTT

[Ano06a] Anonymous. Reviews: A tour of Turing. Scientific American, 294(2):98–100, February 2006. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v294/n2/
full/scientificamerican0206-98.html;   http://www.nature.
com/scientificamerican/journal/v294/n2/pdf/scientificamerican0206-
98.pdf.

Anonymous:2006:TPM

[Ano06b] Anonymous. Turing patterning in the mouse hairs. Science, 314(5804):1349, December 1, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/314/5804/1349.8.full.pdf.

Anonymous:2009:ATP

[Ano09a] Anonymous. Adding a Turing pattern reaction. Science, 324 (5928):687, May 8, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/324/5928/687.5.full.pdf.

Anonymous:2009:ATG

[Ano09b] Anonymous. Alan turing gets belated apology. New Scientist, 203(2726):7, September 16, 2009. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S026240790962443X.

Anonymous:2010:HLB

[Ano10a] Anonymous. HP to liberate Bletchley archive. Web site, June 9, 2010.

Anonymous:2010:TME

[Ano10b] Anonymous. Turing model explained. Science, 329(5999): 1569, September 24, 2010. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/329/5999/1569.5.full.pdf.

Anonymous:2011:PAN

[Ano11a] Anonymous. PILOT ACE — NPL’s legacy. Web site., March

3, 2011. URL http://www.npl.co.uk/news/pilot-ace-npls-

legacy. Today’s ability to multi-task on our computers is taken for granted, but it all started with NPL’s Pilot ACE Computer and the genius of mathematician Alan Turing.

Anonymous:2011:TPS

[Ano11b] Anonymous. Turing papers saved for Bletchley Park. BBC News,

February 25, 2011. URL http://news.bbc.co.uk/today/hi/
today/newsid_9407000/9407249.stm.

Anonymous:2011:TP

[Ano11c] Anonymous. Turing patterns in 3D. Science, 331(6022):1239– 1241, March 11, 2011. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/331/6022/1239.7.full.pdf.

Anonymous:2012:ATYa

[Ano12a] Anonymous. 2012 — the Alan Turing year. Trends in Cognitive Sciences, 16(9):447–448, September 2012. CO-DEN TCSCFK. ISSN 1364-6613 (print), 1879-307X (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136466131200174X.

Anonymous:2012:ATB

[Ano12b] Anonymous. Alan Turing: (1912–1954). Biografie eines Genies. (German) [Alan Turing: (1912–1954). Biography of a genius ]. Spektrum der Wissenschaft (German translation of Scientific American), ??(6):80, ???? 2012. CODEN SPEKDI. ISSN 0170-2971.

Anonymous:2012:ATYb

[Ano12c] Anonymous. Alan Turing Year starts today. Web site, January 1,

2012. URL https://i-programmer.info/news/82-heritage/
3550-alan-turing-year-starts-today.html.

Anonymous:2012:ATA

[Ano12d] Anonymous. Alan Turing’s ACE. Web site, May 28,

2012. URL https://i-programmer.info/history/machines/
11-an-ace-of-a-machine.html.

Anonymous:2012:CCH

[Ano12e] Anonymous. Computer composer honours Turing’s centenary. New Scientist, 215(2872):7, July 7, 2012. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0262407912617217.

Anonymous:2012:KVT

[Ano12f] Anonymous. Kasparov versus Turing. Manchester University Web news story, June 26, 2012.

Anonymous:2012:MM

[Ano12g] Anonymous. Manchester Mark 1. Web encyclopedia article., 2012. Discusses Alan Turing’s role in the design of the Mark 1, and in writing an improved version of a program for finding Mersenne primes.

Anonymous:2012:MNR

[Ano12h] Anonymous. Mathematician Norman Routledge on Alan Turing. Web news story., July 10, 2012. URL http://www.

scientificcomputing.com/news-DA-Mathematician-Norman-
Routledge-on-Alan-Turing-070912.aspx.

Anonymous:2012:T

[Ano12i] Anonymous. Turing at 100. Nature, 482(7386):440, February 22, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v482/n7386/full/482440a.html.

Anonymous:2012:TCB

[Ano12j] Anonymous. Turing centenary: Is the brain a good model for machine intelligence? Nature, 482(7386):462–463, February 22, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v482/n7386/full/482462a.html.

Anonymous:2012:TP

[Ano12k] Anonymous. Turing on parade. New Scientist, 215(2880):4, September 1, 2012. CODEN NWSCAL. ISSN 0262-4079 (print),

1364-8500 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0262407912622325.

Anonymous:2012:TS

[Ano12l] Anonymous. Turing’s sunflowers. Manchester Museum of Science and Industry Web site, 2012. URL http://www.

turingsunflowers.com/about/why. Crowd-sourced project to

grow sunflowers and report their spiral counts online; results are expected to be posted by late summer 2012.

Anonymous:2013:ATP

[Ano13] Anonymous. Alan Turing’s Pilot Ace computer — video.

Web site., April 12, 2013. URL http://www.guardian.co.uk/
science/video/2013/apr/12/alan-turing-pilot-ace-computer-
video.

Anonymous:2014:ATH

[Ano14] Anonymous. Alan Turing: His Work and Impact wins prestigious PROSE Award. ScientificComputing Web site, February 7, 2014.

URL http://www.scientificcomputing.com/news/2014/02/
alan-turing-his-work-and-impact-wins-prestigious-prose-
award.

Anonymous:2015:BCB

[Ano15a] Anonymous. British code breaker Alan Turing’s notebook goes to auction. Scientific Computing, ??(??):??, January 26, 2015. CO-DEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.scientificcomputing.com/news/2015/01/
british-code-breaker-alan-turings-notebook-goes-auction.

Anonymous:2015:TRD

[Ano15b] Anonymous. Turing reaction-diffusion model confirmed. Scientific Computing, ??(??):??, September 22, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.scientificcomputing.com/news/2015/09/
turing-reaction-diffusion-model-confirmed.

Anonymous:2016:RWF

[Ano16] Anonymous. Restoring world’s first recorded computer music. Scientific Computing, ??(??):??, September 2016. URL http:

//www.scientificcomputing.com/news/2016/09/restoring-
worlds-first-recorded-computer-music.

Anonymous:2019:ATB

[Ano19a] Anonymous. Alan Turing banknote concept. Bank of Eng-

land Web site., 2019. URL https://www.bankofengland.co.
uk/banknotes/polymer-50-pound-note.

Anonymous:2019:ATCa

[Ano19b] Anonymous. Alan Turing collection. London Mathematical Society press release., October 18, 2019. URL https://www.lms.

ac.uk/publications/alan-turing-collection-0.

Anonymous:2019:ATCb

[Ano19c] Anonymous. Alan Turing collection. Wiley Web site., October

18, 2019. URL https://londmathsoc.onlinelibrary.wiley.
com/doi/toc/10.1112/(ISSN)1460-244X.alan_turing_collection.

Anonymous:2019:FMC

[Ano19d] Anonymous. Father of modern computing becomes the new face of UK’s 50-pound note: Facts on Alan Turing. India Today,

??(??):??, July 16, 2019. URL https://www.indiatoday.in/
education-today/gk-current-affairs/story/alan-turing-
uk-50-pound-note-father-of-modern-computing-1569994-
2019-07-16.

Anonymous:2020:RTC

[Ano20] Anonymous. Review of two collections of essays about Alan Turing: Copeland, Jack, Jonathan Bowen, Mark Sprevak, Robin Wilson, and others, The Turing Guide, Oxford University Press, Oxford, 2017. 546 pages, Paperback, $29.95. ISBN 978-0-19-874783-3. Floyd, Juliet and Alisa Bokulich (eds.), Philosophical Explorations of the Legacy of Alan Turing: Turing 100, Springer International Publishing, Cham, Switzerland, 2017. 361 pages, Hardcover, $139.99. ISBN 978-3-319-53278-3. Cryptologia, 44 (1):82–86, 2020. CODEN CRYPE6. ISSN 0161-1194 (print),

1558-1586 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/01611194.2019.1650846.

Anonymous:2021:FPN

[Ano21] Anonymous. Fifty pound note with a portrait of Alan Turing in circulation. Web site, June 23, 2021. URL https:

//marketresearchtelecast.com/fifty-pound-note-with-a-
portrait-of-alan-turing-in-circulation/82673/.

Appel:2012:ATS

[App12] Andrew W. Appel, editor. Alan Turing’s systems of logic: the Princeton thesis. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15574-7 (hardcover). xv + 142 pp. LCCN QA9.2 .T86 2012. Reproduction of Turing’s thesis [Tur38c].

ApSimon:1965:IP

[ApS65] H. G. ApSimon. “An impossible program”. The Computer Journal, 8(1):72, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/72.full.pdf+html. Re-

buttal to [Str65].

Arbib:1995:UTM

[Arb95] Michael A. Arbib. From Universal Turing Machines to selfreproduction. In Herken [Her95], pages 161–172. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.

springer.com/book/10.1007/978-3-7091-6597-3.

Alesso:2008:CPD

[AS08a] H. P. Alesso and C. F. (Craig Forsythe) Smith. Connections: patterns of discovery. Wiley-Interscience, New York, NY, USA, 2008. ISBN 0-470-11881-4 (hardcover). xiii + 207 pp. LCCN

T58.5 .A54 2008; for. URL http://www.loc.gov/catdir/
enhancements/fy0741/2007017344-d.html;   http://www.
loc.gov/catdir/enhancements/fy0806/2007017344-b.html;
 http://www.loc.gov/catdir/toc/ecip0716/2007017344.html.

Alesso:2008:CI

[AS08b] H. Peter Alesso and Craig F. Smith. Connecting intelligence. In Connections: patterns of discovery [AS08a], chapter 9, pages 155–173. ISBN 0-470-11881-4, 0-470-19153-8 (e-book). LCCN

T58.5 .A54 2008; for. URL http://www.loc.gov/catdir/
enhancements/fy0741/2007017344-d.html;   http://www.
loc.gov/catdir/enhancements/fy0806/2007017344-b.html;
 http://www.loc.gov/catdir/toc/ecip0716/2007017344.html.

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor. ACM Turing Award Lectures: the first twenty years, 1966–1985. ACM Press anthology series.

ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Aspray:1980:MCC

[Asp80] William F. Aspray. From mathematical constructivity to computer science: Alan Turing, John von Neumann, and the origins of computer science in mathematical logic. Thesis (Ph.D.), University of Wisconsin–Madison, Madison, WI, USA, 1980. v + 443 pp.

Aspray:1984:BRA

[Asp84] William Aspray. Book review: Alan Turing: The Enigma by Andrew Hodges. Isis, 75(3):625–626, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233015.

Avigad:2014:ATH

[Avi14] Jeremy Avigad. Alan Turing: his work and impact [book review, Elsevier Sci., 2013]. Notices of the American Mathematical Society, 61(8):886–890, September 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201408/rnoti-p886.pdf.

Alton:1977:RPA

[AW77] Jeannine Alton and Harriot Weiskittel. Report on the papers of Alan Mathison Turing OBE, FRS (1912–1954) mathematician, 1923–55. Report 77/39 and CSAC 53/7/77, The Royal Commission on Historical Manuscripts, Contemporary Scientific Archives Centre, London, UK, 1977. 19 pp.

Aspray:1988:RCD

[AWL^{+}88] William Aspray, Maurice V. Wilkes, Albert C. Lewis, Greg Mellen, Harold Chucker, Robert V. D. Campbell, Wendy Wilkins, G. J. Tee, Ernest Braun, and Arthur L. Norberg. Reviews: Carpenter and Doran (eds.): A. M. Turing’s ACE Report of 1946 and Other Papers; Masani (ed.): Norbert Wiener: Collected Works with Commentaries; Kozaczuk: Enigma: How the German Machine Cipher Was Broken and How It Was Read by the Allies in World War Two; Worthy: William C. Norris: Portrait of a Maverick; Harvard Computation Laboratory: A Manual of Operation for the Automatic Sequence Controlled Calculator; Proceedings of a Symposium on Large-Scale Digital Calculating Machinery; Gardner: The Mind’s New Science: A History of the Cognitive Revolution; Hartree: Calculating Machines: Recent and Prospective Developments and Their Impact on Mathematical Physics and Calculating Instruments and Machines; McLean and Rowland: The Inmos Saga; Pennings and Buifendam (eds.): New Technology as Organizational Innovation: The Development and Diffusion of Microelectronics; other literature. Annals of the History of Computing, 10(1):80–97, January/March 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1988/pdf/a1080.pdf;   http://
www.computer.org/annals/an1988/a1080abs.htm. See minor

erratum [Ano88]: Hartree as a mathematical physicist, not a physical chemist.

Axelsen:2012:TCT

[Axe12] Holger Bock Axelsen. Time complexity of tape reduction for reversible Turing machines. In Alexis De Vos and Robert Wille, editors, Reversible Computation, volume 7165 of Lecture Notes in Computer Science, pages 1–13. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNCSD9. ISBN 3-642-29516-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-29516-4/;   http:/
/www.springerlink.com/content/k42j417462628m71/.

Bartocci:2011:VMM

[B^{+}11] C. (Claudio) Bartocci et al., editors. Vite matematiche. Mathematical lives: protagonists of the twentieth century from Hilbert to Wiles. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-13605-2. xiii + 238 pp. LCCN QA28 .M38 2011.

Ben-Amram:2005:CTT

[BA05] Amir M. Ben-Amram. The Church–Turing thesis and its lookalikes. ACM SIGACT News, 36(3):113–114, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bacon:1990:DTT

[Bac90] Ben Bacon. In defense of the Turing test. Computers in Physics, 4(2):216–??, March–April 1990. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.1063/1.4822906.

Bacon:2012:CFP

[Bac12] Dave Bacon. Computation and fundamental physics. The Computer Journal, 55(7):826–829, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/826.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Barmpalias:2014:TAM

[BAC14] George Barmpalias, Manindra Agrawal, and S. Barry Cooper. Theory and applications of models of computation at the Turing Centenary in China. Theoretical Computer Science, 560 (part 2)(??):107, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304397514008858.

Bajcsy:2012:CI

[Baj12] Ruzena Bajcsy. Computation and information. The Computer Journal, 55(7):825, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/825.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Barnette:1998:AT

[Bar98] Paul J. Barnette, Jr. Alan Turing (1912–1954). World-Wide

Web site., 1998. URL http://www.geocities.com/Athens/
Acropolis/6681/turing.htm.

Batey:2017:BMP

[Bat17] Mavis Batey. Breaking machines with a pencil. In Copeland et al. [CBSW17], chapter 11, pages 97–108. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Bauer:2000:EGM

[Bau00] Friedrich L. Bauer. Entzifferte Geheimnisse. Methoden und Maximen der Kryptologie. (German) [Deciphering Secrets: Methods and Maxima of Cryptology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third revised and expanded edition, 2000. ISBN 3-540-67931-6, 3-642-63545-8 (print), 3-642-58345-8 (e-book). xiii + 502 pp. LCCN

QA76.9.A25. URL http://www.springerlink.com/content/

978-3-642-58345-2. This work refers to [?].

Bauer:2012:YTT

[Bau12] Craig Bauer. 100 years times two: Alan Turing and the Voynich Manuscript. Cryptologia, 36(2):85–87, 2012. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Berrisford:1994:ROT

[BB94] G. Berrisford and M. Burrows. Reconciling OO with Turing machines. The Computer Journal, 37(10):888–906, ???? 1994. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/37/10/
888.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_10/Vol37_10.body.html#AbstractBerrisford.

Bailey:2012:PCC

[BB12a] David Bailey and Jonathan Borwein. Person or computer: could you pass the Turing Test? The Conversation, ??(??):??,

May 3, 2012. URL https://theconversation.com/person-
or-computer-could-you-pass-the-turing-test-6769.

Borwein:2012:PCC

[BB12b] Jonathan Borwein and David Bailey. Person or computer: could you pass the Turing Test? Web article., May

3, 2012. URL http://theconversation.edu.au/person-or-
computer-could-you-pass-the-turing-test-6769.

Brooks:2016:MTP

[BB16] Heather A. Brooks and Paul C. Bressloff. A mechanism for Turing pattern formation with active and passive transport. SIAM Journal on Applied Dynamical Systems, 15(4):1823–1843, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Bringsjord:2003:CTT

[BBF03] Selmer Bringsjord, Paul Bello, and David Ferrucci. Creativity, the Turing Test, and the (better) Lovelace test. In Moor [Moo03b], pages 215–239. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_12/.

Beckmann:2006:LAC

[BBLT06] Arnold Beckmann, Ulrich Berger, Benedikt Löwe, and John V. Tucker, editors. Logical approaches to computational barriers: Second Conference on Computability in Europe, CiE 2006, Swansea, UK, June 30–July 5, 2006: proceedings, volume 3988 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.59 .C67 2006. URL

http://link.springer.com/book/10.1007/11780342.

Bates:1953:DCA

[BBST53] M. Audrey Bates, B. V. Bowden, C. Strachey, and A. M. Turing. Digital computers applied to games. In Bowden [Bow53a], pages 286–310. LCCN QA76.5 .B66. URL https://archive.org/ details/FasterThanThought. Alan Turing wrote only the part on chess (pages 288–295). The draughts part is due to Christopher Strachey, and the nim part may be due to Audrey Bates.

Bowen:2017:TL

[BC17] Jonathan Bowen and Jack Copeland. Turing’s legacy. In Copeland et al. [CBSW17], chapter 42, pages 463–474. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Beggs:2010:POT

[BCT10] Edwin J. Beggs, José Félix Costa, and John V. Tucker. Physical oracles: The Turing machine and the Wheatstone Bridge. Studia Logica, 95(1–2):279–300, July 2010. CODEN SLOGAP. ISSN 0039-3215 (print), 1572-8730 (electronic). URL http://

www.springerlink.com/content/w480483t57u74v38/. Special

Issue: The Contributions of Logic to the Foundations of Physics.

Bullynck:2015:VWD

[BDD15] Maarten Bullynck, Edgar G. Daylight, and Liesbeth De Mol. Viewpoint: Why did computer science make a hero out of Turing? Communications of the Association for Computing Machinery, 58(3):37–39, March 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.

acm.org/magazines/2015/3/183592/fulltext.

Beaver:1984:BRT

[Bea84] Donald deB. Beaver. Book review: Turing’s Man: Western Culture in the Computer Age by David J. Bolter. Isis, 75(4):770– 771, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
232475.

Beausoleil:1989:MPE

[Bea89] Jean-Roch Beausoleil. The metamathematics–Popperian epistemology connection and its relation to the logic of Turing’s programme. British Journal for the Philosophy of Science, 40(3):307–322, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/40/3/307.full.pdf+html;
http://www.jstor.org/stable/687779.

Becher:2012:TNN

[Bec12] Verónica Becher. Turing’s normal numbers: Towards randomness. In Cooper et al. [CDL12], pages 35–45. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012. URL http://www.springerlink.

com/content/5016568053026532/.

Beeson:1995:CML

[Bee95] Michael J. Beeson. Computerizing mathematics: Logic and computation. In Herken [Her95], pages 173–205. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.springer.

com/book/10.1007/978-3-7091-6597-3.

Beeson:2004:MM

[Bee04] Michael J. Beeson. The mechanization of mathematics. In Teuscher [Teu04a], pages 77–134. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Bennett:1995:LDP

[Ben95] Charles H. Bennett. Logical depth and physical complexity. In Herken [Her95], pages 207–235. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613.

LCCN QA267 .U55 1995. URL http://link.springer.com/
book/10.1007/978-3-7091-6597-3.

Benda:1997:TLI

[Ben97] M. Benda. Turing’s legacy for the Internet. IEEE Internet Computing, 1(6):75–77, November/December 1997. CODEN IICOFX. ISSN 1089-7801 (print), 1941-0131 (electronic).

Bentley:2012:MWA

[Ben12] Peter J. Bentley. Maurice Wilkes on Alan Turing. OUP Blog,

June 2012. URL http://blog.oup.com/2012/06/maurice-
wilkes-on-alan-turing/.

Bernhardt:2016:TVB

[Ber16] Chris Bernhardt. Turing’s Vision: the Birth of Computer Science. MIT Press, Cambridge, MA, USA, 2016. ISBN 0-262-03454-9, 0-262-33380-5 (e-book), 0-262-33381-3 (e-book). xvii + 189 pp. LCCN QA29.T8 A57 2015eb. URL http:/

/ieeexplore.ieee.org/servlet/opac?bknumber=7580019;
http://www.jstor.org/stable/10.2307/j.ctt1c2crt7.

Bielikova:2012:STP

[BFG^{+}12] Mária Bieliková, Gerhard Friedrich, Georg Gottlob, Stefan Katzenbeisser, and György Turán, editors. SOFSEM 2012: Theory and Practice of Computer Science: 38th Conference on Current Trends in Theory and Practice of Computer Science, Špindleruv Mlýn, Czech Republic, January 21–27, 2012. Proceedings, volume 7147 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-27659-0/.

Becher:2007:TUA

[BFP07] Verónica Becher, Santiago Figueira, and Rafael Picchi. Turing’s unpublished algorithm for normal numbers. Theoretical Computer Science, 377(1–3):126–138, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borgqvist:2024:TPF

[BGL24] Johannes G. Borgqvist, Philip Gerlee, and Carl Lundholm. Turing pattern formation on the sphere is robust to the removal of a hole. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00285-023-02034-z.

Booss:2003:MW

[BH03] Bernhelm Booss and Jens Høyrup, editors. Mathematics and war. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2003. ISBN 3-7643-1634-9 , 0-8176-1634-9. viii + 416 pp. LCCN QA10.8 .M38 2003.

Bianco:1979:IFM

[Bia79] Edmond Bianco. Informatique fondamentale: de la machine de Turing aux ordinateurs modernes. (French) [Fundamental Computer Science: from the Turing Machine to Modern Computers], volume 70 of ISR, Interdisciplinary systems research. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1979. ISBN 3-7643-1090-1. 151 + 2 pp. LCCN QA267 .B52. 28.00F.

Biever:2012:BBP

[Bie12] Celeste Biever. Bot with boyish personality wins biggest Turing test. New Scientist Web story., June 25, 2012. URL

http://www.newscientist.com/blogs/onepercent/2012/06/
bot-with-boyish-personality-wi.html.

Barroca:2011:DTI

[BLA^{+}11] Bruno Barroca, Levi Lúcio, Vasco Amaral, Roberto Félix, and Vasco Sousa. DSLTrans: a Turing incomplete transformation language. In Malloy et al. [MBS11], pages 296–305. CO-DEN LNCSD9. ISBN 3-642-19439-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/dg0v556983490629/.

Blank:2014:BRT

[Bla14] Brian E. Blank. Book review: Turing’s Cathedral: The Origins of the Digital Universe. Notices of the American Mathematical Society, 61(7):759–767, August 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.

ams.org/notices/201407/rnoti-p759.pdf.

Bloor:1998:GMA

[Blo98] Robin Bloor. The gene machine: an analysis of a Universal Turing Machine. Bloor Research, Milton Keynes, UK, 1998. ISBN 1-874160-31-7. iv + 131 pp. LCCN ????

Blomer:2012:TKG

[Blö12] Johannes Blömer. Turing und Kryptografie. (German) [Turing and cryptography]. Informatik Spektrum, 35(4):261–270, August 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X

(electronic). URL http://www.springerlink.com/content/

703t016671n87094/. Special Issue: Alan Turing.

Blum:2004:CRW

[Blu04] Lenore Blum. Computing over the reals: where Turing meets Newton. Notices of the American Mathematical Society, 51(9): 1024–1034, October 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blum:2014:ATO

[Blu14] Lenore Blum. Alan Turing and the other theory of computation (expanded). In Downey [Dow14c], chapter 3, pages 48–69. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Baeten:2011:RTM

[BLvT11] Jos C. M. Baeten, Bas Luttik, and Paul van Tilburg. Reactive Turing machines. In Olaf Owe, Martin Steffen, and Jan Arne Telle, editors, Fundamentals of Computation Theory, volume 6914 of Lecture Notes in Computer Science, pages 348– 359. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-22952-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-22952-7/;   http://www.springerlink.com/content/
b12v577614p888xt/.

Baeten:2012:TMM

[BLvT12] Jos C. M. Baeten, Bas Luttik, and Paul van Tilburg. Turing meets Milner. Lecture Notes in Computer Science, 7454: 1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32940-1_1/.

Bodewig:1949:RRE

[Bod49] E. Bodewig. Review of “Rounding-Off Errors in Matrix Processes” by A. M. Turing. Mathematical Reviews, 10:405, 1949. CODEN MAREAR. ISSN 0025-5629.

Boden:2017:PAL

[Bod17] Margaret Boden. Pioneer of artificial life. In Copeland et al. [CBSW17], chapter 33, pages 359–372. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Bolter:1984:TMW

[Bol84] J. David Bolter. Turing’s man: Western culture in the Computer Age. University of North Carolina Press, Chapel Hill, NC, USA, 1984. ISBN 0-8078-1564-0, 0-8078-4108-0 (paperback). xii + 264 pp. LCCN QA76.9.C66B64 1984.

Boone:1952:TMW

[Boo52] William W. Boone. Turing, A. M.. The word problem in semigroups with cancellation. Annals of mathematics, ser. 2, vol. 52 (1950), pp. 491–505. Journal of Symbolic Logic, 17(1):74–76, March 1952. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

Booker:2006:ACT

[Boo06a] Andrew R. Booker. Artin’s conjecture, Turing’s method, and the Riemann hypothesis. Experimental Mathematics, 15(4):385– 407, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://projecteuclid.org/euclid.
em/1175789775.

Booker:2006:TRH

[Boo06b] Andrew R. Booker. Turing and the Riemann Hypothesis. Notices of the American Mathematical Society, 53(10):1208–1211, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/notices/
200610/fea-booker.pdf.

Bowden:1953:FTT

[Bow53a] Baron Bertram Vivian Bowden, editor. Faster Than Thought: a Symposium on Digital Computing Machines. Sir Isaac Pitman and Sons, Ltd., London, UK, 1953. LCCN QA76.5 .B66. URL

https://archive.org/details/FasterThanThought. With a

foreword by the Right Honourable the Earl of Halsbury.

Bowden:1953:TM

[Bow53b] Baron Bertram Vivian Bowden. Türing machine. In Faster Than Thought: a Symposium on Digital Computing Machines [Bow53a], page 414. LCCN QA76.5 .B66. URL https://

archive.org/details/FasterThanThought. Humorous glos-

sary entry.

Brady:1995:BBG

[Bra95] Allen H. Brady. The busy beaver game and the meaning of life. In Herken [Her95], pages 237–254. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.springer.

com/book/10.1007/978-3-7091-6597-3.

Braverman:2013:CRN

[Bra13] Mark Braverman. Computing with real numbers, from Archimedes to Turing and beyond. Communications of the Association for Computing Machinery, 56(9):74–83, September 2013. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brenner:2012:ATI

[Bre12a] Sydney Brenner. Alan Turing II: Kode des Lebens. (German) [Alan Turing II: Life’s code script]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??): ??, ???? 2012. CODEN SPEKDI. ISSN 0170-2971. URL http:

//www.spektrum.de/alias/spezial/alan-turing-ii-kode-
des-lebens/1149656.

Brenner:2012:TCL

[Bre12b] Sydney Brenner. Turing centenary: Life’s code script. Nature, 482(7386):461, February 22, 2012. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v482/n7386/full/
482461a.html.

Bretos:2012:ATP

[Bre12c] Lydia Bretos. Alan Turing: la pensée informatique. (French) [Alan Turing: his computer-science thought]. CRDP de l’académie de Versailles, Paris, France, 2012. ISBN 2-86637-562-9. ISSN 1957-3367. 65 pp. LCCN ????

Breuer:2013:GTP

[Bre13] Thomas Breuer. A Gödel–Turing perspective on quantum states indistinguishable from inside. In Zenil [Zen13], pages 605–616. ISBN 981-4374-29-6. LCCN QA267.7 .C676 2013.

URL http://www.worldscientific.com/worldscibooks/10.

1142/8306. Foreword by Roger Penrose.

Bridger:1990:RTO

[Bri90] Mark Bridger. Reviews: The Turing Omnibus: 61 Excursions in Computer Science, by A. K. Dewdney. American Mathematical Monthly, 97(4):355–357, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bringsjord:1995:IWJ

[Bri95] Selmer Bringsjord. If I were judge. In Epstein et al. [ERB08], pages 89–102. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Bringsjord:2009:IWJ

[Bri09] Selmer Bringsjord. If I were judge. In Epstein et al. [EBR09], pages 89–102. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335

.P35445 2008. URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6710-5_6.pdf.

Brogi:1997:TMC

[Bro97] Antonio Brogi. A Turing machine contest for introducing high school students to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2): 23–27, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brooks:2005:TLC

[Bro05] Frederick P. Brooks, Jr. Turing Lecture 2005: Collaboration and telecollaboration in design. World-Wide Web

site., 2005. URL http://www.bcs.org/BCS/Awards/Events/
TuringLecture/Turing2005/.

Brown:2009:TAT

[Bro09] Prime Minister Gordon Brown. Treatment of Alan Turing was “appalling”. UK Government Web site, September 10, 2009.

URL http://www.number10.gov.uk/Page20571. Issued in re-

sponse to a public appeal.

Brooks:2013:EKK

[Bro13] Richard Brooks. Enigma of Keira Knightley as codebreaker Turing’s lover. The Times [London], June 23,

2013. URL http://www.thesundaytimes.co.uk/sto/news/
uk_news/Arts/article1277723.ece.

Brynjolfsson:2022:TTP

[Bry22] Erik Brynjolfsson. The Turing trap: The promise & peril of human-like artificial intelligence. Dædalus, 151(2):272–287, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print),

1548-6192 (electronic). URL /daed/article-pdf/151/2/272/
2009138/daed_a_01915.pdf.

Blagodatski:2015:DST

[BSK^{+}15] Artem Blagodatski, Anton Sergeev, Mikhail Kryuchkov, Yuliya Lopatina, and Vladimir L. Katanaev. Diverse set of Turing nanopatterns coat corneae across insect lineages. Proceedings of the National Academy of Sciences of the United States of America, 112(34):10750–10755, 2015. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.pnas.

org/content/112/34/10750.abstract.

Boutel:1965:CIP

[BSPI65] B. E. Boutel, C. Strachey, J. H. G. Phillips, and Michael Irish. Correspondence: An impossible program. The Computer Journal, 8(3):215, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_08/Issue_03/tiff/215.tif.

See [Str65, ApS65, HLOS65, AMKM66].

Billock:2012:WUF

[BT12] Vincent A. Billock and Brian H. Tsou. Wahrnehmung Unmögliche Farben. (German) [exercise: Impossible colors]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, ???? 2012. CODEN SPEKDI. ISSN 0170-

2971. URL http://www.spektrum.de/alias/wahrnehmung/
unmoegliche-farben/1056534.

Bradley:2012:IRC

[BTHS12] Jeremy T. Bradley, Nigel Thomas, Richard A. Hayden, and Anton Stefanek. Invited response to Computer Journal Lecture by Prof. Jane Hillston. The Computer Journal, 55(7):882–886, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/55/7/882.full.pdf+html. Special Focus on the Cen-

tenary of Alan Turing.

Buhrmann:2014:TQ

[Buh14] Hurry Buhrmann. Turing in Quantumland. In Downey [Dow14c], chapter 4, pages 70–89. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Bullynck:2015:CPT

[Bul15] Maarten Bullynck. Computing primes (1929–1949): Transformations in the early days of digital computing. IEEE Annals of the History of Computing, 37(3):44–54, July/September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://www.computer.org/csdl/mags/an/2015/
03/man2015030044-abs.html.

Bullynck:2021:GEL

[Bul21] Maarten Bullynck. German encounters of logic and programming (1948–1958), featuring three readings of Turing machines. IEEE Annals of the History of Computing, 43(4):10–26, 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Burke:2011:AMD

[Bur11] Colin Burke. Agnes Meyer Driscoll vs. the Enigma and the Bombe. Report, University of Maryland, Baltimore County, 1000 Hilltop Circle Baltimore, MD 21250, USA, January 7, 2011. 132

pp. URL https://userpages.umbc.edu/~burke/driscoll1-
2011.pdf.

Buzen:2012:CUR

[Buz12] Jeffrey P. Buzen. Computation, uncertainty and risk. The Computer Journal, 55(7):838–847, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/838.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Bansagi:2011:TRD

[BVE11] Tamás Bánsági, Jr., Vladimir K. Vanag, and Irving R. Epstein. Tomography of reaction-diffusion microemulsions reveals three-dimensional Turing patterns. Science, 331(6022):1309– 1312, March 11, 2011. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/331/6022/1309.full.pdf.

Staff:2014:LES

[CAC14a] CACM Staff. Letters to the Editor: On the significance of Turing’s test. Communications of the Association for Computing Machinery, 57(12):8–9, December 2014. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

cacm.acm.org/magazines/2014/12/180771/fulltext.

Staff:2014:NAT

[CAC14b] CACM Staff. News: ACM’s Turing Award prize raised to $1 million. Communications of the Association for Computing Machinery, 57(12):20, December 2014. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.

acm.org/magazines/2014/12/180785/fulltext.

Cai:2012:RFB

[Cai12] Jinhai Cai. Robust filtering-based thinning algorithm for pattern recognition. The Computer Journal, 55(7):887–896, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/887.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Cappuccio:2005:ATU

[Cap05] Massimiliano Cappuccio. Alan Turing: l’uomo, la macchina, l’enigma: per una genealogia dell’incomputabile. (Italian) [Alan Turing: the man, the machine, the Enigma; towards a genealogy of the incomputable], volume 1 of Pragmata. AlboVersorio, Milano, Italy, 2005. ISBN 88-89130-29-6. 342 pp. LCCN QA29.T8 C37 2005.

Carter:2010:TB

[Car10] Frank Carter. The Turing Bombe. Rutherford Journal, 3(??): ??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article030108.html.

Castelfranchi:2001:AAA

[Cas01] Cristiano Castelfranchi. Again on agents’ autonomy: a homage to Alan Turing. Lecture Notes in Computer Science, 1986:339– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1986/19860339.htm;   http://
link.springer-ny.com/link/service/series/0558/papers/
1986/19860339.pdf.

Casselman:2006:BTM

[Cas06a] Bill Casselman. Blueprint for a Turing Machine: About the cover… and a bit more. Notices of the American Mathematical Society, 53(10):1186–1189, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/200610/comm-aboutcov.pdf.

Casselman:2006:MTE

[Cas06b] Bill Casselman. Mathematical theory of the Enigma machine. Notices of the American Mathematical Society, 53 (4):433, April 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.mathaware.

org/;   http://www.nationalarchives.gov.uk/;   http:/
/www.turingarchive.org/browse.php/C/30. The front cover

of this issue displays eight pages of Alan Turing’s description of the Enigma machine. The issue is a special tribute to Kurt Gödel for the centenary of his birth.

Castelfranchi:2013:ATC

[Cas13] Cristiano Castelfranchi. Alan Turing’s “Computing Machinery and Intelligence”. Topoi, 32(2):293–299, October 2013. ISSN 0167-7411 (print), 1572-8749 (electronic). URL http://link.

springer.com/article/10.1007/s11245-013-9182-y.

Cass:2016:DA

[Cas16] Stephen Cass. The Digital Apple. IEEE Spectrum, 53(1):19– 20, January 2016. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Copeland:2017:LW

[CB17] Jack Copeland and Jonathan Bowen. Life and work. In Copeland et al. [CBSW17], chapter 1, pages 3–18. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67

Clausing:2012:ATI

[CBB12] Achim Clausing and Bernhelm Booß-Bavnbek. Alan Turing’s impact in Münster [Discussion of MR 2952222]. Mitteilungen der Deutschen Mathematiker-Vereinigung, 20(2):68–70, 2012. ISSN 0947-4471.

Copeland:2017:TG

[CBSW17] B. Jack Copeland, Jonathan Bowen, Mark Sprevak, and Robin Wilson, editors. The Turing guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). xv + 546 pp. LCCN QA29.T8 C67 2017.

Carpenter:1977:OTM

[CD77] B. E. Carpenter and R. W. Doran. The other Turing machine. The Computer Journal, 20(3):269–279, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/3/269.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/273.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_20/Issue_03/tiff/274.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_03/tiff/275.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/277.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/278.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_20/Issue_03/tiff/279.

tif. Hodges page 318 note 6.1.

Carpenter:1986:MTA

[CD86] B. E. Carpenter and R. W. Doran, editors. A. M. Turing’s ACE Report of 1946 and Other Papers, volume 10 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-03114-0. vii + 141 pp. LCCN QA75 .A185 1986.

Carpenter:2017:TZ

[CD17] Brian Carpenter and Robert Doran. Turing’s Zeitgeist. In Copeland et al. [CBSW17], chapter 22, pages 223–232. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN

QA29.T8 C67 2017. URL https://www.cs.auckland.ac.nz/~
brian/TuringZeitgeistPreprint.pdf.

Cooper:2012:HWC

[CDL12] S. Barry Cooper, Anuj Dawar, and Benedikt Löwe, editors. How the World Computes: Turing Centenary Conference and 8th Conference on Computability in Europe, CiE 2012, Cambridge, UK, June 18–23, 2012. Proceedings, volume 7318 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012.

Chen:2010:TPW

[CEL10] Chao-Nien Chen, Shin-Ichiro Ei, and Ya-Ping Lin. Turing patterns and wavefronts for reaction-diffusion systems in an infinite channel. SIAM Journal on Applied Mathematics, 70(8):2822– 2843, ???? 2010. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ceruzzi:1985:TMW

[Cer85] Paul E. Ceruzzi. Turing’s Man: Western Culture in the Computer Age by J. David Bolter (review). Technology and Culture, 26(2):341–343, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889800/pdf.

Cerqui:2004:TIS

[Cer04] Daniela Cerqui. From Turing to the Information Society. In Teuscher [Teu04a], pages 59–74. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book).

LCCN QA29.T8 A57 2004. URL http://link.springer.com/
chapter/10.1007/978-3-662-05642-4_4. Foreword by Dou-

glas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Craig:1998:REP

[CF98] Edward Craig and Luciano Floridi, editors. Routledge Encyclopedia of Philosophy Online. Routledge, London, UK, version 2.0 edition, 1998. ISBN 0-415-16916-X (CD-ROM), 0-415-19608-6 (user guide), 0-415-16917-8 (10 vol set + CD-ROM). LCCN

B51. URL http://www.rep.routledge.com/.

Ceruzzi:1991:RCK

[CFK^{+}91] Paul Ceruzzi, Kenneth Flamm, Peggy Aldrich Kidwell, Herbert R. J. Grosch, and John A. N. Lee. Reviews: Campbell-Kelly: ICL: A Business and Technical History; Aspray: Computing Before Computers; Watson and Petre: Father, Son & Co.; Asimov and Frenkel: Robots: Machines in Man’s Image; McNeil: An Encyclopedia of the History of Technology; Byte: Fifteenth Anniversary Summit; Deavours and Kruh: The Turing Bombe: Was it Enough?; Pearcey: A History of Australian Computing; Aspray: The Origins of John von Neumann’s Theory of Automata; Crossley and Henry: Thus Spake al-Khwarizmi: a Translation of the text of Cambridge University Library Ms. li.vi.5; Fauvel and Gerdes: African Slave and Calculating Prodigy: Bicentenary of the Death of Thomas Fuller; Marling: Maestro of Many Keyboards [brief biography of Donald Knuth]. Annals of the History of Computing, 13(1):111–117, January/March 1991. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1991/pdf/a1111.pdf;   http://www.computer.
org/annals/an1991/a1111abs.htm.

Cordy:1987:DIE

[CG87] J. R. Cordy and T. C. N. Graham. Design of an interpretive environment for Turing. ACM SIGPLAN Notices, 22(7):199–204, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.org:80/

pubs/citations/proceedings/plan/29650/p199-cordy/.

Capuni:2012:TMR

[ÇG12] Ilir Çapuni and Peter Gács. A Turing machine resisting isolated bursts of faults. In Bieliková et al. [BFG^{+}12], pages 165– 176. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/77156645761111up/.

Copeland:2012:AT

[CGLWVR12] Jack Copeland, Didier Galmiche, Dominique Larchey-Wendling, and Joseph Vidal-Rosset, editors. Alan Turing. Éditions Kimé, Paris, France, 2012. ISBN 2-84174-603-8. 198 pp. LCCN ????

Cordy:1983:TAN

[CH83a] J. R. Cordy and R. C. Holt. Turing: a new general purpose computer language under UNIX. In Software Tools Users Group [Sof83], pages 249–254. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Cordy:1983:TNG

[CH83b] J. R. Cordy and R. C. Holt. Turing: a new general purpose computer language under UNIX. In USENIX Association [USE83], pages 249–254. ISBN ???? LCCN ????

Cooper:2016:OFT

[CH16] S. Barry Cooper and Andrew Hodges, editors. The Once and Future Turing: Computing the World. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-01083-7 (hardcover), 0-521-28250-0 (paperback), 0-511-86319-5 (e-book). xviii + 379 pp. LCCN ????

Chapnick:1994:BRA

[Cha94] Philip Chapnick. Book review: Ad Infinitum: The Ghost in Turing’s Machine and Zero to Lazy Eight: Playing with infinity. Mathematica Journal, 4(1):20–21, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

tronic). URL http://www.mathematica-journal.com/issue/
v4i1/reviews/chapnick/20-21_Chapnick.mj.pdf;   http:/
/www.mathematica-journal.com/issue/v4i1/reviews/chapnick/
index.html.

Chaitin:1995:AEH

[Cha95] Gregory J. Chaitin. An algebraic equation for the halting probability. In Herken [Her95], pages 255–259. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.springer.

com/book/10.1007/978-3-7091-6597-3.

Chan:2016:TMP

[Cha16] Sewell Chan. Thousands of men to be pardoned for gay sex, once a crime in Britain. New York Times, ??(??):A1, A8, October 21, 2016. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://www.nytimes.com/2016/
10/21/world/europe/britain-will-posthumously-pardon-
thousands-of-gay-and-bisexual-men.html. From the story:

“The law providing for the pardons, which could take effect in a matter of months now that it has the support of the Conservative government, is named for Alan Turing, the mathematician who made a major contribution to Britain in World War II by cracking Germany’s Enigma coding machine and was a central figure in the development of the computer.

Turing was convicted on charges of homosexuality in 1952 and committed suicide in 1954. The government apologized in 2009 for its treatment of him, and in 2013, Queen Elizabeth II formally pardoned him. In April, the head of Britain’s signals intelligence agency, GCHQ, also apologized, for its past discrimination against gays.”.

Chesebro:1993:CCC

[Che93] James W. Chesebro. Communication and computability: The case of Alan Mathison Turing. Communication Quarterly, 41 (1):90–121, January 1993. ISSN 0146-3373 (print), 1746-4102 (electronic).

Chomsky:1995:TBG

[Cho95] Noam Chomsky. Turing on the “Imitation Game”. In Epstein et al. [ERB08], pages 103–106. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Chomsky:2009:TIG

[Cho09] Noam Chomsky. Turing on the “Imitation Game”. In Epstein et al. [EBR09], pages 103–106. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_7.pdf.

Chouard:2012:TLU

[Cho12] Tanguy Chouard. Turing at 100: Legacy of a universal mind. Nature, 482(7386):455, February 22, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Christensen:2010:ATF

[Chr10] Chris Christensen. Alan Turing’s first cryptology textbook and Sinkov’s revision of it. Cryptologia, 34(1):27–43, 2010. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Christensen:2013:RBA

[Chr13] Chris Christensen. Review of biographies of Alan Turing. Cryptologia, 37(4):356–367, 2013. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Christensen:2015:BRA

[Chr15] Chris Christensen. Book review: Alan Turing: His Work and Impact, edited by S. Barry Cooper and Jan van Leeuwen. Cryptologia, 39(2):198–202, 2015. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Christensen:2016:CRP

[Chr16] Chris Christensen. Companion review of Prof: Alan Turing Decoded by Dermot Turing. Cryptologia, 40(6):556–562, 2016. CO-DEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Christensen:2021:RCB

[Chr21] Chris Christensen. Review of The Codebreakers of Bletchley Park by Dermot Turing. Cryptologia, 45(6):563–564, 2021. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/01611194.2020.1839814.

Christensen:2022:RRA

[Chr22a] Chris Christensen. Review of Reflections of Alan Turing: a Relative Story by Dermot Turing. Cryptologia, 46(3):287–289, 2022. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/01611194.2021.1969703.

Christensen:2022:RBM

[Chr22b] Chris Christensen. Review of The Bombe: The Machine that Defeated Enigma by Dermot Turing. Arcturus Publishing Limited, London, 2021. 64 pages, Trade paperback A-5, £11.20. ISBN 978-1-3988-1244-4. Cryptologia, 46(4):385–386, 2022. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/01611194.2022.2026840.

Christensen:2024:RFE

[Chr24] Chris Christensen. Review of The First Enigma Codebreaker: The Untold Story of Marian Rejewski Who Passed the Baton to Alan Turing. Cryptologia, 48(5):474–478, 2024. CO-DEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/01611194.2023.2246470.

Churchland:1995:NI

[Chu95] Paul M. Churchland. On the nature of intelligence. In Epstein et al. [ERB08], pages 107–117. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Churchland:2009:NI

[Chu09] Paul M. Churchland. On the nature of intelligence. In Epstein et al. [EBR09], pages 107–117. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_8.pdf.

Church:2013:BCN

[Chu13] Alonzo Church. On Computable Numbers, with an Application to the Entscheidungsproblem by A. M. Turing — review. In Cooper and van Leeuwen [CvL13], pages 117– 119. ISBN 0-12-386980-3 (hardcover). LCCN QA29.T8

C65 2013. URL http://www.sciencedirect.com/science/
article/pii/B9780123869807500034.

Campbell-Kelly:1984:RAT

[CK84] Martin Campbell-Kelly. Review of “Alan Turing: The Enigma”, by Andrew Hodges. Annals of the History of Computing, 6(2):176–178, April/June 1984. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1984/pdf/a2167.pdf;   http://www.computer.
org/annals/an1984/a2167abs.htm.

Campbell-Kelly:2002:BRJ

[CK02] Martin Campbell-Kelly. Book review: Jon Agar, Turing and the Universal Machine: The Making of the Modern Computer. Revolutions in Science. Duxford: Icon Books, 2001. Pp. iv + 153. ISBN 1-84046-250-7. £5.99, $9.95 (paperback). British Journal for the History of Science, 35(4):475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Campbell-Kelly:2012:ATO

[CK12a] Martin Campbell-Kelly. Alan Turing’s other universal machine. Communications of the Association for Computing Machinery, 55(7):31–33, July 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Campbell-Kelly:2012:NCR

[CK12b] Martin Campbell-Kelly. Von Neumann’s computer: [review of Turing’s Cathedral: the Origins of the Digital Universe, George Dyson, 2012 Pantheon Books £25.00 / $29.95 hardcover 423pp]. Physics World, 25(12):44–45, December 2012. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://iopscience.iop.org/pwa/full/pwa-
pdf/25/12/phwv25i12a42.pdf.

Campbell-Kelly:2017:A

[CK17] Martin Campbell-Kelly. ACE. In Copeland et al. [CBSW17], chapter 21, pages 213–222. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Cooper:2002:TDE

[CL02] S. Barry Cooper and Angsheng Li. Turing definability in the Ershov hierarchy. Journal of the London Mathematical Society, 66(3):513–528, 12 2002. CODEN JLMSAK. ISSN

1469-7750. URL http://journals.cambridge.org/article_
S0024610702003691.

Copeland:2017:THC

[CL17a] B. Jack Copeland and Jason Long. Turing and the history of computer music. In Floyd and Bokulich [FB17], chapter 8, pages 189–218. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346. LCCN QA29.T8 P45 2017. URL https://link.

springer.com/chapter/10.1007/978-3-319-53280-6_8.

Copeland:2017:ATH

[CL17b] Jack Copeland and Jason Long. Alan Turing: How his universal machine became a musical instrument. IEEE Spectrum, 54 (10):??, October 26, 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). URL https://spectrum.ieee.

org/tech-history/silicon-revolution/alan-turing-how-
his-universal-machine-became-a-musical-instrument.

Copeland:2017:CM

[CL17c] Jack Copeland and Jason Long. Computer music. In Copeland et al. [CBSW17], chapter 23, pages 233–248. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Clarke:1972:TMM

[Cla72] J. J. Clarke. Turing machines and the mind-body problem. British Journal for the Philosophy of Science, 23(1):1–12, February 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/23/1/1.full.pdf+html;   http://www.jstor.org/
stable/686217.

Clegg:2017:LBT

[Cle17] Brian Clegg. Last byte: Turing’s taxi. Communications of the Association for Computing Machinery, 60(8):104–ff, August 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://dl.acm.org/ft_gateway.cfm?id=
3107917.

Cooper:2007:CLR

[CLS07] S. Barry Cooper, Benedikt Lowe, and Andrea Sorbi, editors. Computation and Logic in the Real World: Third Conference on Computability in Europe, CiE 2007 Siena, Italy, June 18–23, 2007 Proceedings, volume 4497 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-73000-1, 3-540-73001-X (ebook). LCCN QA9.59 .C67 2007. URL http://public.eblib.

com/EBLPublic/PublicView.do?ptiID=337659.

Clark:1996:LAT

[CM96] Andy Clark and P. J. R. (Peter J. R.) Millican, editors. The legacy of Alan Turing: Connectionism, Concepts, and Folk Psychology, volume 2 of Mind Association occasional series. Clarendon Press, Oxford, UK, 1996. ISBN 0-19-823594-1. ix + 281 pp.

LCCN Q335.5 .L44 1996. URL http://www.oup.co.uk/isbn/

0-19-823594-1. See also volume 1 [MC96].

Cogburn:2010:TMP

[CM10] Jon Cogburn and Jason Megil. Are Turing machines platonists? Inferentialism and the computational theory of mind. Minds and Machines, 20(3):423–439, August 2010. CODEN MMACEO.

ISSN 0924-6495 (print), 1572-8641 (electronic). URL http://

www.springerlink.com/content/t627k1mg6576t636/.

Coates:2013:CMW

[Coa13] Sam Coates. Cabinet ministers at war over failure to back Turing pardon. The Times [London], July 20, 2013. URL http://www.

thetimes.co.uk/tto/news/politics/article3821179.ece.

Cockshott:2012:TIM

[Coc12] Paul Cockshott. Turing: The irruption of materialism into thought. Soapbox Science, 2012. URL http://blogs.nature.

com/soapboxscience/2012/06/20/turing-the-irruption-of-
materialism-into-thought.

Conrad:1995:PP

[Con95] Michael Conrad. The price of programmability. In Herken [Her95], pages 261–281. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267

.U55 1995. URL http://link.springer.com/book/10.1007/
978-3-7091-6597-3.

Conery:2012:CSM

[Con12] John S. Conery. Computation is symbol manipulation. The Computer Journal, 55(7):814–816, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/814.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Conner:2022:PTS

[Con22] Thomas H. Conner. Photography from the Turin Shroud to the Turing Machine by Yanai Toister (review). Technology and Culture, 63(3):912–913, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859758.

Cooper:2006:MWK

[Coo06a] S. Barry Cooper. The Man Who Knew Too Much: Alan Turing and the Invention of the Computer — a book review. Notices of the American Mathematical Society, 53(10):1213–1217, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-

9477 (electronic). URL http://www.ams.org/notices/200610/
rev-cooper.pdf.

Cooper:2006:CE

[Coo06b] S. Barry Cooper. Computability and emergence. In Gabbay et al. [GGZ06], pages 193–231. ISBN 0-387-28688-8, 0-387-31072-X (e-book). ISSN 1571-5485 (print), 1574-8944 (electronic).

LCCN QA9.A1 M38 2006. URL http://link.springer.com/
chapter/10.1007/0-387-31072-X_4.

Cooper:2006:HCN

[Coo06c] S. Barry Cooper. How can nature help us compute? In Wiedermann et al. [WTP^{+}06], pages 1–13. CODEN LNCSD9. ISBN 3-540-31198-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://link.springer.com/chapter/10.
1007/11611257_1.

Cooper:2008:NCP

[Coo08] Stuart Barry Cooper, editor. New computational paradigms: changing conceptions of what is computable. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-36033-6 (hardcover), 0-387-68546-4. LCCN

QA9.59. URL http://deposit.ddb.de/cgi-bin/dokserv?
id=2822306&prov=M&dok\_var=1&dok\_ext=htm;   http://
deposit.ddb.de/cgi-bin/dokserv?id=2822306&prov=M&dok_
var=1&dok_ext=htm;   http://digitool.hbz-nrw.de:1801/
webclient/DeliveryManager?pid=2344379&custom_att_2=simple_
viewer;   http://swbplus.bsz-bw.de/bsz255458851cov.htm;
 http://swbplus.bsz-bw.de/bsz255458851inh.htm;   http:
//swbplus.bsz-bw.de/bsz255458851kap.htm;   http://swbplus.
bsz-bw.de/bsz255458851vor.htm;   http://www.gbv.de/dms/
goettingen/513577068.pdf.

Cooper:2012:TMM

[Coo12a] S. Barry Cooper. From Turing Machine to morphogenesis: Forming and informing computation. In Agrawal et al. [ACL12], pages 3–10. CODEN LNCSD9. ISBN 3-642-29951-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-29952-0_2.

Cooper:2012:IAA

[Coo12b] S. Barry Cooper. Incomputability after Alan Turing. Notices of the American Mathematical Society, 59(6):776–784, June/ July 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-

9477 (electronic). URL http://www.ams.org/notices/201206/
rtx120600776p.pdf.

Cooper:2012:IAT

[Coo12c] S. Barry Cooper. The incomputable Alan Turing. arxiv.org, June

2012. URL http://adsabs.harvard.edu/abs/2012arXiv1206.
1706C.

Cooper:2012:PBI

[Coo12d] S. Barry Cooper. Pushing back the incomputable — Alan Turing’s ten big ideas. Asia Pacific Mathematics Newsletter, 2(1): 2–6, 2012. ISSN 2010-3484.

Cooper:2012:TCI

[Coo12e] S. Barry Cooper. Turing centenary: The incomputable reality. Nature, 482(7386):465, February 22, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v482/n7386/
full/482465a.html.

Cooper:2012:TTM

[Coo12f] S. Barry Cooper. Turing’s Titanic machine? Communications of the Association for Computing Machinery, 55(3):74–83, March 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Copeland:2000:NVW

[Cop00a] B. Jack Copeland. Narrow versus wide mechanism: Including a re-examination of Turing’s views on the mind–machine issue. The Journal of Philosophy, 97(1):5, January 2000. ISSN 0022-362X (print), 1939-8549 (electronic).

Copeland:2000:WTM

[Cop00b] B. Jack Copeland. What is a Turing Machine? Web site, July

2000. URL http://www.alanturing.net/turing_archive/
pages/Reference%20Articles/What%20is%20a%20Turing%20Machine.
html.

Copeland:2003:TT

[Cop03] B. Jack Copeland. The Turing Test. In Moor [Moo03b], pages 1– 21. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003.

URL http://link.springer.com/chapter/10.1007/978-94-
010-0105-2_1/.

Copeland:2004:ETS

[Cop04] B. Jack Copeland, editor. The Essential Turing: Seminal Writings in Computing, Logic, Philosophy, Artificial Intelligence, and Artificial Life, plus The Secrets of Enigma. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-825079-7 (hardcover), 0-19-825080-0 (paperback). viii + 613 pp. LCCN QA29.T8 E77 2004. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/toc/fy053/2004275594.html.

Copeland:2005:ATA

[Cop05a] B. Jack Copeland, editor. Alan Turing’s Automatic Computing Engine: the master codebreaker’s struggle to build the modern computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856593-3 (hardcover). xx + 553 pp. LCCN QA75 .A43 2005. URL http://

ukcatalogue.oup.com/product/9780198565932.do;   http:
//www.oxfordscholarship.com/oso/public/content/maths/
9780198565932/toc.html.

Copeland:2005:IGA

[Cop05b] B. Jack Copeland. The imitation game: Artificial intelligence and the human mind. Inaugural Turing Memorial Lecture to be held on the evening of 25 August 2005 at Bletchley Park, Milton

Keynes, August 25, 2005. URL http://www.beds.bcs.org.uk/
events/2005-08-25-Turing.shtml.

Copeland:2006:CSB

[Cop06a] B. Jack Copeland, editor. Colossus: the secrets of Bletchley Park’s codebreaking computers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-284055-X (hardcover), 0-19-957814-1 (paperback). xvi + 462 + 16 pp. LCCN

D810.C88 C66 2006. URL http://www.colossus-computer.
com/.

Copeland:2006:MHC

[Cop06b] B. Jack Copeland. The modern history of computing. In Edward N. Zalta, editor, The Stanford Encyclopedia of Philosophy. Metaphysics Research Lab, Stanford University, Stanford,

CA, USA, June 9, 2006. URL https://plato.stanford.edu/
entries/computing-history/.

Copple:2009:BAL

[Cop09] Kevin L. Copple. Bringing AI to life. In Epstein et al. [EBR09], pages 359–376. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335

.P35445 2008. URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6710-5_22.pdf.

Copeland:2011:MCRa

[Cop11a] B. Jack Copeland. The Manchester computer: a revised history part 1: The memory. IEEE Annals of the History of Computing, 33(1):4–21, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2011:MCRb

[Cop11b] B. Jack Copeland. The Manchester computer: a revised history part 2: The baby computer. IEEE Annals of the History of Computing, 33(1):22–37, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2012:ATE

[Cop12a] B. Jack Copeland, editor. Alan Turing’s Electronic Brain: the Struggle to Build the ACE, the World’s Fastest Computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-960915-2 (paperback). xxi + 553 pp. LCCN ???? UK £14.99.

Copeland:2012:TPI

[Cop12b] B. Jack Copeland. Turing: pioneer of the information age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-963979-5. 192 pp. LCCN QA29.T8 C66 2012.

Copeland:2017:BP

[Cop17a] Jack Copeland. At Bletchley Park. In Copeland et al. [CBSW17], chapter 9, pages 79–84. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:Bb

[Cop17b] Jack Copeland. Baby. In Copeland et al. [CBSW17], chapter 20, pages 199–212. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:CP

[Cop17c] Jack Copeland. Crime and punishment. In Copeland et al. [CBSW17], chapter 4, pages 35–40. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:DES

[Cop17d] Jack Copeland. Delilah — encrypting speech. In Copeland et al. [CBSW17], chapter 18, pages 183–188. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:HHF

[Cop17e] Jack Copeland. Hilbert and his famous problem. In Copeland et al. [CBSW17], chapter 7, pages 57–66. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:IM

[Cop17f] Jack Copeland. Intelligent machinery. In Copeland et al. [CBSW17], chapter 25, pages 265–276. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:THB

[Cop17g] Jack Copeland. Tunny: Hitler’s biggest fish. In Copeland et al. [CBSW17], chapter 14, pages 143–160. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:TGI

[Cop17h] Jack Copeland. Turing’s great invention: the universal computing machine. In Copeland et al. [CBSW17], chapter 6, pages 49–56. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2018:ATL

[Cop18] Jack Copeland. Alan Turing’s lost notebook. Rutherford Journal, 5(??):??, ???? 2018. CODEN ???? ISSN 1177-1380.

URL https://www.math.utah.edu/pub/bibnet/authors/t/
turing-alan-mathison.bib;   http://rutherfordjournal.
org/article050104.html.

Copeland:2023:EAB

[Cop23] B. Jack Copeland. Early AI in Britain: Turing et al. IEEE Annals of the History of Computing, 45(3):19–31, July/September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2025:LSA

[Cop25] Jack Copeland. The lost story of alan turing’s secret “Delilah” project. Web site, February 4, 2025. URL https://spectrum.

ieee.org/alan-turings-delilah.

Corrigan:2007:AT

[Cor07] Jim Corrigan. Alan Turing. Profiles in mathematics. Morgan Reynolds Pub., Greensboro, NC, USA, 2007. ISBN 1-59935-064-

5. ???? pp. LCCN QA29.T8 C67 2007. URL http://www.loc.

gov/catdir/toc/ecip0713/2007011704.html.

Corry:2017:TPW

[Cor17] Leo Corry. Turing’s pre-war analog computers: the fatherhood of the modern computer revisited. Communications of the Association for Computing Machinery, 60(8):50–58, August 2017. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dl.acm.org/ft_gateway.cfm?id=3104032.

Cotogno:2003:HPC

[Cot03] Paolo Cotogno. Hypercomputation and the physical Church– Turing thesis. British Journal for the Philosophy of Science, 54(2):181–223, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/54/2/181.full.pdf+html;
http://www.jstor.org/stable/3541964.

Cowell:2019:ATC

[Cow19] Alan Cowell. Alan Turing, condemned code breaker and computer visionary. New York Times, ??(??):A22, June 5, 2019. URL

https://www.nytimes.com/2019/06/05/obituaries/alan-turing-
overlooked.html.

Copeland:1995:TT

[CP95] Jack Copeland and Diane Proudfoot. Turing’s Test. In Epstein et al. [ERB08], pages 119–138. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Copeland:1996:ATA

[CP96] B. Jack Copeland and Diane Proudfoot. On Alan Turing’s anticipation of connectionism. Synthese, 108(3):361–377, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/

10.1007/BF00413694. See correction [CP23].

Copeland:1999:ATF

[CP99] B. Jack Copeland and Diane Proudfoot. Alan Turing’s forgotten ideas in computer science. Scientific American, 280 (4):98–103, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
98.pdf.

Copeland:2000:WTD

[CP00] B. Jack Copeland and Diane Proudfoot. What Turing did after he invented the Universal Turing Machine. Journal of Logic, Language, and Information, 9(4):491–509, October 2000. CODEN JLLIEN. ISSN 0925-8531 (print), 1572-9583 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A:1008371426608;   http://www.jstor.org/pss/40180239.

Copeland:2001:ANT

[CP01] Jack Copeland and Diane Proudfoot. AlanTuring.net: the Turing Archive for the History of Computing. World-Wide Web

site., 2001. URL http://www.alanturing.net/.

Copeland:2004:CAI

[CP04] B. Jack Copeland and Diane Proudfoot. The computer, artificial intelligence, and the Turing Test. In Teuscher [Teu04a], pages 317–351. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Copeland:2009:TT

[CP09] Jack Copeland and Diane Proudfoot. Turing’s Test. In Epstein et al. [EBR09], pages 119–138. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_9.pdf.

Copeland:2010:DET

[CP10] B. Jack Copeland and Diane Proudfoot. Deviant encodings and Turing’s analysis of computability. Studies in History and Philosophy of Science Part A, 41(3):247–252, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000427.

Ciobanu:2012:CPN

[CP12a] Gabriel Ciobanu and G. Michele Pinna. Catalytic Petri nets are Turing complete. In Dediu and Martin-Vide [DMV12], pages 192–203. ISBN 3-642-28331-4 (soft cover). LCCN ???? URL

http://www.springerlink.com/content/r154v86002n26g75/

.

Copeland:2011:ATF

[CP12b] B. Jack Copeland and Diane Proudfoot. Alan Turing and father of the modern computer. Rutherford Journal, 4(??):??, ???? 2011–2012. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article040101.html.

Copeland:2017:CCF

[CP17a] Jack Copeland and Dani Prinz. Computer chess — the first moments. In Copeland et al. [CBSW17], chapter 31, pages 327–346. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2017:CCN

[CP17b] Jack Copeland and Diane Proudfoot. Connectionism: computing with neurons. In Copeland et al. [CBSW17], chapter 29, pages 309–314. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Copeland:2023:CAT

[CP23] B. Jack Copeland and Diane Proudfoot. Correction to: On Alan Turing’s anticipation of connectionism. Synthese, 201 (2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-04023-z. See [CP96].

Cook:2011:PPT

[CPR11] Byron Cook, Andreas Podelski, and Andrey Rybalchenko. Proving program termination. Communications of the Association for Computing Machinery, 54(5):88–98, May 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crato:2010:FIE

[Cra10a] Nuno Crato. Figuring It Out: Entertaining Encounters with Everyday Math. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-04832-3, 3-642-04833-1. ???? pp. LCCN QA99. URL http://www.

springerlink.com/content/978-3-642-04832-6/.

Crato:2010:TTP

[Cra10b] Nuno Crato. Turing’s test. Part 5. In Figuring It Out: Entertaining Encounters with Everyday Math [Cra10a], pages 183–186. ISBN 3-642-04832-3, 3-642-04833-1. LCCN QA99. URL http:/

/www.springerlink.com/content/m106l6507g659k34/.

Crockett:1994:TTF

[Cro94] Larry Crockett. The Turing test and the frame problem: AI’s mistaken understanding of intelligence. Ablex series in artificial intelligence. Ablex Pub. Corp., Norwood, NJ, USA, 1994. ISBN 0-89391-926-8 (hardcover), 1-56750-030-7. viii + 216 pp. LCCN Q341 .C76 1994.

Cardelli:2011:DCM

[CS11a] Luca Cardelli and William Shih, editors. DNA computing and molecular programming: 17 International Conference, DNA 17, Pasadena, CA, USA, September 19–23, 2011. proceedings, volume 6937 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-23637-5 (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-23637-2/.

Copeland:2011:DAT

[CS11b] B. Jack Copeland and Oron Shagrir. Do accelerating Turing machines compute the uncomputable? Minds and Machines, 21(2):221–239, May 2011. CODEN MMACEO. ISSN 0924-6495 (print), 1572-8641 (electronic). URL http://www.

springerlink.com/content/w60x1951124vl380/.

Clark:2012:RLA

[CS12] Liat Clark and Ian Steadman. The rich legacy of Alan Turing. Wired UK Web site., June 18, 2012. URL http://www.wired.

com/wiredscience/2012/06/alan-turing-legacy/.

Copeland:2019:CTT

[CS19] B. Jack Copeland and Oron Shagrir. The Church–Turing thesis: logical limit or breachable barrier? Communications of the Association for Computing Machinery, 62(1):66–74, January 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL https://cacm.acm.org/magazines/2019/1/
233526/fulltext.

Copeland:2017:WUC

[CSS17] Jack Copeland, Mark Sprevak, and Oron Shagrir. Is the whole universe a computer? In Copeland et al. [CBSW17], chapter 41, pages 445–462. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Cucker:2012:LTN

[Cuc12] Felipe Cucker. The legacy of Turing in numerical analysis. In Bieliková et al. [BFG^{+}12], pages 1–13. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
g6h0144587267827/.

Cuff:2024:CMH

[Cuf24] Thomas Mark Cuff. Computer memories, a history, corrected, revised & expanded. Web document, August 25,

2024. URL https://www.researchgate.net/publication/
355972939_Computer_Memories_A_History_Revision_3x.

Curtis:1965:TMS

[Cur65] M. W. Curtis. A Turing machine simulator. Journal of the ACM, 12(1):1–13, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cooper:2013:AMT

[CV13a] S. Barry Cooper and Jan Van Leeuwen. Alan Mathison Turing by Max Newman. In Cooper and van Leeuwen [CvL13], pages 3–12. ISBN 0-12-386980-3 (hardcover). LCCN QA29.T8

C65 2013. URL http://www.sciencedirect.com/science/
article/pii/B9780123869807500010.

Cooper:2013:TLL

[CV13b] S. Barry Cooper and Jan Van Leeuwen. Turing’s lecture to the London Mathematical Society on 20 February 1947. In Cooper and van Leeuwen [CvL13], pages 481– 497. ISBN 0-12-386980-3 (hardcover). LCCN QA29.T8

C65 2013. URL http://www.sciencedirect.com/science/
article/pii/B9780123869807500216.

Cooper:2013:TTE

[CV13c] S. Barry Cooper and Jan Van Leeuwen. Turing’s treatise on the Enigma (prof’s book). In Cooper and van Leeuwen [CvL13], pages 413–437. ISBN 0-12-386980-3 (hardcover). LCCN

QA29.T8 C65 2013. URL http://www.sciencedirect.com/
science/article/pii/B9780123869807500174.

Copeland:2017:Ba

[CVC17] Jack Copeland, Jean Valentine, and Catherine Caughey. Bombes. In Copeland et al. [CBSW17], chapter 12, pages 109– 128. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Cooper:2013:ATH

[CvL13] S. Barry Cooper and Jan van Leeuwen, editors. Alan Turing — His Work and Impact. Elsevier Science, Inc., Amsterdam, The Netherlands, 2013. ISBN 0-12-386980-3 (hardcover). xxi + 914 pp. LCCN QA29.T8 C65 2013. URL http:/

/amzn.to/VS1tdc;   http://store.elsevier.com/product.
jsp?isbn=9780123869807;   http://www.mathcomp.leeds.
ac.uk/turing2012/give-page.php?300.

Chang:2024:TIL

[CWS^{+}24] Lili Chang, Xinyu Wang, Guiquan Sun, Zhen Wang, and Zhen Jin. A time independent least squares algorithm for parameter identification of Turing patterns in reaction–diffusion systems. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00285-023-02026-z.

Carlucci:2012:NRT

[CZ12] Lorenzo Carlucci and Konrad Zdanowski. A note on Ramsey theorems and Turing jumps. In Cooper et al. [CDL12], pages 89– 95. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012. URL http:

//www.springerlink.com/content/lp57210160h12234/.

Dahlhaus:1995:GPM

[Dah95] Elias Dahlhaus. Gandy’s principles for mechanisms as a model of parallel computation. In Herken [Her95], pages 283–288. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.

springer.com/book/10.1007/978-3-7091-6597-3.

Dalrymple:2012:TBM

[Dal12a] David Dalrymple. Turing: Brain model still incomplete. Nature, 483(7389):275, March 14, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v483/n7389/full/483275c.html.

Daly:2012:BRA

[Dal12b] Peter Daly. Book review: Alan Turing and his Contemporaries, edited by Simon Lavington and others, ISBN-13 978-1-906124-90-8. ITNOW, 54(2):60, June 2012. CODEN ????

ISSN 1746-5710. URL http://itnow.oxfordjournals.org/
cgi/reprint/54/2/60?etoc.

Dasgupta:2014:IBB

[Das14] Subrata Dasgupta. It began with Babbage: the genesis of computer science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-930941-8 (hardcover), 0-19-930942-6 (e-book). x + 328 pp. LCCN QA76.17 .D36 2014.

Davis:1965:UBP

[Dav65] Martin Davis. The undecidable: basic papers on undecidable propositions, unsolvable problems and computable functions. Raven Press, New York, NY, USA, 1965. ISBN 0-911216-01-4. 439 pp. LCCN ????

Davis:1995:IML

[Dav95a] Martin Davis. Influences of mathematical logic on computer science. In Herken [Her95], pages 289–299. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.springer.

com/book/10.1007/978-3-7091-6597-3.

Davis:1995:MLO

[Dav95b] Martin Davis. Mathematical logic and the origin of modern computers. In Herken [Her95], pages 135–158. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.springer.

com/book/10.1007/978-3-7091-6597-3.

Davis:2000:UCR

[Dav00] Martin Davis. The universal computer: the road from Leibniz to Turing. W. W. Norton & Co., New York, NY, USA, 2000. ISBN 0-393-04785-7. xii + 257 pp. LCCN QA76.17. D38 2000; QA76.17 .D38 2000.

Davis:2004:MH

[Dav04] Martin Davis. The myth of hypercomputation. In Teuscher [Teu04a], pages 195–211. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Davis:2006:CTT

[Dav06a] Martin Davis. The Church–Turing thesis: Consensus and opposition. In Beckmann et al. [BBLT06], pages 125–132. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.59 .C67 2006. URL http:/

/link.springer.com/chapter/10.1007/11780342_13.

Davis:2006:WLT

[Dav06b] Martin Davis. What is … Turing reducibility? Notices of the American Mathematical Society, 53(10):1218–1219, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-

9477 (electronic). URL http://www.ams.org/notices/200610/
whatis-davis.pdf.

Davis:2012:UCR

[Dav12] Martin Davis. The universal computer: the road from Leibniz to Turing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, Turing centenary edition, 2012. ISBN 1-4665-0519-2 (paperback). xiv + 224 pp. LCCN QA76.17 .D38 2012.

Davies:2013:ECA

[Dav13] C. Davies. Enigma codebreaker Alan Turing receives royal pardon: Mathematician lost his job and was given experimental ‘chemical castration’ after being convicted for homosexual activity in 1952. The Guardian, ??(??):??, December 24, 2013. ISSN 0261-3077 (print), 1756-3224 (elec-

tronic). URL http://www.theguardian.com/science/2013/
dec/24/enigma-codebreaker-alan-turing-royal-pardon.

Davis:2018:UCR

[Dav18] Martin Davis. The Universal Computer: the Road from Leibniz to Turing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2018. ISBN 0-8153-8402-5 (hardback), 1-138-50208-1 (paperback), 1-315-14472-7 (e-book), 1-351-38481-3 (ePub). xv + 222 pp. LCCN QA76.17 .D38 2018.

Dawes:2016:ALD

[Daw16] Jonathan H. P. Dawes. After 1952: the later development of Alan Turing’s ideas on the mathematics of pattern formation. Historia Mathematica, 43(1):49–64, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086015000397.

Daylight:2012:DSE

[Day12a] Edgar G. Daylight. The Dawn of Software Engineering: from Turing to Dijkstra. Lonely Scholar, Heverlee, Belgium, 2012. ISBN 94-91386-02-6. vi + 239 pp. LCCN QA76.17 .D38 2012. Edited by Kurt De Grave.

Daylight:2012:DTR

[Day12b] Edgar G. Daylight. Deromanticizing Turing’s role in history. In The Dawn of Software Engineering: from Turing to Dijkstra [Day12a], chapter 8, pages 187–198. ISBN 94-91386-02-6. LCCN QA76.17 .D38 2012. Edited by Kurt De Grave.

Daylight:2012:TIP

[Day12c] Edgar G. Daylight. Turing’s influence on programming. In The Dawn of Software Engineering: from Turing to Dijkstra [Day12a], chapter 2, pages 13–42. ISBN 94-91386-02-6. LCCN QA76.17 .D38 2012. Edited by Kurt De Grave.

Daylight:2013:TPF

[Day13] Edgar G. Daylight. Turing’s 1936 paper and the first Dutch computers. Communications of the Association for Computing Machinery, ??(??):??, August 19, 2013. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://cacm.acm.org/blogs/blog-cacm/167012-turings-
1936-paper-and-the-first-dutch-computers/fulltext?mobile=
false.

Daylight:2021:AQW

[Day21] Edgar G. Daylight. Addressing the question What is a Program Text? via Turing scholarship. IEEE Annals of the History of Computing, 43(4):87–91, 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

DeBrosse:2004:SBU

[DB04] Jim DeBrosse and Colin B. Burke. The secret in Building 26: the untold story of America’s ultra war against the Uboat Enigma codes. Random House, New York, NY, USA, 2004. ISBN 0-375-50807-4, 1-58836-353-8, 0-375-75995-6. xxix + 272 pp. LCCN D810.C88 D43 2004. URL http://www.loc.

gov/catdir/samples/random045/2003058494.html;   http:
//www.randomhouse.com/catalog/display.pperl?isbn=9781588363534.

DeAngelis:2005:CPD

[DB05] Gina DeAngelis and David J. Bianco. Computers: processing the data. Innovators. Oliver Press, Minneapolis, MN, USA, 2005. ISBN 1-881508-87-0. 143 pp. LCCN QA76.17 D43 2005.

Pereira:2010:LCP

[dBPZM10] Hernane B. de B. Pereira, Gilney F. Zebende, and Marcelo A. Moret. Learning computer programming: Implementing a fractal in a Turing Machine. Computers and Education, 55(2):767–776, September 2010. CODEN COMEDR. ISSN 0360-1315 (print),

1873-782X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0360131510000898.

daCunha:2011:TMC

[dC11a] Aubrey da Cunha. Turing machines on Cayley graphs. In L. D. Beklemishev and Ruy de Queiroz, editors, Logic, language, information and computation: 18th international workshop, WOLLIC 2011, Philadelphia, PA, USA, proceedings, volume 6642 of Lecture Notes in Computer Science, pages 84–94. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-20919-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-20919-2/;   http:/
/www.springerlink.com/content/d4uktw837297480r/.

Dodig-Crnkovic:2011:SMC

[DC11b] Gordana Dodig-Crnkovic. Significance of models of computation, from Turing model to natural computation. Minds and Machines, 21(2):301–322, May 2011. CODEN MMACEO. ISSN 0924-6495 (print), 1572-8641 (electronic). URL http://www.

springerlink.com/content/613323432j506027/.

Dodig-Crnkovic:2012:ATL

[DC12]  Gordana Dodig-Crnkovic.  Alan Turing’s legacy: Info-

computational philosophy of nature. arxiv.org, July 2012. URL

http://adsabs.harvard.edu/abs/2012arXiv1207.1033D.

Dodig-Crnkovic:2013:ATL

[DC13]  Gordana Dodig-Crnkovic.  Alan Turing’s legacy: Info-

computational philosophy of nature. In Gordana Dodig-Crnkovic and Raffaela Giovagnoli, editors, Computing Nature: Studies in Applied Philosophy, Epistemology and Rational Ethics, volume 7, pages 115–123. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. URL http:/

/link.springer.com/chapter/10.1007/978-3-642-37225-4_
6.

Dormann:2001:FAT

[DDL01] Sabine Dormann, Andreas Deutsch, and Anna T. Lawniczak. Fourier analysis of Turing-like pattern formation in cellular automaton models. Future Generation Computer Systems, 17(7): 901–909, May 2001. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.elsevier.com/gej-
ng/10/19/19/45/34/33/abstract.html.

deFrutosEscrig:2012:ATP

[de 12] David de Frutos Escrig. Alan Turing: a personal assessment of his works. Gac. R. Soc. Mat. Esp., 15(4):675–695, 2012. ISSN 1138-8927.

Kruh:1998:TBW

[Dea98] Louis Kruh C. A. Deavours. The Turing bombe: was it enough? In Deavours et al. [DKK^{+}98], pages 403–421. ISBN 0-89006-862-3. LCCN Z103.S45 1998. US$78.20. URL http://www.

opengroup.com/open/cbbooks/089/0890068623.shtml. Third

volume of selected papers from issues of Cryptologia.

Delvenne:2006:TUD

[Del06] Jean-Charles Delvenne. Turing universality in dynamical systems. In Beckmann et al. [BBLT06], pages 147–152. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.59 .C67 2006. URL http:/

/link.springer.com/chapter/10.1007/11780342_16.

Dennett:2004:CMT

[Den04] Daniel C. Dennett. Can machines think? In Teuscher [Teu04a], pages 295–316. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Denning:2012:CSW

[Den12a] Peter J. Denning. Closing statement: What have we said about computation? The Computer Journal, 55(7):863–865, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/863.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Denning:2012:OSW

[Den12b] Peter J. Denning. Opening statement: What is computation? The Computer Journal, 55(7):805–810, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/805.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Denning:2012:RSC

[Den12c] Peter J. Denning. Reflections on a symposium on computation. The Computer Journal, 55(7):799–802, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/799.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Deutsch:1985:QTC

[Deu85] David Deutsch. Quantum theory, the Church–Turing Principle and the universal quantum computer. Proceedings of the Royal Society of London, A400(1818):97–117, July 8, 1985. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (elec-

tronic). URL http://rspa.royalsocietypublishing.org/
content/400/1818/97.

Dewdney:1989:TOE

[Dew89] A. K. Dewdney. The Turing omnibus: 61 excursions in computer science. Computer Science Press, Rockville, MD, USA, 1989. ISBN 0-7167-8154-9. xiv + 415 pp. LCCN QA76 D45 1989.

Dewdney:1992:TT

[Dew92] A. K. Dewdney. Turing test. Scientific American, 266(1): 30–32, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
30.pdf.

Dewdney:1993:NTO

[Dew93] A. K. Dewdney. The (new) Turing omnibus: 66 excursions in computer science. Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1993. ISBN 0-7167-8271-5. xvi + 455 pp. LCCN DA76.D45 1993; QA76 .D448 1993.

DeGaris:2009:AD

[DH09] Hugo De Garis and Sam Halioris. The artilect debate. In Epstein et al. [EBR09], pages 487–509. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_29.pdf.

Downey:2010:ART

[DH10] Rodney G. Downey and Denis R. Hirschfeldt. Algorithmic randomness and Turing reducibility. Part 2. In Rodney G. Downey and Denis R. Hirschfeldt, editors, Algorithmic Randomness and Complexity, Theory and Applications of Computability, pages 323–401. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-95567-4, 0-387-68441-7 (e-book). LCCN QA267.7 .D67 2010. URL http://www.

springerlink.com/content/978-0-387-95567-4/;   http:/
/www.springerlink.com/content/h4h487l0025m2711/.

Diaz:2012:TCO

[Día12] Josep Díaz. Turing’s Cathedral: The Origins of the Digital Universe, George Dyson. Pantheon Books (2012): Review. Computer Science Review, 6(4):185–186, July 2012. ISSN 1574-0137 (print), 1876-7745 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1574013712000147.

Dick:2013:CT

[Dic13] Stephanie Dick. A celebration of Turing. Endeavour, 37(4): 194–195, 2013. CODEN ENDEAS. ISSN 0160-9327 (print),

1873-1929 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0160932713000598.

Dilao:2005:TIP

[Dil05] Rui Dilão. Turing instabilities and patterns near a Hopf bifurcation. Applied Mathematics and Computation, 164(2):391–414, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300304003261.

Dediu:2011:LAT

[DIMV11] Adrian-Horia Dediu, Shunsuke Inenaga, and Carlos Martin-Vide, editors. Language and automata theory and applications: 5th international conference, LATA 2011, Tarragona, Spain, May 26–31, 2011. Proceedings, volume 6638 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-21253-0 (softcover). LCCN ????

Duque:2012:TPT

[DJ12] David Fernández Duque and Joost J. Joosten. Turing progressions and their well-orders. In Cooper et al. [CDL12], pages 212–221. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012. URL

http://www.springerlink.com/content/hl45rj1160764573/

.

Deavours:1990:TBW

[DK90] C. A. Deavours and Louis Kruh. The Turing bombe: Was it enough? Cryptologia, 14(4):331–349, October 1990. CO-DEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). Reprinted in [DKK^{+}98, pp. 403–421].

Deavours:1998:SCH

[DKK^{+}98] Cipher A. Deavours, David Kahn, Louis Kruh, Greg Mellen, and Brian J. Winkel, editors. Selections From Cryptologia: History, People, And Technology. The Artech House telecommunications library. Artech House Inc., Norwood, MA, USA, February 1998. ISBN 0-89006-862-3. vii + 552 pp. LCCN Z103.S45 1998.

US$78.20. URL http://www.opengroup.com/open/cbbooks/

089/0890068623.shtml. Third volume of selected papers from issues of Cryptologia.

Diez:2024:TPF

[DKM^{+}24] Antoine Diez, Andrew L. Krause, Philip K. Maini, Eamonn A. Gaffney, and Sungrim Seirin-Lee. Turing pattern formation in reaction–cross-diffusion systems with a bilayer geometry. Bulletin of Mathematical Biology, 86(2):??, February 2024. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s11538-023-01237-1.

Durand-Lose:2006:RCR

[DL06] Jérôme Durand-Lose. Reversible conservative rational abstract geometrical computation is Turing-universal. In Beckmann et al. [BBLT06], pages 163–172. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA9.59 .C67 2006. URL http://link.springer.com/
chapter/10.1007/11780342_18.

deLeeuw:1956:CPM

[dLMSS56] K. de Leeuw, E. F. Moore, C. E. Shannon, and N. Shapiro. Computability by probabilistic machines. In Automata studies, Annals of mathematics studies, no. 34, pages 183–212. Princeton University Press, Princeton, NJ, USA, 1956.

Dediu:2012:LAT

[DMV12] Adrian-Horia Dediu and Carlos Martin-Vide, editors. Language and automata theory and applications: 6th International Conference, LATA 2012, Coruña, Spain, March 5–9, 2012. Proceedings, volume 7183 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-28331-4 (soft cover). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-28331-4/.

Donofrio:2001:BIT

[Don01a] Nick Donofrio. BCS / IEE Turing Lecture 2001: Technology, innovation and the new economy. World-Wide Web

site., 2001. URL http://www.bcs.org/BCS/Awards/Events/
TuringLecture/turing2001.htm.

Donofrio:2001:TML

[Don01b] Nick Donofrio. Turing Memorial Lecture. The Computer Journal, 44(2):67–74, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/cgi/reprint/44/2/67;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_44/
Issue_02/440067.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_44/Issue_02/freepdf/440067.
pdf.

Donovan:2014:ATM

[Don14] Peter W. Donovan. Alan Turing, Marshall Hall, and the alignment of WW2 Japanese Naval intercepts. Notices of the American Mathematical Society, 61(3):258–264, March 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://www.ams.org/notices/201403/rnoti-p258.pdf.

Dowek:2012:APC

[Dow12a] Gilles Dowek. Around the physical Church–Turing thesis: Cellular automata, formal languages, and the principles of quantum theory. In Dediu and Martin-Vide [DMV12], pages 21–37. ISBN 3-642-28331-4 (soft cover). LCCN ???? URL http://

www.springerlink.com/content/r3137w88v8806107/.

Dowson:2012:TA

[Dow12b] Mark Dowson. Turing’s ACE. New Scientist, 214(2870):31, June 23, 2012. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0262407912616285.

Dowek:2013:ATR

[Dow13] Gilles Dowek. Alan Turing et la résolution numérique des équations différentielles. (French) [Alan Turing and the numerical solution of differential equations]. Gazette des Mathématiciens, 135:31–33, 2013. ISSN 0224-8999.

Downey:2014:CTA

[Dow14a] Rod Downey. Computability theory, algorithmic randomness and Turing’s anticipation. In Turing’s Legacy: Developments from Turing’s Ideas in Logic [Dow14c], chapter 5, pages 90–123. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Downey:2014:TLDa

[Dow14b] Rod Downey. Turing’s legacy: developments from Turing’s ideas in logic. In Turing’s Legacy: Developments from Turing’s Ideas in Logic [Dow14c], chapter 1, pages vii–x. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Downey:2014:TLDb

[Dow14c] Rod Downey, editor. Turing’s Legacy: Developments from Turing’s Ideas in Logic, volume 42 of Lecture Notes in Logic. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). ???? pp. LCCN ????

Downey:2017:TR

[Dow17] Rod Downey. Turing and randomness. In Copeland et al. [CBSW17], chapter 39, pages 427–436. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Drew:2002:NAS

[DP02] H. R. Drew and S. Pellegrino, editors. New Approaches to Structural Mechanics, Shells and Biological Structures, volume 104 of Solid Mechanics and Its Applications, 0925-0042. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 90-481-6120-7 (print), 94-015-9930-0 (e-book). x + 526 pp. LCCN TA349-359.

Dresner:2010:TCM

[Dre10] Eli Dresner. Turing on computation, memory and behavior. Rutherford Journal, 3(??):??, ???? 2010. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article030104.html.

deSojo:2013:TSD

[dSAL^{+}13] Aurea Anguera de Sojo, Juan Ares, Juan A. Lara, David Lizcano, María A. Martínez, and Juan Pazos. Turing and the serendipitous discovery of the modern computer. Foundations of Science, 18(3):545–557, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-013-9327-x.

Diaz:2012:PAT

[DT12] J. Díaz and C. Torras. A personal account of Turing’s imprint on the development of computer science. Computer Science Review, 6(5–6):225–234, November 2012. ISSN 1574-0137 (print),

1876-7745 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S157401371200041X.

Dutt:2010:TPA

[Dut10] A. K. Dutt. Turing pattern amplitude equation for a model glycolytic reaction-diffusion system. Journal of Mathematical Chemistry, 48(4):841–855, November 2010. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL http://

www.springerlink.com/content/e04jmlj1q418p52k/.

Demchenko:2009:WFW

[DV09] Eugene Demchenko and Vladimir Veselov. Who fools whom? In Epstein et al. [EBR09], pages 447–459. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_26.pdf.

Denning:2012:IWC

[DW12] Peter J. Denning and Peter Wegner. Introduction to what is computation. The Computer Journal, 55(7):803–804, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/803.full.pdf+html. Special Focus on the Centenary of Alan Turing.

DeBenedictis:2016:HWM

[DW16] Erik P. DeBenedictis and R. Stanley Williams. Help wanted: A modern-day Turing. Computer, 49(10):76–79, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL https://www.computer.org/csdl/mags/co/
2016/10/mco2016100076-abs.html.

Dyson:2012:ATG

[Dys12a] George Dyson. Alan Turing I: Der geistige Vater des Computers: Alan Turing gelang der Brückenschlag zwischen Logik und Maschinen; damit legte er die Basis für alle heutigen Computer. (German) [Alan Turing I: The spiritual father of the computer: Alan Turing succeeded in bridging the gap between logic and machinery, so he laid the basis for all of today’s computers]. Spektrum der Wissenschaft (German translation of Scientific American), ??(6):81–83, ???? 2012. CODEN SPEKDI. ISSN 0170-2971.

Dyson:2012:TCD

[Dys12b] George Dyson. Turing centenary: The dawn of computing. Nature, 482(7386):459–460, February 22, 2012. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v482/n7386/
full/482459a.html.

Dyson:2012:TCO

[Dys12c] George Dyson. Turing’s Cathedral: the Origins of the Digital Universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover), 1-4000-7599-8 (paperback). xxii + 401 + 32 pp. LCCN QA76.17 .D97 2012.

Epstein:2009:PTT

[EBR09] Robert Epstein, Grace Beber, and Gary Roberts, editors. Parsing the Turing Test: Philosophical and Methodological Issues in the Quest for the Thinking Computer. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). xxiii + 1 + 517 pp. LCCN Q335 .P35445

2008. URL http://www.springerlink.com/content/978-1-
4020-6710-5.

Edmonds:1995:SEI

[Edm95] Bruce Edmonds. The social embedding of intelligence. In Epstein et al. [ERB08], pages 211–235. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Edmonds:2003:CAI

[Edm03] Bruce Edmonds. The constructibility of artificial intelligence (as defined by the Turing Test). In Moor [Moo03b], pages 145–150. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003.

URL http://link.springer.com/chapter/10.1007/978-94-
010-0105-2_7/.

Edmonds:2009:SEI

[Edm09] Bruce Edmonds. The social embedding of intelligence. In Epstein et al. [EBR09], pages 211–235. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_14.pdf.

Edmonds:2012:LSI

[EG12] Bruce Edmonds and Carlos Gershenson. Learning, social intelligence and the Turing test: Why an “out-of-the-box” Turing machine will not pass the Turing test. In Cooper et al. [CDL12], pages 182–192. ISBN 3-642-30869-4. LCCN QA9.59

.C664 2012. URL http://www.springerlink.com/content/
098331403g670428/.

Eberbach:2004:TIM

[EGW04] Eugene Eberbach, Dina Goldin, and Peter Wegner. Turing’s ideas and models of computation. In Teuscher [Teu04a], pages 159–194. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Endresen:1991:TTA

[EH91] Alexander Endresen and Ivar Håkonsen. TAPE (Turing Application Programming Environment): attribute grammar and language-based editor for the Turing programming language.

Hovedoppgave i informatikk til cand. scient. eksamen [Thesis in computer science for the Cand. Scient. examination], Universitetet i Bergen, Bergen, Norway, 1991. Two volumes.

Elliott:2013:PPA

[Ell13] Francis Elliott. Posthumous pardon for Alan Turing over ‘gay activity’ conviction. The Times [London], Decem-

ber 24, 2013. URL http://www.thetimes.co.uk/tto/news/
politics/article3957457.ece.

Elliott:2019:ATF

[Ell19] Larry Elliott. Alan Turing to feature on new £50 note: Mathematician who cracked enigma code was persecuted for his homosexuality in 1950s. The Guardian, ??(??):??, July 15, 2019. ISSN 0261-3077 (print), 1756-3224 (electronic). URL https://

www.theguardian.com/business/2019/jul/15/alan-turing-
to-feature-on-new-50-note.

Emmer:2013:IMB

[Emm13] Michele Emmer, editor. Imagine math 2: between culture and mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 88-470-2888-4, 88-470-2889-2 (e-book). vii + 262 pp. LCCN QA10.7 .I485 2013.

Epstein:1995:QTC

[Eps95] Robert Epstein. The quest for the thinking computer. In Epstein et al. [ERB08], pages 3–12. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Epstein:2009:QTC

[Eps09] Robert Epstein. The quest for the thinking computer. In Epstein et al. [EBR09], pages 3–12. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_1.pdf.

Evans:1968:CKP

[ER68] Christopher Riche Evans and A. D. J. Robertson, editors. Cybernetics: key papers. University Park Press, Baltimore, MD, USA, 1968. ISBN 0-8391-0015-9. viii + 289 pp. LCCN Q310 .E8.

Epstein:2008:PTT

[ERB08] Robert Epstein, Gary Roberts, and Grace Beber, editors. Parsing the Turing Test. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). xxiii + 517 pp. LCCN Q335 .P35445 2008.

Erion:2003:CTA

[Eri03] Gerald J. Erion. The Cartesian test for automatism. In Moor [Moo03b], pages 241–251. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_13/.

Erskine:1984:ATE

[Ers84] Ralph Erskine. Alan Turing: The Enigma — book review. Cryptologia, 8(4):332–336, October 1984. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http://

www.informaworld.com/smpp/content~content=a741902615~
db=all~order=page.

Evans:1981:MMH

[Eva81] Christopher Evans. The Making of the Micro: a History of the Computer. Victor Gollanz, London, UK, 1981. ISBN 0-575-02913-7. 118 pp. LCCN ????

Etz:2017:JBH

[EW17] Alexander Etz and Eric-Jan Wagenmakers. J. B. S. Haldane’s contribution to the Bayes factor hypothesis test. Statistical Science, 32(2):313–329, May 2017. CODEN STSCEP.

ISSN 0883-4237. URL http://projecteuclid.org/euclid.
ss/1494489818.

Faizrahmanov:2010:CNF

[Fai10a] M. Kh. Faizrahmanov. Computable numberings of families of low sets and Turing jumps in the Ershov hierarchy. Siberian mathematical journal, 51(6):1135–1138, November 2010. CODEN SMTJAW. ISSN 0037-4466 (print), 1573-9260

(electronic). URL http://www.springerlink.com/content/
8117v360802216g7/.

Faizrakhmanov:2010:DLC

[Fai10b] M. Kh. Faizrakhmanov. Decomposability of low 2-computably enumerable degrees and Turing jumps in the Ershov hierarchy. Russian Mathematics (Iz VUZ), 54(12):51–58, December 2010. CODEN ???? ISSN 1066-369X (print), 1934-810X

(electronic). URL http://www.springerlink.com/content/
6r5wwj4341270374/.

Faizrakhmanov:2011:TJE

[Fai11] M. Kh. Faizrakhmanov. Turing jumps in the Ershov hierarchy. Algebra and Logic, 50(3):279–289, July 2011. CODEN ALL0A6. ISSN 0002-5232 (print), 1573-8302 (electronic). URL http://

www.springerlink.com/content/c1082qn15l118858/.

Fairhead:2012:BRA

[Fai12] Harry Fairhead. Book review: Alan Turing’s Electronic Brain, B. Jack Copeland, Oxford University Press, 592 Pages ISBN 978-0-19-960915-4. Web site, 2012. URL https:

//i-programmer.info/bookreviews/24-history/5091-alan-
turings-electronic-brain.html.

Floyd:2017:PEL

[FB17] Juliet Floyd and Alisa Bokulich, editors. Philosophical Explorations of the Legacy of Alan Turing: Turing 100, volume 324 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346. xvii + 361 pp. LCCN QA29.T8 P45 2017. URL https:/

/link.springer.com/book/10.1007/978-3-319-53280-6.

Feferman:1995:TLZ

[Fef95] Solomon Feferman. Turing in the land of O(z). In Herken [Her95], pages 103–134. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267

.U55 1995. URL http://link.springer.com/book/10.1007/
978-3-7091-6597-3.

Feffer:1999:BRT

[Fef99] Loren Butler Feffer. Book review: Turing’s Legacy: A History of Computing at the National Physical Laboratory, 1945–1995 by David Yates. Isis, 90(2):390–391, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237112.

Feferman:2006:TT

[Fef06] Solomon Feferman. Turing’s thesis. Notices of the American Mathematical Society, 53(10):1200–1206, November 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/notices/200610/fea-feferman.
pdf.

Fenstad:1995:LC

[Fen95] Jens Erik Fenstad. Language and computations. In Herken [Her95], pages 301–321. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267

.U55 1995. URL http://link.springer.com/book/10.1007/
978-3-7091-6597-3.

Feldman:1963:CTC

[FF63] Julian Feldman and Edward A. Feigenbaum, editors. Computers and thought: a collection of articles. McGraw-Hill, New York, NY, USA, 1963. xvi + 535 pp. LCCN Q 335.5 .F4 SMC.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

FragaDelfinoKunz:2024:NAP

[FGG^{+}24] Camile Fraga Delfino Kunz, Alf Gerisch, James Glover, Denis Headon, Kevin John Painter, and Franziska Matthäus. Novel aspects in pattern formation arise from coupling Turing reaction– diffusion and chemotaxis. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01225-5.

Fabrizio:2015:LAT

[FH15] Doug Fabrizio and Andrew Hodges. The life of Alan Turing. RadioWest interview., January 21, 2015. URL http://radiowest.

kuer.org/post/life-alan-turing.

Fokina:2014:CMT

[FHM14] Ekaterina B. Fokina, Valentina Harizanov, and Alexander Melnikov. Computable model theory. In Downey [Dow14c], chapter 6, pages 124–194. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Fienberg:2006:WDB

[Fie06] Stephen E. Fienberg. When did Bayesian inference become “Bayesian”? Bayesian Analysis, 1(1):1–40, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue01/fienberg.pdf;   http://projecteuclid.org/euclid.
ba/1340371071.

Fields:2015:CAP

[Fie15] Chris Fields. Co-authorship proximity of A. M. Turing Award and John von Neumann Medal winners to the disciplinary boundaries of computer science. Scientometrics, 104(3):809– 825, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-015-1575-9.

Finkelstein:1995:FP

[Fin95] David Finkelstein. Finite physics. In Herken [Her95], pages 323–347. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55

1995. URL http://link.springer.com/book/10.1007/978-
3-7091-6597-3.

Fisher:2015:NGB

[Fis15] Lawrence M. Fisher. News: Google boosts ACM’s Turing Award prize to $1 million. Communications of the Association for Computing Machinery, 58(1):31, January 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

cacm.acm.org/magazines/2015/1/181631/fulltext.

Fisher:2017:NTL

[Fis17] Lawrence M. Fisher. News: Turing laureates celebrate award’s 50th anniversary. Communications of the Association for Computing Machinery, 60(9):20–23, September 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://cacm.acm.org/magazines/2017/9/220435/fulltext.

Floyd:2017:TCS

[Flo17] Juliet Floyd. Turing on “common sense”: Cambridge resonances. In Floyd and Bokulich [FB17], chapter 5, pages 103–149.

ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346.

LCCN QA29.T8 P45 2017. URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-6_5.

Flynn:2002:CS

[Fly02] Roger R. Flynn, editor. Computer sciences, volume 1. Macmillan Publishing Company, New York, NY, USA, 2002. ISBN 0-02-865567-2 (Volume 1: Foundations: Ideas and People (hardcover)), 0-02-865568-0 (Volume 2: Software and Hardware), 0-02-865569-9 (Volume 3: Social Applications), 0-02-865570-2 (Volume 4: Electronic Universe). ???? pp. LCCN QA76 .C572 2002.

Feldmann:1971:EUT

[FOO71] H. Feldmann, H. Oberquelle, and C.-P. Ortlieb. Eine einfache universelle Turingmaschine in ALGOL 60 Simulation. (German) [A simple universal Turing machine in ALGOL 60 simulation]. Computing, 8(3–4):241–249, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fortnow:2012:ELT

[For12] Lance Fortnow. The enduring legacy of the Turing machine. The Computer Journal, 55(7):830–831, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/830.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Franzen:2006:WDI

[Fra06] Torkel Franzén. What does the incompleteness theorem add to the unsolvability of the halting problem? In Beckmann et al. [BBLT06], page 198. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA9.59 .C67 2006. URL http://link.springer.com/
chapter/10.1007/11780342_21.

Frailey:2012:CP

[Fra12] Dennis J. Frailey. Computation is process. The Computer Journal, 55(7):817–819, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/817.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Freeman:1986:ATC

[Fre86] W. J. Freeman. Alan Turing: The chemical basis of morphogenesis. In Palm and Aertsen [PA13], pages 235–236. ISBN

3-642-70913-3. LCCN ???? URL http://link.springer.com/
chapter/10.1007/978-3-642-70911-1_16.

Freeman:2012:CQW

[Fre12a] Peter A. Freeman. Consideration of the question “What Is Computation” considered harmful. The Computer Journal, 55(7): 861–862, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/55/7/861.full.pdf+html. Special Focus on the

Centenary of Alan Turing.

French:2012:DTT

[Fre12b] Robert M. French. Dusting off the Turing test. Science, 336(6078):164–165, April 13, 2012. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/336/6078/164.full.pdf.

French:2012:MBT

[Fre12c] Robert M. French. Moving beyond the Turing test. Communications of the Association for Computing Machinery, 55(12): 74–77, December 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frith:2005:AT

[Fri05] Holden Frith. Alan Turing. The Times [London], January 25,

2005. URL http://www.thetimes.co.uk/tto/life/fashion/
mensstyle/article1761995.ece.

Freer:2014:TCS

[FRT14] Cameron E. Freer, Daniel M. Roy, and Joshue B. Tenenbaum. Towards common-sense reasoning via conditional simulation: legacies of Turing in Artificial Intelligence. In Downey [Dow14c], chapter 7, pages 195–252. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Fu:2012:NTR

[Fu12] Bin Fu. NE is not NP Turing reducible to nonexponentially dense NP sets. In ????, editor, LATIN 2012: Theoretical Informatics, volume 7256 of Lecture Notes in Computer Science, pages 375– 386. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-29343-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-29343-6/;   http://www.springerlink.com/content/
d237503641vh0686/.

Furbach:2012:TKI

[Fur12] Ulrich Furbach. Turing und Künstliche Intelligenz. (German) [Turing and artificial intelligence]. Informatik Spektrum, 35(4):280–286, August 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://www.

springerlink.com/content/v85626771126985k/. Special Is-

sue: Alan Turing.

Galton:2006:CTT

[Gal06] Antony Galton. The Church–Turing thesis: Still valid after all these years? Applied Mathematics and Computation, 178(1):93– 102, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300305008374.

Garzon-Alvarado:2011:BHF

[GAM11] Diego A. Garzón-Alvarado and Angelica M. Ramírez Martinez. A biochemical hypothesis on the formation of fingerprints using a Turing patterns approach. Theoretical Biology and Medical Modelling, 8(1):24, December 2011. CODEN

???? ISSN 1742-4682. URL http://www.springerlink.com/
content/7911j562012827t6/.

Gams:2013:ATT

[Gam13] Matjaž Gams. Alan Turing, Turing machines and stronger. Informatica (Ljubljana, Slovenia), 37(1):9–14, 2013. CODEN IN-FOFF. ISSN 0350-5596 (print), 1854-3871 (electronic).

Gandy:1954:DMT

[Gan54] R. O. Gandy. Dr. A. M. Turing, O.B.E., F.R.S. Nature, 174(4429):535–536, September 18, 1954. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v174/n4429/pdf/
174535a0.pdf.

Gandy:1995:CI

[Gan95] Robin Gandy. The confluence of ideas in 1936. In Herken [Her95], pages 51–102. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55

1995. URL http://link.springer.com/book/10.1007/978-
3-7091-6597-3.

Garner:1995:THS

[Gar95] Robby Garner. The Turing Hub as a standard for Turing Test interfaces. In Epstein et al. [ERB08], pages 319–324. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Garner:2009:THS

[Gar09] Robby Garner. The Turing Hub as a standard for Turing Test interfaces. In Epstein et al. [EBR09], pages 319–324. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445

2008. URL http://link.springer.com/content/pdf/10.
1007/978-1-4020-6710-5_19.pdf.

Gasarch:2016:RTC

[Gas16] William Gasarch. Review of: Turing Computability: Theory and Applications by Robert Soare. ACM SIGACT News, 47(4):6–8, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Graham-Cumming:2012:ATCb

[GC12a] John Graham-Cumming. Alan Turing: Codebreaking and code-making. New Scientist, 214(2867):iv–v, June 2, 2012. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0262407912613773.

Graham-Cumming:2012:ATCa

[GC12b] John Graham-Cumming. Alan Turing: Computation. New Scientist, 214(2867):ii–iii, 2012. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0262407912613761.

Graham-Cumming:2012:ATI

[GC12c] John Graham-Cumming. Alan Turing: Intelligence & life. New Scientist, 214(2867):vi–vii, June 2, 2012. CODEN NWSCAL.

ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0262407912613785.

Graham-Cumming:2012:ATL

[GC12d] John Graham-Cumming. Alan Turing: Life, interrupted. New Scientist, 214(2867):viii, June 2, 2012. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0262407912613797.

Graham-Cumming:2012:EAY

[GC12e] John Graham-Cumming. Exploring the ancestry of your desktop PC: [book review:] Turing’s Cathedral: The origins of the digital universe by George Dyson, Allen Lane/ Pantheon. New Scientist, 213(2854):52, March 3–9, 2012. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0262407912605855.

Grabchak:2017:PTF

[GC17a] Michael Grabchak and Victor Cosme. On the performance of Turing’s formula: A simulation study. Communications in Statistics: Simulation and Computation, 46(6):4199–4209, 2017. CODEN CSSCDB. ISSN 0361-0918.

Greenish:2017:TM

[GC17b] Jonathan Bowen Simon Greenish and Jack Copeland. Turing’s monument. In Copeland et al. [CBSW17], chapter 19, pages 189–196. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Gee:2011:BPA

[Gee11] Sue Gee. Bletchley Park awarded funds for historic restoration. Web site, October 5, 2011. URL https://i-

programmer.info/news/82-heritage/3154-bletchley-park-
awarded-funds-for-historic-restoration.html.

Gee:2012:BRA

[Gee12a] Sue Gee. Book review: Alan Turing and His Contemporaries, Chris Burton, Martin Campbell-Kelly, Simon Lavington and Roger Johnson, British Computer Society, 100 pages, ISBN 978-1-906124-90-8. Web site, 2012. URL https://i-programmer.

info/bookreviews/24-history/4031-alan-turing-and-his-
contemporaries.html.

Gee:2012:CSA

[Gee12b] Sue Gee. Commemorative stamp for Alan Turing. Web site,

January 3, 2012. URL https://i-programmer.info/news/
82-heritage/3557-commemorative-stamp-for-alan-turing.
html.

Gelenbe:2012:NC

[Gel12] Erol Gelenbe. Natural computation. The Computer Journal, 55(7):848–851, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/55/7/848.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Goranzon:1991:DTA

[GF91] Bo Göranzon and Magnus Florin, editors. Dialogue and Technology: Art and Knowledge, The Springer Series on Artificial Intelligence and Society. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-19574-2, 1-4471-1731-X (e-book). ISSN 1431-0856. LCCN TJ210.2-211.495; Q334-342.

Granstrom:2012:TMD

[GG12] Helena Granström and Bo Göranzon. Turing’s man: a dialogue. AI and Society, ??(??):????, ???? 2012. CODEN AISCEM. ISSN 0951-5666 (print), 1435-5655 (electronic). URL http://www.

springerlink.com/content/x67x6577u6123153/.

Grattan-Guinness:2013:MAT

[GG13] I. Grattan-Guinness. The mentor of Alan Turing: Max Newman (1897–1984) as a logician. The Mathematical Intelligencer, 35 (3):54–63, September 2013. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://link.springer.

com/article/10.1007/s00283-013-9387-3.

Gratton-Guinness:2017:TMM

[GG17] Ivor Gratton-Guinness. Turing’s mentor, Max Newman. In Copeland et al. [CBSW17], chapter 40, pages 437–442. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Gabbay:2006:MPA

[GGZ06] Dov M. Gabbay, Sergei S. Goncharov, and Michael Zakharyaschev, editors. Mathematical problems from applied logic I: logics for the XXIst century, volume 4 of International mathematical series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-28688-8, 0-387-31072-X (e-book). ISSN 1571-5485 (print), 1574-8944 (electronic). xxviii + 348 pp. LCCN

QA9.A1 M38 2006. URL http://www.loc.gov/catdir/
enhancements/fy0663/2005931830-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2005931830-t.html.

Gherardi:2011:ATF

[Ghe11] Guido Gherardi. Alan Turing and the foundations of computable analysis. Bulletin of Symbolic Logic, 17(3):394–430, 2011. CO-DEN BSLOF3. ISSN 1079-8986 (print), 1943-5894 (electronic).

Ginoux:2019:BRJ

[Gin19] Jean-Marc Ginoux. Book review: Juliet Floyd; Alisa Bokulich, eds. Philosophical Explorations of the Legacy of Alan Turing: Turing 100. Isis, 110(4):851–852, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goranzon:1995:JAG

[GKO95] Bo Göranzon, Anders Karlqvist, and Eva Obenfeldner. Jenseits aller Gewissheit: die Begegnung zwischen Alan Turing und Ludwig Wittgenstein. (German) [Beyond all certainty: the meeting between Alan Turing and Ludwig Wittgenstein]. Haymon-Verlag, Innsbruck, Austria, 1995. ISBN 3-85218-203-4. 64 pp. LCCN ???? Translated from the Swedish and English by Eva Obenfeldner. With a foreword by Allan Janik.

Gladwin:2001:ATV

[Gla01] Lee A. Gladwin. Alan Turing’s visit to Dayton. Cryptologia, 25 (1):11–17, January 2001. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Gladwin:2003:AMT

[Gla03] Lee A. Gladwin. Alan M. Turing’s “Critique of Running Short Cribs on the U. S. Navy Bombe”. Cryptologia, 27(1):50–54, January 2003. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Gladwin:2004:AMT

[Gla04] Lee A. Gladwin. Alan M. Turing’s contributions to co-operation between the UK and the US. In Teuscher [Teu04a], pages 463– 473. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Glaschick:2012:ATW

[Gla12] Rainer Glaschick. Alan Turings Wirkung in Münster. (German) [Alan Turing’s influence in Münster]. Mitteilungen der Deutschen Mathematiker-Vereinigung, 20(1):42–48, 2012. ISSN 0947-4471.

Gambin:2012:PWD

[GMC12] Anna Gambin and Anna Marciniak-Czochra. Preface: Watching the daisies grow: from biology to biomathematics and bioinformatics — Alan Turing centenary special issue. Fundamenta Informaticae, 118(4):i–ii, 2012. CODEN FUINE8. ISSN 0169-2968 (print), 1875-8681 (electronic).

Good:2012:BTC

[GMT^{+}12] Irving John Good, Donald Michie, G. (Geoffrey) Timms, James A. Reeds, Whitfield Diffie, and Judith Veronica Field, editors. Breaking teleprinter ciphers at Bletchley Park: general report on Tunny with emphasis on statistical methods (1945). Wiley, New York, NY, USA, 2012. ISBN 0-470-46589-1 (hardcover). cxi + 673 pp. LCCN D810.C88 G66 2015.

Goldreich:1995:RIP

[Gol95] Oded Goldreich. Randomness, interactive proofs, and zeroknowledge — a survey. In Herken [Her95], pages 349–375. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http:

//link.springer.com/book/10.1007/978-3-7091-6597-3.

Gollifer:2012:ASA

[Gol12] Sue Gollifer. Art space: Alan Turing Year 2012. Digital Creativity, 23(3-4):307–311, December 2012.

Goncalves:2022:TTT

[Gon22] Bernardo Gonçalves. The Turing Test is a thought experiment. Minds and Machines, 33(1):1–31, November 2022. CODEN MMACEO. ISSN 1572-8641 (print), 1572-8641 (electronic).

Goncalves:2023:GRR

[Gon23a] Bernardo Gonçalves. Galilean resonances: the role of experiment in Turing’s construction of machine intelligence. Annals of Science, 81(3):359–389, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00033790.2023.
2234912.

Goncalves:2023:IPA

[Gon23b] Bernardo Gonçalves. Irony with a point: Alan Turing and his intelligent machine utopia. Philosophy & Technology, 36(3):50:1– 50:31, July 2023. ISSN 2210-5433 (print), 2210-5441 (electronic).

Goncalves:2024:LLO

[Gon24a] Bernardo Gonçalves. Lady Lovelace’s objection: The Turing– Hartree disputes over the meaning of digital computers, 1946– 1951. IEEE Annals of the History of Computing, 46(1):6– 18, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Goncalves:2024:TTB

[Gon24b] Bernardo Gonçalves. Turing’s test, a beautiful thought experiment. IEEE Annals of the History of Computing, 46(3):36–49, 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Good:1979:EWC

[Goo79a] I. J. Good. Early work on computers at Bletchley. Annals of the History of Computing, 1(1):38–48, July/September 1979. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1979/pdf/a1038.pdf;   http://
www.computer.org/annals/an1979/a1038abs.htm.

Good:1979:SHP

[Goo79b] I. J. Good. Studies in the history of probability and statistics. XXXVII. A. M. Turing’s statistical work in World War II. Biometrika, 66(2):393–396, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2335677.

Good:1984:TC

[Goo84] I. J. Good. Turing and the computer. Nature, 307(5952):663– 664, February 16, 1984. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v307/n5952/pdf/307663a0.pdf.

Good:1992:IRA

[Goo92] I. J. Good. Introductory remarks for the article in Biometrika 66 (1979), “A. M. Turing’s Statistical Work in World War II”. In Pure mathematics [Tur92b], pages 211–223. ISBN 0-444-88059-3. LCCN ???? Edited and with an introduction and postscript by J. L. Britton and Irvine John Good. With a preface by P. N. Furbank.

Good:2000:TAE

[Goo00] I. J. Good. Turing’s anticipation of empirical Bayes in connection with the cryptanalysis of the naval Enigma. Journal of Statistical Computation and Simulation, 66(2):101–111, 2000. CODEN JSCSAT. ISSN 0094-9655 (print), 1563-5163 (electronic). 50th Anniversary of the Department of Statistics, Virginia Tech, Part II (Blacksburg, VA, 1999).

Goranzon:1991:TP

[Gör91] Bo Göranzon. Turing’s paradox. In Göranzon and Florin [GF91], pages 85–92. ISBN 3-540-19574-2, 1-4471-1731-X (ebook). ISSN 1431-0856. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-1-4471-1731-5_10.

Goranzon:1995:STE

[Gör95a] Bo Göranzon, editor. Skill, Technology and Enlightenment: On Practical Philosophy, Artificial Intelligence and Society, 1431-0856. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-19920-9, 1-4471-3001-4 (ebook). ISSN 1431-0856. LCCN QA75.5-76.95.

Goranzon:1995:TP

[Gör95b] Bo Göranzon. Turing’s paradox. In Skill, Technology and Enlightenment: On Practical Philosophy [Gör95a], pages 85–92. ISBN 3-540-19920-9, 1-4471-3001-4 (e-book). ISSN 1431-0856. LCCN QA75.5-76.95.

Gottfried:1996:ATA

[Got96] Ted Gottfried. Alan Turing: the architect of the computer age. An impact biography. Franklin Watts, New York, NY, USA and London, UK, 1996. ISBN 0-531-11287-X. 128 + 16 pp. LCCN QA29.T8 G68 1996.

Goutefangea:1999:ATP

[Gou99] Patrick Goutefangea. Alan Turing: la “pensé” de la machine et l’idée de pratique. (French) [Alan Turing: machine thought and practical idea]. Thesis (doctoral), Département de philosophie, Université de Nantes, Nantes, France, 1999. 243 pp. Published in 2001 by Presses universitaires du Septentrion, Villeneuve d’Ascq, France.

Glaschick:2012:ATD

[GR12] Rainer Glaschick and Norbert Ryska. Alan Turing und Deutschland: Berührungspunkte. (German) [Alan Turing and Germany: points of contact]. Informatik Spektrum, 35(4):295–300, August 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-012-0624-5;   http://www.springerlink.
com/content/e740173816206188/. Special Issue: Alan Turing.

Grcar:2011:MGE

[Grc11] Joseph F. Grcar. Mathematicians of Gaussian elimination. Notices of the American Mathematical Society, 58(6):782–792, June/July 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Greenberg:2017:EM

[Gre17] Joel Greenberg. The Enigma machine. In Copeland et al. [CBSW17], chapter 10, pages 85–96. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Gheorghe:2012:MSM

[GS12] Marian Gheorghe and Mike Stannett. Membrane system models for super-Turing paradigms. Natural Computing, 11(2):253–259, ???? 2012. CODEN ???? ISSN 1567-7818 (print), 1572-9796

(electronic). URL http://www.springerlink.com/content/
k77jup2086985084/.

Gubb:1986:TF

[Gub86] D. Gubb. Turing’s fly. Nature, 323(6090):675, October 23, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v323/
n6090/pdf/323675c0.pdf.

Gurevich:1995:AWB

[Gur95] Yuri Gurevich. Algorithms in the world of bounded resources. In Herken [Her95], pages 377–385. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613.

LCCN QA267 .U55 1995. URL http://link.springer.com/
book/10.1007/978-3-7091-6597-3.

Goldstine:1951:NIM

[GvN51] Herman H. Goldstine and John von Neumann. Numerical inverting of matrices of high order. II. Proceedings of the American Mathematical Society, 2:188–202, 1951. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http:/

/www.jstor.org/view/00029939/di970628/97p0185x/0. See

[vNG47] for Part I. Reprinted in [Tau63b, Paper 15, pp. 558– 572].

Greenberg:2014:GWB

[GW14] Joel Greenberg and Rosamond Welchman. Gordon Welchman: Bletchley Park’s architect of ultra intelligence. Frontline Books, Barnsley, UK, 2014. ISBN 1-84832-752-8 (hardcover), 1-4738-3463-5 (e-book). xvi + 286 + 16 pp. LCCN TK5102.94 .G744

2014. URL http://lib.myilibrary.com?id=943722.

Hull:1985:NT

[HAC^{+}85] T. E. Hull, A. Abraham, M. S. Cohen, A. F. X. Curley, C. B. Hall, D. A. Penny, and J. T. M. Sawchuk. Numerical Turing. ACM SIGNUM Newsletter, 20(3):26–34, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Haeusler:2012:CAT

[Hae12] Edward Hermann Haeusler. A celebration of Alan Turing’s achievements in the year of his centenary. International Transactions in Operational Research, 19(3):487–491, 2012. CODEN ITORF9. ISSN 0969-6016 (print), 1475-3995 (electronic).

Haigh:2014:ATD

[Hai14] Thomas Haigh. Actually, Turing did not invent the computer. Communications of the Association for Computing Machinery, 57(1):36–41, January 2014. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Haigh:2016:BRT

[Hai16] Thomas Haigh. Book reviews: Turing: Pioneer of the Information Age. Jack Copeland. Oxford: Oxford University Press, 2013. 224 pp. $17.95 (paperback). (ISBN 978-0-19-871918-2). Journal of the Association for Information Science and Technology, 67 (7):1787–1789, July 2016. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haigh:2017:HRC

[Hai17] Thomas Haigh. Historical reflections: Colossal genius: Tutte, Flowers, and a bad imitation of Turing. Communications of the Association for Computing Machinery, 60(1):29–35, January 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://cacm.acm.org/magazines/2017/1/
211102/fulltext.

Hales:2013:MAT

[Hal13] T. Hales. Mathematics in the age of the Turing Machine.

arxiv.org, February 2013. URL http://adsabs.harvard.edu/
abs/2013arXiv1302.2898H.

Hales:2014:MAT

[Hal14] Thomas C. Hales. Mathematics in the age of the Turing machine. In Downey [Dow14c], chapter 8, pages 253–298. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Hamilton:1990:TT

[Ham90] Russell J. Hamilton. Turing test. Computers in Physics, 4 (3):224–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822908.

Hamer:2016:RPA

[Ham16] David H. Hamer. Review of Prof: Alan Turing Decoded by Dermot Turing. Cryptologia, 40(6):553–555, 2016. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Hanlon:2012:TFT

[Han12] Michael Hanlon. Turing’s flower theory blossoms. The Times [London], October 28, 2012. URL http://www.

thesundaytimes.co.uk/sto/news/uk_news/Environment/article1156602.
ece.

Hartree:1947:MTL

[Har47] D. R. Hartree. A. M. Turing’s lectures on ACE. World-Wide

Web document, 1947. URL http://turing.ecs.soton.ac.uk/
browse.php/B/2.

Harnad:2003:MMT

[Har03] S. Harnad. Minds, machines and Turing. In Moor [Moo03b], pages 253–273. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN

Q341 .T87 2003. URL http://link.springer.com/chapter/
10.1007/978-94-010-0105-2_14/.

Harnad:2012:ATH

[Har12a] Stevan Harnad. Alan Turing and the “hard” and “easy” problem of cognition: Doing and feeling. arxiv.org, June 2012. URL http:

//adsabs.harvard.edu/abs/2012arXiv1206.3658H.

Hartmanis:2012:TMI

[Har12b] Juris Hartmanis. Turing machine-inspired computer science results. In Cooper et al. [CDL12], pages 276–282. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012. URL http://www.

springerlink.com/content/5559v05340530r12/.

Hasslacher:1995:BTM

[Has95] Brosl Hasslacher. Beyond the Turing Machine. In Herken [Her95], pages 387–402. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267

.U55 1995. URL http://link.springer.com/book/10.1007/
978-3-7091-6597-3.

Hauser:2003:LWM

[Hau03] Larry Hauser. Look who’s moving the goal posts now. In Moor [Moo03b], pages 185–195. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_10/.

Hawking:2005:GCI

[Haw05] Stephen Hawking. God created the integers: the mathematical breakthroughs that changed history. Running Press Book Publishers, Philadelphia, PA and London, UK, 2005. ISBN 0-7624-1922-9 (hardcover). xiii + 1160 pp. LCCN

???? US$29.95. URL http://www.perseusbooksgroup.com/
runningpress/book_detail.jsp?isbn=0762419229.

Hayton:2017:BRT

[Hay17] Dan Hayton. Book review: The Turing Guide. Resurrection: The Journal of the Computer Conservation Society, ??(79):??, Autumn 2017. ISSN 0958-7403. URL https:

//computerconservationsociety.org/resurrection/res79.
htm#i.

Holt:1987:TPL

[HC87] Richard C. Holt and James R. Cordy. The Turing programming language. Technical Report 87-200, Department of Computing and Information Science, Queen’s University at Kingston, Kingston, Ontario, Canada, 1987.

Holt:1988:TPLa

[HC88] R. C. Holt and J. R. Cordy. The Turing programming language. Communications of the Association for Computing Machinery, 31(12):1410–1423, 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heath:2015:HNS

[Hea15] Nick Heath. Hacking the Nazis: The secret story of the women who broke Hitler’s codes. TechRepublic, ??(??):??, March

26, 2015. URL http://www.techrepublic.com/article/the-
women-who-helped-crack-nazi-codes-at-bletchley-park/.

Hejhal:2007:TBB

[Hej07] Dennis A. Hejhal. Turing: a bit off the beaten path. The Mathematical Intelligencer, 29(1):27–35, Winter 2007. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL

http://link.springer.com/article/10.1007/BF02984757.

Hellman:2017:TLC

[Hel17] Martin E. Hellman. Turing Lecture: Cybersecurity, nuclear security, Alan Turing, and illogical logic. Communications of the Association for Computing Machinery, 60(12):52–59, December 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL https://cacm.acm.org/magazines/2017/
12/223042-cybersecurity-nuclear-security-alan-turing-
and-illogical-logic.

Henderson:2011:ATC

[Hen11] Harry Henderson. Alan Turing: computing genius and wartime code breaker. Makers of modern science. Chelsea House, New York, NY, 2011. ISBN 0-8160-6175-0. xx + 133 pp. LCCN QA29.T8 H46 2011.

Herken:1988:UTM

[Her88] Rolf Herken, editor. The Universal Turing Machine: a halfcentury survey. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853741-7. xiv + 661 pp. LCCN QA267 .U55 1988.

Herken:1995:UTM

[Her95] Rolf Herken, editor. The Universal Turing Machine: a halfcentury survey, volume 2 of Computerkultur. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. xvi + 611 pp.

LCCN QA267 .U55 1995. URL http://link.springer.com/
book/10.1007/978-3-7091-6597-3.

Hertel:1998:QTM

[Her98] Joachim Hertel. Quantum Turing Machine simulator. Mathematica Journal, 8(3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hewitt:2013:WCA

[Hew13] Carl Hewitt. What is computation? Actor model versus Turing’s model. In Zenil [Zen13], pages 159–185. ISBN 981-4374-29-6. LCCN QA267.7 .C676 2013. URL http://www.

worldscientific.com/worldscibooks/10.1142/8306. Fore-

word by Roger Penrose.

Hayes:1995:TTC

[HF95] P. Hayes and K. Ford. Turing test considered harmful. In ????, editor, Proceedings of the 14th International Joint Conference on Artificial Intelligence (IJCAI 95), pages 972–977. ????, ????, 1995.

Hawkes:1998:BTL

[HfB98] Peter Hawkes and Association for Biometrics. Biometrics and Turing’s legacy. Information Security Technical Report, 3(1):

95–97, ???? 1998. CODEN ISTRFR. ISSN 1363-4127 (print),

1873-605X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1363412798800265.

Hochhuth:1989:AT

[HG89] Rolf Hochhuth and Ya‘akov Gotshlak. Alan Turing. Sifriyat po‘alim, Tel Aviv, Israel, 1989. ISBN 965-04-2049-5. 158 pp. LCCN ????

Holt:1984:ICS

[HH84] R. C. (Richard C.) Holt and J. N. P. Hume. Introduction to computer science using the Turing programming language. Reston Publishing Co. Inc., Reston, VA, USA, 1984. ISBN 0-8359-3168-4, 0-8359-3167-6 (paperback). xii + 404 pp. LCCN QA76 .H623 1984.

Hume:1990:ICS

[HH90] J. N. P. Hume and R. C. (Richard C.) Holt. Introduction to computer science using the Turing programming language. Holt Software Associates Inc., Toronto, ON, Canada, second edition, 1990. ISBN 0-921598-06-8. viii + 389 pp. LCCN QA 76 .H62 1990.

Husbands:2008:MMH

[HHW08] Philip Husbands, Owen Holland, and Michael Wheeler, editors. The mechanical mind in history. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-08377-9 (hardcover). viii + 458 pp. LCCN Q335 .M3956 2008.

Hicks:2008:RTH

[Hic08] Marie Hicks. Repurposing Turing’s ‘human brake’. IEEE Annals of the History of Computing, 30(4):108, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hidrogenesse:2012:GBD

[Hid12] Hidrogenesse. Un dígito binario dudoso: recital para Alan Turing: todas las canciones, letra y música, Segarra y Ballesteros . (Spanish) [A bit dubious recital for Alan Turing: all the songs, lyrics and music, Segarra and Ballesteros]. Austrohúngaro, D.L., Barcelona, Spain, 2012. One audio CD-ROM.

Hilton:1991:WAT

[Hil91] Peter Hilton. Working with Alan Turing. The Mathematical Intelligencer, 13(4):22–23, Fall 1991. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://link.

springer.com/article/10.1007/BF03028336.

Hill:1993:ATM

[Hil93] Chris Hill. Alan Turing: a mathematical genius. Altrincham History Society Occasional Papers, 6:10, 1993.

Hilton:2000:RRC

[Hil00a] Peter Hilton. Reminiscences and reflections of a codebreaker. In Joyner [Joy00], pages 1–8. ISBN 3-540-66336-3 (softcover), 3-642-59663-0 (e-book). LCCN QA268 .C67 1999.

UK£44.50. URL http://link.springer.com/chapter/10.

1007/978-3-642-59663-6_1. Proceedings of the Conference on Coding Theory, Cryptography and Number Theory held at the U.S. Naval Academy during October 25–26, 1998.

Hilton:2000:BGC

[Hil00b] Peter John Hilton. Breaking German codes. VHS video tape, Western Washington University., 2000.

Hilton:2017:MG

[Hil17] Peter Hilton. Meeting a genius. In Copeland et al. [CBSW17], chapter 3, pages 31–34. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Hinsen:2017:DSS

[Hin17] Konrad Hinsen. A dream of simplicity: Scientific computing on Turing machines. Computing in Science and Engineering, 19 (3):78–85, May/June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https://www.computer.

org/csdl/mags/cs/2017/03/mcs2017030078-abs.html.

Hammerstein:2002:TBA

[HL02] Peter Hammerstein and Olof Leimar. Theoretical biology: Ants on a Turing trail. Nature, 418(6894):141–142, July 11, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v418/
n6894/full/418141a.html.

Higman:1965:CIP

[HLOS65] Bryan Higman, W. F. Lunnon, C. F. J. Outred, and C. Strachey. Correspondence: An impossible program. The Computer Journal, 8(2):175–176, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_08/Issue_02/tiff/175.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/

Issue_02/tiff/176.tif. See [Str65, ApS65, BSPI65, AMKM66].

Harrison:1992:TON

[HM92] Harry Harrison and Marvin Lee Minsky. The Turing option: a novel. Warner Books, New York, NY, USA, 1992. ISBN 0-446-51565-5. 422 pp. LCCN PS3558.A667 T88 1992. US$21.95 (US$26.95 Can.).

Hofer:1996:TPF

[HM96] Thomas Höfer and Philip K. Maini. Turing patterns in fish skin? Nature, 380(6576):678, April 25, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v380/n6576/
pdf/380678a0.pdf.

Harman:2002:CSM

[HM02] P. M. (Peter Michael) Harman and Simon Mitton, editors. Cambridge scientific minds. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78100-0, 0-521-78612-6 (paperback). viii + 343 pp. LCCN Q141 .C1283

2002.  URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
cam021/2001025664.html;   http://www.loc.gov/catdir/
samples/cam033/2001025664.html;   http://www.loc.gov/
catdir/toc/cam021/2001025664.html.

Holt:1988:TPLb

[HMRC88] R. C. (Richard C.) Holt, Philip A. Matthews, J. Alan Rosselet, and James R. Cordy, editors. The Turing programming language: design and definition. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-933136-0 (paperback). viii + 325 pp. LCCN QA76.73.T85 T87 1988.

Hochhuth:1987:ATE

[Hoc87] Rolf Hochhuth. Alan Turing: Erzählung (German) [Alan Turing: Narration]. Rowohlt, Reinbek bei Hamburg, Germany, 1987. ISBN 3-498-02879-0. 188 pp. LCCN PT2668.O3 A64 1987.

Hodges:1983:ATEa

[Hod83a] Andrew Hodges. Alan Turing: the enigma. Burnett Books, London, UK, 1983. ISBN 0-09-152130-0. 587 + 8 pp. LCCN QA29.T8 H63 1983.

Hodges:1983:ATEb

[Hod83b] Andrew Hodges. Alan Turing: the enigma. Simon and Schuster, New York, NY, USA, 1983. ISBN 0-671-49207-1, 0-671-52809-2. x + 587 + [8] pp. LCCN QA29.T8 H63 1983.

Hodges:1985:ATE

[Hod85] Andrew Hodges. Alan Turing: the enigma of intelligence. Unwin Paperbacks, London, UK, 1985. ISBN 0-04-510060-8. 586 pp. LCCN ????

Hodges:1988:ATO

[Hod88] Andrew Hodges. Alan Turing ou l’énigme de l’intelligence. (French) [Alan Turing, or the enigma of intelligence]. Payot, Paris, France, 1988. ISBN 2-228-88081-7. 437 pp. LCCN ????

Hodges:1989:ATEa

[Hod89a] Andrew Hodges. Alan Turing, enigma. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1989. ISBN 3-211-82627-0. ???? pp. LCCN ????

Hodges:1989:ATEb

[Hod89b] Andrew Hodges. Alan Turing, enigma, volume 1 of Computerkultur. Kammerer & Unverzagt, Berlin, Germany, 1989. ISBN 3-9801050-5-9. 662 pp. LCCN ????

Hodges:1992:ATE

[Hod92] Andrew Hodges. Alan Turing: the enigma. Vintage, London, UK, 1992. ISBN 0-09-911641-3. xix + 586 + 8 pp. LCCN ????

Hodges:1994:ATEa

[Hod94a] Andrew Hodges. Alan Turing, Enigma, volume 1 of Computerkultur. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1994. ISBN 3-211-82627-0. 662 pp. LCCN ????

Hodges:1994:ATEb

[Hod94b] Andrew Hodges. Alan Turing, Enigma, volume 1 of Computerkultur. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1994. ISBN 3-7091-9381-8, 3-7091-5832-X. vi + 662 + 22 pp. LCCN TJ210.2-211.495; Q334-342.

Hodges:1994:UGS

[Hod94c] Andrew Hodges. Am Ufer. (German) [On the shore]. In Alan Turing, Enigma [Hod94b], pages 526–608. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-3-7091-9381-5_9.

Hodges:1994:AGC

[Hod94d] Andrew Hodges. Aufholjagd. (German) [Catching up]. In Alan Turing, Enigma [Hod94b], pages 299–361. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-342. URL

http://link.springer.com/chapter/10.1007/978-3-7091-
9381-5_6;   http://link.springer.com/content/pdf/bfm%
3A978-3-7091-9381-5/3/1.pdf.

Hodges:1994:ECG

[Hod94e] Andrew Hodges. Esprit de Corps. (German) [Team spirit]. In Alan Turing, Enigma [Hod94b], pages 3–54. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-3-7091-9381-5_1.

Hodges:1994:GWG

[Hod94f] Andrew Hodges. Geist der Wahrheit. (German) [Spirit of truth]. In Alan Turing, Enigma [Hod94b], pages 55–130. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-3-7091-9381-5_2.

Hodges:1994:NGE

[Hod94g] Andrew Hodges. Nachwort. (German) [Epilogue]. In Alan Turing, Enigma [Hod94b], pages 610–621. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-342. URL http:/

/link.springer.com/chapter/10.1007/978-3-7091-9381-5_
11.

Hodges:1994:NMG

[Hod94h] Andrew Hodges. Neue Männer. (German) [New men]. In Alan Turing, Enigma [Hod94b], pages 131–186. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-3-7091-9381-5_3.

Hodges:1994:PGP

[Hod94i] Andrew Hodges. Postskriptum. (German) [Postscript]. In Alan Turing, Enigma [Hod94b], page 609. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-342. URL http:/

/link.springer.com/chapter/10.1007/978-3-7091-9381-5_
10.

Hodges:1994:RRG

[Hod94j] Andrew Hodges. Relais–Rennen. (German) [Relay races]. In Alan Turing, Enigma [Hod94b], pages 187–280. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-3-7091-9381-5_4.

Hodges:1994:UGR

[Hod94k] Andrew Hodges. Überleitung. (German) [Reconciliation]. In Alan Turing, Enigma [Hod94b], pages 281–295. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-

342. URL http://link.springer.com/chapter/10.1007/
978-3-7091-9381-5_5.

Hodges:1994:VGD

[Hod94l] Andrew Hodges. Verzögerung. (German) [Delay]. In Alan Turing, Enigma [Hod94b], pages 362–449. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-342. URL http:/

/link.springer.com/chapter/10.1007/978-3-7091-9381-5_
7.

Hodges:1994:VGO

[Hod94m] Andrew Hodges. Vogelfrei. (German) [Outlaw]. In Alan Turing, Enigma [Hod94b], pages 450–525. ISBN 3-7091-9381-8, 3-7091-5832-X. LCCN TJ210.2-211.495; Q334-342. URL http://link.

springer.com/chapter/10.1007/978-3-7091-9381-5_8.

Hodges:1995:ATTa

[Hod95a] Andrew Hodges. Alan Turing and the Turing Machine. In Herken [Her95], pages 3–14. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267

.U55 1995. URL http://link.springer.com/book/10.1007/
978-3-7091-6597-3.

Hodges:1995:ATTb

[Hod95b] Andrew Hodges. Alan Turing and the Turing Test. In Epstein et al. [ERB08], pages 13–22. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Hodges:1997:ATHa

[Hod97a] A. Hodges. The Alan Turing home page. World-Wide Web site.,

1997. URL http://www.turing.org.uk/turing/.

Hodges:1997:ATHb

[Hod97b] Andrew Hodges. Alan Turing home page. World-Wide Web site.,

1997. URL http://www.wadham.ox.ac.uk/~ahodges/Turing.
html.

Hodges:1997:TNP

[Hod97c] Andrew Hodges. Turing: a natural philosopher, volume III of The great philosophers. Phoenix, London, UK, 1997. ISBN 0-7538-0192-2. 58 pp. LCCN QA29.T8 H62993 1997. URL http:

//www.turing.org.uk/philosophy/book.html.

Hodges:1999:T

[Hod99] Andrew Hodges. Turing, volume 3 of The great philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1999. ISBN 0-415-92378-6. 58 pp. LCCN QA29.T8 H632 1999.

Hodges:2000:ATE

[Hod00] Andrew Hodges. Alan Turing: the enigma. Walker, New York, NY, USA, 2000. ISBN 0-8027-7580-2 (paperback). xvii + 587 + 8 pp. LCCN QA29.T8 H63 2000.

Hodges:2001:ATO

[Hod01] Andrew Hodges. Alan Turing ou l’énigme de l’intelligence. (French) [Alan Turing, or the enigma of intelligence]. Payot, Lausanne, Switzerland, 2001. ISBN ???? ???? pp. LCCN ????

Hodges:2002:AT

[Hod02a] Andrew Hodges. Alan Turing. In Harman and Mitton [HM02], pages 253–268. ISBN 0-521-78100-0, 0-521-78612-6 (paperback). LCCN Q141 .C1283 2002. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/description/cam021/2001025664.html;   http://
www.loc.gov/catdir/samples/cam033/2001025664.html;
http://www.loc.gov/catdir/toc/cam021/2001025664.html.

Hodges:2002:EZSa

[Hod02b] Andrew Hodges. Enigma: Życie i śmierć Alana Turinga. (Polish) [Enigma: the life and death of Alan Turing]. Prószyński i S-ka, Warszawa, Poland, 2002. ISBN 83-7255-087-5. 460 + 1 + 8 pp. LCCN ???? Translation of [Hod00] to Polish by Wiktor Bartol.

Hodges:2003:MUA

[Hod03a] Andrew Hodges. Military use of Alan Turing. In Booss and Høyrup [BH03], pages 312–325. ISBN 3-7643-1634-9 , 0-8176-1634-9. LCCN QA10.8 .M38 2003. URL http://link.

springer.com/chapter/10.1007/978-3-0348-8093-0_16.

Hodges:2003:SEV

[Hod03b] Andrew Hodges. Storia di un enigma. Vita di Alan Turing (1912–1954). (Italian) [Story of an Enigma. Life of Alan Turing (1912–1954)]. Bollati Boringhieri, Torino, Italy, 2003. ISBN 88-339-1501-8. 762 pp. LCCN ????

Hodges:2004:ATI

[Hod04a] Andrew Hodges. Alan Turing: an introductory biography. In Teuscher [Teu04a], pages 3–8. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Hodges:2004:WWA

[Hod04b] Andrew Hodges. What would Alan Turing have done after 1954? In Teuscher [Teu04a], pages 43–58. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book).

LCCN QA29.T8 A57 2004. URL http://link.springer.com/
chapter/10.1007/978-3-662-05642-4_3. Foreword by Dou-

glas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Hodges:2006:BRB

[Hod06a] Andrew Hodges. Book reviews: B. Jack Copeland (ed.), The Essential Turing: The Ideas that Gave Birth to the Computer Age. Oxford: Clarendon Press, 2004. Pp. viii + 613. ISBN 0-19-825079-7. £50.00 (hardback). ISBN 0-19-825080-0. £14.99 (paperback). British Journal for the History of Science, 39(3): 470–471, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4028508.

Hodges:2006:ETB

[Hod06b] Andrew Hodges. The Essential Turing — a book review. Notices of the American Mathematical Society, 53(10):1190–1199, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/notices/
200610/rev-hodges.pdf.

Hodges:2008:ATL

[Hod08a] Andrew Hodges. Alan Turing, logical and physical. In Cooper [Coo08], pages 3–15. ISBN 0-387-36033-6 (hardcover), 0-387-

68546-4. LCCN QA9.59. URL http://link.springer.com/
chapter/10.1007/978-0-387-68546-5_1.

Hodges:2008:WDA

[Hod08b] Andrew Hodges. What did Alan Turing mean by “machine”? In Husbands et al. [HHW08], pages 75–90. ISBN 0-262-08377-9 (hardcover). LCCN Q335 .M3956 2008.

Hodges:2009:ATT

[Hod09] Andrew Hodges. Alan Turing and the Turing test. In Epstein et al. [EBR09], pages 13–22. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_2.pdf.

Hodges:2012:TTT

[Hod12a] A. Hodges. Turing and the test of time [spectral lines]. IEEE Spectrum, 49(6):8, June 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hodges:2012:ATI

[Hod12b] Andrew Hodges. Alan Turing: an introductory biography. Boletim da Sociedade Portuguesa de Matemática, 67:1–8, 2012. ISSN 0872-3672. Translated from the English [see MR2106942] by Ana Cristina Ferreira.

Hodges:2012:MHM

[Hod12c] Andrew Hodges. Alan Turing IV: Der Mann hinter der Maschine: Alan Turing ist heute für viele Leistungen berühmt; doch es dauerte lange, bis seine Arbeiten Anerkennung fanden. (German) [Alan Turing IV: The man behind the machine: Alan Turing is today famous for many services, but it was not until his work was recognized]. Spektrum der Wissenschaft (German translation of Scientific American), ?? (6):87–88, ???? 2012. CODEN SPEKDI. ISSN 0170-

2971. URL http://www.spektrum.de/alias/spezial/alan-
turing-iv-der-mann-hinter-der-maschine/1149658.

Hodges:2012:ATE

[Hod12d] Andrew Hodges. Alan Turing: the enigma. Princeton University Press, Princeton, NJ, USA, centenary edition, 2012. ISBN 0-691-15564-X (paperback), 1-4008-4497-5 (e-book). xxxi + 586 + 8 pp. LCCN QA29.T8 H63 2012.

Hodges:2012:BTM

[Hod12e] Andrew Hodges. Beyond Turing’s machines. Science, 336(6078): 163–164, April 13, 2012. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/336/6078/163.full.pdf.

Hodges:2012:TCR

[Hod12f] Andrew Hodges. Turing in context — response. Science, 336 (6089):1639, June 29, 2012. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/336/6089/1639.1.full.pdf.

Hodges:2014:ATE

[Hod14] Andrew Hodges. Alan Turing: The enigma: The book that inspired the film The Imitation Game. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-16472-X (paperback), 0-09-911641-3, 0-691-15564-X, 1-4481-3781-0, 1-78470-008-8, 1-4008-6512-3 (e-book). xxx + 736 + 8 pp. LCCN QA29.T8

H63 2012. URL http://press.princeton.edu/titles/10413.

html. With a foreword by Douglas Hofstadter and a new preface by the author.

Hofstadter:1983:BRM

[Hof83] Douglas R. Hofstadter. [book review:] mind, body and machine: Alan Turing, The Enigma, by Andrew Hodges. New York Times, pages BR1, BR24–BR25, November 13, 1983. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/docview/122176300.

Hofstadter:1985:RAT

[Hof85] Douglas R. Hofstadter. Review of “Alan Turing: The Enigma”. In Metamagical Themas: Questing for the Essence of Mind and Pattern, pages 483–491. Bantam Books, New York, NY, USA, 1985.

Holt:1986:DGT

[Hol86] Richard C. Holt. Design goals for the Turing programming language. Technical Report CSRI 187, Computer Systems Research Institute, University of Toronto, Toronto, Ontario, Canada, 1986. 44 pp.

Holt:1990:ICS

[Hol90] R. C. (Richard C.) Holt. Introduction to computer science using the Turing programming language. Holt Software Associates Inc., Toronto, Ontario, Canada, second edition, 1990. ISBN 0-921598-06-8. ???? pp. LCCN ????

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (ebook). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Hopcroft:1984:TM

[Hop84] John E. Hopcroft. Turing machines. Scientific American, 250(5):86–92, 94–98 (Intl. ed. 70–80), May 1984. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v250/n5/pdf/scientificamerican0584-86.pdf.

Hopcroft:2012:ITM

[Hop12] John Hopcroft. On the impact of Turing Machines. In Agrawal et al. [ACL12], pages 1–2. CODEN LNCSD9. ISBN 3-642-29951-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Horn:1995:TT

[Hor95] Robert E. Horn. The Turing Test. In Epstein et al. [ERB08], pages 73–88. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Horn:2009:TT

[Hor09] Robert E. Horn. The Turing Test. In Epstein et al. [EBR09], pages 73–88. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335

.P35445 2008. URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6710-5_5.pdf.

House-of-Lords:2012:ATS

[Hou12] House of Lords. Alan Turing (Statutory Pardon) Bill [HL]. Parliament of Great Britain, London, UK, 2012. ISBN 0-10-844204-

7. 1 pp. LCCN ???? URL http://services.parliament.uk/
bills/2012-13/alanturingstatutorypardon.html. A Bill to

give a statutory pardon to Alan Mathison Turing for offences under section 11 of the Criminal Law Amendment Act 1885 of which he was convicted on 31 March 1952. [First reading 25 July 2012; not yet passed].

Holt:1988:CPO

[HP88a] R. C. Holt and D. A. Penny. The concurrent programming of operating systems using the Turing Plus language. Technical Report ????, Computer Systems Research Institute, University of Toronto, Toronto, Ontario, Canada, 1988. 300 (est.) pp.

Holt:1988:CPU

[HP88b] R. C. Holt and D. A. Penny. Concurrent programming using the Turing Plus language. Technical Report ????, Computer Systems Research Institute, University of Toronto, Toronto, Ontario, Canada, 1988. 108 pp.

Hodges:2000:ATA

[HP00] Andrew Hodges and Kimmo Pietiläinen. Alan Turing, arvoitus. (Finnish) [Alan Turing, enigma]. Terra cognita, Helsinki, Finland, 2000. ISBN 952-5202-14-3. 604 + 8 pp. LCCN ????

Hey:2015:CUJ

[HP15] Anthony J. G. Hey and Gyuri Pápay. The computing universe: a journey through a revolution. Cambridge University Press, Cambridge, UK, 2015. ISBN 0-521-76645-1 (hardcover), 0-521-15018-3 (paperback) 1-316-12976-4 (e-book). 416 pp. LCCN

QA76.17 .H49 2015. URL http://alltitles.ebrary.com/
Doc?id=10992514.

Haigh:2020:HRN

[HP20] Thomas Haigh and Mark Priestley. Historical reflections: von Neumann thought Turing’s universal machine was ‘simple and neat.’: but that didn’t tell him how to design a computer. Communications of the Association for Computing Machinery, 63 (1):26–32, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3372920.

Harris:1982:TSS

[HS82] R. A. Harris and Leo Stodolsky. Two state systems in media and ‘Turing’s Paradox’. Physics Letters, B116(6):464–468, October 28, 1982. CODEN PYLBAJ. ISSN 0031-9163 (print), 1873-2410 (electronic).

Hinsley:1993:CIS

[HS93] F. H. Hinsley and Alan Stripp, editors. Codebreakers: the inside story of Bletchley Park. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-820327-6, 0-19-285304-X. xxi + 321 pp. LCCN D810.C88 M46 1993.

Homer:2014:TDC

[HS14] Steven Homer and Alan L. Selman. Turing and the development of computational complexity. In Downey [Dow14c], chapter 9, pages 299–328. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Horvath:2009:EDM

[HSD09] Judit Horváth, István Szalai, and Patrick De Kepper. An experimental design method leading to chemical Turing patterns. Science, 324(5928):772–775, May 8, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/324/5928/772.full.pdf.

Hillston:2012:SPA

[HTG12] Jane Hillston, Mirco Tribastone, and Stephen Gilmore. Stochastic process algebras: from individuals to populations. The Computer Journal, 55(7):866–881, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/866.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Humphrys:1995:HMP

[Hum95] Mark Humphrys. How my program passed the Turing Test. In Epstein et al. [ERB08], pages 237–260. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Humphrys:2009:HMP

[Hum09] Mark Humphrys. How my program passed the Turing Test. In Epstein et al. [EBR09], pages 237–260. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_15.pdf.

Humphries:2014:NLP

[Hum14] Will Humphries. ‘now let’s put Turing on a banknote’. The Times [London], October 9, 2014. URL http://www.thetimes.

co.uk/tto/arts/film/article4231364.ece.

Huskey:1991:MED

[Hus91] Harry D. Huskey. Memoir: The early days. Annals of the History of Computing, 13(3):290–306, July/September 1991. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1991/pdf/a3285.pdf;   http://www.computer.
org/annals/an1991/a3285abs.htm.

Hutchinson:1984:SNH

[Hut84] Alan Hutchinson. Short notes: the halting problem does not matter. The Computer Journal, 27(4):376, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/4/376.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_27/Issue_04/tiff/376.
tif.

Hutchens:1995:CSS

[Hut95] Jason L. Hutchens. Conversation simulation and sensible surprises. In Epstein et al. [ERB08], pages 325–342. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Hutchens:2009:CSS

[Hut09] Jason L. Hutchens. Conversation simulation and sensible surprises. In Epstein et al. [EBR09], pages 325–342. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008. URL http://link.

springer.com/content/pdf/10.1007/978-1-4020-6710-5_20.
pdf.

Hyman:2012:HAT

[Hym12] Paul Hyman. In honor of Alan Turing. Communications of the Association for Computing Machinery, 55(9):20–23, September 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Miller:2014:TMW

[III14] Charles F. Miller III. Turing machines to word problems. In Downey [Dow14c], chapter 10, pages 329–385. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Istrail:2013:ATJ

[IM13] Sorin Istrail and Solomon Marcus. Alan Turing and John von Neumann — their brains and their computers. Lecture Notes in Computer Science, 7762:26–35, 2013. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-36751-9_2.

Ireland:2017:WWW

[Ire17] Eleanor Ireland. We were the world’s first computer operators. In Copeland et al. [CBSW17], chapter 15, pages 161–166. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Irvine:2004:MNC

[Irv04] Andrew Irvine. MI5 neither confirms nor denies allegations concerning Alan Turing. The Mathematical Intelligencer, 26(2):57, Spring 2004. CODEN MAINDC. ISSN 0343-6993 (print), 1866-

7414 (electronic). URL http://link.springer.com/article/
10.1007/BF02985655.

Isaacson:2014:IHG

[Isa14] Walter Isaacson. The Innovators: How a Group of Hackers, Geniuses and Geeks Created the Digital Revolution. Simon and Schuster, New York, NY, USA, 2014. ISBN 1-4711-3879-8 (hardcover), 1-4767-0869-X (cloth), 1-4711-3897-6 (paperback), 1-4104-7497-6 (cloth), 1-4767-0870-3. viii + 542 pp. LCCN QA76.2.A2.

Ito:2010:PTM

[IST^{+}10] Takao Ito, Makoto Sakamoto, Ayumi Taniue, Tomoya Matsukawa, Yasuo Uchida, Hiroshi Furutani, and Michio Kono. Parallel Turing machines on four-dimensional input tapes. Artificial Life and Robotics, 15(2):212–215, September 2010. CODEN ???? ISSN 1433-5298 (print), 1614-7456 (electronic). URL http://

www.springerlink.com/content/j76634441j5445w7/.

Ibarra:2012:WSS

[IT12] Oscar H. Ibarra and Nicholas Q. Tran. Weak synchronization and synchronizability of multitape pushdown automata and Turing machines. In Dediu and Martin-Vide [DMV12], pages 337–350. ISBN 3-642-28331-4 (soft cover). LCCN ???? URL http://

www.springerlink.com/content/n485213l08032735/.

Ivey:2015:LVU

[Ive15] Prudence Ivey. Little Venice: A unique waterside location with high-end independent shops and cafe culture. Web site, Octo-

ber 29, 2015. URL http://www.hamhigh.co.uk/polopoly_fs/
1.4288666!/image/image.jpg_gen/derivatives/landscape_
630/image.jpg;   http://www.hamhigh.co.uk/property/little_
venice_a_unique_waterside_location_with_high_end_independent_
shops_and_cafe_culture_1_4288678. The Web site has a pho-

tograph of a blue plaque at the Colonnade Hotel in Westminster, labeled “English Heritage: Alan Turing 1912–1954 Code-breaker and Pioneer of Computer Science was born here.”.

Jacobs:2011:CWQ

[Jac11] Bart Jacobs. Coalgebraic walks, in quantum and Turing computation. In Martin Hofmann, editor, Foundations of software science and computational structures: 14th international conference, FOSSACS 2011, held as part of the joint European conference on theory and practice of software, ETAPS 2011, Saarbrucken, Germany, March 26–April 3, 2011. proceedings, volume 6604 of Lecture Notes in Computer Science, pages 12– 26. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-19804-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-19804-5/;   http://www.springerlink.com/content/
mq563u37r2114637/.

Jackson:2012:HAT

[Jac12] Joab Jackson. How Alan Turing set the rules for computing: The Turing Machine gave the world a model for how computers could operate. ComputerWorld Online, June

22, 2012. URL http://www.computerworld.com/s/article/
9228397/How_Alan_Turing_set_the_rules_for_computing.

James:2006:ASH

[Jam06] I. M. (Ioan Mackenzie) James. Asperger’s syndrome and high achievement: some very remarkable people. Jessica Kingsley, London, UK, 2006. ISBN 1-84310-388-5 (paperback). 224 pp. LCCN RC553.A88 J35 2006.

Jeandel:2012:ICT

[Jea12] Emmanuel Jeandel. On immortal configurations in Turing machines. In Cooper et al. [CDL12], pages 334–343. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012. URL http://www.

springerlink.com/content/0p3754hj27504621/.

Jefferson:1949:MMM

[Jef49] Geoffrey Jefferson. The mind of mechanical man. British Medical Journal, 1(4616):1105–1110, June 25, 1949. CODEN BMJOAE. ISSN 1468-5833. URL https://www.jstor.org/

stable/25372573.

Johnstone:2015:TPI

[Joh15] Adrian Johnstone. Turing: Pioneer of the Information Age by B. Jack Copeland (review). Technology and Culture, 56(3):772– 773, July 2015. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/593085.

Jones:2004:FBB

[Jon04] Allan Jones. Five 1951 BBC broadcasts on automatic calculating machines. IEEE Annals of the History of Computing, 26 (2):3–15, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://csdl.computer.

org/comp/mags/an/2004/02/a2003abs.htm;   http://csdl.
computer.org/dl/mags/an/2004/02/a2003.htm;   http://
csdl.computer.org/dl/mags/an/2004/02/a2003.pdf.

Jones:2016:RMC

[Jon16] Matthew L. (Matthew Laurence) Jones. Reckoning with Matter: Calculating Machines, Innovation, and Thinking About Thinking from Pascal to Babbage. University of Chicago Press, Chicago, IL, USA and London, UK, 2016. ISBN 0-226-41146-X (hardcover), 0-226-41163-X (e-book). 331 pp. LCCN QA75 .J66 2016.

Jones:2017:BRT

[Jon17] Cliff B. Jones. Book review: The Turing Guide, By Jack Copeland, Jonathan Bowen, Mark Sprevak, Robin Wilson and others. Oxford University Press, Oxford, UK, 26 January 2017, xv + 576 pp, 246 × 189 mm, ISBN: 978-0-19-874782-6 (Hardback, $75.00), ISBN: 978-0-19-874783-3 (Paperback, $19.99). Formal Aspects of Computing, 29(6):1121–1122, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00165-017-0446-y.

Jorge:2007:TWB

[Jor07] Martinez-Gil Jorge. Thinking on the Web: Berners-Lee, Gödel and Turing. The Computer Journal, 50(3):371–372, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/cgi/
content/full/50/3/371-a;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/3/371-a.

Joyner:2000:CTC

[Joy00] David Joyner, editor. Coding theory and cryptography: from Enigma and Geheimschreiber to quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66336-3 (softcover), 3-642-59663-0 (ebook). LCCN QA268 .C67 1999. UK£44.50. URL http:/

/frode.home.cern.ch/frode/pubs/cryptoday.pdf. Proceed-

ings of the Conference on Coding Theory, Cryptography and Number Theory held at the U.S. Naval Academy during October 25–26, 1998.

Jastrow:1997:GGE

[JTS97] Robert Jastrow, Alan M. Turing, and John Searle. Giza garuna eta ordenadorea: (garunaren eboluzioaz). (Basque) [The human brain and the computer: (the evolution of the brain)]. Gaiak, Donostia, Poland, 1997. ISBN 84-87203-98-1. 292 pp. LCCN ????

Kondo:1996:TPF

[KA96] Shigeru Kondo and Rihito Asai. Turing patterns in fish skin? Nature, 380(6576):678, April 25, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v380/n6576/
pdf/380678b0.pdf.

Knauff:1999:CCM

[KAB99] Bob Knauff, Isaac Asimov, and Harry Blairy. The colorful characters of mathematics. Carolina Mathematics, Burlington, NC, USA, 1999. ISBN ???? LCCN ???? Set of posters on mathematicians through history, each with a portrait and biographical information.

Kahn:1984:COS

[Kah84] David Kahn. Cryptology and the origins of spread spectrum: Engineers during World War II developed an unbreakable scrambler to guarantee secure communications between Allied leaders; actress Hedy Lamarr played a role in the technology. IEEE Spectrum, 21(9):70–80, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahan:2012:WMA

[Kah12] W. Kahan. What might Alan Turing say about the inevitable fallibility of software? Report, Department of Mathematics and Department of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, CA, USA, June 15,

2012. 1–12 pp. URL https://people.eecs.berkeley.edu/~

wkahan/15June12.pdf. Prepared for the ACM’s Celebration in San Francisco, 15–16 June 2012, of the Centennial of Alan Turing’s Birth.

Kanan:2012:TBO

[Kan12] Christopher Kanan. Turing: Beyond the original concept. Nature, 483(7389):275, March 14, 2012. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v483/n7389/full/
483275d.html.

Karlqvist:1995:LTL

[Kar95] Anders Karlqvist. The legacy of Turing — on the limits of the calculable. In Göranzon [Gör95a], pages 167–181. ISBN 3-540-19920-9, 1-4471-3001-4 (e-book). ISSN 1431-0856. LCCN

QA75.5-76.95. URL http://link.springer.com/chapter/10.
1007/978-1-4471-3001-7_16.

Kenner:1989:RDT

[Ken89] ?. Kenner. Review of Dewdney, The Turing Omnibus: 61 Excursions in Computer Science (1989). Byte Magazine, 14(??):??, ?? 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kennedy:2017:TGB

[Ken17] Juliette Kennedy. Turing, Gödel and the “bright abyss”. In Floyd and Bokulich [FB17], chapter 3, pages 63–91. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346. LCCN QA29.T8 P45

2017. URL https://link.springer.com/chapter/10.1007/
978-3-319-53280-6_3.

Krause:2024:TIE

[KGJ^{+}24] Andrew L. Krause, Eamonn A. Gaffney, Thomas Jun Jewell, Václav Klika, and Benjamin J. Walker. Turing instabilities are not enough to ensure pattern formation. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-023-01250-4.

Kjos-Hanssen:2019:OHB

[KH19] Bjørn Kjos-Hanssen. Only human [book review of The Turing guide, MR3618873]. Notices of the American Mathematical Society, 66(4):556–561, April 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kidwell:1996:CWM

[Kid96] Peggy Kidwell. Collected works of A. M. Turing — morphogenesis. IEEE Annals of the History of Computing, 18(4):69, October–December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://ieeexplore.ieee.

org/iel4/85/11673/00539923.pdf.

Kidwell:2006:ATA

[Kid06] Peggy Aldrich Kidwell. Alan Turing’s Automatic Computing Engine: The Master Codebreaker’s Struggle to Build the Modern Computer (review). Technology and Culture, 47(2):460– 462, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200857.

Kiefer:2012:AT

[Kie12] Dorothea Kiefer. Alan Turing (1912–1954). Mitteilungen der Deutschen Mathematiker-Vereinigung, 20(1):41, 2012. ISSN 0947-4471.

Kilov:2014:RUCa

[Kil14a] Haim Kilov. Review of The universal computer: the road from Leibniz to Turing by Martin Davis. ACM SIGACT News, 45 (3):29–31, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kilov:2014:RUCb

[Kil14b] Haim Kilov. Review of The Universal Computer. The Road from Leibniz to Turing by Martin Davis. ACM SIGACT News, 45(4): 17–20, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kurzweil:2009:WTT

[KK09] Ray Kurzweil and Mitchell Kapor. A wager on the Turing Test. In Epstein et al. [EBR09], pages 463–477. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_27.pdf.

Kleene:1995:TAC

[Kle95] Stephen C. Kleene. Turing’s analysis of computability, and major applications of it. In Herken [Her95], pages 15–49. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book).

ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.

springer.com/book/10.1007/978-3-7091-6597-3.

Kondo:2012:TPF

[Kon12] Shigeru Kondo. Turing pattern formation without diffusion. In Cooper et al. [CDL12], pages 416–421. ISBN 3-642-30869-4.

LCCN QA9.59 .C664 2012. URL http://www.springerlink.
com/content/452061701l8520j6/.

Koppel:1995:S

[Kop95] Moshe Koppel. Structure. In Herken [Her95], pages 403–419. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http:

//link.springer.com/book/10.1007/978-3-7091-6597-3.

Korner:1996:PC

[Kör96] T. W. (Thomas William) Körner. The Pleasures of Counting. Cambridge University Press, Cambridge, UK, 1996. ISBN 1-107-05056-1 (e-book), 0-521-56823-4, 0-521-56087-X, 1-316-02367-2. x + 534 pp. LCCN QA93 .K65 1996. URL http://catdir.loc.

gov/catdir/description/cam029/97108334.html;   http:/
/catdir.loc.gov/catdir/toc/cam027/97108334.html.

Kovac:2003:TLC

[Kov03] Carol Kovac. Turing Lecture 2003: Computing in the age of the genome. World-Wide Web site., 2003. URL http://www.bcs.

org/BCS/Awards/Events/TuringLecture/Turing2003/.

Kalos:2002:BRM

[KP02] Malvin H. Kalos and Douglass E. Post. Book review: Martin Davis, The Universal Computer: The Road From Leibniz to Turing. New York: W. W. Norton and Company, 2000, xii + 237 pages. $25.95 (cloth). Physics in Perspective (PIP), 4(1):118– 119, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kruh:2005:RTCa

[Kru05] Louis Kruh. Reviews and things cryptologic: Bamford, James. A Pretext For War: 9/11, Iraq, and the Abuse of America’s Intelligence Agencies. Doubleday. 2004. 420 pp. $26.95; Goldreich, Oded. Foundations of Cryptography: Volume II: Basic Applications. Cambridge University Press, 40 West 20th Street, New York NY 10011-4211 USA. 2004. 798 pp. $75.00; McBain. Ed. Hark!. Simon & Schuster, 1230 Avenue of the Americas, New York NY 10020 USA. 2004. 293 pp. $24.95; Spillman, Richard J. Classical and Contemporary Cryptology. Pearson Prentice Hall, Pearson Education, Inc., Upper Saddle River NJ 07458 USA. 2005. 285 pp. $54.00; Teuscher, Christof (Ed.) Alan Turing: Life and Legacy of a Great Thinker. Springer-Verlag, 175 Fifth Ave., New York NY 10010 USA. 2004. 542 pp. $69.95; Balliett, Blue and Brett Helquist, Illustrator. Chasing Vermeer. Scholastic Press, 557 Broadway, New York NY 10012 USA. 2004. 254 pp. $16.96; Caldwell, Ian and Dustin Thomason. The Rule of Four. The Dial Press/Random House, Inc., 1745 Broadway NY 10019 USA. 2004. 372 pp. $24.00/$34 Canada. Cryptologia, 29(1):88–93, January 2005. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL

http://www.findarticles.com/p/articles/mi_qa3926/is_
200501/ai_n13244737;   http://www.findarticles.com/p/
articles/mi_qa3926/is_200501/ai_n13244743;   http://
www.findarticles.com/p/articles/mi_qa3926/is_200501/
ai_n13244747;   http://www.findarticles.com/p/articles/
mi_qa3926/is_200501/ai_n13244749;   http://www.findarticles.
com/p/articles/mi_qa3926/is_200501/ai_n13244806;   http:
//www.informaworld.com/smpp/content~content=a748639586~
db=all~order=page.

Kurzweil:2004:LAR

[Kur04] Ray Kurzweil. The law of accelerating returns. In Teuscher [Teu04a], pages 381–416. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Katajainen:1988:FST

[KvLP88] Jyrki Katajainen, Jan van Leeuwen, and Martti Penttonen. Fast simulation of Turing machines by random access machines. SIAM Journal on Computing, 17(1):77–88, February 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kealy:2012:NSA

[KW12] Bonni J. Kealy and David J. Wollkind. A nonlinear stability analysis of vegetative Turing pattern formation for an interaction–diffusion plant–surface water model system in an arid flat environment. Bulletin of Mathematical Biology, 74(4):803–833, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/s11538-011-9688-7;
 http://link.springer.com/content/pdf/10.1007/s11538-
011-9688-7.pdf.

Lin:2012:AAA

[LA12] Jie Lin and Don Adjeroh. All-against-all circular pattern matching. The Computer Journal, 55(7):897–906, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/897.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Laplante:1996:GPC

[Lap96] Phillip Laplante, editor. Great Papers in Computer Science. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). iv + 717 pp. LCCN

QA76 .G686 1996. URL http://bit.csc.lsu.edu/~chen/
GreatPapers.html.

Lassgue:1995:DJI

[Las95] Jean Lassègue. Doing justice to the Imitation Game. In Epstein et al. [ERB08], pages 151–169. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Lassegue:1998:T

[Las98] Jean Lassègue. Turing. Belles Lettres, Paris, France, 1998. ISBN 2-251-76014-8. 210 pp. LCCN ????

Lassegue:2009:DJI

[Las09] Jean Lassègue. Doing justice to the Imitation Game. In Epstein et al. [EBR09], pages 151–169. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_11.pdf.

Lavington:1980:EBC

[Lav80] Simon Hugh Lavington. Early British Computers: the Story of Vintage Computers and the People Who Built Them. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1980. ISBN 0-932376-08-8 (paperback). 139 pp. LCCN

TK7885.A5L38. URL http://www.bitsavers.org/pdf/dec/_
Books/_Digital_Press/Lavington_Early_British_Computers_
1980.pdf;   https://ed-thelen.org/comp-hist/EarlyBritish.
html.

Lavington:2012:SBA

[Lav12] Simon Lavington. A synopsis of the book Alan Turing and his Contemporaries: Building the World’s First Computers. The Computer Journal, 55(7):779–787, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/779.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Labinger:2001:OCC

[LC01] Jay A. Labinger and H. M. (Harry M.) Collins, editors. The one culture?: a conversation about science. University of Chicago Press, Chicago, IL, USA and London, UK, 2001. ISBN 0-226-46722-8 (hardcover), 0-226-46723-6 (paper). xi + 329 pp. LCCN Q175.55.O54 2001.

Lavington:2012:ATH

[LCKBJ12] S. H. (Simon Hugh) Lavington, Martin Campbell-Kelly, Christopher P. Burton, and Roger Johnson, editors. Alan Turing and his contemporaries: building the world’s first computers. British Computer Society, London, UK, 2012. ISBN 1-906124-90-6 (paperback), 1-78017-105-6 (PDF e-book), 1-78017-106-4 (ePub ebook), 1-78017-107-2 (Kindle e-book). xiv + 111 pp. LCCN QA76.17 .A423 2012. UK£11.69.

Lengyel:1991:MTS

[LE91] István Lengyel and Irving R. Epstein. Modeling of Turing structures in the chlorite–iodide–malonic acid–starch reaction system. Science, 251(4994):650–652, February 8, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/251/4994/650.full.pdf.

Leavitt:2005:MWK

[Lea05] David Leavitt. The man who knew too much: Alan Turing and the invention of the computer. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2005. ISBN 0-393-05236-2 (hardcover). 319 pp. LCCN QA29.T8 L43 2005. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/toc/ecip0514/2005018034.html.

Leavitt:2007:ATH

[Lea07] David Leavitt. Alan Turing: l’homme qui inventa l’informatique. (French) [Alan Turing: the man who invented computer science]. Dunod, Paris, France, 2007. ISBN 2-10-050357-X. 276 pp. LCCN ???? French translation of [Lea05] by Julien Famonet.

Leavitt:2012:HQS

[Lea12] David Leavitt. El hombre que sabía demasiado: Alan Turing y la invención de la computadora. (Spanish) [The Man Who Knew Too Much: Alan Turing and the invention of the computer ]. Antoni Bosch, D.L., Barcelona, Spain, 2012. ISBN 84-95348-30-6. 304 pp. Spanish translation by Federico Corriente Basús of [Lea05].

Leavitt:2017:TP

[Lea17] David Leavitt. Turing and the paranormal. In Copeland et al. [CBSW17], chapter 32, pages 347–356. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Leatherdale:2019:BRA

[Lea19] Dik Leatherdale. Book review : Alan Turing’s Manchester. Resurrection: The Journal of the Computer Conservation Society, ??(87):??, Autumn 2019. ISSN 0958-

7403. URL https://computerconservationsociety.org/
resurrection/res87.htm#f.

Lehmer:1956:RRZ

[Leh56] D. H. Lehmer. On the roots of the Riemann zeta-function. Acta Mathematica, 95(1):291–298, December 1956. CODEN AC-MAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Lehman:1970:DZR

[Leh70] R. S. Lehman. On the distribution of zeros of the Riemann zetafunction. Proceedings of the London Mathematical Society. First Series, 3(20):303–320, ???? 1970. ISSN 0024-6115 (print), 1460-244X (electronic). This paper corrects several errors in [Tur53]. See also [Tru11].

Leiber:2001:TFI

[Lei01] Justin Leiber. Turing and the fragility and insubstantiality of evolutionary explanations: A puzzle about the unity of Alan Turing’s work with some larger implications. Philosophical Psychology, 14(1):83–94, March 2001. ISSN 0951-5089 (print), 1465-394X (electronic).

Lemire:2004:ATH

[Lem04] Laurent Lemire. Alan Turing: l’homme qui a croqué la pomme. (French) [Alan Turing: The man who broke the apple]. Hachette, Paris, France, 2004. ISBN 2-01-235618-4. 191 pp. LCCN QA29.T8.

Lemire:2012:ATH

[Lem12] Laurent Lemire. Alan Turing: l’homme qui a croquée la pomme. (French) [Alan Turing: the man who crunched the apple]. Fayard, Paris, France, 2012. ISBN 2-213-67196-6. 191 pp. LCCN ????

Lenat:1995:BMS

[Len95] Douglas B. Lenat. Building a machine smart enough to pass the Turing Test. In Epstein et al. [ERB08], pages 261–282. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Lenat:2009:BMS

[Len09] Douglas B. Lenat. Building a machine smart enough to pass the Turing Test. In Epstein et al. [EBR09], pages 261–282. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445

2008. URL http://link.springer.com/content/pdf/10.
1007/978-1-4020-6710-5_16.pdf.

Levy:1988:CCC

[Lev88] David N. L. Levy. Computer chess compendium. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-91331-9. 440 pp. LCCN GV1449.3 .L47 1988.

Levin:2006:MDT

[Lev06] Janna Levin. A madman dreams of Turing machines. Alfred A. Knopf, New York, NY, USA, 2006. ISBN 1-4000-4030-2. 230 pp. LCCN PS3612.E9238 M33 2006. URL http://www.loc.gov/

catdir/enhancements/fy0643/2005037124-s.html;   http:

//www.loc.gov/catdir/enhancements/fy0661/2005037124-
b.html;   http://www.loc.gov/catdir/enhancements/fy0661/
2005037124-d.html.

Levesque:2017:CST

[Lev17] Hector J. Levesque. Common sense, the Turing test, and the quest for real AI. MIT Press, Cambridge, MA, USA, 2017. ISBN 0-262-03604-5 (hardcover). xv + 172 pp. LCCN Q335 .L4634 2017.

Lewin:1978:UGW

[Lew78] Ronald Lewin. Ultra Goes to War: The Secret Story. Hutchinson, London, UK, 1978. ISBN 0-09-134420-4. 398 pp. LCCN ????

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Lee:2011:DTP

[LGB11] S. Seirin Lee, E. A. Gaffney, and R. E. Baker. The dynamics of Turing patterns for morphogen-regulated growing domains with cellular response delays. Bulletin of Mathematical Biology, 73(11):2527–2551, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:/

/link.springer.com/article/10.1007/s11538-011-9634-8;
 http://link.springer.com/content/pdf/10.1007/s11538-
011-9634-8.pdf.

Liu:2022:DPS

[LGS22] Haicheng Liu, Bin Ge, and Jihong Shen. Dynamics of periodic solutions in the reaction–diffusion glycolysis model: Mathematical mechanisms of Turing pattern formation. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300322003988.

Lehmann-Haupt:1983:BTA

[LH83] Christopher Lehmann-Haupt. Books of the times: Alan Turing: The Enigma. By Andrew Hodges. 587 pages. Illustrated. Simon & Schuster. $22.50. New York Times, page C25, November 10, 1983. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
docview/122159956.

Liebig:2011:KZE

[Lie11] Hans Liebig. Konrad Zuse, Erfinder des Computers — im Vergleich mit Alan Turing und John v. Neumann. (German) [Konrad Zuse, inventor of the computer — in comparison with Alan Turing and John von Neumann]. Informatik Spektrum, 34(6):553–564, December 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://www.

springerlink.com/content/h1j0r5m668715865/. Special is-

sue: Konrad Zuse.

Lipton:2011:ATG

[Lip11] Richard J. Lipton. The annotated Turing: a guided tour through Alan Turing’s historic paper on computability and the Turing machine [book review, Wiley, Indianapolis, IN, 2008]. Notices of the American Mathematical Society, 58(8):1120–1121, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lipton:2012:MTW

[Lip12] Richard J. Lipton. Might Turing have won a Turing Award? Computer, 45(6):96–97, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liskov:2012:KPP

[Lis12] Barbara Liskov. Keynote presentation: Programming the Turing machine. ACM SIGADA Ada Letters, 32(3):23–24, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Livesley:2002:EMW

[Liv02] R. K. Livesley. Elastic minimum-weight design: An encounter with Alan Turing. In Drew and Pellegrino [DP02], pages 155– 163. ISBN 90-481-6120-7 (print), 94-015-9930-0 (e-book). LCCN TA349-359.

Lindsay:1997:BC

[LJWH97] Charles Lindsay, Derek Jacobi, Hugh Whitemore, and Andrew Hodges. Breaking the code, 1997. ISBN 1-56442-662-9. Based on the play of the same title by Hugh Whitemore, and on the book, “Alan Turing: the enigma”, by Andrew Hodges. Originally broadcast as an episode of the PBS television series, Mobil masterpiece theatre Credits: Director of photography, Robin Vidgeon ; editor, Laurence Mery-Clark ; introduced by Russell Baker Performers: Derek Jacobi, Alun Armstrong, Richard Johnson, Harold Pinter, Amanda Root, Prunella Scales The story of Alan Turing, British mathematical genius and designer of the computer that broke the German Enigma code during World War II, whose admittance to homosexuality at a time when it was illegal presented problems for him, for his family, for his colleagues, and for the State’s preoccupation with national security.

Lengyel:1993:TTS

[LKE93] István Lengyel, Sándor Kádár, and Irving R. Epstein. Transient Turing structures in a gradient-free closed system. Science, 259(5094):493–495, January 22, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/259/5094/493.full.pdf.

Lassegue:2012:WTC

[LL12] Jean Lassègue and Giuseppe Longo. What is Turing’s comparison between mechanism and writing worth? In Cooper et al. [CDL12], pages 450–461. ISBN 3-642-30869-4. LCCN QA9.59

.C664 2012. URL http://www.springerlink.com/content/
2lg0174160715017/.

Lloyd:2012:TTF

[Llo12] Seth Lloyd. A Turing test for free will. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 370(1971):3597–3610, July 28, 2012. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2012RSPTA.
370.3597L.

Loebner:1995:HHT

[Loe95] Hugh Loebner. How to hold a Turing Test contest. In Epstein et al. [ERB08], pages 173–179. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Loebner:2009:HHT

[Loe09] Hugh Loebner. How to hold a Turing Test contest. In Epstein et al. [EBR09], pages 173–179. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_12.pdf.

Lolli:2013:AMT

[Lol13] Gabriele Lolli. Alan M. Turing (1912–1954). In Emmer [Emm13], pages 247–254. ISBN 88-470-2888-4, 88-470-2889-2 (e-book). LCCN QA10.7 .I485 2013.

Li:2001:TTP

[LOM^{+}01] Yong-Jun Li, Julia Oslonovitch, Nadia Mazouz, Florian Plenge, Katharina Krischer, and Gerhard Ertl. Turing-type patterns on electrode surfaces. Science, 291(5512):2395–2398, March 23, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/291/
5512/2395.full.pdf.

Lombardi:2005:LML

[Lom05] Gabriel Lombardi. L’aventure mathématique: liberté et rigueur psychotiques, Cantor, Gödel, Turing. (French) [Mathematical adventure: freedom and psychotic rigor: Cantor, Gödel, Turing]. In progress. Champ lacanien, Paris, France, 2005. ISBN 2-914332-10-6. 227 pp. LCCN ????

Longo:2009:LTI

[Lon09] Giuseppe Longo. Laplace, Turing and the “Imitation Game” impossible geometry. In Epstein et al. [EBR09], pages 377–411. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445

2008. URL http://link.springer.com/content/pdf/10.
1007/978-1-4020-6710-5_23.pdf.

Love:2004:BRL

[Lov04] Peter Love. Book reviews: The legacy of Alan Turing: Alan Turing: Life and Legacy of a Great Thinker, edited by Christof Teuscher. Computing in Science and Engineering, 6(4):97–99, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://csdl.

computer.org/comp/mags/cs/2004/04/c4097.pdf;   http:/
/csdl.computer.org/dl/mags/cs/2004/04/c4097.htm.

Lowe:2016:BCE

[Löw16] Benedikt Löwe. Barry Cooper (1943–2015): The engine of Computability in Europe. Computability, 5(1):3–11, ???? 2016. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Lakin:2011:MSV

[LP11] Matthew R. Lakin and Andrew Phillips. Modelling, simulating and verifying Turing-powerful strand displacement systems. In Cardelli and Shih [CS11a], pages 130–144. CODEN LNCSD9. ISBN 3-642-23637-5 (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/p02hj1m07hv83w20/.

Lara:2022:PMC

[LPAA22] Juan A. Lara, Juan Pazos, Aurea Anguera de Sojo, and Shadi Aljawarneh. The paternity of the modern computer. Foundations of Science, 27(3):1029–1040, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-021-09797-y.

Lighthill:1951:MCM

[LTM^{+}51] M. J. Lighthill, G. C. Tootill, J. C. P. Miller, A. M. Turing, and E. A. Newman. Manchester computing machine: general topics. Technical report, Manchester University, Manchester, UK, 1951. ???? pp. Reproduced in [WCK89, p. 194–196].

Leyshon:2021:DPD

[LTS^{+}21] Thomas Leyshon, Elisa Tonello, David Schnoerr, Heike Siebert, and Michael P. H. Stumpf. The design principles of discrete Turing patterning systems. Journal of Theoretical Biology, 531(??):Article 110901, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0022519321003209.

Lucas:1995:CTC

[Luc95] John Lucas. Commentary on Turing’s “Computing Machinery and Intelligence”. In Epstein et al. [ERB08], pages 67–70. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Lucas:2009:CTC

[Luc09] John Lucas. Commentary on Turing’s “Computing Machinery and Intelligence”. In Epstein et al. [EBR09], pages 67–70. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008. URL http://link.

springer.com/content/pdf/10.1007/978-1-4020-6710-5_4.
pdf.

Lupkowski:2011:TIG

[ŁW11] Paweł Łupkowski and Andrzej Wiśniewski. Turing interrogative games. Minds and Machines, 21(3):435–448, August 2011. CODEN MMACEO. ISSN 0924-6495 (print), 1572-8641

(electronic). URL http://www.springerlink.com/content/
441u757u26372n41/.

Macintyre:2012:ATW

[Mac12a] Ben Macintyre. Alan Turing was more than just a gay victim. The Times [London], June 22, 2012. URL http://www.

thetimes.co.uk/tto/opinion/columnists/benmacintyre/article3452827.
ece.

Macintyre:2012:RPB

[Mac12b] Ben Macintyre. The review: Pioneer broken by society’s code: The new exhibition that explores the life and work of Alan Turing — genius and martyr. The Times [London], July

4, 2012. URL http://www.thetimes.co.uk/tto/science/
eureka/article3464335.ece.

Madden:2012:JTD

[Mad12] Niall Madden. John Todd and the development of modern numerical analysis. Irish Mathematical Society Bulletin, 69:11–24, Summer 2012. ISSN 0791-5578. URL http://www.maths.tcd.

ie/pub/ims/bull69/Madden.pdf.

Mahon:2010:NEH

[Mah10] A. P. Mahon. Naval Enigma: the history of Hut Eight 1939–1945, volume 7 of Bletchley archive. Military Press, Milton Keynes, UK, 2010. ISBN 0-85420-443-1. 119 (est.) pp. LCCN ????

Mairs:2006:TLL

[Mai06] Chris Mairs. Turing Lecture 2006: Lifestyle access for the disabled — adding positive drift to the random walk with technology. World-Wide Web site., January 26, 2006. URL http://www.

bcs.org/BCS/Awards/Events/TuringLecture/Turing2006/.

Mairs:2007:IED

[Mai07] Chris Mairs. Inclusion and exclusion in the digital world: Turing Lecture 2006. The Computer Journal, 50(3):274–280, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/cgi/
content/abstract/50/3/274;   http://comjnl.oxfordjournals.
org/cgi/content/full/50/3/274;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/3/274.

Makowsky:1995:MIA

[Mak95] Johann A. Makowsky. Mental images and the architecture of concepts. In Herken [Her95], pages 421–432. ISBN 3-211-82637-8 (paperback), 3-211-82628-9, 3-7091-6597-0 (e-book). ISSN 0946-9613. LCCN QA267 .U55 1995. URL http://link.springer.

com/book/10.1007/978-3-7091-6597-3.

Malitz:1987:TM

[Mal87] I. Malitz. The Turing machine. Byte Magazine, 12(11):345–357, November 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mangel:1990:CTB

[Man90] Marc Mangel, editor. Classics of theoretical biology: from material presented at a meeting held on 5 July 1988 in Oxford, UK. Bulletin of mathematical biology. Pergamon, New York, NY, USA, 1990. ISBN ???? 334 (vol. 1), 3326 (vol. 2) pp. LCCN ???? Two volumes. Printed in Bulletin of mathematical biology (ISSN 0092-8240) 52(1/2) 1990 and 53(1/2) 1991.

Maruoka:2011:CCB

[Mar11a] Akira Maruoka. Computational complexity based on Turing machines. Part 4. In Concise guide to computation theory [Mar11b], pages 185–199. ISBN 0-85729-534-9 (hardcover).

LCCN ???? URL http://www.springerlink.com/content/
m18757l776854340/.

Maruoka:2011:CGC

[Mar11b] Akira Maruoka. Concise guide to computation theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-85729-534-9 (hardcover). ???? pp. LCCN

???? URL http://www.springerlink.com/content/978-0-
85729-534-7/.

Maruoka:2011:TMP

[Mar11c] Akira Maruoka. Turing machine. Part 3. In Concise guide to computation theory [Mar11b], pages 133–159. ISBN 0-85729-534-

9 (hardcover). LCCN ???? URL http://www.springerlink.
com/content/v41j818v370k66q8/.

Maruoka:2011:UTM

[Mar11d] Akira Maruoka. Universality of Turing machine and its limitations. Part 3. In Concise guide to computation theory [Mar11b], pages 161–181. ISBN 0-85729-534-9 (hardcover).

LCCN ???? URL http://www.springerlink.com/content/
k3325rx644280304/.

Margenstern:2013:BTM

[Mar13a] Maurice Margenstern. Bacteria, Turing Machines and hyperbolic cellular automata. In Zenil [Zen13], pages 209–230. ISBN 981-4374-29-6. LCCN QA267.7 .C676 2013. URL http://www.

worldscientific.com/worldscibooks/10.1142/8306. Fore-

word by Roger Penrose.

Margenstern:2013:CQT

[Mar13b] Maurice Margenstern. Ce qu’Alan Turing nous a laissé. (French) [What Alan Turing left us]. Gazette des Mathématiciens, 135: 17–31, 2013. ISSN 0224-8999.

Marton:2013:CGG

[Mar13c] Yuval Marton. Cylons, gaylons and gay grammar. AI and Society, October 2013. CODEN AISCEM. ISSN 0951-5666 (print), 1435-5655 (electronic). URL http://link.springer.

com/article/10.1007/s00146-013-0503-x.

Mason:2012:ATT

[Mas12] Caroline Mason. An artistic Turing test. ITNOW, 54(4): 58–59, December 2012. CODEN ???? ISSN 1746-5710.

URL http://itnow.oxfordjournals.org/content/54/4/58.
full.pdf?etoc.

Mauldin:2009:GUC

[Mau09] Michael L. Mauldin. Going under cover: Passing as human. In Epstein et al. [EBR09], pages 413–429. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_24.pdf.

Mays:1952:CMT

[May52] W. Mays. Can machines think? Philosophy, 27(101):148–162, April 1952. ISSN 0031-8191 (print), 1469-817X (electronic).

May:1961:RPA

[May61] Kenneth O. May. Recent publications: Alan M. Turing, by Sara Turing. American Mathematical Monthly, 68(8):827, October 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mays:2001:MRT

[May01] Wolfe Mays. My reply to Turing: Fiftieth anniversary. Journal of the British Society for Phenomenology, 32(1):4–23, January 2001. ISSN 0007-1773 (print), 2332-0486 (electronic).

Maini:2006:TMC

[MBC06] Philip K. Maini, Ruth E. Baker, and Cheng-Ming Chuong. The Turing model comes of molecular age. Science, 314(5804):1397– 1398, December 1, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/314/5804/1397.full.pdf.

Malloy:2011:SLE

[MBS11] Brian Malloy, Mark Brand, and Steffen Staab, editors. Software Language Engineering: Third International Conference, SLE 2010, Eindhoven, The Netherlands, October 12–13, 2010, Revised Selected Papers, volume 6563 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-19439-7. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
978-3-642-19439-9/.

Millican:1996:LAT

[MC96] P. J. R. (Peter J. R.) Millican and Andy Clark, editors. The legacy of Alan Turing: Machines and Thought, volume 1 of Mind Association occasional series. Clarendon Press, Oxford, UK, 1996. ISBN 0-19-823593-3. x + 297 pp. LCCN Q335.5 .L44 1996. See also volume 2 [CM96].

Mainzer:2012:ILT

[MC12a] Klaus Mainzer and Leon Chua. Introduction: Leibniz, Turing, Zuse, and beyond. In The Universe as Automaton: From Simplicity and Symmetry to Complexity [MC12b], pages 1–16. ISBN 3-642-23476-3. LCCN ???? URL http://www.springerlink.

com/content/w971m8857498w085/.

Mainzer:2012:UAH

[MC12b] Klaus Mainzer and Leon Chua, editors. The Universe as Automaton: From Simplicity and Symmetry to Complexity. SpringerBriefs in Complexity. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-23476-3. LCCN ????

McGrayne:2011:TWH

[McG11] Sharon Bertsch McGrayne. The theory that would not die: how Bayes’ rule cracked the Enigma code, hunted down Russian submarines, and emerged triumphant from two centuries of controversy. Yale University Press, New Haven, CT, USA, 2011. ISBN 0-300-16969-8. xiii + 320 pp. LCCN QA279.5 .M415 2011.

McGinnes:2012:NCD

[McG12] Jamie McGinnes. News: Codebreaker ‘did not kill himself‘: New riddle over death of Enigma cracker as biographer claims mathematician may have accidentally poisoned himself with toxic chemical. The Times [London], June 24,

2012. URL http://www.thesundaytimes.co.uk/sto/news/
uk_news/Science/article1067749.ece.

Mckinstry:1995:MS

[Mck95] Chris Mckinstry. Mind as space. In Epstein et al. [ERB08], pages 283–299. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Mckinstry:2009:MS

[Mck09] Chris Mckinstry. Mind as space. In Epstein et al. [EBR09], pages 283–299. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445

2008. URL http://link.springer.com/content/pdf/10.
1007/978-1-4020-6710-5_17.pdf.

Martin-Delgado:2011:ATO

[MD11] Miguel-Angel Martin-Delgado. Alan Turing and the origins of complexity. arxiv.org, October 2011. URL http://adsabs.

harvard.edu/abs/2011arXiv1110.0271M.

Meier:2012:ATG

[Mei12a] Christian Meier. Alan Turing: Der Geist in der universellen Rechenmaschine. (German) [Alan Turing: The ghost in the universal computing machine]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, ???? 2012. CODEN SPEKDI. ISSN 0170-2971. URL http:

//www.spektrum.de/alias/alan-turing/der-geist-in-der-
universellen-rechenmaschine/1155170.

Meinhardt:2012:MBM

[Mei12b] Hans Meinhardt. Modelle zur biologischen Musterbildung: Turings Theorie und die spätere Entdeckung der Rolle von lokaler Selbstverstärkung und lang-reichweitiger Inhibition. (German) [Models for biological pattern formation: Turing’s theory and the later discovery of the role of local self-enhancement and longrange inhibition]. Informatik Spektrum, 35(4):287–294, August 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X

(electronic). URL http://www.springerlink.com/content/

w612722x6002n503/. Special Issue: Alan Turing.

Metz:2019:TBP

[Met19] Cade Metz. Tony Brooker, pioneer of computer programming, dies at 94. New York Times, ??(??):??, December 13, 2019. URL

https://en.wikipedia.org/wiki/Tony_Brooker;   https:
//www.nytimes.com/2019/12/13/technology/tony-brooker-
dead.html.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Michie:1980:TOC

[Mic80] Donald Michie. Turing and the origins of the computer. New Scientist, 85(195):580–583, 1980. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Michie:1984:LMG

[Mic84] D. Michie. A loner, a misfit, a genius. New Scientist, ??(??): 36–37, ???? 1984. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Michie:2008:ATM

[Mic08] Donald Michie. Alan Turing’s mind machines. In Husbands et al. [HHW08], pages 61–74. ISBN 0-262-08377-9 (hardcover). LCCN Q335 .M3956 2008.

Micali:2015:VWI

[Mic15] Silvio Micali. Viewpoint: What it means to receive the Turing award. Communications of the Association for Computing Machinery, 58(1):52–53, January 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.

acm.org/magazines/2015/1/181611/fulltext.

Minsky:1967:CFI

[Min67] Marvin Lee Minsky. Computation: Finite and Infinite Machines. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1967. xvii + 317 pp. LCCN QA267 .M55.

Minsky:1972:CFI

[Min72] Marvin L. Minsky. Computation: Finite and Infinite Machines. Prentice-Hall Series in Automatic Computation; Open University set book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1972. ISBN 0-13-165449-7. xvii + 317 pp. LCCN QA76.

Mishra:2009:TPW

[Mis09] Bud Mishra. Technical perspective: Where biology meets computing. Communications of the Association for Computing Machinery, 52(3):96, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mitchell:2012:BC

[Mit12] Melanie Mitchell. Biological computation. The Computer Journal, 55(7):852–855, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/55/7/852.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Morris:1984:EPP

[MJ84] F. L. Morris and C. B. Jones. An early program proof by Alan Turing. Annals of the History of Computing, 6(2): 139–143, April/June 1984. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1984/
pdf/a2139.pdf;   http://www.computer.org/annals/an1984/
a2139abs.htm.

Mumford:2009:CIC

[MJ09] Christine L. Mumford and L. C. Jain, editors. Computational Intelligence: Collaboration, Fusion and Emergence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-01798-3 (print), 3-642-01799-1 (e-book).

LCCN Q342 .C66 2009. URL http://link.springer.com/
openurl?genre=book&isbn=978-3-642-01798-8.

Meltzer:1969:MI

[MM69a] Bernard Meltzer and Donald Michie, editors. Machine intelligence, volume 5. Edinburgh University Press, Edinburgh, Scotland, 1969. CODEN ???? ISBN 0-85224-176-3. ISSN 0076-2032, 0541-6418. vii + 588 pp. LCCN Q336. With a previously unpublished report by A. M. Turing. Annual machine Intelligence Workshop.

Michie:1969:MI

[MM69b] Donald Michie and Bernard Meltzer, editors. Machine intelligence 5. Edinburgh University Press, Edinburgh, Scotland, 1969. ISBN 0-85224-176-3. 596 pp. LCCN Q336.

Meltzer:1972:MI

[MM72] Bernard Meltzer and Donald Michie, editors. Machine intelligence 7. Edinburgh University Press, Edinburgh, UK, 1972. ISBN 0-85224-234-4. LCCN ????

Mladenic:2013:EIS

[MMB13] Dunja Mladenić, Stephen Muggleton, and Ivan Bratko. Editors’ introduction to the Special Issue on “100 years of Alan Turing and 20 years of SLAIS”. Informatica (Ljubljana, Slovenia), 37 (1):1, 2013. CODEN INFOFF. ISSN 0350-5596 (print), 1854-3871 (electronic).

Montalban:2019:MCC

[Mon19] Antonio Montalbán. Martin’s conjecture: a classification of the naturally occurring Turing degrees. Notices of the American Mathematical Society, 66(8):1209–1215, September 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moor:2003:SFT

[Moo03a] James H. Moor. The status and future of the Turing Test. In The Turing Test: the elusive standard of artificial intelligence [Moo03b], pages 197–213. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_11/.

Moor:2003:TTE

[Moo03b] James H. Moor, editor. The Turing Test: the elusive standard of artificial intelligence, volume 30 of Studies in cognitive systems. Kluwer Academic Publishers, Dordrecht, The Netherlands, 2003. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. vi + 273 pp.

LCCN Q341 .T87 2003. URL http://www.springerlink.com/
content/978-94-010-0105-2.

Moody:2014:DMW

[Moo14] Oliver Moody. Death of man who cracked Hitler’s code. The Times [London], March 27, 2014. URL http://www.thetimes.

co.uk/tto/news/uk/defence/article4046291.ece.

Moody:2015:TLT

[Moo15] Oliver Moody. Turing’s last theory is back in favour. The Times

[London], March 6, 2015. URL http://www.thetimes.co.uk/
tto/science/article4373607.ece.

Mundici:2017:TM

[MS17] Daniele Mundici and Wilfried Sieg. Turing, the mathematician. In Floyd and Bokulich [FB17], chapter 2, pages 39–62. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346.

LCCN QA29.T8 P45 2017. URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-6_2.

Muhlenbein:2009:CIL

[Müh09] Heinz Mühlenbein. Computational intelligence: The legacy of Alan Turing and John von Neumann. In Mumford and Jain [MJ09], pages 23–43. ISBN 3-642-01798-3 (print), 3-642-01799-1 (e-book). LCCN Q342 .C66 2009. URL http://link.springer.

com/chapter/10.1007/978-3-642-01799-5_2.

Murray:1993:MB

[Mur93] J. D. (James Dickson) Murray. Mathematical Biology, volume 19 of Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1993. ISBN 3-540-57204-X. xiv + 767 pp. LCCN QH323.5 .M88 1993.

Murray:2012:ATM

[Mur12] James D. Murray. After Turing: Mathematical modelling in the biomedical and social sciences: From animal coat patterns to brain tumours to saving marriages. In Cooper et al. [CDL12], pages 517–527. ISBN 3-642-30869-4. LCCN QA9.59

.C664 2012. URL http://www.springerlink.com/content/
15884x52041l8xu4/.

Nemeti:2006:CGR

[NA06] István Németi and Hajnal Andréka. Can general relativistic computers break the Turing barrier? In Beckmann et al. [BBLT06], pages 398–412. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA9.59 .C67 2006. URL http://link.springer.com/
chapter/10.1007/11780342_42.

Nanjundiah:2003:ATB

[Nan03] Vidyanand Nanjundiah. Alan Turing and “The Chemical Basis of Morphogenesis”. In Sekimura et al. [SNUM03], pages 33– 44. ISBN 4-431-65960-9 (print), 4-431-65958-7 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/
978-4-431-65958-7_3.

Naur:1986:TTT

[Nau86] Peter Naur. Thinking and Turing’s test. BIT, 26(2): 175–187, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=26&issue=2&spage=175.

Naur:1993:UTU

[Nau93] P. Naur. Understanding Turing’s universal machine — personal style in program description. The Computer Journal, 36(4):351–372, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/36/4/351.full.pdf+html;
http://www3.oup.co.uk/computer_journal/Volume_36/Issue_
04/Vol36_04.body.html#AbstractNaur.

Naughton:2009:PMG

[Nau09] Philippe Naughton. [Prime Minister] Gordon Brown issues apology for ‘inhumane’ treatment of Alan Turing. The Times [Lon-

don], September 11, 2009. URL http://www.thetimes.co.uk/
tto/news/uk/article1945078.ece.

Nerode:2014:MTT

[Ner14] Anil Nerode. Musings on Turing’s thesis. In Downey [Dow14c], chapter 11, pages 386–396. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Newman:1949:NEA

[New49] M. H. A. Newman. A note on electric automatic computing machines. British Medical Journal, 1(4616):1133, June 25, 1949. CODEN BMJOAE. ISSN 1468-5833. URL https://www.jstor.

org/stable/25372598.

Newman:1955:AMT

[New55] M. H. A. (Maxwell Herman Alexander) Newman. Alan Mathison Turing. 1912–1954. Biographical Memoirs of Fellows of the Royal Society, 1:253–263, November 1955. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1955.0019.

Newton:2003:ATS

[New03] David E. Newton. Alan Turing: a study in light and shadow. Xlibris, Philadelphia, PA, USA, 2003. ISBN 1-4010-9080-X. 122 pp. LCCN ????

Newman:2012:ATR

[New12] William Newman. Alan Turing remembered. Communications of the Association for Computing Machinery, 55(12):39–40, December 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nichols:2017:BRC

[Nic17] Tiffany Nichols. Book review: Chris Bernhardt, Turing’s Vision: The Birth of Computer Science. Cambridge, MA: MIT Press, 2016. Pp. 189. ISBN 978-0-262-03454-8. £19.95 (cloth). British Journal for the History of Science, 50(2):366–368, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL https://www.cambridge.org/core/product/
6FBE184804281FFF83D386E3D7902756.

Nofre:2017:BRG

[Nof17] David Nofre. Book review: Giovanni Sommaruga; Thomas Strahm, eds. Turing’s Revolution: The Impact of His Ideas about Computability. Isis, 108(2):486–487, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Normann:2014:HGT

[Nor14] Dag Normann. Higher generalizations of the Turing Model. In Downey [Dow14c], chapter 12, pages 397–433. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Newman:1942:FTC

[NT42] M. H. A. Newman and A. M. Turing. A formal theorem in Church’s theory of types. Journal of Symbolic Logic, 7:28–33, 1942. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886

(electronic). URL https://projecteuclid.org/euclid.jsl/
1183389307.

Numerico:2005:ATI

[Num05] Teresa Numerico. Alan Turing e l’intelligenza delle macchine. (Italian) [Alan Turing and machine intelligence]. FrancoAngeli, Milan, Italy, 2005. ISBN 88-464-6136-3. 208 pp. LCCN

???? URL http://www.loc.gov/catdir/toc/casalini01/
05052807.pdf.

Neary:2012:CSU

[NW12] Turlough Neary and Damien Woods. The complexity of small universal Turing machines: a survey. In Bieliková et al. [BFG^{+}12], pages 385–405. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
f0mk02258j01h22h/.

Odifreddi:2012:MES

[Odi12] Piergiorgio Odifreddi. Un matematico eclettico e stravagante: conferenza su Alan Turing. (Italian) [An eclectic and quirky mathematician: lectures on Alan Turing]. Casagrande, Bellinzona, Switzerland, 2012. ISBN 88-7713-638-3. 46 pp. LCCN ????

OConnell:2003:DAT

[OF03] Henry O’Connell and Michael Fitzgerald. Did Alan Turing have Asperger’s syndrome? Irish Journal of Psychological Medicine, 20(1):28–31, March 2003. ISSN 0790-9667 (print), 2051-6967 (electronic).

Ocasio-Gonzalez:2012:TCE

[OG12] Víctor A. Ocasio-González. Turing computable embeddings and coding families of sets. In Cooper et al. [CDL12], pages 539–548. ISBN 3-642-30869-4. LCCN QA9.59 .C664 2012. URL http://

www.springerlink.com/content/h7l723170r88070l/.

ONeill:2023:RAT

[O’N23] Laura O’Neill. 6 reasons Alan Turing, the father of modern computing and war hero, is a Manchester icon. Web site, June

23, 2023. URL https://secretmanchester.com/icons-alan-
turing/.

ORegan:2012:AT

[O’R12] Gerard (Cornelius Gerard) O’Regan. Alan Turing. In Gerard (Cornelius Gerard) O’Regan, editor, A Brief History of Computing, pages 219–228. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4471-2358-1. LCCN ???? URL http://www.

springerlink.com/content/978-1-4471-2358-3/;   http:/
/www.springerlink.com/content/k243537142501334/.

Ord-Smith:1965:BRB

[OS65] R. J. Ord-Smith. Book review: Irascible Genius (Charles Babbage). The Computer Journal, 7(4):277, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

7/4/277.full.pdf+html. See comments and corrections [Phi65].

Ouyang:1991:TUS

[OS91] Q. Ouyang and Harry L. Swinney. Transition from a uniform state to hexagonal and striped Turing patterns. Nature, 352(6336):610–612, August 15, 1991. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v352/n6336/pdf/
352610a0.pdf.

Orlitsky:2003:AGT

[OSZ03] Alon Orlitsky, Narayana P. Santhanam, and Junan Zhang. Always good Turing: Asymptotically optimal probability estimation. Science, 302(5644):427–431, October 17, 2003. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/302/5644/427.
full.pdf.

Olderog:2012:TVG

[OW12] Ernst-Rüdiger Olderog and Reinhard Wilhelm. Turing und die Verifikation. (German) [Turing and verification]. Informatik Spektrum, 35(4):271–279, August 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://

www.springerlink.com/content/d3v152650r317367/. Special

Issue: Alan Turing.

Owens:2012:ATC

[Owe12] Barbara Boucher Owens. ACM Turing Centenary celebration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):8, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prank:2011:ULT

[PA11a] Rein Prank and Mart Anton. Using a learner- and teacherfriendly environment for Turing machine programming and testing. In Patrick Blackburn, Hans van Ditmarsch, Maria Manzano, and Fernando Soler-Toscano, editors, Tools for Teaching Logic, volume 6680 of Lecture Notes in Computer Science, pages 198–206. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-21349-9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/
content/601p154683474vn1/;   http://www.springerlink.
com/content/978-3-642-21349-6/.

Putchala:2011:MVA

[PA11b] Santosh Putchala and Nikhil Agarwal. Machine vision: an aid in reverse Turing test. AI and Society, 26(1):95–101, February 2011. CODEN AISCEM. ISSN 0951-5666 (print), 1435-5655

(electronic). URL http://www.springerlink.com/content/
h18u3m2766n53751/.

Palm:2013:BTP

[PA13] Günther Palm and Ad Aertsen, editors. Brain theory: proceedings of the first Trieste meeting on brain theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-70913-3. LCCN ????

Papadimitriou:2003:TNA

[Pap03] Christos H. Papadimitriou. Turing: a novel about computation. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-16218-0 (hc.). 284 pp. LCCN PS3616.A58 T87 2003.

Papadimitriou:2012:APA

[Pap12] Christos H. Papadimitriou. Alan and I: a personal account of Alan Turing’s life and impact. Communications of the Association for Computing Machinery, 55(9):42–43, September 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parry:2012:ECD

[Par12] Marc Parry. Early computing’s ‘deal with the devil’. The Chronicle of Higher Education, ??(??):??, July 9, 2012. ISSN 0009-5982 (print), 1931-1362 (electronic). URL http://chronicle.com/

article/article-content/132763/.

Parry:2014:WAT

[Par14] Hannah Parry. Was Alan Turing murdered? New book claims Enigma machine genius didn’t kill himself after all. Daily Mail (UK), December 11, 2014. URL http://www.

dailymail.co.uk/news/article-2870499/Was-Alan-Turing-
murdered-New-book-claims-Enigma-machine-genius-didn-
t-kill-all.html.

Parikh:2017:TCT

[Par17] Rohit Parikh. Is there a Church–Turing Thesis for social algorithms? In Floyd and Bokulich [FB17], chapter 15, pages 339–357. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346. LCCN QA29.T8 P45 2017. URL https://link.

springer.com/chapter/10.1007/978-3-319-53280-6_15.

Patera:2004:AAH

[Pat04] Valeria Patera. Alan’s apple: Hacking the Turing Test. In Teuscher [Teu04a], pages 9–41. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Patera:2007:MAH

[Pat07] Valeria Patera. La mela di Alan: hacking the Turing test: cybertragicommedia. Teatro. Di Renzo, Roma, Italy, 2007. ISBN 88-8323-170-8. 118 pp. LCCN ????

Paulos:1991:BNR

[Pau91] John Allen Paulos. Beyond Numeracy: Ruminations of a Numbers Man. Alfred A. Knopf, New York, NY, USA, 1991. ISBN 0-394-58640-9, 0-685-48163-8. xiii + 285 pp. LCCN QA5 .P38 1991.

Pavlus:2017:NTT

[Pav17] John Pavlus. The new Turing tests. Scientific American, 316 (3):61–62, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
61.html;   http://www.nature.com/scientificamerican/
journal/v316/n3/pdf/scientificamerican0317-61.pdf.

PazSoldan:2003:DTS

[Paz03] Edmundo Paz Soldán. El delirio de Turing. [(Spanish)] The delirium of Turing. Alfaguara. Santillana de Ediciones, La Paz, Bolivia, 2003. ISBN 99905-2-283-9. 313 pp. LCCN MLCM 2007/41947 (P).

Perelgut:1988:TPC

[PC88] S. Perelgut and J. R. Cordy. Turing Plus: a comparison with C and Pascal. ACM SIGPLAN Notices, 23(1):137–143, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

PazSoldan:2006:TD

[PC06] Edmundo Paz Soldán and Lisa Carter. Turing’s delirium. Houghton-Mifflin, Boston, MA, USA, 2006. ISBN 0-618-54139-X. 291 pp. LCCN PQ7820.P39 D4513 2006. URL http://www.

loc.gov/catdir/enhancements/fy0623/2005024726-d.html;

 http://www.loc.gov/catdir/enhancements/fy0668/2005024726-
s.html;   http://www.loc.gov/catdir/enhancements/fy0736/
2005024726-b.html.

Peachey:2019:NFB

[Pea19] Kevin Peachey. New face of the Bank of England’s £50 note is revealed as Alan Turing. BBC News Web site., July 15, 2019. URL

https://www.bbc.com/news/business-48962557. Includes 40

sec video Turing tribute by Bank of England governor Mark Carney.

Pellen:2009:HIH

[Pel09] Luke Pellen. How not to imitate a human being. In Epstein et al. [EBR09], pages 431–446. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_25.pdf.

Petzold:2008:ATG

[Pet08] Charles Petzold. The annotated Turing: a guided tour through Alan Turing’s historic paper on computability and the Turing Machine. Wiley, New York, NY, USA, 2008. ISBN 0-470-22905-5 (paperback). xii + 372 pp. LCCN QA267 .P48 2008.

Petrocelli:2018:BRT

[Pet18] Carla Petrocelli. Book review: The Turing Guide, by Jack Copeland, Jonathan Bowen, Mark Sprevak, and Robin Wilson. Nuncius, 33(1):166–168, ???? 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/journals/10.
1163/18253911-03301015.

Phillips:1965:IG

[Phi65] William Phillips. “Irascible genius”. The Computer Journal, 8 (1):56, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/56.full.pdf+html. See [OS65].

Piccinini:2003:ATM

[Pic03a] Gualtiero Piccinini. Alan Turing and the mathematical objection. Minds and Machines, 13(1):23–48, February 2003. CO-DEN MMACEO. ISSN 0924-6495 (print), 1572-8641 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021348629167.

Piccinini:2003:TRI

[Pic03b] Gualtiero Piccinini. Turing’s rules for the Imitation Game. In Moor [Moo03b], pages 111–120. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.

springer.com/chapter/10.1007/978-94-010-0105-2_5/.

Piccinini:2011:PCT

[Pic11] Gualtiero Piccinini. The physical Church–Turing thesis: Modest or bold? British Journal for the Philosophy of Science, 62(4):733–769, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/62/4/733.full.pdf+html.

Pilous:2012:IWG

[Pil12] Roland Pilous. Die Informationierung der Welt. (German) [The informatization of the world]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, ???? 2012. CODEN SPEKDI. ISSN 0170-2971. URL http://www.

spektrum.de/alias/die-information/die-informationierung-
der-welt/1152086.

Piper:2004:TLC

[Pip04] Fred Piper. Turing Lecture 2004: Cyberworld security — the good, the bad and the ugly. World-Wide Web

site., 2004. URL http://www.bcs.org/BCS/Awards/Events/
TuringLecture/Turing2004/.

Piper:2005:TLC

[Pip05] Fred Piper. Turing Lecture: Cyberworld security — the good, the bad and the ugly. The Computer Journal, 48(2):145–156, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
cgi/reprint/48/2/145;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_48/Issue_02/bxh076.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_48/
Issue_02/pdf/bxh076.pdf.

Pitogo:2014:WAT

[Pit14] Heziel Pitogo. Was Alan Turing murdered? Author Roger Bristow says “yes”. Web site., December 14, 2014. URL https://

www.warhistoryonline.com/war-articles/was-alan-turing-
murdered-author-roger-bristow-says-yes.html.

Pitt:2023:TTT

[Pit23] Lenny Pitt. Turing Tumble is Turing-Complete. Theoretical Computer Science, 948(??):??, February 28, 2023. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304397523000476.

Platt:2009:GT

[Pla09] Charles Platt. The Gnirut Test. In Epstein et al. [EBR09], pages 479–485. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445

2008. URL http://link.springer.com/content/pdf/10.
1007/978-1-4020-6710-5_28.pdf.

Pool:1991:DTD

[Poo91] R. Pool. Did Turing discover how the leopard got its spots? Science, 251(4994):627, February 8, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/251/4994/627.full.pdf.

Pool:1992:DTD

[Poo92] Robert Pool. Did Turing Discover How the Leopard Got Its Spots?: Understanding the way a simple chemical system produces patterns may offer insights into animal development. ACM SIGPLAN Notices, 27(12):28, December 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Porter:2019:TBC

[Por19] Christopher P. Porter. Three books on computability, with a special focus on Turing’s legacy. Historia Mathematica, 46(??):88– 95, February 2019. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086018301265.

Possati:2023:CTP

[Pos23a] Luca M. Possati. Correction: From Turing to Peirce. a semiotic interpretation of computation. Foundations of Science, 28 (4):1175, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-022-09884-8. See [Pos23b].

Possati:2023:TPS

[Pos23b] Luca M. Possati. From Turing to Peirce. a semiotic interpretation of computation. Foundations of Science, 28(4):1085–1110, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09878-6. See correction [Pos23a].

Potgieter:2010:OCA

[PR10] Petrus H. Potgieter and Elemér E. Rosinger. Output concepts for accelerated Turing machines. Natural Computing, 9(4):853– 864, December 2010. CODEN ???? ISSN 1567-7818 (print),

1572-9796 (electronic). URL http://www.springerlink.com/
content/g752243r4v0127ur/.

Parikh:2017:JTB

[PR17] Rohit Parikh and Adriana Renero. Justified true belief: Plato, Gettier, and Turing. In Floyd and Bokulich [FB17], chapter 4, pages 93–102. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346. LCCN QA29.T8 P45 2017.

URL https://link.springer.com/chapter/10.1007/978-3-
319-53280-6_4.

Prawitz:1995:TWT

[Pra95] Dag Prawitz. Turing and Wittgenstein — two perceptions of reality. In Göranzon [Gör95a], pages 187–192. ISBN 3-540-19920-9, 1-4471-3001-4 (e-book). ISSN 1431-0856. LCCN QA75.5-76.95.

Prager:2001:T

[Pra01] John Prager. On Turing. Wadsworth philosophers series. Wadsworth/Thomson Learning, Belmont, CA, USA, 2001. ISBN 0-534-58364-4. 83 pp. LCCN QA29.T8 P73 2001.

Priestley:2010:SOM

[Pri10] Mark Priestley. A Science of Operations: Machines, Logic and the Invention of Programming. History of computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-554-4 (hardcover), 1-84882-555-2 (ebook). ix + 341 pp. LCCN QA76.6 .P737 2010.

Price:2021:GWB

[Pri21] David A. (David Andrew) Price. Geniuses at War: Bletchley Park, Colossus, and the Dawn of the Digital Age. Alfred A.

Knopf, New York, NY, USA, 2021. ISBN 0-525-52154-2 (hardcover), 0-525-52155-0 (e-book). 243 pp. LCCN D810.C88 P75 2021.

Priestley:2024:WSS

[Pri24] Mark Priestley. Working in the scientific state: John Womersley’s early career. IEEE Annals of the History of Computing, 46(3):86–93, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Proudfoot:2004:RRF

[Pro04] Diane Proudfoot. Robots and rule-following. In Teuscher [Teu04a], pages 359–379. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Proudfoot:2005:NIT

[Pro06] Diane Proudfoot. A new interpretation of the Turing Test. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article010113.html.

Proudfoot:2015:WTH

[Pro15] D. Proudfoot. What Turing himself said about the imitation game. IEEE Spectrum, 52(7):42–47, July 2015. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Proudfoot:2017:CM

[Pro17a] Diane Proudfoot. Child machines. In Copeland et al. [CBSW17], chapter 30, pages 315–326. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Proudfoot:2017:TFW

[Pro17b] Diane Proudfoot. Turing and free will: A new take on an old debate. In Floyd and Bokulich [FB17], chapter 13, pages 305–321. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346.

LCCN QA29.T8 P45 2017. URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-6_13.

Proudfoot:2017:TTE

[Pro17c] Diane Proudfoot. The Turing test from every angle. In Copeland et al. [CBSW17], chapter 27, pages 287–300. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Proudfoot:2017:TCI

[Pro17d] Diane Proudfoot. Turing’s concept of intelligence. In Copeland et al. [CBSW17], chapter 28, pages 301–308. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Patitz:2011:EST

[PSS11] Matthew J. Patitz, Robert T. Schweller, and Scott M. Summers. Exact shapes and Turing universality at temperature 1 with a single negative glue. In Cardelli and Shih [CS11a], pages 175–189. CODEN LNCSD9. ISBN 3-642-23637-5 (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/8gj7xm27v4vw4277/.

Qian:2011:ETU

[QSW11] Lulu Qian, David Soloveichik, and Erik Winfree. Efficient Turing-universal computation with DNA polymers. In Yasubumi Sakakibara and Yongli Mi, editors, DNA computing and molecular programming: 16th international conference, DNA 16, Hong Kong, China, June 14–17, 2010, revised selected papers, volume 6518 of Lecture Notes in Computer Science, pages 123– 140. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-18304-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.com/content/978-3-
642-18304-1/;   http://www.springerlink.com/content/
l35415v031r1w021/.

Rakus-Andersson:2003:BBE

[RA03] Elisabeth Rakus-Andersson. The brains behind the Enigma code breaking before the Second World War. In Booss and Høyrup [BH03], pages 83–102. ISBN 3-7643-1634-9 , 0-8176-1634-9.

LCCN QA10.8 .M38 2003. URL http://link.springer.com/
chapter/10.1007/978-3-0348-8093-0_3.

Rakus-Andersson:2004:PBB

[RA04] Elisabeth Rakus-Andersson. The Polish brains behind the breaking of the Enigma code before and during the Second World War. In Teuscher [Teu04a], pages 419–439. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Ross:1995:EET

[RAM95] John Ross, Adam P. Arkin, and Stefan C. Mueller. Experimental evidence for Turing structures. The Journal of Physical Chemistry, 99(25):10417–10419, 1995. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic). URL http://pubs.

acs.org/doi/abs/10.1021/j100025a051.

Randell:1972:ATOa

[Ran72a] Brian Randell. On Alan Turing and the origins of digital computers. Technical report CS-TR 33, Computing Laboratory, University of Newcastle upon Tyne, Newcastle upon Tyne,

UK, 1972. 36 pp. URL http://www.cs.ncl.ac.uk/research/
pubs/books/papers/126.pdf;   http://www.cs.ncl.ac.uk/
research/pubs/trs/abstract.php?number=33.

Randell:1972:ATOb

[Ran72b] Brian Randell. On Alan Turing and the origins of digital computers. In Meltzer and Michie [MM72], pages 3–20. ISBN 0-

85224-234-4. LCCN ???? URL http://www.cs.ncl.ac.uk/
research/pubs/books/papers/126.pdf.

Randell:1976:C

[Ran76] Brian Randell. The COLOSSUS. Technical Report 90, Computing Laboratory, University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1976. 48 pp. URL http://www.

cs.ncl.ac.uk/publications/trs/papers/90.pdf;   http:/
/www.ncl.ac.uk/computing/research/publication/160056.

Randell:2000:TML

[Ran00] Brian Randell. Turing Memorial Lecture: Facing up to faults. The Computer Journal, 43(2):95–106, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www.cs.ncl.ac.uk/research/pubs/articles/
papers/245.pdf;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_43/Issue_02/430095.pdf;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_43/Issue_02/430095.
sgm.abs.html.

Randell:2012:TE

[Ran12] Brian Randell. A Turing enigma. Lecture Notes in Computer Science, 7454:23–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32940-1_3/.

Randell:2017:TOD

[Ran17a] Brian Randell. Turing and origins of digital computers. In Copeland et al. [CBSW17], chapter 8, pages 67–76. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Randell:2017:UR

[Ran17b] Brian Randell. Ultra revelations. In Copeland et al. [CBSW17], chapter 17, pages 175–182. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Rapaport:2003:HPT

[Rap03] William J. Rapaport. How to pass a Turing Test. In Moor [Moo03b], pages 161–184. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_9/.

Regan:2023:WPM

[Reg23] Matthew Regan. When Pac-Man met Turing: The game runs on an emulated 6502 CPU. IEEE Spectrum, 60(6):16–18, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reinitz:2012:TCP

[Rei12] John Reinitz. Turing centenary: Pattern formation. Nature, 482 (7386):464, February 22, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v482/n7386/full/482464a.html.

Rescorla:2017:OTB

[Res17] Michael Rescorla. From Ockham to Turing — and back again. In Floyd and Bokulich [FB17], chapter 12, pages 279–304. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346.

LCCN QA29.T8 P45 2017. URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-6_12.

Richards:2005:TRM

[Ric06] Bernard Richards. Turing, Richards and morphogenesis. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article010109.html.

Richards:2017:RVT

[Ric17] Bernard Richards. Radiolaria: validating the Turing theory. In Copeland et al. [CBSW17], chapter 35, pages 383–388. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Rider:1984:BRM

[Rid84] Robin E. Rider. Book review: A mathematician: Alan Turing: the enigma, by Andrew Hodges. Science, 223(4638):807, February 24, 1984. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
223/4638/807.1.full.pdf.

Rigamonti:1991:TGS

[Rig91] Gianni Rigamonti. Turing: il genio e lo scandalo. (Italian) [Turing: the genius and the scandal]. Flaccovio, Palermo, Italy, 1991. ISBN 88-7804-055-X. 123 pp. LCCN ????

Raphael:2000:GP

[RM00a] Frederic Raphael and Ray Monk, editors. The great philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-92817-6. 469 pp. LCCN B29 .G677 2000.

Raphael:2000:GPS

[RM00b] Frederic Raphael and Ray Monk, editors. The Great Philosophers: From Socrates to Turing. Weidenfeld and Nicolson, London, UK, 2000. ISBN 0-297-64590-0. 469 pp. LCCN B72 .G742

Raphael:2001:GPS

[RM01] Frederic Raphael and Ray Monk, editors. The Great Philosophers: From Socrates to Turing. Phoenix, London, UK, 2001. ISBN 0-7538-1136-7. 570 pp. LCCN ????

Rodrigues:2011:PFL

[RMP11] Luiz Alberto Díaz Rodrigues, Diomar Cristina Mistro, and Sergei Petrovskii. Pattern formation, long-term transients, and the Turing–Hopf bifurcation in a space- and time-discrete predator–prey system. Bulletin of Mathematical Biology, 73 (8):1812–1840, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/s11538-010-9593-5;
 http://link.springer.com/content/pdf/10.1007/s11538-
010-9593-5.pdf.

Robinson:1997:GIP

[Rob97] Daniel N. Robinson. The great ideas of philosophy: Lecture 49: Breaking the code, Alan Turing in the forest of wisdom. Lecture on audio cassette., 1997.

Robinson:2012:SED

[Rob12] Andrew Robinson, editor. The scientists: an epic of discovery. Thames and Hudson, New York, NY, USA, 2012. ISBN 0-500-25191-6. 304 pp. LCCN Q141 .S3712 2012.

Roberts:2017:TBH

[Rob17] Jerry Roberts. The Testery: breaking Hitler’s most secret code. In Copeland et al. [CBSW17], chapter 16, pages 167–174. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Rocchi:2012:WIB

[Roc12] Paolo Rocchi. What is information: Beyond the jungle of information theories. The Computer Journal, 55(7):856–860, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/55/7/856.full.pdf+html. Special Focus on the Cen-

tenary of Alan Turing.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Rosenbloom:2012:CC

[Ros12] Paul S. Rosenbloom. Computing and computation. The Computer Journal, 55(7):820–824, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/820.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Rouly:2018:AIU

[Rou18] Ovi Chris Rouly. Artificial intelligence using P-type unorganised machines. Rutherford Journal, 5(??):??, ???? 2018. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article050107.html.

Ratz:2012:TIM

[RR12] Andreas Rätz and Matthias Röger. Turing instabilities in a mathematical model for signaling networks. Journal of Mathematical Biology, 65(6–7):1215–1244, December 2012. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0495-4.pdf.

Ronald:2003:IES

[RS03] Edmund M. A. Ronald and Moshe Sipper. Intelligence is not enough: On the socialization of talking machines. In Moor [Moo03b], pages 151–160. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_8/.

Restrepo:2004:ISR

[RTM04] Hector Fabio Restrepo, Gianluca Tempesti, and Daniel Mange. Implementation of a self-replicating Universal Turing Machine. In Teuscher [Teu04a], pages 241–269. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Ruelle:2007:MB

[Rue07] David Ruelle. The Mathematician’s Brain: a Personal Tour Through the Essentials of Mathematics and Some of the Great Minds Behind Them. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12982-7. ix + 160 pp.

LCCN QA8.4 .R84 2007. URL http://www.loc.gov/catdir/
enhancements/fy0726/2006049700-d.html;   http://www.
loc.gov/catdir/enhancements/fy0731/2006049700-t.html;
 http://www.loc.gov/catdir/enhancements/fy0734/2006049700-
b.html.

Russ:1989:BRR

[Rus89] Steve Russ. Book review: Rolf Herken (ed.). The Universal Turing Machine: A Half-Century Survey. Oxford: Oxford University Press, 1988. Pp. xiv + 661. ISBN 0-19-853741-7. £55.00. British Journal for the History of Science, 22(4):451–452, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4026930.

Ramm:2012:CTD

[RV12] A. G. Ramm and V. Volpert. Convergence of time-dependent Turing structures to a stationary solution. Acta Applicandae Mathematicae, ??(??):????, ???? 2012. CODEN AAMADV. ISSN 0167-8019 (print), 1572-9036 (electronic). URL http://

www.springerlink.com/content/e0186258417vr605/.

Sale:2004:ATB

[Sal04] Tony Sale. Alan Turing at Bletchley Park in World War II. In Teuscher [Teu04a], pages 441–462. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Saler:2012:ATT

[Sal12] Michael Saler. Alan Turing in three words [reviews of B. Jack Copeland Turing: Pioneer of the information age, 320pp. Oxford University Press. £14.99 (US $21.95). 978-0-19-963979-

3. Andrew Hodges Alan Turing The enigma — The centenary edition, 632pp. Princeton University Press. Paperback, £15.59. (US $24.95). 978-0-691-15564-7. George Dyson, Turing’s Cathedral: The origins of the digital universe, 432pp. Allen Lane. £25. 978-0-713-99750-7, US: Vintage. Paperback, $16.95. 978-1-4000-7599-7]. The Times Literary Supplement, Decem-

ber 28, 2012. URL http://www.the-tls.co.uk/tls/public/
article1179883.ece.

Santini:2005:WSI

[San05] Simone Santini. We are sorry to inform you . . .. Computer, 38 (12):128, 126–127, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saunders:1993:ATB

[Sau93] P. T. Saunders. Alan Turing and biology. IEEE Annals of the History of Computing, 15(3):33–36, July–September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://dlib.computer.org/an/books/an1993/
pdf/a3033.pdf;   http://www.computer.org/annals/an1993/
a3033abs.htm.

Savage:2024:NBT

[Sav24] Neil Savage. News: Beyond Turing: Testing LLMs for intelligence. Communications of the Association for Computing Machinery, 67(9):10–12, September 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.

acm.org/doi/10.1145/3673427.

Saygin:2000:TTY

[SCA00] Ayse Pinar Saygin, Ilyas Cicekli, and Varol Akman. Turing Test: 50 years later. Minds and Machines, 10(4):463–518, November 2000. CODEN MMACEO. ISSN 0924-6495 (print), 1572-

8641 (electronic). URL http://link.springer.com/article/
10.1023/A:1011288000451.

Saygin:2003:TTY

[SCA03] Ayse Pinar Saygin, Ilyas Cicekli, and Varol Akman. Turing Test: 50 years later. In Moor [Moo03b], pages 23–78. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003.

URL http://link.springer.com/chapter/10.1007/978-94-
010-0105-2_2/.

Schnelle:1988:TNN

[Sch88] Helmut Schnelle. Turing naturalized: von Neumann’s unfinished project. In Herken [Her88], pages 539–559. ISBN 0-19-853741-7. LCCN QA267 .U55 1988.

Scheutz:2002:CND

[Sch02] Matthias Scheutz, editor. Computationalism: new directions. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-19478-3. xiii + 209 pp. LCCN QA76 .C54747 2002. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/toc/fy034/2002019570.html.

Schmidhuber:2004:TWW

[Sch04a] Jürgen Schmidhuber. Turing’s war work counts for more than computers. Nature, 429(6991):501, June 3, 2004. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v429/n6991/
full/429501c.html.

Schnelle:2004:NES

[Sch04b] Helmut Schnelle. A note on enjoying strawberries with cream, making mistakes, and other idiotic features. In Teuscher [Teu04a], pages 353–358. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Schmidhuber:2012:TC

[Sch12a] Jürgen Schmidhuber. Turing in context. Science, 336(6089): 1638–1639, June 29, 2012. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/336/6089/1638.3.full.pdf.

Schmidhuber:2012:TKH

[Sch12b] Jürgen Schmidhuber. Turing: Keep his work in perspective. Nature, 483(7391):541, March 28, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v483/n7391/
full/483541b.html.

Schweizer:2012:EFT

[Sch12c] Paul Schweizer. The externalist foundations of a truly total Turing test. Minds and Machines, ??(??):????, ???? 2012. CODEN MMACEO. ISSN 0924-6495 (print), 1572-8641

(electronic). URL http://www.springerlink.com/content/
n25g2468432445m1/.

Simmons:2023:MTM

[SCPC23] Emily S. G. Simmons, Arielle M. Cooley, Joshua R. Puzey, and Gregory D. ConradiSmith. A multigenerational Turing model reproduces transgressive petal spot phenotypes in hybrid Mimulus. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s11538-023-01223-7.

Schoenick:2017:MBT

[SCT^{+}17] Carissa Schoenick, Peter Clark, Oyvind Tafjord, Peter Turney, and Oren Etzioni. Moving beyond the Turing Test with the Allen AI Science Challenge. Communications of the Association for Computing Machinery, 60(9):60–64, September 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://cacm.acm.org/magazines/2017/9/220439/fulltext.

Searle:1995:TTY

[Sea95] John R. Searle. The Turing Test: 55 years later. In Epstein et al. [ERB08], pages 139–150. ISBN 1-4020-6708-9, 1-4020-6710-0 (ebook). LCCN Q335 .P35445 2008.

Searle:2009:TTY

[Sea09] John R. Searle. The Turing Test: 55 years later. In Epstein et al. [EBR09], pages 139–150. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_10.pdf.

Sen:2021:EFH

[Sen21] Paul Sen. Einstein’s fridge: how the difference between hot and cold explains the universe. Scribner, New York, NY, USA, 2021. ISBN 1-5011-8130-0 (hardcover), 1-5011-8132-7 (e-book). xii + 305 pp. LCCN QC311 .S5118 2021.

Settle:2017:ATC

[Set17] Amber Settle. ACM Turing 50th Celebration Conference in China. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3):2–3, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Severance:2012:ATB

[Sev12] Charles Severance. Alan Turing and Bletchley Park. Computer, 45(6):6–8, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Soni:2017:MPH

[SG17] Jimmy Soni and Rob Goodman. A mind at play: how Claude Shannon invented the information age. Simon and Schuster, New York, NY, USA, 2017. ISBN 1-4767-6668-1 (hardcover), 1-4767-6669-X (paperback), 1-4767-6670-3 (e-book). xv + 366 pp. LCCN QA29.S423 S66 2017.

Szpankowski:2018:FSI

[SG18] Wojciech Szpankowski and Ananth Grama. Frontiers of science of information: Shannon meets Turing. Computer, 51 (1):28–38, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.computer.

org/csdl/mags/co/2018/01/mco2018010028-abs.html.

Schonhage:1994:FAM

[SGV94] Arnold Schönhage, Andreas F. W. Grotefeld, and Ekkehart Vetter. Fast algorithms: a multitape Turing machine implementation. B.I. Wissenschaftsverlag, Mannheim, Germany, 1994. ISBN 3-411-16891-9. x + 297 pp. LCCN QA76.9.A43 S34 1994.

Shannon:1954:UTM

[Sha54] Claude E. Shannon. A universal Turing Machine with two internal states. Memorandum 54-114-38, Bell Laboratories, Murray Hill, NJ, USA, May 15, 1954. Published in [dLMSS56].

Shapiro:2000:TLH

[Sha00] Stuart S. Shapiro. Turing’s Legacy: A History of Computing at the National Physical Laboratory, 1945–1995 (review). Technology and Culture, 41(1):172–174, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33499.

Shallit:2009:SCF

[Sha09a] Jeffrey Outlaw Shallit. A second course in formal languages and automata theory. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-86572-7 (hardcover), 0-511-43622-X (e-book), 0-511-43835-4 (e-book), 0-511-43700-5 (e-book), 0-511-80887-9 (e-book). xi + 240 pp. LCCN QA267.3 .S53 2009. URL http:/

/catdir.loc.gov/catdir/toc/ecip0822/2008030065.html.

Shallit:2009:TM

[Sha09b] Jeffrey Outlaw Shallit. Turing machines. In A second course in formal languages and automata theory [Sha09a], chapter 6, pages 174–201. ISBN 0-521-86572-7 (hardcover), 0-511-43622-X (e-book), 0-511-43835-4 (e-book), 0-511-43700-5 (e-book), 0-511-80887-9 (e-book). LCCN QA267.3 .S53 2009. URL http:/

/catdir.loc.gov/catdir/toc/ecip0822/2008030065.html.

Shahrestani:2012:DNP

[Sha12] Mohammad Reza Vaez Shahrestani. Davidson’s no-priority thesis in defending the Turing Test. Procedia — Social and Behavioral Sciences, 32:456–461, 2012. ISSN

1877-0428. URL http://www.sciencedirect.com/science/
article/pii/S1877042812000705. The 4th International Con-

ference of Cognitive Science.

Shah:2014:EAT

[Sha14] Huma Shah. The emotions of Alan Turing: The boy who explained Einstein’s theory of relativity aged 15\frac{1}{2} for his mother. International Journal of Synthetic Emotions, 5(1):23–30, ???? 2014. CODEN ???? ISSN 1947-9093 (print), 1947-9107 (elec-

tronic). URL https://dblp.org/db/journals/ijse/ijse5.
html#Shah14.

Sherratt:2012:TPD

[She12] Jonathan A. Sherratt. Turing patterns in deserts. In Cooper et al. [CDL12], pages 667–674. ISBN 3-642-30869-4. LCCN

QA9.59 .C664 2012. URL http://www.springerlink.com/
content/013717v153771512/.

Strachey:1981:AIP

[SHH81] C. Strachey, Velma R. Huskey, and Harry D. Huskey. Anecdotes: An impossible program; Charles Babbage and Lady Lovelace. Annals of the History of Computing, 3(4):414–415, October/December 1981. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1981/
pdf/a4414.pdf.

Shieber:2004:TTV

[Shi04] Stuart M. Shieber, editor. The Turing test: verbal behavior as the hallmark of intelligence. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-69293-7 (paperback). xi + 346 pp. LCCN Q341 .T874 2004.

Shipley:2012:TCC

[Shi12] Henry Shipley. Turing: Colossus computer revisited. Nature, 483 (7389):275, March 14, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v483/n7389/full/483275b.html.

Shiu:2014:BRA

[Shi14] Peter Shiu. Book review: Alan Turing’s system of logic: the Princeton thesis, Andrew W. Appel (Ed.) Pp. 142. £16.95 (hbk). 2012. ISBN: 978-0-691-15574-6 (Princeton University Press). The Mathematical Gazette, 98(541):186–187, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shute:1987:ATE

[Shu87] Malcolm Shute. Alan Turing — The Enigma of Intelligence, Andrew Hodges. Unwin Paperbacks (Counterpoint), London (1985). Microelectronics Journal, 18(5):45, 1987. CO-DEN MICEB9. ISSN 0026-2692 (print), 1879-2391 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0026269287804378.

Siegelmann:1995:CBT

[Sie95] Hava T. Siegelmann. Computation beyond the Turing limit. Science, 268(5210):545–548, April 28, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/268/5210/545.full.pdf.

Siegfried:2012:MMA

[Sie12] Tom Siegfried. A mind from math: Alan Turing foresaw machines’ potential to mimic brains. Science News (Washington, DC), 181(13):26–28, 2012. CODEN SCNEBK. ISSN 1943-0930.

Siegelmann:2013:TST

[Sie13] Hava T. Siegelmann. Turing on Super-Turing and adaptivity. Progress in Biophysics and Molecular Biology, 113(1):117–126, September 2013. CODEN PBIMAC. ISSN 0079-6107 (print),

1873-1732 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079610713000278. ¡ce:title¿Can Bi-

ology Create a Profoundly New Mathematics and Computation?¡/ce:title¿ ¡ce:subtitle¿Special Theme Issue on Integral Biomathics¡/ce:subtitle¿.

Sieg:2014:SRS

[Sie14] Wilfried Sieg. Step by recursive step: Church’s analysis of effective calculability. In Downey [Dow14c], chapter 13, pages 434–466. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Simpson:2017:BRD

[Sim17a] Edward Simpson. Banburismus revisited: depths and Bayes. In Copeland et al. [CBSW17], chapter 38, pages 415–426. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Simpson:2017:IB

[Sim17b] Edward Simpson. Introducing Banburisms. In Copeland et al. [CBSW17], chapter 13, pages 129–142. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Sebag-Montefiore:2007:EBC

[SM07] Hugh Sebag-Montefiore. Enigma: the battle for the code. Barnes and Noble, New York, NY, USA, 2007. ISBN 0-7607-9118-X. x + 422 + 16 pp. LCCN D810.C88 S43 2007.

Smillie:2002:BRT

[Smi02] K. Smillie. Book review: Turing and the universal machine: the making of the modern computer. IEEE Annals of the History of Computing, 24(2):95, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smith:2005:TMS

[Smi05] S. W. Smith. Turing is from Mars, Shannon is from Venus: computer science and computer engineering. IEEE Security & Privacy, 3(2):66–69, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//ieeexplore.ieee.org/iel5/8013/30742/01423965.pdf;
http://ieeexplore.ieee.org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423965&count=15&index=11.

Smiley:2010:MWI

[Smi10] Jane Smiley. The man who invented the computer: the biography of John Atanasoff, digital pioneer. Doubleday, New York, NY, USA, 2010. ISBN 0-385-52713-6, 0-385-53372-1 (e-book), 1-299-11995-6 (e-book). 246 + 8 pp. LCCN QA76.2.A75 S64 2010. US$25.95.

Smith:2014:HJD

[Smi14] Alvy Ray Smith. His just deserts: a review of four books [book reviews of MR2963548, MR2919681, MR3185259, and Alan Turing’s electronic brain, Oxford University Press, 2012]. Notices of the American Mathematical Society, 61(8):891–895, September 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-

9477 (electronic). URL http://www.ams.org/notices/201408/

rnoti-p891.pdf. Reviews of: Alan Turing: The Enigma: The Centenary Edition, Andrew Hodges, Princeton University Press, May 2012 US$24.95, 632 pages, ISBN-13 978-0-691-15564-7; Alan M. Turing: Centenary Edition, Sara Turing, Cambridge University Press, April 2012, US$31.99, 193 pages, ISBN-13 978-1-107-02058-0; Alan Turing’s Electronic Brain: The Struggle to Build the ACE, the World’s Fastest Computer, B. Jack Copeland and others, Oxford University Press, May 2012, US$33.97, 592 pages, ISBN-13 978-0-19-960915-4; Turing: Pioneer of the Information Age, Jack Copeland, Oxford University Press, January 2013, US$21.95, 224 pages, ISBN-13 978-0-19-963979-3.

Smith:2015:HHB

[Smi15a] Christopher Smith. The hidden history of Bletchley Park: a social and organisational history, 1939–1945. Palgrave Macmillan, New York, NY, USA, 2015. ISBN 1-137-48492-6. vii + 238 pp. LCCN D810.C88 C653 2015. URL http://www.loc.gov/

catdir/enhancements/fy1608/2015015176-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1608/2015015176-
t.html.

Smith:2015:DBP

[Smi15b] Michael Smith. The Debs of Bletchley Park and other stories. Aurum Press, London, UK, 2015. ISBN 1-78131-387-3 (hardcover), 1-78131-388-1. 298 + 8 pp. LCCN D810.S7 S65 2015.

Sekimura:2003:MPF

[SNUM03] Toshio Sekimura, Sumihare Noji, Naoto Ueno, and Philip K. Maini, editors. Morphogenesis and pattern formation in biological systems: experiments and models. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 4-431-65960-9 (print), 4-431-65958-7 (e-book). ???? pp. LCCN ????

Soare:2007:CI

[Soa07] Robert I. Soare. Computability and incomputability. In Cooper et al. [CLS07], pages 705–715. ISBN 3-540-73000-1, 3-540-73001-X (e-book). LCCN QA9.59 .C67 2007. URL http://link.

springer.com/chapter/10.1007/978-3-540-73001-9_75.

Soare:2014:TDC

[Soa14] Robert Irving Soare. Turing and the discovery of computability. In Downey [Dow14c], chapter 14, pages 467–492. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Soare:2016:TCT

[Soa16] Robert I. Soare. Turing computability: theory and applications. Theory and applications of computability. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016 .S63 2016. ISBN 3-642-31932-7, 3-642-31933-5 (ebook). ISSN 2190-619X, 2190-6203. xxxvi + 263 pp. LCCN

QA9.59. URL http://www.springerlink.com/content/978-
3-642-31933-4.

STUG:1983:PUA

[Sof83] Software Tools Users Group, editor. Proceedings: USENIX Association [and] Software Tools Users Group Summer Conference, Toronto 1983, July 1983, Toronto, Ontario, Canada. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1983. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Solomonides:1987:ATS

[Sol87] Tony Solomonides. Alan Turing on stage. Science as Culture, 1(1):43–50, January 1987. ISSN 0950-5431 (print), 1470-1189 (electronic).

Sorenson:2005:HCA

[Sor05] Jonathan P. Sorenson. An honors course on Alan M. Turing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):103–106, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL https://www.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Springer:2012:CWW

[Spr12] Michael Springer. Computer wie wir. (German) [computer like us]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, ???? 2012. CODEN SPEKDI. ISSN

0170-2971. URL http://www.spektrum.de/alias/springers-
einwuerfe/computer-wie-wir/1064593.

Sprevak:2017:TMM

[Spr17] Mark Sprevak. Turing’s model of the mind. In Copeland et al. [CBSW17], chapter 26, pages 277–286. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Siegelmann:1991:TCN

[SS91] Hava T. Siegelmann and Eduardo D. Sontag. Turing computability with neural nets. Applied Mathematics Letters, 4(6):77–80, 1991. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Sommaruga:2015:TRI

[SS15] Giovanni Sommaruga and Thomas (Thomas Adrian) Strahm, editors. Turing’s revolution: the impact of his ideas about computability. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2015. ISBN 3-319-22155-8, 3-319-22156-6 (e-book). xxiv + 329 pp. LCCN QA29.T8 T87 2015. URL

http://link.springer.com/10.1007/978-3-319-22156-4;
http://swbplus.bsz-bw.de/bsz44537344xcov.htm.

Schoning:2012:TAB

[ST12] Uwe Schöning and Wolfgang Thomas. Turings Arbeiten über Berechenbarkeit — eine Einführung und Lesehilfe. (German) [Turing’s work on computability — an introduction and reading help]. Informatik Spektrum, 35(4):253–260, August 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X

(electronic). URL http://www.springerlink.com/content/

74k69jw454g8725t/. Special Issue: Alan Turing.

Salo:2023:PUT

[ST23] Ville Salo and Ilkka Törmä. A physically universal Turing machine. Journal of Computer and System Sciences, 132(??):16– 44, March 2023. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022000022000617.

Stannett:2004:HM

[Sta04] Mike Stannett. Hypercomputational models. In Teuscher [Teu04a], pages 135–157. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Stewart:1990:DTM

[Ste90] I. A. Stewart. The demise of the Turing Machine in complexity theory. Technical report 310, Computing Laboratory, University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1990. 11 pp.

Stewart:1994:MRS

[Ste94] Ian Stewart. Mathematical recreations: a subway named Turing. Scientific American, 271(3):104–107, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-104.pdf.

Sterrett:2000:TTT

[Ste00] Susan G. Sterrett. Turing’s two tests for intelligence*. Minds and Machines, 10(4):541–559, November 2000. CODEN MMACEO. ISSN 0924-6495 (print), 1572-8641 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1011242120015.

Sterrett:2003:TTT

[Ste03] Susan G. Sterrett. Turing’s two tests for intelligence. In Moor [Moo03b], pages 79–97. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.springer.

com/chapter/10.1007/978-94-010-0105-2_3/.

Steinicke:2012:ENO

[Ste12a] Wolfgang Steinicke. EVZI NUPT ODFL. Spektrum der Wissenschaft (German translation of Scientific American), ??(??): ??, ???? 2012. CODEN SPEKDI. ISSN 0170-2971. URL http:

//www.spektrum.de/alias/verschluesselte-botschaften/
evzi-nupt-odfl/1158190.

Sterrett:2012:BTC

[Ste12b] Susan G. Sterrett. Bringing up Turing’s ‘child-machine’. In Cooper et al. [CDL12], pages 703–713. ISBN 3-642-30869-4.

LCCN QA9.59 .C664 2012. URL http://www.springerlink.
com/content/2482525281q47604/.

Sterrett:2017:TIH

[Ste17] Susan G. Sterrett. Turing on the integration of human and machine intelligence. In Floyd and Bokulich [FB17], chapter 14, pages 323–338. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346. LCCN QA29.T8 P45 2017.

URL https://link.springer.com/chapter/10.1007/978-3-
319-53280-6_14.

Strachey:1965:LEI

[Str65] C. Strachey. Letter to the Editor: An impossible program. The Computer Journal, 7(4):313, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Strathern:1999:TCB

[Str99] Paul Strathern. Turing and the computer: The Big Idea. The big idea. Anchor Press/Doubleday, Garden City, NY, USA, 1999. ISBN 0-385-49243-X (paperback). 105 pp. LCCN QA76.17 .S77 1999.

Strick:2011:J

[Str11] Heinz Klaus Strick. Juni 2011. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, ???? 2011. CODEN SPEKDI. ISSN 0170-2971. URL http://www.

spektrum.de/alias/der-mathematische-monatskalender/juni-
2011/1072866.

Strawn:2015:AT

[Str15] George Strawn. Alan Turing. ComputingEdge, 1(3):45–47, March 2015. ISSN 2376-113X. URL http://www.computer.

org/computingedge.

Sumner:2014:DCV

[Sum14] James Sumner. Defiance to compliance: Visions of the computer in postwar Britain. History and Technology, 30(4):309– 333, 2014. CODEN HITEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sutherland:1985:RTM

[Sut85] R. P. Sutherland. Review of Turing’s Man: Western Culture in the Computer Age, by J. David Bolter. Dr. Dobb’s Journal of Software Tools, 10(3):122–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:2012:TC

[Sut12] Ivan Sutherland. The tyranny of the clock. Communications of the Association for Computing Machinery, 55(10):35–36, October 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sutner:2013:UTI

[Sut13] Klaus Sutner. Universality, Turing incompleteness and observers. In Zenil [Zen13], pages 435–449. ISBN 981-4374-29-6. LCCN

QA267.7 .C676 2013. URL http://www.worldscientific.com/
worldscibooks/10.1142/8306. Foreword by Roger Penrose.

Shah:2010:HIM

[SW10] Huma Shah and Kevin Warwick. Hidden interlocutor misidentification in practical Turing tests. Minds and Machines, 20(3):441– 454, August 2010. CODEN MMACEO. ISSN 0924-6495 (print),

1572-8641 (electronic). URL http://www.springerlink.com/
content/ul03135815l85553/.

Swade:2013:CAT

[Swa13a] Doron Swade. Codebreaker: Alan Turing’s Life and Legacy: 21 June 2012–21 October 2013, Science Museum, London. Technology and Culture, 54(4):957–962, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530519.

Swade:2013:ODC

[Swa13b] Doron Swade. Origins of digital computing: Alan Turing, Charles Babbage, and Ada Lovelace. In Zenil [Zen13], pages 23–43. ISBN 981-4374-29-6. LCCN QA267.7 .C676 2013.

URL http://www.worldscientific.com/worldscibooks/10.

1142/8306. Foreword by Roger Penrose.

Swade:2017:TLB

[Swa17] Doron Swade. Turing, Lovelace, and Babbage. In Copeland et al. [CBSW17], chapter 24, pages 249–262. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Swinton:2004:WDG

[Swi04] Jonathan Swinton. Watching the daisies grow: Turing and Fibonacci phyllotaxis. In Teuscher [Teu04a], pages 477–498. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Swinton:2019:ATM

[Swi19] Jonathan Swinton. Alan Turing’s Manchester. Infang Publishing, Manchester, UK, 2019. ISBN 0-9931789-2-8 (paperback), 0-9931789-3-6 (e-book). 204 pp. LCCN ????

Sykes:1992:BHS

[Syk92] Christopher Sykes. BBC Horizon: The strange life and death of Dr. Turing. Documentary video (48m21s)., 1992. URL http://

www.youtube.com/watch?v=Z-sTs2o0VuY. Edited by Jana Ben-

nett.

Szepietowski:1994:TMS

[Sze94] Andrzej Szepietowski. Turing machines with sublogarithmic space, volume 843 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58355-6 (New York), 0-387-58355-6 (Berlin). viii + 114 pp. LCCN QA267 .S987 1994.

Szudzik:2012:TTC

[Szu12] Matthew P. Szudzik. Is Turing’s thesis the consequence of a more general physical principle? In Cooper et al. [CDL12], pages 714–722. ISBN 3-642-30869-4. LCCN QA9.59

.C664 2012. URL http://www.springerlink.com/content/
a52542772l7j5377/.

Turing:2005:MCM

[T^{+}06] Alan Turing et al. The mind and the computing machine. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article010111.html.

Taussky:1956:AMT

[Tau56] Olga Taussky. Alan Mathison Turing 1912–1954. Mathematical Tables and Other Aids to Computation, 10(55):180–181, 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon, New York, NY, USA, 1961–1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Taylor:1998:MCT

[Tay98] R. Gregory Taylor. Motivating the Church–Turing thesis in the twenty-first century. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):228–231, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turing:2012:RSS

[TB12] Alan M. Turing and D. Bayley. Report on speech secrecy system DELILAH, a technical description compiled by A. M. Turing and Lieutenant D. Bayley REME, 1945–1946. Cryptologia, 36 (4):295–340, 2012. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Tan:2018:PMN

[TCP^{+}18] Zhe Tan, Shengfu Chen, Xinsheng Peng, Lin Zhang, and Congjie Gao. Polyamide membranes with nanoscale Turing structures for water purification. Science, 360(6388):518–521, May 4, 2018. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See news story [Zas18].

Tropp:1984:RPC

[TDCKW84] Henry S. Tropp, Philip H. Dorn, Martin Campbell-Kelly, and Eric A. Weiss. Reviews: P. E. Ceruzzi, Reckoners: The Prehistory of the Digital Computer from Relays to the Stored Program Concept, 1935–1945; E. A. Feigenbaum and P. McCorduck, The Fifth Generation; A. Hodges, Alan Turing: The Enigma; S. S. Husson, 25th Anniversary Issue, IBM Journal of Research and Development; M. Ledger, The Case of the ENIAC and The ENIAC’s Muddled History; capsule reviews. Annals of the History of Computing, 6(2):167–187, April/June 1984. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1984/pdf/a2167.pdf;   http://www.computer.
org/annals/an1984/a2167abs.htm.

Tedre:2015:SCS

[Ted15] Matti Tedre. The Science of Computing: Shaping a Discipline. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4822-1770-8, 1-4822-1769-4. xii + 280 pp. LCCN QA76.

Termini:2011:LDM

[Ter11] Settimo Termini. The life, death and miracles of Alan Mathison Turing. In Bartocci et al. [B^{+}11], pages 91–96. ISBN 3-642-

13605-2. LCCN ???? URL http://www.springerlink.com/
content/q73703k262v6780l/.

Teuscher:2004:ATL

[Teu04a] Christof Teuscher, editor. Alan Turing, life and legacy of a great thinker. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Teuscher:2004:TC

[Teu04b] Christof Teuscher. Turing’s connectionism. In Alan Turing: life and legacy of a great thinker [Teu04a], pages 499–529. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. URL http:/

/link.springer.com/chapter/10.1007/978-3-662-05642-4_

21. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Teuscher:2012:FSI

[Teu12] Christof Teuscher. Foreword: Special issue on Alan Turing. Evolutionary Intelligence, 5(1):1–2, March 2012. CODEN ???? ISSN 1864-5909 (print), 1864-5917 (electronic). URL http://www.

springerlink.com/content/1nq052q0glw60q8l/.

Turing:1995:MT

[TG95] Alan Turing and Jean-Yves Girard. La machine de Turing. Seuil, Paris, France, 1995. ISBN 2-02-013571-X. 174 pp. LCCN ???? Translated from English by Julien Basch et Patrice Blanchard.

Therkildsen:1987:GAA

[The87] Tom Therkildsen. En gjennomgang av Alan M. Turing’s artikkel: “On computable numbers, with an application to the Entscheidungsproblem” (1936), og en drøfting av dens datafaglige aspekter. (Norwegian) [A review of Alan M. Turing’s article “On computable numbers, with an application to the Entscheidungsproblem” (1936), and a discussion of its computational aspects]. ???? ????, ????, ????, 1987. Prosjekt ‘UTM’: en universell Turingmaskin.

Thomas:2018:BRC

[Tho18] Wolfgang Thomas. Book review: Chris Bernhardt. Turing’s Vision: The Birth of Computer Science. Isis, 109(1):213–214, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tomayko:1988:AAT

[THWV88] James E. Tomayko, Peter Hilton, Richard Louis Weis, and Alfred Van Sinderen. Anecdotes: Alan Turing in the Home Guard; Overcoming Murphy’s Law; Babbage and the Scheutz Machine at Dudley Observatory. Annals of the History of Computing, 10(2):133–139, April/June 1988. CO-DEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1988/pdf/a2133b.pdf;   http:/
/www.computer.org/annals/an1988/a2133babs.htm.

Tian:2011:TPC

[Tia11] Canrong Tian. Turing patterns created by cross-diffusion for a Holling II and Leslie–Gower type three species food chain model. Journal of Mathematical Chemistry, 49(6):1128–1150, June 2011. CODEN JMCHEG. ISSN 0259-9791 (print),

1572-8897 (electronic). URL http://www.springerlink.com/
content/e34638060128n170/.

Timpson:2004:QCC

[Tim04] Christopher G. Timpson. Quantum computers: the Church– Turing Hypothesis versus the Turing Principle. In Teuscher [Teu04a], pages 213–240. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Tofts:2003:PCI

[TJC03] Darren Tofts, Annemarie Jonson, and Alessio Cavallaro, editors. Prefiguring cyberculture: an intellectual history. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-20145-3. xiv + 322 pp. LCCN T173.8 .P688 2002b.

Turing:1949:DMC

[TNPY49] Alan M. Turing, Max H. A. Newman, Michael Polanyi, and J. Z. Young. Discussion on the mind and the computing machine, 27 October 1949. Web site, 1949. URL https://www.turing.org.

uk/sources/wmays1.html.

Turing:2006:CA

[TP06] Alan Turing and Parabola. A conversation with Alan. Parabola, 31(3):71–73, ???? 2006. ISSN 0362-1596.

Turing:1985:MM

[TPD85] A. M. Turing, H. Putnam, and D. Davidson. Mentes y maquinas. (Spanish) [Minds and machines]. Tecnos, Madrid, Spain, 1985. ISBN ???? 126 pp. LCCN ????

Traiger:2003:MRI

[Tra03] Saul Traiger. Making the right identification in the Turing Test. In Moor [Moo03b], pages 99–110. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://link.

springer.com/chapter/10.1007/978-94-010-0105-2_4/.

Traub:2012:WRC

[Tra12] Joseph Traub. What is the right computational model for continuous scientific problems? The Computer Journal, 55(7):836– 837, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/55/7/836.full.pdf+html. Special Focus on the

Centenary of Alan Turing.

Tropp:1993:CQD

[Tro93] Henry S. Tropp. Comments, queries, and debate: Turing’s visit to the United States. IEEE Annals of the History of Computing, 15(1):10, January–March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://dlib.

computer.org/an/books/an1993/pdf/a1007.pdf. See correc-

tion [Tro95].

Tropp:1995:CQD

[Tro95] Henry S. Tropp. Comments, queries, and debate: Correction: Turing’s visit to the United States. IEEE Annals of the History of Computing, 17(3):6, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://ieeexplore.

ieee.org/iel4/85/8988/00397054.pdf. See [Tro93].

Trudgian:2011:ITM

[Tru11] Timothy Trudgian. Improvements to Turing’s method. Mathematics of Computation, 80(276):2259–2279, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2011-80-
276/S0025-5718-2011-02470-1/home.html;   http://www.
ams.org/journals/mcom/2011-80-276/S0025-5718-2011-02470-
1/S0025-5718-2011-02470-1.pdf;   http://www.ams.org/
mathscinet-getitem?mr=2813359. See [Tur53, Leh70].

Tsang:2019:ATW

[Tsa19] Amie Tsang. Alan Turing will be new face of £50 note, Bank of England says. New York Times, ??(??):??, July 15, 2019. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL https://www.nytimes.com/2019/07/
15/business/alan-turing-50-pound-note.html.

Taussky:1956:CBT

[TT56] Olga Taussky and John Todd. Commuting bilinear transformations and matrices. Journal of the Washington Academy of Sciences, 46(12):373–375, December 1956. CODEN JWASA3. ISSN

0043-0439. URL http://www.jstor.org/stable/24533908.

Turing:1935:ELR

[Tur35a] A. M. Turing. Equivalence of left and right almost periodicity. Journal of the London Mathematical Society, 10:284–285, 1935. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750

(electronic). URL http://turing.ecs.soton.ac.uk/browse.
php/B/10.

Turing:1935:GEF

[Tur35b] A. M. Turing. On the Gaussian error function. Fellowship dissertation, King’s College, Cambridge University, Cambridge, UK,

Turing:1936:CNA

[Tur36] A. M. Turing. On computable numbers, with an application to the Entscheidungsproblem. Proceedings of the London Mathematical Society. Second Series, 42:230–265, 1936. CODEN PLM-TAL. ISSN 0024-6115 (print), 1460-244X (electronic). This is the paper that introduced what is now called the Universal Turing Machine. See correction [Tur37b]. Reprinted in [Haw05].

Turing:1937:CD

[Tur37a] A. M. Turing. Computability and λ-definability. Journal of Symbolic Logic, 2:153–163, 1937. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL http://turing.ecs.

soton.ac.uk/browse.php/B/11.

Turing:1937:CNA

[Tur37b] A. M. Turing. On computable numbers, with an application to the Entscheidungsproblem. A correction. Proceedings of the London Mathematical Society. Second Series, 43:544–546, 1937. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (elec-

tronic). URL http://turing.ecs.soton.ac.uk/browse.php/

B/12. See [Tur36].

Turing:1937:FC

[Tur37c] A. M. Turing. The p-function in λ − K-conversion. Journal of Symbolic Logic, 2(4):164, December 1937. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL http://

homepage.mac.com/a.eppendahl/work/others/pflkc/.

Turing:1938:EG

[Tur38a] A. M. Turing. The extensions of a group. Compositio Mathematica, 5(3):357–367, 1938. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic). URL http://turing.ecs.

soton.ac.uk/browse.php/B/14;   http://www.numdam.org/
item?id=CM_1938__5__357_0.

Turing:1938:FAL

[Tur38b] A. M. Turing. Finite approximations to Lie groups. Annals of Mathematics (2), 39(1):105–111, January 1938. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL

http://turing.ecs.soton.ac.uk/browse.php/B/13.

Turing:1938:SLB

[Tur38c] Alan Mathison Turing. Systems of logic based on ordinals: a dissertation. Ph.D. dissertation, Cambridge University, Cambridge,

UK, 1938. 161–228 + 8 pp. URL http://turing.ecs.soton.

ac.uk/browse.php/B/15. Published by Hodgson & Son, London, UK.

Turing:1939:SLB

[Tur39] A. M. Turing. Systems of logic based on ordinals. Proceedings of the London Mathematical Society. Second Series, 45:161–228, 1939. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic).

Turing:1940:TE

[Tur40] Alan M. Turing. Turing’s treatise on Enigma [the Prof’s book]. Report, ????, Bletchley Park, UK, 1940. 140 pp. URL http:/

/frode.home.cern.ch/frode/crypto/Turing/index.html;
http://www.turing.org.uk/publications/profsbook.html.

Retypeset by Ralph Erskine, Philip Marks, and Frode Weierud from previously-secret material released in April 1996 by the US National Security Agency as reference number NR 964, Box 201, RG 457. The correct title was later found from British sources to be Mathematical Theory of ENIGMA Machine.

Turing:1941:APC

[Tur41a] Alan M. Turing. The applications of probability to cryptography. Report, GCHQ, Cheltenham, UK, 1941. URL http:

//cacm.acm.org/news/148812-turings-rapid-nazi-enigma-
code-breaking-secret-revealed/fulltext;   http://www.
gchq.gov.uk/Press/Pages/turing-papers-released.aspx;
 http://www.nationalarchives.gov.uk/documentsonline/
details-result.asp?Edoc_Id=8835615;   http://www.theregister.
co.uk/2012/04/23/turing_papers_released/;   http://www.
wired.co.uk/news/archive/2012-04/20/turing-papers;
https://arxiv.org/abs/1505.04714. Unclassified and re-

leased 23 April 2012. Date uncertain, but believed to be between April 1941 and April 1942. National Archives of the UK document HW 25/37.

Turing:1941:SR

[Tur41b] Alan M. Turing. On statistics of repetitions. Report, GCHQ, Cheltenham, UK, 1941. URL http://cacm.acm.org/

news/148812-turings-rapid-nazi-enigma-code-breaking-
secret-revealed/fulltext;   http://www.gchq.gov.uk/Press/
Pages/turing-papers-released.aspx;   http://www.nationalarchives.
gov.uk/documentsonline/details-result.asp?Edoc_Id=8835614;
 http://www.theregister.co.uk/2012/04/23/turing_papers_
released/;   http://www.wired.co.uk/news/archive/2012-
04/20/turing-papers;   https://arxiv.org/abs/1505.04715.

Unclassified and released 23 April 2012. Date uncertain, but believed to be between April 1941 and April 1942. National Archives of the UK document HW 25/38.

Turing:1942:UDB

[Tur42a] A. M. Turing. The use of dots as brackets in Church’s system. Journal of Symbolic Logic, 7(4):146–156, December 1942. CO-DEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

URL http://turing.ecs.soton.ac.uk/browse.php/B/16.

Turing:1942:ATR

[Tur42b] Alan M. Turing. Alan Turing’s report from Washington, DC, November 1942. Transcription of the first section of Turing’s report dated 28 November 1942, in the (British) National Archives, box HW 57/10, released from secrecy in October 2004., 1942.

URL http://www.turing.org.uk/sources/washington.html.

Turing:1943:MCZ

[Tur43] A. M. Turing. A method for the calculation of the zeta-function. Proceedings of the London Mathematical Society. Second Series, 48:180–197, 1943. CODEN PLMTAL. ISSN 0024-6115 (print),

1460-244X (electronic). URL http://turing.ecs.soton.ac.
uk/browse.php/B/17.

Turing:1945:PDM

[Tur45] A. M. Turing. Proposal for development in the Mathematics Division of an Automatic Computing Engine (ACE). Report E.882, Executive Committee, National Physical Laboratory, Teddington, UK, 1945. ???? pp. Reprinted in [CD86, pp. 20-105] and [Tur92c, pp. 1–86]. See also [CD77] for an analysis of the design.

Turing:1946:PEC

[Tur46] A. M. Turing. Proposed electronic calculator. Report E882, National Physical Laboratory, Teddington, UK, 1946. 48 pp.

URL http://www.emula3.com/docs/Turing_Report_on_ACE.

pdf. Reprinted in [Tur72].

Turing:1947:LMF

[Tur47] A. M. Turing. Lecture to London Mathematical Society, February 20, 1947. Turing Digital Archive, 1947. URL http://

turing.ecs.soton.ac.uk/browse.php/B/1.

Turing:1948:IM

[Tur48a] A. M. Turing. Intelligent machinery. Report, National Physical Laboratory, Teddington, UK, 1948. ???? pp. Reprinted in [ER68] and [MM69b, pp. 3–23].

Turing:1948:PFT

[Tur48b] A. M. Turing. Practical forms of type theory. Journal of Symbolic Logic, 13:80–94, 1948. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

Turing:1948:REM

[Tur48c] A. M. Turing. Rounding-off errors in matrix processes. Quarterly Journal of Mechanics and Applied Mathematics, 1 (part 3) (1):287–308, September 1948. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic). URL http://turing.ecs.

soton.ac.uk/browse.php/B/18. Reprinted in [Tur92b] with

summary and notes (including corrections). According to Niall Madden [Mad12, page 18], this paper was the first to introduce the term condition number of a matrix, although von Neumann and Goldstine had used a phrase figure of merit for a similar idea in their 1947 paper [vNG47].

Turing:1949:CLR

[Tur49] Alan M. Turing. Checking a large routine. In Anonymous [Ano49], pages 67–69. LCCN ???? URL http://www.

turingarchive.org/browse.php/B/8. A corrected version is

printed in [MJ84]. The original is reprinted in [WCK89, pp. 70– 72].

Turing:1950:CMI

[Tur50a] A. M. Turing. Computing machinery and intelligence. MIND: A Quarterly Review of Pyschology and Philosophy, 59(236): 433–460, October 1950. CODEN MNDDAB. ISSN 0026-4423 (print), 1460-2113 (electronic). URL http://turing.

ecs.soton.ac.uk/browse.php/B/19;   http://turing.ecs.
soton.ac.uk/browse.php/B/9. Reprinted in [FF63, Tur92c].

Turing:1950:PHM

[Tur50b] A. M. Turing. Programmers’ handbook for Manchester electronic computer. University of Manchester, Manchester, UK, 1950.

Turing:1950:WPS

[Tur50c] A. M. Turing. The word problem in semi-groups with cancellation. Annals of Mathematics (2), 52:491–505, 1950. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL

http://turing.ecs.soton.ac.uk/browse.php/B/31.

Turing:1951:LPM

[Tur51a] A. M. Turing. Local programming methods and conventions. In Anonymous [Ano51], page ?? LCCN ???? Reproduced in Part III of the Mathematical Logic volume of the Collected Works

[Tur01b] and in [WCK89, p. 178].

Turing:1951:PHM

[Tur51b] A. M. Turing. Programmers’ handbook for Manchester electronic computer. Mark II. University of Manchester, Manchester, UK,

1951. URL http://turing.ecs.soton.ac.uk/browse.php/B/
32;   http://www.alanturing.net/programmers_handbook/
;   https://spectrum.ieee.org/tech-history/silicon-revolution/
alan-turing-how-his-universal-machine-became-a-musical-
instrument.

Turing:1952:CBM

[Tur52] A. M. Turing. The chemical basis of morphogenesis. Philosophical transactions of the Royal Society of London Series B, Biological sciences, B 237(641):37–72, August 14, 1952. CO-DEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (elec-

tronic). URL http://turing.ecs.soton.ac.uk/browse.php/

B/22. Reprinted in [Man90, pp. 153–197] and [Tur90].

Turing:1953:SCR

[Tur53] A. M. Turing. Some calculations of the Riemann zeta-function. Proceedings of the London Mathematical Society. Third Series, 3(3):99–117, 1953. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://turing.ecs.

soton.ac.uk/browse.php/B/21. See corrections and improve-

ments [Leh70], and [Tru11]. The latter comments: “Turing’s Method has become the standard technique used in modern verification of the Riemann hypothesis.” See also [Leh56].

Turing:1954:SUP

[Tur54] A. M. Turing. Solvable and unsolvable problems. Science News (Penguin Books), 31(??):7–13, February 1954.

Turing:1959:AMT

[Tur59] Ethel Sara Stoney Turing. Alan M. Turing. W. Heffer and Sons, Cambridge, UK, 1959. xiv + 157 pp. LCCN QA29.T8 T8.

Turing:1960:KMD

[Tur60a] A. M. Turing. Kann eine Maschine denken? [Mozet li masina myslit’?] Mit einem Anhang von J. von Neumann: Allgemeine und logische Theorie der Automaten. [Obscaja i logiceskaja teorija avtomatov.]. (German) [Can a machine think? With an appendix by J. von Neumann: General and Logical Theory of Automata]. Staatsverlag für physikalisch-mathematische Literatur, Moscow, USSR, 1960. 112 pp. LCCN ????

Turing:1960:MLM

[Tur60b] Alan M. Turing. Mozhet li mashina myslit? With a supplement by J. von Neumann. Translated from the English by Ju. A. Danilov; edited and with a foreword by S. A. Janovskaja. Gosudarstv. Izdat. Fiz.-Mat. Lit., Moscow, USSR, 1960. 112 pp. LCCN ????

Turing:1965:SLB

[Tur65] Alan M. Turing. Systems of logic based on ordinals. In The undecidable: basic papers on undecidable propositions, unsolvable problems and computable functions [Dav65], pages 154–222. ISBN 0-911216-01-4. LCCN ????

Turing:1969:IM

[Tur69] A. M. Turing. Intelligent machinery. Machine Intelligence, 5: 3–23, 1969. CODEN MACIBX. ISSN 0076-2032, 0541-6418.

Turing:1972:MTO

[Tur72] A. Turing. A. M. Turing’s original proposal for the development of an electronic computer. Report, Division of Computer Science, National Physical Laboratory, Teddington, UK, April 1972. Reprint, with a foreword by D. W. Davies. The National Archives of the UK: DSIR 30/163, page 2.

Turing:1974:PPM

[Tur74] A. M. Turing. ¿Puede pensar una máquina? Departamento de Logica y Filosofia de la Ciencia, Universidad de Valencia, Valencia, Spain, 1974. ISBN 84-600-6348-8. 64 pp. LCCN ???? Translation by Manuel Garrido and Amador Anton from the English original (Mind 59 (1950), 433–460), Quadernos Teorema, No. 1.

Turing:1987:ISS

[Tur87] Alan M. Turing. Intelligence Service: Schriften. (German) [Intelligence Service: Writings]. Brinkmann und Bose, Berlin, Germany, 1987. ISBN 3-922660-22-3. 239 pp. LCCN ???? Edited by Bernhard Dotzler and Friedrich Kittler.

Turing:1990:CBM

[Tur90] A. M. Turing. The chemical basis of morphogenesis. Bulletin of Mathematical Biology, 52(1–2):153–197, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02459572. Reprinted from Philosophical Transactions of the Royal Society (Part B), Vol 237, Pg 37–72, 1953 [Tur52].

Turing:1991:CMT

[Tur91] Alan Turing. Can a machine think? In Ferris and Fadiman [FF91], pages 492–?? ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Turing:1992:M

[Tur92a] A. M. Turing. Morphogenesis. Collected Works of A. M. Turing. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88486-6. xxvi + 131 pp. LCCN ???? With a preface by P. N. Furbank and edited by P. T. Saunders.

Turing:1992:PM

[Tur92b] A. M. Turing. Pure mathematics. Collected Works of A. M. Turing. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88059-3. xxii + 287 pp. LCCN ???? Edited and with an introduction and postscript by J. L. Britton and Irvine John Good. With a preface by P. N. Furbank.

Turing:1992:MI

[Tur92c] Alan Mathison Turing. Mechanical intelligence. Collected works of A. M. Turing. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88058-5. xix + 227 pp. LCCN Q335.5 .T87 1992. Edited and with an introduction by D. C. (Darrel C.) Ince, and with a preface by P. N. Furbank.

Turing:1994:IM

[Tur94] Alan M. Turing. Intelligenza meccanica, volume 268 of Serie Scientifica [Science Series]. Bollati Boringhieri, Turin, Italy, 1994. ISBN 88-339-0880-1. 160 pp. LCCN ???? Translated from the 1992 English original by Gabriele Lolli and Nino Dazzi with the collaboration of Luisa Conte, Translation edited and with an introduction by Lolli.

Turing:1995:LLM

[Tur95a] A. M. Turing. Lecture to the London Mathematical Society on 20 February 1947. M.D. computing: computers in medical practice, 12(5):390–397, ???? 1995. CODEN MDCOE7. ISSN 0724-6811. Reprinted from the Charles Babbage Institute Reprint Series for the History of Computing, Vol 10, Pg 106, 1986 [TWCD86].

Turing:1995:CMI

[Tur95b] Alan M. Turing. Computing machinery and intelligence. In Epstein et al. [ERB08], pages 23–65. ISBN 1-4020-6708-9, 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008.

Turing:1996:IMH

[Tur96] A. M. Turing. Intelligent machinery, a heretical theory. Philosophia Mathematica. Series III, 4(3):256–260, 1996. CO-DEN ???? ISSN 0031-8019 (print), 1744-6406 (electronic).

Turing:1999:TTE

[Tur99] Alan Turing. Turing’s treatise on Enigma. Technical report, CERN, Geneva, Switzerland, 1999. URL http://home.cern.

ch/~frode/crypto/Turing/index.html. This document is re-

typed from the original (undated??) Turing typescript by the editors Ralph Erskine, Philip Marks and Frode Weierud. Chapters 1, 2, and 6 (of 8) are available; the remainder are in preparation.

Turing:19xx:NNN

[Turxx] A. M. Turing. A note on normal numbers. Unpublished, undated, and reprinted first in [Tur92b, pp. 117–119]; handwritten on the back of part of the draft typescript of the only known surviving original of [Tur36]. See [BFP07] for a detailed review and extension of this work., 19xx. URL http://www.turingarchive.

org/viewer/?id=131&title=01a.1.

Turing:2000:ATM

[Tur00] Alan M. Turing. Alan Turing’s Manual for the Ferranti Mk. I, February 13, 2000. 96 pp. URL http://curation.cs.

manchester.ac.uk/computer50/www.computer50.org/kgill/
mark1/progman.aux.html;   http://curation.cs.manchester.
ac.uk/computer50/www.computer50.org/kgill/mark1/RobertTau/

turing.pdf. Manual written in 1951. Transcribed and typeset by Robert S. Thau.

Turing:2001:CWM

[Tur01a] A. M. Turing. Collected works of A. M. Turing. Mathematical logic. Elsevier, Amsterdam, The Netherlands, 2001. ISBN ???? xii + 293 pp. LCCN ???? US$116.00. Edited by R. O. Gandy and C. E. M. Yates. Including prefaces by Solomon Feferman.

Turing:2001:ML

[Tur01b] A. M. Turing. Mathematical logic. Collected Works of A. M. Turing. North-Holland, Amsterdam, The Netherlands, 2001. ISBN 0-444-50423-0. xii + 293 pp. LCCN ???? Edited by the late R. O. Gandy and C. E. M. Yates, Including prefaces by P. N. Furbank, Yates, Solomon Feferman, Andrew Hodges, Jack Good and Martin Campbell-Kelly.

Turing:2001:VNC

[Tur01c] Alan M. Turing. Visit to National Cash Register Corporation of Dayton, Ohio. Cryptologia, 25(1):1–10, January 2001. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). See

[DB04] for the NCR story and brief comments about Turing’s visit there.

Turing:2003:CRS

[Tur03] Alan M. Turing. Critique of “Running Short Cribs on the U. S. Navy Bombe”. Cryptologia, 27(1):44–49, January 2003. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Turing:2004:BS

[Tur04] Alan M. Turing. Bombe and Spider. In Copeland [Cop04], pages 313–335. ISBN 0-19-825079-7 (hardcover), 0-19-825080-0 (paperback). LCCN QA29.T8 E77 2004. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/fy053/2004275594.html. Text

prepared by Ralph Erskine and Philip Marks and Frode Weierud from the only two surviving copies of Turing’s typescript.

Turing:2005:NM

[Tur05a] Alan M. Turing. Notes on memory (1945). In Copeland [Cop05a], page ?? ISBN 0-19-856593-3 (hardcover). LCCN QA75

.A43 2005. URL http://ukcatalogue.oup.com/product/
9780198565932.do;   http://www.oxfordscholarship.com/
oso/public/content/maths/9780198565932/toc.html.

Turing:2005:PEC

[Tur05b] Alan M. Turing. Proposed electronic calculator (1945). In Copeland [Cop05a], page ?? ISBN 0-19-856593-3 (hardcover). LCCN QA75 .A43 2005. URL http://

ukcatalogue.oup.com/product/9780198565932.do;   http:
//www.oxfordscholarship.com/oso/public/content/maths/
9780198565932/toc.html.

Turing:2009:CMI

[Tur09] Alan M. Turing. Computing machinery and intelligence. In Epstein et al. [EBR09], pages 23–65. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_3.pdf.

Turing:2012:AMT

[Tur12] Sara Turing. Alan M. Turing. Cambridge University Press, Cambridge, UK, centenary edition, 2012. ISBN 1-107-02058-1 (hardcover). xxiv + 169 pp. LCCN QA29.T8

T8 2012. URL http://www.cambridge.org/us/academic/
subjects/mathematics/recreational-mathematics/alan-m-
turing-centenary-edition.

Turing:2015:PAT

[Tur15a] Dermot Turing. Prof: Alan Turing decoded: a biography. The History Press, Stroud, Gloucestershire, UK, 2015. ISBN 1-84165-643-7 (print), 0-7509-6524-X (e-book). 319 pp. LCCN QA29.T8 T78 2015.

Turing:2015:AMT

[Tur15b] Sara Turing. Alan M. Turing. Cambridge University Press, Cambridge, UK, centenary paperback edition edition, 2015. ISBN 1-107-52422-9 (paperback). xxiv + 169 pp. LCCN QA29.T8 T8 2015.

Turing:2017:MTT

[Tur17] Sir John Dermot Turing. The man with the terrible trousers. In Copeland et al. [CBSW17], chapter 2, pages 19–30. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Turing:2018:XYZ

[Tur18] Dermot Turing. X, Y and Z: the Real Story of How Enigma Was Broken. The History Press, Gloucestershire, UK, 2018. ISBN 0-7509-8782-0 (hardcover), 0-7509-8967-X (ePub). 319 + 1 pp. LCCN D810.C88 T87 2018.

Turing:2020:CBP

[Tur20] Dermot Turing. The Codebreakers of Bletchley Park: the Secret Intelligence Station That Helped Defeat the Nazis. Arcturus Publishing Limited, London, UK, 2020. ISBN 1-78950-621-2, 1-83857-650-9 (paperback). 256 pp. LCCN ???? With an introduction by Christopher M. Andrew.

Turing:2021:CNA

[Tur21a] Alan Mathison Turing. On computable numbers, with an application to the Entscheidungsproblem (1936). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 6, pages 51–60. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Turing:2021:BMD

[Tur21b] Dermot Turing. The Bombe: The Machine that Defeated Enigma. Arcturus Publishing Limited, London, UK, 2021. ISBN 1-398-81244-7. 64 pp.

Turing:2021:RAT

[Tur21c] Dermot Turing. Reflections of Alan Turing: a relative story. The History Press, Gloucestershire, UK, 2021. ISBN 0-7509-9609-9 (hardcover), 0-7509-9707-9 (ePub e-book). 207 pp. LCCN QA29.T8 T87 2021.

Turing:1999:MLM

[TvN99] A. Turing and J. von Neumann. Mozhet li mashina myslit’?. (Russian) [Can the machine think?]. Nauchno-Izdatel’skij Tsentr “Regulyarnaya i Khaoticheskaya Dinamika”, Izhevsk, Russia, 1999. ISBN ???? 104 pp. LCCN ????

Turing:2005:TWL

[TW05] Alan M. Turing and James H. Wilkinson. The Turing–Wilkinson lecture series (1946–1947). In Copeland [Cop05a], page ?? ISBN 0-19-856593-3 (hardcover). LCCN QA75 .A43 2005. URL

http://ukcatalogue.oup.com/product/9780198565932.do;
 http://www.oxfordscholarship.com/oso/public/content/
maths/9780198565932/toc.html. Edited with an introduction

by B. Jack Copeland.

Turing:2012:TWL

[TW12] Alan M. Turing and James H. Wilkinson. The Turing–Wilkinson lecture series (1946–7) (edited with an introduction by B. Jack Copeland). In Copeland [Cop12a], pages 459–527. ISBN 0-19-960915-2 (paperback). LCCN ???? UK £14.99.

Turing:1986:MTA

[TWCD86] Alan Mathison Turing, Michael Woodger, B. E. Carpenter, and R. W. Doran. A. M. Turing’s ACE report of 1946 and other papers. Charles Babbage Institute reprint series for the history of computing ; 10. Tomash Publishers, Los Angeles, CA, USA, 1986. ISBN 0-262-03114-0. 140 + 1 pp. LCCN QA75 .A1851 1986.

Underwood:2013:ATY

[Und13] Sarah Underwood. The Alan Turing Year leaves a rich legacy. Communications of the Association for Computing Machinery, 56(10):24–25, October 2013. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Unknown:1984:ATE

[Unk84] Unknown. Alan Turing: The Enigma. Cryptologia, 8(4):??, October 1984. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

USENIX:1983:UCPb

[USE83] USENIX Association, editor. USENIX Conference Proceedings, Winter, 1983. San Diego, CA. USENIX, San Francisco, CA, USA, Winter 1983. ISBN ???? LCCN ????

Uchida:2010:SPF

[UST^{+}10] Yasuo Uchida, Makoto Sakamoto, Ayumi Taniue, Ryuju Katamune, Takao Ito, Hiroshi Furutani, and Michio Kono. Some properties of four-dimensional parallel Turing machines. Artificial Life and Robotics, 15(4):385–388, December 2010. CODEN ???? ISSN 1433-5298 (print), 1614-7456 (electronic). URL http:

//www.springerlink.com/content/d72327q6027pj420/.

Vardi:2014:ELW

[Var14] Moshe Y. Vardi. Editor’s letter: Would Turing have passed the Turing Test? Communications of the Association for Computing Machinery, 57(9):5, September 2014. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2017:VIW

[Var17] Moshe Y. Vardi. Vardi’s insights: Would Turing have won the Turing award? Communications of the Association for Computing Machinery, 60(11):7, November 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

cacm.acm.org/magazines/2017/11/222163/fulltext.

Villani:2015:RLQ

[VB15] Cédric Villani and Edmond Baudoin. Les rêveurs lunaires: quatre génies qui ont changé l’histoire. (French) [The lunar dreamers: four geniuses who changed history]. Gallimard, Paris, France, 2015. ISBN 2-07-066593-3 (paperback). 179 + 12 pp. LCCN ????

Boas:2012:TMD

[vEB12] Peter van Emde Boas. Turing machines for dummies: Why representations do matter. In Bieliková et al. [BFG^{+}12], pages 14– 30. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/0623661278802w34/.

Vaizey:2012:TT

[VFR^{+}12] Ed Vaizey, David Ferrucci, Justin Ratner, Nigel Shadbolt, Ian Livingstone, Ariel Eckstein, Kate Russell, and Alan Bundy. Turing test. ITNOW, 54(2):51, June 2012. CODEN ????

ISSN 1746-5710. URL http://itnow.oxfordjournals.org/
cgi/content/abstract/54/2/51?etoc.

Vincenzi:2013:ATP

[Vin13] Massimo Vincenzi. Alan Turing and the poisoned apple. In Emmer [Emm13], pages 255–262. ISBN 88-470-2888-4, 88-470-2889-2 (e-book). LCCN QA10.7 .I485 2013. URL http://link.

springer.com/chapter/10.1007/978-88-470-2889-0_27.

vonLunen:2013:BRA

[vL13] Alexander von Lünen. Book reviews: Andrew W. Appel (ed.), Alan Turing’s System of Logic: The Princeton Thesis. Princeton and Oxford: Princeton University Press, 2012. Pp. xv + 142. ISBN 978-0-691-15574-6. £16.95 (hardback). British Journal for the History of Science, 46(4):728–729, December 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

vonNeumann:1947:NIM

[vNG47] John von Neumann and Herman H. Goldstine. Numerical inverting of matrices of high order. Bulletin of the American Mathematical Society, 53(11):1021–1099, November 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X

(electronic). URL http://projecteuclid.org/euclid.bams/

1183511222. See [GvN51] for Part II. Reprinted in [Tau61a, vol. 5, pp. 479–557].

Voss:2013:IBE

[Vos13] G. S. Voss. ‘It is a beautiful experiment’: queer(y)ing the work of Alan Turing. AI and Society, page ??, November 2013. CODEN AISCEM. ISSN 0951-5666 (print), 1435-5655 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00146-013-0517-4.

Wallace:1995:AC

[Wal95] Richard S. Wallace. The anatomy of A.L.I.C.E. In Epstein et al. [ERB08], pages 181–210. ISBN 1-4020-6708-9, 1-4020-6710-0 (ebook). LCCN Q335 .P35445 2008.

Wallace:2009:AC

[Wal09] Richard S. Wallace. The anatomy of A.L.I.C.E. In Epstein et al. [EBR09], pages 181–210. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335

.P35445 2008. URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6710-5_13.pdf.

Warwick:2012:ALT

[War12] Kevin Warwick. Not another look at the Turing test! In Bieliková et al. [BFG^{+}12], pages 130–140. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/
content/p6w42015w2t04858/.

Watt:1995:CPT

[Wat95] Stuart Watt. Can people think? Or machines? In Epstein et al. [ERB08], pages 301–318. ISBN 1-4020-6708-9, 1-4020-6710-0 (ebook). LCCN Q335 .P35445 2008.

Watt:2009:CPT

[Wat09] Stuart Watt. Can people think? Or machines? In Epstein et al. [EBR09], pages 301–318. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book).

LCCN Q335 .P35445 2008. URL http://link.springer.com/
content/pdf/10.1007/978-1-4020-6710-5_18.pdf.

Watson:2012:CGP

[Wat12a] Ian Watson. The computer gets personal. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 125–159. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.

springer.com/chapter/10.1007/978-3-642-28102-0_7.

Watson:2012:CBB

[Wat12b] Ian Watson. Computers and big business. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 89–103. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.

springer.com/chapter/10.1007/978-3-642-28102-0_5.

Watson:2012:CGW

[Wat12c] Ian Watson. Computers go to war. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 51–88. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.

springer.com/chapter/10.1007/978-3-642-28102-0_4.

Watson:2012:DCa

[Wat12d] Ian Watson. The dawn of computing. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 9–40. ISBN 3-642-28101-X, 3-642-28102-8 (ebook). LCCN QA76.17 .W38 2012eb. URL http://link.

springer.com/chapter/10.1007/978-3-642-28102-0_2.

Watson:2012:DPH

[Wat12e] Ian Watson. Deadheads and propeller heads. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 105–123. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.

springer.com/chapter/10.1007/978-3-642-28102-0_6.

Watson:2012:DCb

[Wat12f] Ian Watson. Digital consciousness. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 307–329. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.springer.

com/chapter/10.1007/978-3-642-28102-0_14.

Watson:2012:DU

[Wat12g] Ian Watson. Digital underworld. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 259–283. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.springer.

com/chapter/10.1007/978-3-642-28102-0_12.

Watson:2012:D

[Wat12h] Ian Watson. Dotcom. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 183–200. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17

.W38 2012eb. URL http://link.springer.com/chapter/10.
1007/978-3-642-28102-0_9.

Watson:2012:I

[Wat12i] Ian Watson. Introduction. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 1– 7. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17

.W38 2012eb. URL http://link.springer.com/chapter/10.
1007/978-3-642-28102-0_1.

Watson:2012:MLG

[Wat12j] Ian Watson. Machines of loving grace. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 285–306. ISBN 3-642-28101-X, 3-642-28102-8 (e-book).

LCCN QA76.17 .W38 2012eb. URL http://link.springer.

com/chapter/10.1007/978-3-642-28102-0_13.

Watson:2012:MM

[Wat12k] Ian Watson. Marvelous machines. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 41–49. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.springer.

com/chapter/10.1007/978-3-642-28102-0_3.

Watson:2012:SC

[Wat12l] Ian Watson. The second coming. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 201–234. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.springer.

com/chapter/10.1007/978-3-642-28102-0_10.

Watson:2012:UMD

[Wat12m] Ian Watson. The universal machine: from the dawn of computing to digital consciousness. Copernicus Books, New York, NY, USA, 2012. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). xiv + 353 pp. LCCN QA76.17 .W38 2012eb.

Watson:2012:WW

[Wat12n] Ian Watson. Weaving the Web. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 161–181. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17 .W38 2012eb. URL http://link.springer.

com/chapter/10.1007/978-3-642-28102-0_8.

Watson:2012:W

[Wat12o] Ian Watson. Web 2.0. In The universal machine: from the dawn of computing to digital consciousness [Wat12m], pages 235–258. ISBN 3-642-28101-X, 3-642-28102-8 (e-book). LCCN QA76.17

.W38 2012eb. URL http://link.springer.com/chapter/10.
1007/978-3-642-28102-0_11.

Witzany:2012:TFC

[WB12] Guenther Witzany and František Baluška. Turing: a formal clash of codes. Nature, 483(7391):541, March 28, 2012. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v483/n7391/
full/483541c.html.

Baker:2017:TTM

[WBM17] Thomas Woolley, Ruth Baker, and Philip Maini. Turing’s theory of morphogenesis. In Copeland et al. [CBSW17], chapter 34, pages 373–382. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Williams:1989:EBC

[WCK89] Michael R. Williams and Martin Campbell-Kelly, editors. The Early British Computer Conferences, volume 14 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-23136-0. xvi + 508 pp. LCCN QA76.17 .E171 1989.

Webster:2012:ATU

[Web12] Craig S. Webster. Alan Turing’s unorganized machines and artificial neural networks: his remarkable early work and future possibilities. Evolutionary Intelligence, 5(1):35–43, March 2012. CODEN ???? ISSN 1864-5909 (print), 1864-5917

(electronic). URL http://www.springerlink.com/content/
7n140j28g88847pw/.

Wegner:2012:EC

[Weg12] Peter Wegner. The evolution of computation. The Computer Journal, 55(7):811–813, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/55/7/811.full.pdf+

html. Special Focus on the Centenary of Alan Turing.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a4348.pdf;   http://www.computer.
org/annals/an1988/a4348abs.htm.

Welland:2002:TLS

[Wel02] Mark E. Welland. Turing Lecture 2002: Smaller, faster, better — but is it nanotechnology? World-Wide Web

site., 2002. URL http://www.bcs.org/BCS/Awards/Events/
TuringLecture/Turing2002/.

Wells:2004:CST

[Wel04] Andrew J. Wells. Cognitive science and the Turing Machine: an ecological perspective. In Teuscher [Teu04a], pages 271–292. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Welch:2006:NDH

[Wel06a] P. D. Welch. Non-deterministic halting times for Hamkins-Kidder Turing Machines. In Beckmann et al. [BBLT06], pages 571–574. ISBN 3-540-35466-2, 3-540-35468-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.59 .C67

2006. URL http://link.springer.com/chapter/10.1007/
11780342_58.

Wells:2006:RCC

[Wel06b] Andrew Wells. Rethinking cognitive computation: Turing and the science of the mind. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, England, 2006. ISBN 1-4039-1161-4 (hardcover), 1-4039-1162-2 (paperback). xx + 265 pp. LCCN Q342 .W46 2006. URL http://

catdir.loc.gov/catdir/enhancements/fy0625/2005051170-
b.html;   http://catdir.loc.gov/catdir/enhancements/
fy0625/2005051170-d.html;   http://catdir.loc.gov/catdir/
enhancements/fy0625/2005051170-t.html.

Welch:2012:SRA

[Wel12] P. D. Welch. Some reflections on Alan Turing’s centenary. European Mathematical Society. Newsletter, 85:32–38, 2012. ISSN 1027-488X.

Welch:2014:TMM

[Wel14] P. D. Welch. Transfinite machine models. In Downey [Dow14c], chapter 15, pages 493–529. ISBN 1-107-04348-4 (hardcover), 1-107-63858-5 (paperback), 1-107-33857-3 (e-book). LCCN ????

Whitemore:1987:BCa

[WH87a] Hugh Whitemore and Andrew Hodges. Breaking the code. Amber Lane, Oxford, UK, 1987. ISBN 0-906399-80-7. 80 pp. LCCN PR6073.H577 B73 1987. Play based on the book Alan Turing, the Enigma, by Andrew Hodges.

Whitemore:1987:BCb

[WH87b] Hugh Whitemore and Andrew Hodges. Breaking the Code. Fireside Theatre, Garden City, NY, USA, 1987. ISBN ???? 112 + 4 pp. LCCN ???? Play based on the book Alan Turing, the Enigma, by Andrew Hodges.

Whitemore:1988:BC

[WH88a] Hugh Whitemore and Andrew Hodges. Breaking the code. S. French, New York, NY, USA, 1988. ISBN 0-573-69030-8. 114 + 2 pp. LCCN PR6073.H577 B74 1988.

Whitemore:1988:BCP

[WH88b] Hugh Whitemore and Andrew Hodges. Breaking the code: a play. French, London, UK, 1988. ISBN 0-573-01656-9. 64 pp. LCCN ????

Whalen:2009:CBT

[Wha09] Thomas E. Whalen. A computational behaviorist takes Turing’s Test. In Epstein et al. [EBR09], pages 343–357. ISBN 1-4020-9624-0 (paperback), 1-4020-6708-9 (hardcover), 1-4020-6710-0 (e-book). LCCN Q335 .P35445 2008. URL http://link.

springer.com/content/pdf/10.1007/978-1-4020-6710-5_21.
pdf.

Whitemore:1987:BCS

[Whi87] Hugh Whitemore. Breaking the code: The story of Alan Turing,

1987.

Whitemore:1991:WAA

[Whi91] Hugh Whitemore. Writing about Alan Turing. The Mathematical Intelligencer, 13(4):26–27, Fall 1991. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://link.

springer.com/article/10.1007/BF03028338.

Whittle:2012:NCC

[Whi12] P. Whittle. A natural channel coding for the finite and infinite time axes. The Computer Journal, 55(7):788–798, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/

55/7/788.full.pdf+html. Special Focus on the Centenary of Alan Turing.

Whitty:2017:DE

[Whi17] Robin Whitty. Decidability and the Entscheidungsproblem. In Copeland et al. [CBSW17], chapter 37, pages 405–414. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Wiedermann:2012:TCM

[Wie12] Jiří Wiedermann. Towards computational models of artificial cognitive systems that can, in principle, pass the Turing test. In Bieliková et al. [BFG^{+}12], pages 44–63. CODEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
253264926413825k/.

Wilkinson:1971:SCN

[Wil71] J. H. Wilkinson. Some comments from a numerical analyst. Journal of the ACM, 18(2):137–147, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1980:TWN

[Wil80] J. H. Wilkinson. Turing’s work at the National Physical Laboratory and the construction of Pilot ACE, DEUCE, and ACE. In Metropolis et al. [MHR80], pages 101–114. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Williams:1997:HCT

[Wil97] Michael R. Williams. A History of Computing Technology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, second edition, 1997. ISBN 0-8186-7739-2 (hardcover). xi + 426 pp. LCCN QA76.17 .W55 1997.

URL http://catalogue.bnf.fr/ark:/12148/cb37514972q.

Wilson:2010:BRA

[Wil10] Kevin A. Wilson. Book review: The annotated Turing by Charles Petzold, Publisher Wiley, 2008. ACM SIGACT News, 41(2): 22–26, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea makers: personal perspectives on the lives and ideas of some notable people. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (e-book), 1-57955-011-8. 250 (est.) pp. LCCN

Q141 .W678562 2016. URL http://www.wolfram-media.com/
products/idea-makers.html.

Wolfram:2017:CT

[Wol17] Stephen Wolfram. A century of Turing. In Copeland et al. [CBSW17], chapter 5, pages 43–48. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Wright:2016:RST

[Wri16] John Wright. A recursive solution for Turing’s H-M factor. Cryptologia, 40(4):327–347, 2016. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Wollkind:2000:CTP

[WS00] David J. Wollkind and Laura E. Stephenson. Chemical Turing pattern formation analyses: Comparison of theory with experiment. SIAM Journal on Applied Mathematics, 61(2): 387–431, 2000. CODEN SMJMAP. ISSN 0036-1399 (print),

1095-712X (electronic). URL http://epubs.siam.org/sam-
bin/dbq/article/32621.

Warwick:2016:TIG

[WS16] Kevin Warwick and Huma Shah. Turing’s imitation game: conversations with the unknown. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-05638-1 (hardcover), 1-107-29723-0 (e-book). vii + 195 pp. LCCN Q341 .W37 2016.

Wiedermann:2006:STP

[WTP^{+}06] Jiří Wiedermann, Gerard Tel, Jaroslav Pokorný, Mária Bieliková, and Július Štuller, editors. SOFSEM 2006: Theory and Practice of Computer Science: 32nd Conference on Current Trends in Theory and Practice of Computer Science, Merin, Czech Republic, January 21–27, 2006. Proceedings, volume 3831 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-31198-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.springerlink.com/openurl.asp?
genre=issue&issn=0302-9743&volume=3831.

Whitty:2017:ITM

[WW17] Robin Whitty and Robin Wilson. Introducing Turing’s mathematics. In Copeland et al. [CBSW17], chapter 36, pages 391– 404. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Wei:2012:TSS

[WWG12] Meihua Wei, Jianhua Wu, and Gaihui Guo. Turing structures and stability for the 1-D Lengyel–Epstein system. Journal of Mathematical Chemistry, 50(9):2374–2396, October 2012. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/
10.1007/s10910-012-0037-3;   http://www.springerlink.
com/content/4865417834u11n37/.

Yang:2012:AIE

[Yan12] Xin-She Yang. Artificial intelligence, evolutionary computing and metaheuristics: in the footsteps of Alan Turing, volume 427 of Studies in computational intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-29693-9, 3-642-29694-7 (e-book). ISSN 1860-949X (print), 1860-9503 (electronic). xix + 794 pp. LCCN Q335 .A78 2013.

Yao:2003:CPC

[Yao03] Andrew Chi-Chih Yao. Classical physics and the Church–Turing Thesis. Journal of the ACM, 50(1):100–105, January 2003. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yapp:2012:TAW

[Yap12] Chris Yapp. The Turing AI wine taster. ITNOW, 54(4): 52–53, December 2012. CODEN ???? ISSN 1746-5710.

URL http://itnow.oxfordjournals.org/content/54/4/52.
abstract.html?etoc.

Zabell:1995:ATC

[Zab95] S. L. Zabell. Alan Turing and the Central Limit Theorem. American Mathematical Monthly, 102(6):483–494, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zabell:2012:CAM

[Zab12] Sandy Zabell. Commentary on Alan M. Turing: The applications of probability to cryptography. Cryptologia, 36(3):191–214, 2012. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). See [Tur41a].

Zabell:2017:ATA

[Zab17] Sandy Zabell. Alan Turing and the applications of probability to cryptography. 60-minute video of lecture at Rutgers Univer-

sity., April 4, 2017. URL https://www.youtube.com/watch?v=
vQKMdHnCgrs.

Zastrow:2018:WFI

[Zas18] Mark Zastrow. Water filter inspired by Alan Turing passes first test. Membrane’s structure predicted in mathematician’s lone biology paper. Nature News Web site., May 3,

2018. URL https://www.nature.com/articles/d41586-018-

05055-7. See [TCP^{+}18].

Zdenek:2003:PLT

[Zde03] Sean Zdenek. Passing Loebner’s Turing Test: a case of conflicting discourse functions. In Moor [Moo03b], pages 121–144. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003.

URL http://link.springer.com/chapter/10.1007/978-94-
010-0105-2_6/.

Zenil:2013:CUU

[Zen13] Hector Zenil, editor. A computable universe: understanding and exploring nature as computation. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2013. ISBN 981-4374-29-6. xliv + 810 pp. LCCN

QA267.7 .C676 2013. URL http://www.worldscientific.com/
worldscibooks/10.1142/8306. Foreword by Roger Penrose.

Ziegler:2009:PRC

[Zie09] Martin Ziegler. Physically-relativized Church–Turing Hypotheses: Physical foundations of computing and complexity theory of computational physics. Applied Mathematics and Computation, 215(4):1431–1447, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300309004226.

Ziliak:2008:CSS

[ZM08] Stephen Thomas Ziliak and Deirdre N. McCloskey. The Cult of Statistical Significance: How the Standard Error Costs Us Jobs, Justice, and Lives. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 2008. ISBN 0-472-07007-X (cloth), 0-472-05007-9 (paperback). xxiii + 321 pp.

LCCN HB137 .Z55 2008. URL http://www.loc.gov/catdir/
enhancements/fy0809/2007035401-d.html.